A Complete Bibliography of Publications in Experimental Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

26 February 2025 Version 1.33

Title word cross-reference

(((X^{2} − P)^{2} − Q)^{2} − R)^{2} − S^{2} [Bre08a]. (−2, 3, n) [GK12]. (1, 2) [EF18]. (1 − q)(1 − q^{2})(1 − q^{3}) ⋯ (1 − q^{n−1})(1 − q^{n}) [BU23]. (2, 3, 6) [Sch06]. (4k + 2) [AJK08]. {−1, 0, 1} [SS06]. 0 [Cum04, Har15]. 1 [BR18, BM00, CF95, Cum04, Fin92]. 1/2 [Car03]. 1/π [Gui06a, Gui06b]. 1/π^{2} [AG12, Gui06b]. 100 [BCVZ02]. 10651 [MR95]. 10^{19} [AD07]. 10^{27} [RT16]. 11 [BLP09, BGKT05, Lep93, LM19]. 1132 [KP08]. 12 [GP94, LM19]. 128 [JM14, Elk01]. 15 [BGKT05, DL92, KO95]. 16 [EMRSÉ14, MTC14]. 19 [CHLM97]. 191 [HPP08]. 1 < n < 10 [OM08]. 1: 4 [Kra94]. 2 [AT00, AMM08, BLP09, CD09, Elk01, EL14, KdC05, KV17, Kat10b, Man18, Nic18, RW11, Van17]. 23 [DGKMY15]. 24 [CP03]. 2n [Kie21]. 2 × 2 [CILO19, LR17]. 3 [ABL17b, AT00, Ban06, BFG^{+}13, BL00, Bur04, But05, CLLM07, CLT06, CGHN00, Der15, DBG19, EM02, FMP04, Fuk11, GHL19, HS15, HIK^{+}16, Kat10b, KY17, LR07, Sha19, Tan09, Ume09, Web97]. 32 [BN97, CH08a]. 3630 [BCVZ02]. 3n + 1 [LSW99]. 3x + 1 [AL95, BM98, Hew21, Nic18]. 3x + d [BM98]. 4 [Con06, JK01, RT00, SZ04]. 47 [Tak14]. 4 × 4 [CD21]. 5 [De 15a, HS07, Kat10b]. 6 [ALT^{+}12, FP92, FP93, FPDH98, HP08]. 6, 6 [Nil03]. 7 [ALT^{+}12, BGKT05]. 71 [Tak14]. 8 [ALT^{+}12, BN97, Bur07, FJ96]. 8.875 · 10^{30} [HP13]. 89 [Tak14]. 9 [MP01]. > 1 [ABG^{+}21]. [0.5, 1) [Hew21]. ^{n} [BDGS23]. _{2} [GM18]. _{3} [GMY21]. A [Esc14, GK12]. a^{3} + b^{3} = c^{p} [Kra98]. A_{3}(2) [AB98]. A_{5} [FP92]. A_{6} [FP93]. abc [Nit93]. Aut(F_{2}) [– Dok00]. ax + by ≡ 1 [BHK05]. B_{2}[g] [HP18]. modℓ [Die04a]. modp [Muk21]. C^{2} [BFK^{+}92]. D [Esc14, BGKMI01, HK02b, Har15]. D_{4} [CCEK07]. Δ_{2,n} [Yun22]. DW (5, 2) [Pra05]. E_{6(−26)} [JN06]. E_{6(6)} [JN06]. E_{8} [GMR20]. ℓ [DJP^{+}05]. ℓ > 5 [Die04a]. End(A) = Z [Die02]. F [HHRS21]. f(2, 3, 4) [DR08]. F (2, 9) [Hol95]. F_{11} [MSS15]. F_{2} [– Dok00]. F_{4} [Pur18]. Fi_{23} [GHL19]. g [GPS18]. Γ [Oud11]. Γ_{0}(4) [BFM13]. GL(3, Z) [AAC98]. GL_{2}(F_{p}) [Lip21]. GL_{24}(Q) [Neb96]. H(q, n) [KO95]. h^{⋆} [BL21]. H_{4} [Alv08]. j(z) [BK03]. J_{0}(N) [Web97]. J_{3} [SW95, Lee04]. K [ST05, Tan24]. K3 [EOT11, EJ22, Shi16, Sta23, Jai09]. K = 2 [Tos22]. K^{3} [SK11]. L [ABG^{+}21, ACF^{+}21, BDM^{+}18, BBW17, Boy98, DFK04, DD03a, DG21, DPVZ22, Dok04, Dum01, FK10, FK12, Fer92, Hir00, Mil06, Oma01, Ots15, PT08, Tan09, OB19]. L(1/2, χ_{d}) [AR12]. L^{2} [GGM11]. L^{p} [Joh08]. Λ(n) [PR17b]. Λ(n)/n [PR17b]. log 2 [Bri01]. M_{11} [BDL96]. M_{12} [CEM06]. M_{13} [CEM06]. M_{2,n} [BCGY23]. M_{22} [SW93]. M_{24} [Gra96, Pfe97, EOT11]. C^{2} [BZ16]. F_{2} [Lom02]. F_{q} [GS05]. P^{2} [BDM07]. P^{m} × P^{n} [AB09]. Q [Del01]. T^{3} [GMK19]. Z_{5}^{3} [Bir22]. Z_{2} [AD21]. Z_{29} [KM20]. Z_{31} [KM20]. C0 [BA02]. R^{2} [AL03]. K [Tar92]. M_{g} [BRTC23]. O(1, 2) [AB09]. ∫↕_{3} [GL15]. μ [War11]. μ ≥ τ [Sha19]. n [BRTC23, BMRS19, CC07, GSS16, KO95, OM08, Rei96]. n + 3 [BDP16]. N = 14 [MT15]. N = 4 [HM21]. n ≤ 6 [Oha08]. n × 2 [LR17]. O(3, 1) [FWW02]. O_{6} [OS22a]. Out(F_{n}) [Oha08]. p [AAC98, BCM21b, BBF12, Boy94, BGMR06, CGM15, FI18, GM18, PR03, WS98]. p^{8} [LVL22]. PGL(3, C) [BFG^{+}13, FPUP22]. π [AKP03, BBC^{+}12, Chu05, Gan14]. π(x) [Gar02]. π/\sqrt{3} [Bri01]. ± [Kat10b]. ±1 [BC99]. PSL(2, Z) [Ave10b]. PSL(2, Z) [HR92a, CP03, Cum04]. Q [Fer96, APS04, FK09, KS01, Chi24, vB22, BU23, Hik03, JST16, MGO22, XZ22]. Q(\sqrt{5}) [CD09, Dem05]. Q(T) [Gra96]. q − 1 [Har15]. R [GP06, Rim92]. R^{3} [CLW15]. S [Wil06, CFH^{+}15]. S^{4} [BB22]. S_{5} [FPDH98]. S_{n} [War11, Yun22]. SL(2, C) [FGT10]. SL(3, C) [FGK^{+}16]. SL(3, Z) [LR11]. SL(3, Z) [AM92, vGvdKTV97]. SL(4, R) [CLT06]. SL(4, Z) [LT18]. SL_{2} [LR92]. SL_{2}[Z/pZ] [RS19]. SL_{3}(F_{2}) [APS04]. SL_{3}(Z) [KMN24]. sM_{⟨3⟩} [Con21]. Sp_{2n}((C)) [EGP^{+}12]. \sqrt{d} [Duj01]. \sqrt{|[Cyrillic − sha]|} [Qua06]. ⋆ mod ō(h̄^{4}) [BK22]. SU(2, 1) [GK21]. SU(3) [Coq20]. SU(N + 1) [GdlI07]. T [HHRS21, HM12, NT15, She09]. θ [DdLV14]. × [DH22]. u [GMK19]. W [KO95, Fuk11]. W (2, 6) [KP08]. W (2n + 1) [WPW97]. X_{0}^{+}(N) [CB09]. X_{0}^{+}(p) [Gal99]. x_{1} + x_{2} + ⋯ + x_{k} + c = x_{k+1} [BMRS19]. X_{E}(7) [HK03]. Xndép(11) [Hal98]. xy [BHK05]. xy + yz + xz = n [Pet04]. xy + yz + zx [BC00]. Y^{2} = X^{5} + k [Bre08b]. Z^{2} [GG97]. Z_{2} [FK09]. Z_{p} [KW23, FK02]. ζ(1/2) [Vin01]. ζ(2n + 2) [BBB06]. ζ(4n + 3) [AG99, BB97]. ζ(\frac{1}{2}) [Bax02].

|L′/L(1, χ)| [LL23].

-additive [CF95, Fin92]. -Adic [ABL17b, AMM08, BCM21b, BGMR06, CGM15, GM18, RW11, Boy94, WS98]. -ary [GPS18]. -Associahedra [Man18]. -block [CHLM97]. -Braids [HS07]. -Class [DJP^{+}05]. -Clue [MTC14]. -Coefficient [War11]. -Color [BMRS19]. -Complex [Car03]. -Convergence [Oud11]. -Crossing [LM19]. -cube [Rei96]. -Cycles [AJK08]. -Deformed [MGO22]. -Dimensional [BGKMI01, EM02, HK02b, BN97, DBG19]. -Equivariant [Yun22]. -Expansions [JST16]. -Extension [FK09]. -Extensions [APS04, FK02, KM20, KW23]. -Functions [ACF^{+}21, BBW17, DFK04, DPVZ22, Dok04, Dum01, FK10, FK12, Mil06, Oma01, Ots15, Boy98]. -Gonal [BRTC23]. -Graph [HM12]. -graphs [KO95]. -Groups [BBF12, ST05]. -Homology [GL15]. -Integral [GP06]. -Invariant [Wil06]. -Invariants [Fuk11, DG21]. -Iterations [She09]. -Loop [GSS16]. -Manifold [BFG^{+}13, Bur04]. -Manifolds [Ban06, CLLM07, Con06, CLT06, Der15, FMP04, HIK^{+}16, But05, CGHN00, RT00]. -Measures [GMK19]. -Modular [GHL19, OB19]. -Number [DdLV14, Lip21]. -periodic [KdC05]. -Polynomial [GK12]. -Polynomials [BL21]. -Polytopes [SZ04]. -problem [LSW99]. -rang [Lep93]. -Regular [SW10]. -Representations [BFG^{+}13]. -Series [BU23, Hik03, XZ22, DD03a, PT08]. -singularities [Tar92]. -Space [KY17, Ume09, Tan09, ABG^{+}21]. -Spaces [JK01]. -Sphere [HS15, BL00]. -Spheres [Fuk11, Tan09]. -String [OM08]. -Sufficiency [BA02]. -Symmetric [AD21]. -Tetrahedron [Bur07]. -trees [Rim92]. -Tuples [Tan24]. -Type [Sha19, GM18]. -Types [EL14]. -Unit [CFH^{+}15]. -values [BDM^{+}18, Hir00].

1 [BDM07]. 11-Tetrahedron [BB22]. 12-Crossing [KA24]. 16-vertex [BL00]. 1997 [ELT^{+}97].

2-Adic [Chi24]. 2-class [BB21]. 2-Compact [Sem23]. 2-Forms [BLS22]. 24-Cell [RT23].

3-Manifolds [BB22, Zha22]. 3-Sets [Pul23].

4-Manifolds [RT23, ST22].

Abel [Sze98]. abelian [Lou98, Ben23, CMZ21, Die02, DN21, ET14, GL22, GM18, HM02, MSS18, Sch19]. Absolute [BU23, BDF^{+}22, Fla08]. Abstract [SSS03]. Abundant [Bri06, Del98]. Acceleration [CS22a, CVZ00]. Accessibility [SP23]. Accumulation [GW18]. Accuracy [MV10]. Accurate [Pau17]. Achille [FWW02]. across [CKL^{+}23]. Action [GL22]. Actions [JR02]. Acylindrical [Zha22]. Addenda [HMM99]. additive [CF95, Fin92]. Addressing [ACG^{+}21]. Adelic [Gre10]. Adic [ABL17b, Chi24, AMM08, BCM21b, Boy94, BGMR06, CGM15, GM18, RW11, WS98]. Adjoint [Wei14, Hir00]. Advanced [XKA14]. Adventures [ACNL^{+}23]. Affine [BS99, CO02, EHPN20, KL21, SS06, HSV94]. after [BCVZ02]. Agashe [LS02]. AGM [BCF04, BC04]. ago [PZHC09]. Aharonov [BNH11]. Ahmadi [DKRV21]. Aided [BLS22]. al [Gan17, Wan01]. Alan [CLLM07]. Alexander [DFJ12, Hor14, Koh18]. Algebra [Alv08, BH04, HV19, HP06, HP08, HHS21, MSS18, MPY17, Pur18, Roo08]. Algebraic [BCM21a, BHM^{+}13, BHI^{+}11, ÇFM23, CR01, Dum13, EF18, GMR20, HST22, Kid04, Kre10, MR10, ST19b, Sta21, Ver00, Wal94]. Algebras [BEN12, BZ17, BH04, EM17, GdlI07, Hum02, JN06, KW08, KY20, KS22, LL21, LS03, LS07, Rit07, Soh21, dG05, BJS^{+}21, KO95, Sch05, dG02b]. algébrique [Wal94, Wal95]. algébriques [Ber95]. Algorithm [Dem08, DN04, DJP^{+}05, FIS^{+}14, GG03, GJN23, Har02, IZ14, Kap07, KB19, MM01, MN14, NSW23, Opg01, OS22b, Sch04, Sch07, Wad06, Bri00, BJP94, DU95, HR92b]. Algorithmic [BBvBK15, FGJ22, JR02, BB92]. algorithmique [BB92]. Algorithms [Abr22, DG02a, DFH20, EM17, ERSÉ19, Klü02, LR17, Nit93, Sca09, HMM98, HMM99]. Aliquot [BCVZ02, CGJM20, SS11]. Almkvist [Riv04]. Almkvist-Granville [Riv04]. Almost [Bur10, EL09, EGSC18, Har02, Kob01, Ter22, LL94]. Almost-Crystallographic [Ter22]. almost-free [LL94]. Alon [RSS18]. Along [KKSV16]. Alt [HH22b]. Alternating [DHST08, DERZ18, OS22b, SZJ04, CVZ00]. Alternative [DZ11, HP06, HP08, Hol95]. Amalgam [IJK^{+}23]. Amenability [ER19, HHRS21]. Amicable [SS11]. Ammann [DEFM23]. Amod [LS02]. Amoebas [The02]. Among [AD21]. Ample [LM17b]. Amplitude [Olv01]. Analogues [Hum02, Nic18]. Analysis [BBKL17, BNH11, CGJM20, DdLV14, DSV09, GIKW23, JR02, Sev18, XKA14, LR92]. Analytic [BA02, Qua06, DS95, FG99]. Analytical [BDGS23, CPS^{+}01]. analyticity [GMSB96]. Andrews [BM06]. Angenent [BK21]. Angled [Ino22]. Anisotropic [KCK22]. anneau [Cou94]. Anosov [Bri04]. Anosovs [KKT09]. Ansonov [HS07]. Anti [KW23]. Anti-cyclotomic [KW23]. Any [IT19]. Apéry [AG99, BBB06, BB97, Gor21, Riv04]. Apéry-Like [Gor21, Riv04, AG99, BBB06, BB97]. Apollonian [CJK^{+}19, FS11]. Appendix [De 06a, CLLM07, GM08, LS02, Mil06, Mil93]. Application [BNH11, De 06a, De 06b, AB98, Boo00]. Applications [CMOS13, Gor21, HH22a, IT19, LPU11, MSV03, DE02]. Applied [BCGY23]. Approach [BBC^{+}12, BM15, DH22, HJL22, IKMF16, Soh21, Vor20, Zon22, BJS^{+}21, CPS^{+}01, Kob94, Sch05]. Approaches [EGP^{+}12, GKR01]. Approximate [LS04, Lam18]. Approximately [BK21]. Approximating [Bra21, iIS04]. Approximation [BM15, BHS09, BCPV21, GMRST14, LvF03, Mat01, Oud11, Tom96a]. Approximations [BCM21a, BCM21b, Lan98]. Arbitrary [BPS12, ER15, Kou19, dC02]. Area [Ano20a, LRZ20]. Arising [Ano20b, Bon21]. Arithmetic [Aga13, APS04, BST20, GST06, GPS18, HR14a, JST16, JST20, LZ10, LR07, Mac08, MR21, MNT23, PR17b, RCD21, SW12, Bre99, CGHN00, Jun00, PR03]. Arithmetic-Geometric [LR07]. Arithmetical [Lag01]. Arithmétique [PR03]. Arrangement [Roo08, Sne05]. Arrangements [BGBVS20, Esc14, HR14b, Caz97]. Arrays [HWW05]. Artin [Boo06, Oma01]. ary [GPS18]. Asai [DPVZ22]. aspects [DE02, LR02]. Assisted [BK22]. Associahedra [Man18]. Associated [BCLM17, BHI^{+}11, CLLM07, CL04, DEFM23, DLN23, EGP^{+}12, HM21, JN06, KB17, Mer20, Shi05, Sta23, ST05]. Associative [EM17]. Asymptotic [ArE11, Alm98, AGS18, Alv08, BT18, CS22a, DG93, GN22, GWW16, Hen23, HRSS17, Hik03, MRRTS16, Pat03, Pat05, SZJ04, DE93]. Asymptotics [GM08, KKSV16, ST19b, JRS92]. asymptotiques [DE93]. Atkin [FHL^{+}10]. Atlas [CN13]. ATR [GM13]. Attached [Die02, Lam18, Hir00]. Attracting [Jeo15]. Attractors [BDM07, De 15b, Rei96]. author [Mil93]. Authors [Ano92a, Ano93]. Automata [BL07, DG93, MPR94]. Automated [Adz17]. Automatic [KH19, Soh21, EIFZ96]. Automorphic [DS21, Boo00]. Automorphism [Lut13]. Automorphisms [DK21, KB19, Shi16, Soh21, GR00]. Average [DL22]. Averaged [May01]. Averages [CHMW15]. Avoiding [BCG10, dW98].

Back [Ano01a, Ano01b]. backmatter [Ano11c]. Bad [Kou19, Sha21]. Bailey [Gan17]. Balanced [CC07, Pec19]. Ball [Cra14, Muk17]. Band [BTZK23, OTZK22, KK99]. band-gap [KK99]. Band-Limited [BTZK23, OTZK22]. Banded [EGSC18]. Bar [TTS17]. barycentre [Jen00]. Barycentric [BKN^{+}12, IC16]. Base [Bar06, Har15, OM08]. Base- [Har15]. Base-Tangle [OM08]. Based [DRT11, BDL96, Sch05]. Bases [AFMS17, BCOS24, BHP12, BHI^{+}11, Gaá01, GOP02, GR19, Ola05, dG02b]. Basic [DSV09]. basis [HMM98, HMM99]. Baumslag [MV00]. Bayesian [GW18]. Beauville [BBF12]. Beenker [DEFM23]. Behavior [BT18, CL07a, CMOS13, ER19, Hen23, MR21, DG93]. Bell [WPW97]. below [Gun00]. Bennequin [HIK22]. Bernhard [BH21]. Bernoulli [CGM15, HS18]. Bernstein [BP21]. Bers [KSWY06]. Bessel [BS08a, GP92]. Best [PE06]. Bestvina [Bri00]. Betti [Sne05]. Between [Mat01, LL21, LS03, PdlL02]. Beurling [LR02]. Bhargava [Lip21]. Bhatia [Ber17]. Bianchi [Şen11]. Bias [FS10, GKL19, RS94]. Biases [ACF^{+}21, FM00]. Bicolored [CG14]. Bicycle [LT09]. Bieberbach [CS01]. bifurcation [Kra94, Ros05]. Bijective [BP07]. Billiard [MRRTS16, ST19a]. Billiards [BMS21, HPR22, Jeo15, New16, Sch06, Sch09, KdC05]. Binary [Aic04, BS17, BÖW18, EH04, Fer22, KN22, Pet04, Wag01]. Binet [CS22a]. Binomial [Abl17a, BBK01]. Birational [PPSF17]. Birationality [Gui18]. Birch [Ble11, vB22]. Birkhoff [KdC05]. Birth [PZHC09]. bistellar [BL00]. Bit [Hew21]. Bits [Wag01]. Blasius [Boy99]. Block [Hul95, CHLM97]. Blocki [ÅCL18]. Blow [He21]. Blow-Ups [He21]. Blown [BDP16]. Blown-up [BDP16]. Blowup [ALT^{+}12, GIKW23]. Blowups [ALP^{+}11, SK11]. blueprints [AB98]. BMR [Tsu20]. Board [Ano12b, Ano13, Ano14, Ano15, Ano17, Ano18, Ano11b]. Bodies [DLM13, Ger03, LP14]. Bogomolov [Fab09, JM14]. Bohm [BNH11]. Bolza [KKSV16]. Bombs [New16]. Boolean [CC07]. Boppana [RSS18]. Borcherds [May10]. Border [HIL13, LR17]. Borel [GMR20]. Borwein [AG99, Bai17]. Bott [Da 19]. Bound [AM01, AJK08, Cd18, HS18, PZ04]. Boundaries [DPPS05, LN03]. Boundary [CV03, FMP04, SP23, Zha22, Jen00]. Bounded [AL03, CC07, Lip21]. Bounding [BCG19]. Bounds [Ano20a, AF06, BY13, Boo21, BÖW18, BS11, Del98, ES22, Fer22, GKT20, HP18, HIL13, IC16, LRZ20, MV10, MVG05, RSS18, Saw23, TS06, BC98, Dre93, DE93, JLW95]. Box [BCC10]. BPS [Cha23]. BPS-Series [Cha23]. Braces [KSV22, KSV21]. Bradley [AG99]. Braid [BM04, MSV03, HdC02]. Braids [HS07]. Branching [PR13, CD10]. Brauer [AKN^{+}16, CHLM97, Poo06]. Breaker [ST21]. Bredon [BG95]. Brennan [Tos22]. Briques [BF16]. Broken [LM06, Pec19]. Broken-Cycle-Free [LM06]. Bruno [Car03]. Bubbles [ÁCWB03]. Buildings [KV17]. Bundles [Els11, KK12, MMT00]. Burau [FS19].

Cable [Cha23]. Calabi [AP23, BKZ22, Die04b]. Calcul [RV99]. calculate [DU95]. Calculating [CO02, Gaá15, LM19]. Calculation [HLMZ01, May10, Hir00, NR99]. Calculations [CMZ21, GL23, HV22, LOS20, KSV97]. Calculator [CKL^{+}22]. Calculus [GPHH^{+}12, HHdC^{+}15, Knu01, LOS20, RSSS06, Sot00]. Can [Bre08a, Raw03]. Cannon [BGSS22]. Canonical [BGM22, BKZ22, CO02, GHSV23, Hal14, JST20, EW98, dG02b]. Cantor [ABL17b, JRS92, Jeo15]. capacity [Ros97]. Capillary [BT18]. Carathéodory [Bru07, GMSB96]. Cardinal [GGM11]. Case [BBBR17, BCF04, BC04, Gan17, GW18, GK21, PT08, Sch13a, SdW94]. cases [IL00]. Casorati [DPV15]. Casse [BF16]. Casse-Briques [BF16]. Castelnuovo [CLMR15]. Catalan [HLM23, MR95]. catastrophes [PS92]. Category [LOS20]. Cauchy [DL04, LSV98, RV99]. Caustics [kAL01]. Caustique [JR99]. Cayley [DdLV14, RS19]. Cell [RT23, BC98]. Cells [HOY12]. Cellular [BL07, DY07, DG93]. Census [BB22, Bur07, Bur14, DPP11, FGG^{+}16, Gar21b]. censuses [But05]. Center [kAL01]. Central [BGBVS20, BBK01, Mil06, PT08, AAB^{+}21]. centre [Jor99]. Certain [Bar06, BP15, CGKS20, DD03a, EF18, JST16, JST20, Yab09, BSTV93, CRY00, DU95, DL92, LL94, Thu94, VS00, VV99]. certificates [Pfe21]. Certified [BL12, HLLM15]. CF [EF09]. Chain [CR22, LW12, Mar19]. Chains [GPS18]. Challenging [BHI^{+}11]. Chamber [ADGH20]. champions [ORW99]. Chan [Coo08]. Change [Hew21]. Chaos [kAL01, RS19, WS98]. chaotic [Lan98, Pin00]. Character [BC01, CG09, FGK^{+}16, Sto09, BG99, Neb00, TW99]. Characteristic [BS21, BLS22, GG03, KL16, MSS15, Nic18, PR08]. Characterizing [Pul23]. Characters [GHL19]. Charactistic [Elk01]. Chebotarev [KZ12].

Chebyshev [FS10, Har11, HH22a, RS94]. Checkings [RT16]. Chen [Tha95]. Chern [MMO^{+}02]. Chinburg [JRS03]. Chromatic [LM06]. Chua [Coo08]. Chyzak [Wan01]. Circle [BTZK23, BH01, BL21, DS95, FS11, HS14, OTZK22, PdlL02]. circles [GR00, Nur00]. Circulant [EVV16]. Circumcenters [HPR22]. Class [BBBC07, BR18, DSS07, DG02a, DJP^{+}05, ET14, FI18, FK09, Gui06a, HJK^{+}19, KM20, LVL22, Lee23, LPU11, Mal10, Wil23, AT00, BB21, CF95, KSV97, Lou98, Tah00]. Classes [ABG09, Sca09, War11, AT00, LL94]. Classical [Koh18]. Classification [BN97, BT18, BP07, Soh21, dG05, Sch05]. Clausen [BBK01]. Clebsch [DM06, PPS09]. Clifford [BJS^{+}21, DSV09]. Closed [BLS22, CLM13, DS19, EGSC18, Pul23, HW94, Tah00]. Closure [KSY08, MMW07]. Closures [Gal14]. Clouds [HR04]. Clue [MTC14]. Clues [MTC14]. Cluster [KY20]. Clusters [Kap07]. CM [Chi24, DW10]. CMC [HS15]. Cobordism [BCG19]. Coclass [EF09, EM17]. Cocycle [AER^{+}08]. Codes [BÖW18, Wan09]. Coefficient [Van17, War11, KdC05]. Coefficients [ACF^{+}21, Ave10a, Ave10b, BK03, BU23, BFT16, CKRS06, DM06, GJU12, HHS21, PR08, AAC98, Boo00, BC99]. cofinite [Tah00]. Cogrowth [ER19]. Cohen [KLNP20, WtR03]. Coherence [GN22]. Cohomological [KMN24, Gun00]. Cohomology [AP23, APS04, BGSS22, CH13, DR17, FGT10, GMY21, Kho03, MPY17, Şen11, Web09, AAC98, AM92, Tah00, vGvdKTV97]. Coincidences [PRS22]. Coincidental [Hop22]. Collatz [Hew21]. Color [BMRS19]. Colored [BHP12]. Colorings [EFMPR18]. Combinatorial [Ban06, Caz97, KM18, Lom20, SK11]. Combinatorics [Gar21a, Wil06]. Commensurability [BST20]. Commensurable [Cum04]. Commensurators [GHH08]. Comments [KY20]. Commutator [Wal13]. Commutators [LV15]. Compact [GR22, GBP97, Sem23, BJS^{+}21]. Compactified [BCGY23]. compacts [Ber95]. Comparison [CJUE04, BJL05]. Complement [Mil01, MR00, Ran99]. Complements [Gör15, AR97]. Complete [ACG^{+}21, CLW15, RCD21]. Completely [PR17a]. Complex [Abr22, Ano20b, Bon21, BC04, Car03, DPP11, Der22, DS14, EJ22, Jon10, JJJ02, Kaw18, Kre10, LvF03, Roo08, SSS03, BFK^{+}92, Boy99, Cra99b, – Dok00, FG99]. Complex-Parameter [BC04]. Complexes [EVV16, EM02, KLNP20, Kre10, SZ04, LM97]. Complexity [Con06, HHMP10, MP01, Nao21]. Component [Aga13]. Composite [DERZ18, GOP02, GR19, PT08]. composition [ABT96]. Compression [LP14]. Computation [Bad16, BB18, BPW95, BS12, Cho94, CS01, DM06, DJP^{+}05, GSS16, GHL19, Gra02, GM13, LSY03, LPU11, Lou98, MSV03, Par21, Jor99]. Computational [BBBR17, DSS07, DE02, EGP^{+}12, Gal14, Gan17, HHRS21, HJL22, IKMF16, KK12, KO95, Lau21, Lom20, MRS22, TTS17, BJS^{+}21]. Computations [Ave10a, Ave10b, BQS95, BDdG07, BH18, BEGY20, BHI^{+}11, But12, Dem05, HS05, HIK^{+}16, Sta23, WtR03, LS02, MN05, Rog00]. Compute [BK03, Pfe97]. Computer [BBKL17, Boo21, BLS22, BL07, BK22, But11, GP92, HLMZ01, LM97, Mat98, Zon22, BD09, Guc95, Sch05]. Computer-Aided [BLS22]. Computer-Assisted [BK22]. computer-based [Sch05]. Computers [Raw03]. Computing [BP92, BdJS23, BS01, CDS02, CH13, CLT06, CGHN00, Cre97, CD09, DG21, DR96, Die04b, Dok04, ET14, Fir15, FK19, Gun00, Hor14, HT00, IZ14, JLY16, Kou19, LZ10, LS03, LS07, PT08, PP93, Poo01, Ran99, Ros97, ST02a, Ste95, The02, Wat02, Wik03, dC02, GPP21, RV99]. Concavity [LM06]. Concerning [Aga13, AR12, EH04, WtR03, Gar02, TW99]. Condensation [LN01]. Conditional [Boe23]. Conditions [CF96, Tse22]. Conductor [Lip21]. Cone [BDP16, DD03b]. Cones [BHI^{+}11, Opg01]. Configuration [MPY17, OS22a]. Configurations [BBC^{+}09b, BCGY23, CG14, OS22a, BC98]. Confined [ES22]. Confirming [Tos22]. Confluence [NSYY08]. Conformally [BH01]. Congruence [BR18, BHK05, CL04, CP03, Cum04, GMY21, KT16, KKSV16, Tak14, vGvdKTV97]. Congruences [BDM^{+}18, Gor21, GM08, IT19, Tan24]. congruent [Rog00]. Conjectural [TS06]. Conjecturally [Nur00]. Conjecture [ÅCL18, AFMS17, BHK^{+}18, Bad16, BO11, BY12, Bir14, Bir22, Ble11, BZ16, Boo06, BS21, BH21, Car03, CKL^{+}23, Coo08, De 15b, Dou05, Dou07, DH22, DLM13, DKRV21, DHJ15, EM17, EM09, Fab09, GPHH^{+}12, GN22, GXB15, HHdC^{+}15, HP13, Hik03, Kie21, KM18, LT09, LM06, Mal22, MMO^{+}02, Oud11, RW11, RT16, Sha19, SZ06, Sne05, Tos22, Wan01, vB22, BM98, BG95, GMSB96, MN05, Pau96, Sot00, Ver00, AK23, BH12, CMPR21, HOKMS24, Tos22, Win22]. Conjectured [Gui06a]. Conjectures [Aga13, AR12, BDdG07, Ber24, BL07, BEGY20, DLN23, DPV15, Han09, Kat19, MR21, RSSS06, She20, Svi18, Nit93, Rob00, Mar22]. Conjecturing [Adz17]. conjugacies [PdlL02]. Conjugacy [AT00, GHK20, RD04]. conjugates [PP93]. Conjuncture [JRS03]. Connected [Bax02, NR00]. Connectedness [SS06]. Connection [DS14]. Consecutive [MU22, JW00]. Constant [BCIO01, BC01, BBC09a, BÖW18, CV03, CIO07, DGJ22, PE06, RS98, Ume09, Cra99a, GBP97, GB97, KMS00]. Constant-Dimension [BÖW18]. Constant-Mean-Curvature [Ume09]. Constantes [HMT10]. Constants [FK19, HR14a, IC16, Lag01, LZ10, HMT10]. Constellations [BCG10]. Constrained [ACPR11]. constructed [BB92, VS00]. Constructing [DHST08, EF09, KSW01, PS96, PR05, Roe07, dG02b]. Construction [BBvBK15, BDSS18, DSS07, Gra96, IT19, Lee23, PPSF17, SW10, SZ04, SW93, Tha08, Zha22, CPS^{+}01, KO95, Gra96]. Constructions [LM17b]. construits [BB92]. Contained [Fir15]. Contents [Ano03a, Ano03b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b]. context [BM98]. Continuation [JLY16]. Continued [Abr22, AK23, BCM21b, BCLM17, CHMW15, Duj01, FIS^{+}14, HV22, MU22]. continues [PZHC09]. Continuous [MO07]. contour [FS98]. Contraction [Jeo15]. contrast [KK99]. Control [LN10]. Convergence [CVZ00, HK02b, Oud11, She09, KK97]. Convergent [CCG21, Har02]. Convex [BZ16, BS11, DL22, DLM13, HT19, KSY08]. Convolution [Ben23, Ste21]. Convolutions [HS18]. Conway [Pin00]. Coordinates [Bur10]. Copies [Fir15]. Corank [Sha19]. Corks [Nao21]. corps [Lep93]. Correct [EL09]. Correction [Ano20a, Ano20b, Ano23]. Corrections [AAB^{+}21]. Correspondence [CD09, LL21]. correspondences [BP94]. Corresponding [GdlI07]. Corrigendum [Ano11a, Ano12a, FKS19]. Cosets [Els14]. cotangent [MMT00]. Could [FKS15, FKS19]. Count [ABB22]. Counterexample [ÅCL18, BH21, Tos22, Mar95]. Counterexamples [Ale19, DKRV21]. Counting [CC07, CGP23, DS15, GG03, HLLM15, LS13, Mel22, PS00]. courbe [Hal98, HK03]. Courbes [Lep93, Fer92, Fer96, PR03]. cousin [Pin00]. cover [SW93]. Covered [EVV16]. Covering [BRS^{+}18, GKS22]. Coverings [LRSST23, Nur00]. Covers [AAHwaabAR22, BCOS24, CHRR07, LGRS23]. Cox [DBG19]. Coxeter [BKM24, dC02]. CP [BDGS23]. CR [Ale21, Der15]. created [PS92]. Crepant [BSW24, DBG19]. critère [LR02]. Criteria [KALP22, LR02]. Criterion [AKN^{+}16, KL16, WPW97]. Critical [BK12, BG22, BH18, DLS22, FK10, FK12, GLP16, NR00, RS22, WZ21, Mil00, PdlL02]. Critical-Orbit-Relation [RS22]. Critically [DS13]. cross [He02]. Crossing [ÇFM23, KA24, LM19]. Crossings [BKM24]. crucible [BD09]. Cryptography [SP05, Sch19]. Crystal [GG06]. Crystallographic [Lam18, Lut13, Ter22, PS00]. Cube [BGKMI01, Rei96]. cubes [JK95]. Cubic [BB21, Coo08, CKL^{+}22, DLN23, DS13, DGKMY15, Duq01, EJ10, Gaá01, LSY03, Sta23, Mil92, SL00]. Cubical [SZ04]. Cubics [Chi02, Wil06]. Curious [DP15]. Currents [Kap07]. Curtis [BM06]. Curvature [CV03, CKL^{+}22, EM02, GIKW23, May01, Tra08, Ume09, CS93, Cho94, GBP97, GB97, HKS92, KMS00, RS98]. Curvatures [CKL^{+}22]. Curve [CL22, DG21, EHPN20, Mel22, Mil06, Wat02, EW98, GPP21, SdW94]. Curve-Shortening [EHPN20]. Curves [Ano20b, ADGH20, BPS12, BLP09, BMS21, BvBK09, BDM07, Bon21, BBW17, BS08b, BHH01, ÇFM23, CR22, CP10, CP12, Cha14, CL07b, CM16, Cul19, DGJ22, DG02a, DFK04, DLN23, DS13, Del01, DD03a, Dem08, DLS22, DGKMY15, DT02, Dum02, Duq01, Fab09, FG23, FS12, GST06, GHSV23, GJN23, Gre10, GdlI07, GM13, GHY13, HS13, Hye17, JLY16, KKP^{+}19, Kat10b, Kob01, Kou19, KH22, KS01, LM22, LL21, Leh22, LRR^{+}19, Lip21, LH23, Lom02, MR10, Muk21, Pio07, Poo01, Poo06, Qua06, RS01, Rub13, SZJ04, Sch19, SS11, Tse22, Wat08, vB22, BG97, BHvdP95, Bre99, Cre97, Fer92, Fer96, PR03, Rog00, Ros00, RS00, VV99]. Cusp [BV22, RT23, Cre97, GH96]. Cusped [GHH08]. Cuspidal [Aga13, Pio07]. Cusps [Pio07]. Cut [iIS04, ST02a]. CW [Kre10]. CY [BR17]. CY-type [BR17]. Cycle [LM06]. Cycles [ACG^{+}21, AJK08, Hay18, Joh11, SS11, ST95]. Cyclic [BB21, EHT01, ES14, HR03, LSY03, Tha08, Wal13]. Cyclically [MW22]. Cycling [Knu01]. Cyclotomic [Fie07, FI18, FK09, IJK^{+}23, KM20, KMP22, PRS22, BQS95, BC99, KW23]. Cylinder [GY07, OS22a].

D [TTS17]. D-Bar [TTS17]. Damped [AL03]. Darmon [GM13]. Data [BH12, BHM^{+}13, KKM05, XKA14]. Data-Mining [XKA14]. Database [Bro07, KKM^{+}01]. Daubechies [Wan01]. Decic [DJ10]. Deciding [BS08b, ER19, KL16]. Decimal [Gan14]. Decision [GHK20]. Decomposable [Chi02, PR17a]. Decomposing [LGRS23]. Decomposition [Ame08, ADGH20, BEN12, Con21]. Decompositions [CILO19, KM18, LS07, OM08]. Dedekind [Alm98, HNR20]. Defect [HIK22]. Defined [Chi24, Cul19, Del01, Duq01, Kob01, NTTY22, NSYY08]. Defining [LH23, Muk21]. Deformation [BK22, GK21, LS02]. Deformations [HS15, Ste95]. Deformed [MGO22]. Deforming [HW02]. defy [SdW94]. degeneracy [MRS22]. Degenerations [DDG^{+}14]. Degree [BA02, BvBK09, CH08a, CC07, EMRSÉ14, HLMZ01, HOY12, HP08, Hop22, KKP^{+}19, LRSST23, MS11, RSÉ03, Wat02, Web09, BSTV93, FP92, FP93, FPDH98]. Degree- [EMRSÉ14]. degree-two [BSTV93]. Degrees [Alu15, Fou21, Sch09]. Del [CS13]. Delaunay [Ros05]. Deligne [FPUP22]. Deligne-Mostow [FPUP22]. Delone [MVG05]. dendritic [Kob94]. Dense [DFH20, LT18]. Densest [Zon22]. Densité [Wal95, Wal94]. Densities [STZ95]. Density [DEFM23, Fer22, MVG05, TS06, Del98, Wal94]. d’entiers [Cou94]. Dependent [BPL22]. Depth [IZ14, IKMF16]. Derivation [MSS18]. Derivatives [BB18, DD03a, FK10, FK12]. Derived [HV19, Mac08, LL94]. Descartes [FKS19, FKS15]. Descent [ACPR11, Knu01]. Descent-Cycling [Knu01]. Description [EF18]. Destabilizing [BMN22]. detect [DH97]. Detecting [KB19, MN14]. Determinant [vdP01, FJ93]. Determinants [BHvdP95, DPV15, FJ96]. Determination [BBB06, Die02, PE06, Wad06]. Determine [Zon22]. Determined [Sha19, BB97]. Determining [GJN23, ST22]. Deviation [Xu04]. Deviations [AAB^{+}21]. Diagonal [Ber24, EJ10]. Diagrams [BCL^{+}22]. Diameter [BS11, CS22b]. Diamond [KSW01]. Diffeomorphisms [GMK19]. Difference [Ano20a, LRZ20, Neu06]. Different [AFMS17, HS13]. Differential [Dur11, Gar17, GdlI07, NSYY08, Rom18, Sne05]. Differentials [CMZ21]. Differentiation [HK21, Wei14]. diffstrata [CMZ21]. Diffuse [TTS17]. Digit [CHMW15]. Digital [GG06]. Digits [BBBR17, CFH^{+}21, Gan17, Har15, BCVZ02]. Digraphs [CM21]. dihedral [GPP94]. Dilatation [CH08b, HS07]. Dilation [DM06]. Dimension [BÖW18, De 15b, GY10, HS18, HOKMS24, Gun00, Web97]. Dimensional [BCC10, BW07, BGKMI01, BK16, CS01, CGKS20, EM02, HK02b, Har02, LS03, LS07, Mat01, Sch07, Soh21, Ter22, AM92, BN97, Caz97, – Dok00, Kob94, Now97, Tom96a, DBG19]. Dimensions [BNHLR11, BDdG07, BGKT05, DS14, LvF03, LN03, Cra01, JR99, PS00, VS00]. diophantine [ST99, DP08, GP06, LvF03, Pet04, Riv10, Shi05, SdW94, dW98]. Dirac [ER15, HY22, SSS03]. Direct [DHJ15, LN01]. Dirichlet [AF06, ACF^{+}21, BM15, Der05, DS14]. Dirty [GS05]. Disc [Fer22, GR22]. Disc-Packings [GR22]. Discontinuities [DEFM23]. Discover [Raw03]. Discovering [Abl17a, Abl22]. Discovery [BBKL17, PZHC09]. Discrepancies [LS13]. Discrete [BL14, DSV09, Eng09, GY10, GK21, HY22, KKM05, LLT03, LR05, MO22, PPS09, DS95, GG97, Lan98, PP93]. Discreteness [Wad06]. discretization [BS99]. Discretizations [PPS09]. Discretized [Vor20]. Discriminant [DGKMY15, FP92, FP93, FPDH98]. Discriminant- [DGKMY15]. Discriminants [AS12, ABZ07, DJ10]. Discs [BMN22]. Disk [dVV22]. Disks [GY07, OS22b]. Dissections [Ano20a, LRZ20]. Distance [BY13, LLS95]. distance-transitive [LLS95]. Distinct [Tan24]. Distinguished [CO02]. Distributing [KS03]. Distribution [ABB22, Bax02, Bor14, CLM13, GJU12, LMO13, Mal04, Mal10, MNS08, Pat03, Pat05, Qua06, RCD21, Van17, AL95]. Distributions [LRR^{+}19]. Disturbing [SZ06]. Divide [ÇFM23]. divided [BCPV21]. Divisibility [Yab09]. Division [DT02]. Divisor [McN17]. Divisorial [Tse22]. Divisors [CJUE04, GJL03, KB17, Tse22, Yab09, CtR96, CtR97]. Do [BU23, Ger03]. Does [DH97]. Domain [AL03, HOY12]. Domains [AD21, Der05, SW22b, NR00]. Dominant [DS15, ER19]. Don [GM08]. Donaldson [She09]. Doppelgängers [Hop22]. Dorfman [KS22]. Double [ÁCWB03, BCR22, BLS22]. Double-generic [BCR22]. Doubly [CP12, Con18, HJL22, CR99, RTW00]. Doubting [Bla22]. down [Hop22]. down-Degree [Hop22]. Doyle [BH01]. Drawing [KSWY06, MPR94]. Dream [He21]. Drinfeld [BV22]. Dual [CS21, Gar21a, GHSV23, Opg01, SZ04, Bra95, FH99]. Duality [KK12, BS99]. Dueñez [Mil06]. d’une [Gra96]. Duplicate [Bur14]. Dyadic [Hew21]. Dyer [Ble11, FHL^{+}10, vB22]. Dynamical [AKN^{+}16, AD21, BCH^{+}14, CMOS13, MO22, EW98]. Dynamics [BMS21, BSTV93, Cra14, HK21, Kaw18, ST19a, CR93, Cra99b, LSW99, Mil93, BDM07]. Dyson [SZ06].

Earthquakes [Gar23]. Edge [BS11, MT16]. Edge-Graph [BS11]. Edge-Marked [MT16]. Editorial [Ano04a, Ano04b, Ano13, Ano14, Ano15, Ano17, Ano18, Ano11b, Ano12b]. editors [Ano92b, Ano97, EL92]. Eduardo [Mil06]. Effect [DM06]. Effective [BDP16, Coc05, Leh22]. Efficiency [CHRR07]. Eigenforms [Van17, vGvdKTV97]. Eigenfunctions [CS21, MS01]. Eigenvalue [AF06, Bar06, HJL22, MM01]. Eigenvalues [BNH11, EGSC18, FIK20, MNS08, Oka02, GH96, Gun00]. Eight [FGK^{+}16, Mil01, MP22, MR00, Ran99]. Einstein [BGKT05]. Eisenstein [BDM^{+}18, GMY21]. either [MR95]. EKG [LRS02]. Electron [Sch13a]. élémentaires [DE93]. Elementary [DE93]. Elements [BDS16, HP08, CM00]. Eleventh [Boe23]. Elimination [PR08]. Elite [Mül06]. Elkies [MM01]. Ellipses [AD21]. ellipsoid [BFK^{+}92]. ellipsoïdale [JR99]. Ellipsoids [BZ16]. Elliptic [BK03, BPS12, BDM07, CL22, CL07b, DGJ22, DG02a, DFK04, DLN23, Del01, DD03a, DG21, Dem08, DGKMY15, DT02, Dum02, Duq01, FS12, FK10, FK12, GST06, GJN23, Gre10, GM13, GHY13, Jai09, Kat05, Kat10b, Kob01, Kou19, KS01, Lip21, Mil06, Neu06, PPSF17, Qua06, RS01, Rub13, Sch19, SS11, Wat02, Wat08, Yab09, Bre99, Cre97, EW98, Fer92, Fer96, PR03, Rog00, Ros00, RS00, SdW94, ST99, dW98]. elliptiques [Fer92, Fer96, PR03]. Embedded [AB09, CLW15, RTW00, Tra08]. Embedding [BM98]. Embeddings [AER^{+}08, BB22, KSWY06]. Empirical [BBC^{+}12, DH22, EFHY21]. Empirically [BB97]. Encryption [HWW05]. ends [BCVZ02, RS98]. Energy [BBC^{+}09b, He02, KK93]. Energy-Minimizing [BBC^{+}09b]. Enrico [PZHC09]. Enriques [Dol18, MRS22]. Entropy [AJK08, BK21, DK21, Hal14, KKT09]. Enumerating [AB98, ADGH20, LN01, MP15]. Enumeration [ArE11, Ame08, MTC14, Sil17]. Enumerative [GM08]. Enveloping [CG14, Hum02, dG02b]. EOV [Ano11b, Ano12b, Ano13, Ano14, Ano15, Ano17, Ano18]. Equal [CP03, Nur00]. Equality [ABZ07]. Equation [BBW17, BMRS19, Bre08b, Gar17, GHK03, NSYY08, Pet04, Kra98, MR95, Sze98, Kra98]. Equations [Dur11, GP06, Gaá15, HIL13, MR10, Neu06, OS16, RS22, Rom18, GPP94, JW00, SdW94, ST99, dW98]. equator [LP97]. Equidistant [IC16]. Equidistribution [Kat14]. equilateral [Nur00]. equilibrium [BPW95]. Equisymmetric [BRTC23]. Equivalence [BST20, War11]. Equivalent [BZ17, Cha14]. Equivariant [Ble11, Mal22, RW11, Yun22]. Ermolaev [Pur18]. Errata [CtR97, Wal95, HMM99]. Erratum [KSV22]. errors [Viv94]. Escaping [HPR22]. Eschenburg [But12]. essential [NR99]. Essentially [BP15]. Estimate [BA02]. Estimates [Joh08, PR17b]. Estimating [LV15]. Estimation [Olv01]. Eta [HNR20]. etc [Ano97, CHS96]. étude [HMT10, BB92, Fer96]. Euclid [KS08]. Euclidean [HR04, dVV22]. Euler [BBG94, CB94, Eel07, FS98, MSS15]. Eulerian [CGM15]. Eutactic [AG22]. evaluating [SW99]. Evaluation [BB18, Boo21, CS22a, JST20, BBG94, CB94]. evaluations [Dil00]. Even [Muk22, OB19]. Every [De 15a, Die04a]. Everywhere [Har02]. Evidence [AKN^{+}16, AAB^{+}21, AD21, Ble11, DG02a, GW18, RW11, Ver00]. Evolutes [AST22]. Evolution [Adz17]. Evolutionary [Soh21]. evolver [Bra92, BPW95]. Evolving [DLS22]. Exact [BHM^{+}13, GSS16, Kat10a, Rah21, Rom18, Wan09]. Exactly [ST19b]. Example [May01, Mil04]. Examples [BGM22, LM17b, LRSST23, NP23, PR13, Sha19, Nit93]. Exceptional [BM04, CLLM07, HPR22, Sil95, IL00, IKM02]. Exchange [KY20]. excitable [DG93]. Existence [BS08b, Die04a, LM22, RAG18]. Exotic [BGKT05, Nao21]. Exp [Ano04a, Ano04b]. Exp. [Ano03a, Ano03b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a]. Expansion [BK22, Duj01, Gan14, Hik03, dVV22]. Expansions [AK23, BC01, Gaá01, GPS18, Har15, JST16, MU22, XZ22]. Expansivity [GLP16]. Expectations [Hop22]. Experiment [BS08b, Sch22a, Sto09, ABG^{+}21, PZHC09]. Experiment. [Ano00, Ano01a, Ano01b, Ano01c, Ano01d, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano11c]. Experimental [AM92, BH12, BBG94, BBB06, Bai17, BB18, BBC^{+}09b, BEZ03, Bel21, GKR01, GW18, LN10, PPS09, PZHC09, Wor20, XKA14, BD09, DRR99, PdlL02, Dre93, Fer96]. expérimentale [Fer96]. expérimentales [Dre93]. Experimentation [RSSS06]. Experimenting [BCH04, DFH20, DFH23]. Experiments [AR12, CLM13, FS11, HLLM15, KZ08, Riv16, dCIL22, DS95, JRS92, LR92]. Explicit [BM04, Dem05, Die02, GL22, LR07, PR17b, RSS18, SW10, Sta23]. Exploiting [Cd18]. Exploration [Lom20]. Explorations [BL07, HHRS21]. Exploring [Ino22, Tra08]. exponent [MR95]. Exponential [CL07a, Kat19, Pat03, Pat05]. Exponentially [MRRTS16]. Exponents [Fou21, HK02b, SY01]. Extended [ALT^{+}12, HMM98, HMM99]. Extending [DM22]. Extension [BTZK23, CEM06, FK09, OTZK22, FP92, FP93, Gra96]. Extensions [AST17, BLS22, DW10, FK02, Kat14, KM20, DE02, FPDH98, KW23, APS04]. Extremal [Chu05, GR22, OB19, OS22a, Sca09]. Extremality [Riv10]. Extreme [CIO07]. extremizing [KK93]. Extrinsic [HY22].

Facets [BS11]. Factoring [BtR96]. Factorisations [Hol04]. Factorization [Shl21]. Factorizations [GXB15, Muk21, Rah21, VRS03]. Factors [Bre08a, Web97]. Failure [KW08]. False [DW10]. Families [ABL17b, ACF^{+}21, Hal14, Kat10b, LH23, MNS08, Pul23, RS22, Tha08, Fer96, RS00, Thu94]. familles [Fer96]. Family [AGS18, BG22, Ber17, BP21, BNL17, Das23, GLP16, HPR22, MR10, RTY21, Sto17a, Tsu20, CHK93, HK02a, Ola05, SL00]. Fan [Man18]. Fano [CGKS20, Els11, JK01]. Fans [JLS20, HT00]. Farey [BHH01, LM17a]. Farthest [Sch22b]. Fast [BM06, LR92]. Fat [HP14]. Feigenbaum [Thu94]. Feit [DS21]. Fekete [MNT23]. Fermat [HS13]. Fermi [PZHC09]. Feuilletages [CD10]. Few [BS11, BCPV21, KH19, BP92]. Fewer [KH19]. fiber [Cos02]. Fibonacci [DRT11, Hol95, Wei16]. Fibrations [HLM23]. fibré [Cos02]. Fibred [But05]. Field [ABZ07, DGKMY15, FS12, FI18, GHY13, KM20, Kou19, LSY03, Van17, EH96, LR92, SW99]. Fields [BCIO01, BH12, BPS12, Bir14, BB21, CDD^{+}16, CIO07, DG02a, Dem08, DJP^{+}05, DJ10, DT02, Duq01, Fie07, FK02, GOP02, GP06, GR19, GMRST14, GXB15, HM02, HJK^{+}19, KN22, Kob01, LRR^{+}19, Lip21, Mal10, Oka02, Oma01, SP05, FP99, GPP94, Guc95, JLW95, Klü02, Lou98, Ola05, PS92, Ste93, ST05]. Fifth [GHY13, OS16]. Figure [FGK^{+}16, Mil01, MR00, Ran99]. Figure-Eight [Mil01, MR00, Ran99]. Filled [SW22b]. film [Bra95]. Find [OS22b]. Finding [BCG10, CL07b, EMRSÉ14, ERSÉ19, MM01, Sch04, Wan09, Nit93]. Fingering [AD14]. Finite [Abr22, Alm02, BPS12, BNHLR11, BM15, BMS21, DdLV14, Der22, EM17, EGP^{+}12, FS12, GMRST14, GXB15, HWW05, HM02, HRSS17, Kir11, Kob01, KZ12, LvT05, LRR^{+}19, LM20, Lut13, LS03, LS07, Neb96, PR05, RAG18, Tra08, VRS03, LR92, MPR94, Pfe97, PS96, PS92, Ter96]. Finite-Dimensional [LS07]. Finitely [But11, DHJ15, ERvR15, Gal14, Hul01, Roo08, Sha19, HR96]. Finiteness [DHJ15, LM22, Now97]. Fintushel [Fuk11]. First [Hor14, IKMF16, BP92, Jor99, KdC05, Ros05, Tah00, Wil02]. Fischer [GHL19, LLS95]. Five [CS01, Sch13a]. Five- [CS01]. Five-Electron [Sch13a]. Fixed [FG23, FS12, Gra02, Wag01]. Flag [RSSS06]. Flat [Lut09, Wal03, Thu94]. flat-top [Thu94]. flips [BL00]. Flow [CS93, EHPN20, GIKW23, LOS20, May01, BGSS22, Cho94, RS05]. Foam [Boo21, BPW95]. Fold [HdJF06]. folding [CR93]. Foliations [CV03, CR01]. Folkman [DR08]. Fonction [Dre93, DE93]. fonctions [Fer92]. Forbidden [GWW16]. forced [MP94]. Forcing [Col05, GHK03, HdC02]. Foreword [Kon22]. Form [EGSC18, GPP94, HMM98, HMM99]. Formalizing [BL22, DKAP22]. formation [CS93]. Formats [BKZ22]. forme [BB92]. Formes [CNJ96]. Forms [Aic04, AG22, BBBC07, BV22, BP15, BP01, Bir21, BLS22, BS12, CKRS06, CIO07, CD09, Dem05, Dum13, DS21, FHL^{+}10, GHY13, HV19, Kat05, Kat10a, Lau21, Pet04, Tak14, BJS^{+}21, Boo00, Bre98, CNJ96, Cre97, Hir00, HKS92, Jor99, Now97]. Formula [CS22a, CG09, Duj01, Dum13, EL09, Hum02, JST16, LR07, Rit07]. Formulae [Riv04, AG99, BB97]. Formulas [AKP03, Dur11, MRRTS16, Alm98]. Foulkes [MN05]. Fountain [Mar19]. Four [BSW13, BK16, CGKS20, FS19, Ger03, GW18, KKP^{+}19, Nao21, Sch22c]. Four-Dimensional [BK16, CGKS20]. Four-Manifolds [Nao21]. Four-Step [BSW13]. Four-Strand [FS19]. fourfold [SW93]. Fourfolds [Sta23]. Fourier [Ave10a, Ave10b, BTZK23, BDHZ09, Boo00, CKRS06, DdLV14, Eng09, GJU12, GP92, JRS92, LR92, OTZK22]. Fourth [BR17]. Fractal [AGS18, BHS09, DS14, HS01, KSW01, LvF03, Mer20, Sev18]. Fractals [Bar06, Hew21, BGSS22]. Fraction [Abr22, AK23, BCF04, BC04, CHMW15, Duj01, FIS^{+}14]. Fractions [BCM21b, BCLM17, HV22, LM17a, MU22]. Frame [BDGS23]. Frames [HWW05]. Framework [DSV09]. Frederick [ELT^{+}97]. Free [CJUE04, EFHY21, HP06, HP08, HR14b, HR96, LM06, MSS18, Sil17, Ste21, Wal13, CF96, LL94]. Freeness [Tsu20]. French [BB92, BM94, CNJ96, CD10, Cos02, Dre93, DE93, DO99, Fer92, Fer96, Gra96, HMT10, Kra98, LR02, PR03, RV99, Wal94]. Frequencies [ABG09, ABGS11, DERZ18, RS01]. Frequency [Jen00]. Frequent [McN17]. Fresnel [Goe93]. friables [HMT10]. Frieze [Cun17]. Fringe [GWW16]. Frobenius [Adz17, ABG09, BEZ03, HS05, LRR^{+}19, Van17]. Front [Ano01c, Ano01d, dVV22]. Fuchsian [Ave10a, Mac08]. Fuglede [Bir22]. Fully [KB19]. Function [Ave10a, Ave10b, BK03, BH12, BK12, BM15, BH18, BFFvdW07, BFM13, Car03, De 15b, HNR20, KB17, KvdL04, LSY03, McN17, Rub13, SY01, Sha21, SW12, Wik03, BP92, Boy99, CtR96, CtR97, DR96, Dre93, DE93, DS95, FP99, Klü02, SW99]. Functional [BBW17, HLZ10, MSSS19, WZ21, NR00]. Functions [ACF^{+}21, BCH^{+}14, Ber17, BP21, BA02, BBW17, CC07, CLP15, Coq20, DFK04, DPVZ22, Dok04, Dum01, Eng09, FK10, FK12, GP06, GY10, HS01, JST16, Kat05, Kim22, KKM05, KH19, Mer21, Mil06, Oma01, Ots15, Wan01, Boy98, EIFZ96, Fer92, IKM02, Mar95, Tar92]. Functors [DM22]. Fundamental [BC01, BGBVS20, HOY12, LSY03, JW00].

G [Neb00]. G-modules [Neb00]. Gackstatter [Tha95]. Gain [OB19]. gallery [BP94]. Galois [AAC98, AM92, BL22, CNJ96, Chi24, Die02, Die04a, Dou05, Dou07, Gra96, Gre10, GL22, Hul95, LGRS23, Mal04]. galoisiennes [CNJ96]. Games [ST21]. Gap [AST17, KMN24, KK99, MSV03]. Gasket [DL04, HPR22, OSS03]. Gaskets [CJK^{+}19]. Gauss [HK02b, Har02]. Gaussian [CLM13, CKL^{+}23, GS97, RWW01]. GCD [HMM99, HMM98]. Gelfand [KS22]. Genera [LM19, PR17a, BN97]. General [LM20, Olv01, RD04]. Generalization [BH01, CMPR21, Dou05, Dou07, EM09, Koh18, Svi18]. Generalizations [XZ22]. Generalized [AM22, AS16, BOSS19, Cul19, De 15a, Har11, MNT23, Alm98, BJP94]. Generating [KKM05, BP92]. Generation [Riv04, WS98]. Generations [DRT11]. Generator [Lin04]. Generators [BC02, JST20, MOY20, RW94, SW95]. Generic [Frö22, Leh22, Sne05, BCR22]. Genetic [ERSÉ19]. Genus [ALT^{+}12, AAHwaabAR22, BLP09, BCG19, CH08b, Con18, CP03, Cum04, CD09, Fab09, FG23, JST20, KV17, KH22, LR07, Leh22, LH23, Mac08, ST22, vB22, AT00, GBP97, HW02, Tha95]. Genus- [BLP09, CD09]. genus-one [HW02]. Genus-Two [CH08b]. Geodesic [AR97, FMP04, Mil01, NP23, BGSS22]. Geodesics [CLM13, Kus09, LP14, Lin04, LR05, CS93]. Geometric [Fab09, LR07, SY01, Sin03, HIS16]. Geometrical [ABB11]. Geometries [CMST22, KCK22, Lee04, BDL96]. Geometry [BHM^{+}13, CL07a, CILO19, CS21, CHMWS22, DS13, GM08, LR17, LT09, Mil93, Wil06, Cra99b, DF00, Ver00]. Germain [FJ93]. Germs [Sha19]. Gibbs [GMK19, GP92]. Ginzburg [NR00]. Given [CLMR15, CL07b, Tan09, BP92]. Glimpse [CDS02]. Global [CL07a, GP06, EW98]. Goldbach [HP13, Mar22, RT16]. Goldfeld [BH12]. Golomb [GP94]. Golyshev [BGM22]. Gonal [BRTC23]. Gonality [FG23]. Good [CL07b, Sha21, Nit93, SZ06]. Good-to-Bad [Sha21]. Göpel [FM19]. Gordan [DM06]. Gorenstein [BCR22, BKZ22, HHS21, KW08]. Goresky [Coc05, CMPR21]. Gorshkov [Har11]. Gottschling [HOY12]. Gould [Koh18]. Gradient [ACPR11]. Gram [Sha21]. Grammatical [Adz17]. Granville [Riv04]. Graph [AS12, BS11, CKL^{+}22, HM12, KSY08, LRSST23, ACL01, ST95]. Graphic [BBG^{+}13]. graphics [GP92]. Graphs [ACG^{+}21, ABB22, BCGY23, Chi24, CLP15, CKL^{+}22, DdLV14, EVV16, GKT20, Hay18, HHMP10, IJK^{+}23, KY20, MSSS19, MNS08, MN14, Neu06, OS14, RS19, ST21, ACL01, BB93, Coc05, KO95, KK99, MS05]. Grassmannian [Ano97, CHS96, DHST08]. Greater [BR18]. Green [Ave10a, Ave10b, BS21, Wik03]. Grid [EHPN20]. Gröbner [BHP12, HT00, MS11]. Grothendieck [BPL22]. Group [BPS12, BBG^{+}13, CEM06, DM22, EHT01, ES14, EOT11, FS12, GMR20, GHL19, HHRS21, HR03, JR02, Kob01, KZ12, KS01, LW12, LPU11, LM20, Oha08, RD04, Sch04, Sem23, AT00, BDL96, CHLM97, CM00, DE02, Geb00, Gra96, Hol95, Hul95, KSV97, Pfe97, RW94, SW93, Tah00, Wal94]. groupe [Gra96, Wal94, Wal95]. groupes [Ber95, MSWZ93]. Groups [Aga13, Ale21, AP23, Ave10a, BHK^{+}18, BST20, BBF12, BGBVS20, BCH04, Ben23, BFG^{+}13, BM04, BKM24, Bon21, But11, CHRR07, CH08a, CH13, CH21a, CFH^{+}15, CO02, CS01, Cum04, DdLV14, Del01, DPP11, Der22, DFH23, DJP^{+}05, Dum01, Dum02, DHJ15, EF09, ERvR15, ESTB10, GKS22, GK21, Hol04, HJK^{+}19, Hul01, IJK^{+}23, JM14, JST16, JST20, Kat19, Kid04, Kir11, LVL22, LvT05, Lut13, Mac08, Mal04, Mal10, MOY20, MW22, NSW23, NW05, PR05, RAG18, Rah21, Şen11, Sil17, Sin15, Ter22, VRS03, Wal13, Wil23, dC02, Alp98, EIFZ96, HR92b, HR96, LL94, LLS95, MV00, MPR94, PS96, PS00, ST05, Ano20b]. Growth [AD14, Bri04, DW10, EIFZ96, GJN23, GG06, GG97, Sze98]. Günter [Ano20a]. Gushel [Sta23]. Gyroids [GB97].

Hadamard [AM01, HWW05]. Half [CKRS06]. Half-Integral-Weight [CKRS06]. Hall [EM09, PW01]. Hamiltonian [BCPV21, Hay18, Jor99, ST21]. Hamiltonians [BNH11]. Handel [Bri00]. Harada [RW94]. Hard [BCL^{+}22]. Hardy [ABT96, PE06, Rit07]. Harmonic [BS12, CLW15, GY10, HLLM15, Boy94, IKM02]. Harmonics [Ber24]. Hasse [BG97]. Hassett [ADGH20]. Hatchet [LT09]. Hausdorff [De 15b, KK97]. Having [De 15a, Fla08, MP01, Sta21, Jun00, Ste93]. Hecke [AAC98, Alv08, Gun00, HV19, KO95, KW08, LS02, Oka02, Ots15, vGvdKTV97]. Heegner [Hop10]. Heel [FWW02]. Heesch [MT16]. Height [BGKMI01, DH04, BM00, EW98]. Heights [BCH^{+}14, BdJS23]. Heisenberg [DS19]. helicoid [HW02]. Hénon [Ara07, FG99, JJJ02]. Hermite [HMM99, BCIO01, HMM98]. Hermitian [DS21, KN22]. Heron [GJL03]. Hesselink [Lee23]. Heuristic [Pau17]. Heuristics [BB21, Del01, GL23, HMM05, KLNP20, KW23, Lip21, Poo06, Wat08, WtR03]. Hexacarpet [BKN^{+}12]. hexagon [Smi00]. Hidden [BW07, Ste17]. Hierarchical [BS17]. High [AAHwaabAR22, JZ05, KS01, MV10, RSÉ03, Rub13, BJL05, KK99, Tom96a]. High-Accuracy [MV10]. high-dimensional [Tom96a]. High-Order [JZ05]. high-precision [BJL05]. Higher [BCC10, CH13, Con18, Duk23, Dur11, ER15, FK12, Hop10, Hor14, Tha95, vB22, JK95, ST05]. Higher-Dimensional [BCC10]. Higher-genus [Tha95]. Higher-Order [Dur11, Hor14]. Higher-Weight [Hop10]. Highly [CH21a]. Hilbert [BCOS24, Bir21, BHI^{+}11, CLMR15, CD09, Dem05, Frö22, HS14, HHS21, HM12, IZ14, Lee23, May10, MS11, Rob00]. Hirota [PPS09]. Hirschhorn [XZ22]. Hitting [MTC14]. Hives [Lom20]. Hodge [BdJS23, DN21]. HOL [DKAP22, KALP22]. Holds [Bir22]. Hole [BRS^{+}18]. Holes [MVG05]. Holomorphic [CR01, JST16, Kaw18, LSW99]. Holonomy [Gui18]. Homeomorphisms [BS01, Bri00]. HOMFLY [CG23]. HOMFLY-PT [CG23]. Homoclinic [Col05, GN09]. Homologies [DGR06]. Homology [BCGY23, CG23, Fuk11, GOR13, GL15, HLMZ01, Jon10, KLNP20, LRSST23, MPS^{+}18, Muk22, Oha08, Tan09, Yun22, BL00]. homometric [GP94]. Homomorphism [LS03]. homomorphisms [DU95]. homophones [MSWZ93]. Homotopy [BL12, EL14, JLY16, LOS20, Roo08]. Honey [Lom20]. Hook [Han09]. Horn [BP21]. Horseshoe [Col05, HdC02]. House [Sta21]. huge [PS92]. HUM [LN10]. Hundred [Sch09]. Hurwitz [BN97, BBvBK15, HV22]. Hyper [Coq20]. Hyper-Roots [Coq20]. Hyperbolic [Ara07, Bor14, CDV22, CH21a, CLLM07, CLM13, CV03, Coi22, CLT06, DPP11, Der22, DFJ12, FGG^{+}16, FMP04, GR22, GMK19, GHH08, GK21, HHMP10, HIK^{+}16, Ino22, KV17, KM18, NSYY08, RT23, Roe07, SYY08, Zha22, But05, GH96, HW94, RT00, RS98]. Hyperboloid [ST02b]. Hyperelliptic [BBW17, Web97, vB22, AT00]. Hypergeometric [BBBC07, BST20, DPVZ22, DFH23, Fou21, MS01, NSYY08, SY01, SYY08, Sin15]. Hypergeometric-Type [MS01]. Hypergraph [SW10]. Hyperkähler [Saw23]. Hypermetric [DD03b]. Hyperplane [Roo08, Sne05]. Hyperplanes [Leh22, Pra05]. Hypersurfaces [BK16, GIKW23, JK01, Lee23, Tse22]. hypothèse [LR02]. Hypotheses [BGM22]. Hypothesis [Boo06, Bri06, Kaw18, Vor20, Bla22, LR02].

Icosahedral [JRS03, Tsu20]. Ideal [CV03, Raw03]. Ideals [AM22, BCR22, Frö22, KSV22, HT00, KSV21]. identifying [Boo00]. Identities [Abl17a, Abl22, BBB06, BHP12, KR15, KS22, OS15]. IdentityFinder [KR15]. Igusa [ALT^{+}12]. II [ABL17b, BTZK23, BC04, Bra21, CL07a, Cha23, CMPR21, FI18, GP06, GIKW23, Mal04, MOY20, OS22a, Pat05, Sch09]. III [Elk01]. Im [LRSST23]. Images [Chi24, Die02]. Imaginary [FK02, HNR20, HJK^{+}19, SY01, Lou98]. immersed [Ran99]. implementation [Bri00, EH96, HR92b]. Imprimitive [CH21a, DJ10]. improvement [ST99]. Improving [Cd18]. incommensurably [VS00]. Indecomposable [CH13, Pet04]. Independence [EVV16]. Independent [RT16]. Index [Der22, Kaw18, GPP94]. indications [AM92]. Inequalities [CHMW15, Gar02]. Inequality [BTZK23, Boe23, DP15, HIK22, PE06, OTZK22]. Inexact [BHM^{+}13]. Infimum [Bad16]. Infinite [Abl17a, Abl22, ABL17b, BCH04, HH22a, Joh08, KALP22, Law17, Smi00, Hol95, Tha95]. Infinitely [RR21]. Infinity [RS22]. Inflection [CL22]. Informal [Lan98]. information [Mur93]. Ingleton [Boe23]. Inhomogeneous [PW01]. Initial [BCR22]. inspire [PZHC09]. Instructions [Ano92a, Ano93]. Integer [EMRSÉ14, Har11, HH22a, Shl21, Dil00]. Integers [GKS22, GPS18, GMY21, Shp16, Sta21, Svi18, BtR96, Del98]. Integrability [PPS09, ST19a]. Integrable [PPSF17]. Integral [Bru07, CKRS06, DF00, DL04, Duq01, FS11, Gaá01, GOP02, GP06, GR19, MSS15, Şen11, Web09, FS98, GPP21, HKS92, Ola05, VV99]. Integrals [BBBC07, BCC10, BSW13, Jor99]. Integrated [DEFM23]. Integration [EJ22, GM18, JZ05]. Interactive [Kon22]. Interior [HP14]. Internal [Tan24]. Interpolation [Coi22, GMRST14, IC16]. interpretation [EW98]. Intersection [CP10, CP12, Cha14, HLM23, HLZ10]. Intersections [ABL17b, ST95]. Interval [GMRST14, CPS^{+}01]. Intervals [Bra21, GL23]. Intrinsic [HY22]. Intrinsically [MN14]. Introducing [Bur04]. introduction [BD09]. Invariant [BMS21, BH04, HV22, KKM^{+}01, Koh18, KZ12, MRS22, Wil06, CHK93, FH99, GR00, Tom96a, Tom96b]. Invariants [AER^{+}08, Ano20b, BNHLR11, Bon21, DG21, Fuk11, GSS16, JN06, JST20, KSV97, KM18, MMO^{+}02, Neb00, PR05, BG97, CGHN00, IKM02, ST05, Wil02]. Inversion [CO20]. Inversions [KSY08]. Investigating [Wan09]. Investigation [MW22, JLW95]. Investigations [DD03a, Mil06]. Involutions [CO02]. Involving [Svi18]. Irrational [GM05, GW18]. Irrationalities [BCM21a]. Irrationality [Bri01, KALP22]. Irreducibility [BBB^{+}18, GS05]. Irreducible [BH04, ES14, EMRSÉ14, HR03, KB19, – Dok00]. Isabelle [DKAP22, KALP22]. Isabelle/HOL [DKAP22, KALP22]. Ising [BBBC07]. Ising-Class [BBBC07]. Islands [BCPV21]. Isogeneous [EF18]. Isogeny [Chi24]. Isogeny-Torsion [Chi24]. Isolated [Sch19]. Isometries [BP07, HW94]. Isomorphism [LL94]. Isospectral [Lau21]. iss [Ano00]. Issue [Kon22]. Iterated [Das23, De 15b, BP94, Mil92]. Iterates [GXB15]. Iterating [CtR96, CtR97]. iteration [Cra99b, Cra01]. Iterations [She09, Chu05]. Iwasawa [ST05].

J [ELT^{+}97]. Jablan [BH21]. Jacobi [AGS18, Ots15, Sch07, Sin03]. Jacobi-Sum [Ots15]. Jacobian [PR17a]. Jacobians [BLP09, CLP15, Cul19, EF18, HM02, LGRS23]. Jacquet [CD09]. Jain [Ber17]. Jean [Ano20a]. Jean-Philippe [Ano20a]. John [PZHC09]. Johnson [ACG^{+}21]. Jointly [BL07]. Jonathan [Bai17]. Jones [DH97]. Journal [ELdlL92]. Jr [ELT^{+}97]. Julia [Mil04, SW22b]. Jumping [ORW99].

K3 [Bro07, CT14, FLST23]. Kac [Hen23]. Kanenobu [Sto00]. Kaplansky [DHJ15]. Kashaev [MMO^{+}02]. Kazhdan [FK19, NT15, Tah00, War11, dC02]. Kedlaya [GG03]. Keiper [Vor20]. Keiper/Li [Vor20]. Kelvin [Oud11]. Kepler [DS19]. kernel [DU95]. Khovanov [GOR13, LOS20, MPS^{+}18, Muk22]. Kim [KT16]. Kimura [PPS09]. Kimura-Type [PPS09]. Kind [Gui03]. Kirillov [KM18]. Kissing [Cd18, MV10]. Klapper [CMPR21]. Klein [BJS^{+}21, PR21]. Kleinian [MOY20, MPR94]. Kloosterman [Boo00]. Knot [ACPR11, BKM24, BCG19, Cha23, DERZ18, DGR06, FGK^{+}16, Kho03, KS05, Mil01, AR97, MR00, Ran99]. Knots [Ano23, BNHLR11, BDHZ09, CN13, DERZ18, DFJ12, GK12, GSS16, GOR13, GL15, HIK22, Hor14, KA24, Lam21, LM19, Mil01, MMO^{+}02, OS22b, Raw03, RR21, SZJ04, ABG^{+}21, KK93, Sto00]. Knotted [MN14]. Known [EGSC18, FKS15, FKS19]. Knuth [Pau96]. Kobayashi [BFK^{+}92]. Koecher [Riv04]. Kohnert [AS22]. Kolakoski [She20]. Kontsevich [BK22]. Koszul [DZ11]. Kottwitz [Coc05]. Kronecker [HR14a, JST16]. Kronheimer [Boo21]. Kubilius [HMT10]. Kummer [RSSS13]. Kuttner [GKR01].

Labbé [Ano20a]. ladder [CLZ92]. Ladders [Pec19]. Lagrange [BJP94]. Lagrangians [AP23]. Laguerre [Cul19]. Lamé [Joh08]. Landau [NR00]. Langlands [CD09]. Laplacian [Bar06, BHS09, CS21, GH96, KMN24, KSW01, NR99]. Laplacians [DEFM23, SW22b, Ter96]. Large [Ale19, ABB22, AAB^{+}21, BBKL17, But11, CS22b, LN01, Mül06, Nao21, OS14, dVV22, BtR96, Ros00]. large-prime [BtR96]. Large-Time [dVV22]. Largest [McN17, MNS08]. Last [Sin03]. Latin [FJ96]. Lattice [All05, AG22, AAHwaabAR22, CCG21, Der22, Elk01, Hen18, HM21, Yi19, HMM98, HMM99]. Lattice-Packing [AG22]. Lattices [Coq20, Der05, DS21, Elk01, FGT10, Gar21a, HKPP13, IKMF16, KN22, OB19, BB92, BN97, BM94, BQS95, Chu05, FPUP22]. Laurent [HHS21]. law [ACL01]. LCM [IKMF16]. Leading [CFH^{+}21]. Lean [BHL^{+}22]. Learning [GHK20]. Least [Oud11, Xu04, Web97]. Lebesgue [IC16]. Leech [All05]. Legendre [Kat10b, Sto17a]. Legendrian [CN13]. LEGO [ArE11]. Lehmer [Sto17b]. Lei [Mil93]. Length [BO17, CLM13, Cha14, Hew21, MRRTS16, MW22, VRS03, Wal13, dVV22, HW94]. Length-Equivalent [Cha14]. Lengths [GJL03, Han09]. Lens [Lau21, Tan09]. Lenstra [CIO07, KLNP20, WtR03]. Less [CHRR07, CDV22, CP03, GP94, MR95]. Level [Die04b]. Levels [PT08]. Lex [BCOS24]. Li [Vor20]. Library [BL14]. libre [Cou94]. libres [MSWZ93]. lié [Cos02, BJS^{+}21, BLS22, BH04, JN06, LVL22, MSS18, Rit07, Roo08, Sch05, Soh21, dG05]. Life [GM05, BC98]. Lift [KT16]. Lifts [IT19]. Like [CS22a, Gor21, KL16, Riv04, AG99, AG12, BBB06, BB97]. Limit [AAB^{+}21, BM05, ERSÉ19, Joh11, JST16, KK97, MPR94]. Limited [BTZK23, OTZK22]. Limits [BdJS23]. Line [BK12, BH18, Rub13, STZ95]. Linear [AST17, Bre08a, CL22, HS14, LN10, LM20, RD04, HR92b, Ros97]. Lines [Esc14, Sch22c, Ano97, CHS96]. Link [Gör15]. Links [BR18, HIK22, MMO^{+}02, He02, Koh18]. Liquid [Sch22a]. Lissajous [BDHZ09]. List [Ino22]. Littlewood [Bad16, PE06]. Ljunggren [BOSS19]. Local [BY12, BFG^{+}13, CL07a, Fou21, ST19a, Mar95]. Loci [Alu15, Chi04, iIS04, LM22, SS06, ST02a]. locking [Jen00]. Löffler [BG95].

Log [Hye17, LM06]. Log-Concavity [LM06]. logarithm [ST99]. Logarithmic [CJUE04, DJP^{+}05, LvT05, RAG18, Ros97]. Logarithms [BCLM17, dW98]. Loki [ST02a]. longest [ST95]. Loop [GSS16]. Loops [SW22a]. Lorenz [Win22]. Lovász [DdLV14]. Low [EL14, LRSST23, MS11, Ter22, Web09, GBP97, PS00]. Low-Degree [MS11, Web09]. Lower [AJK08, ACF^{+}21, BGBVS20, BÖW18, Fer22, GKT20, HS18, HIL13, MVG05, Soh21, TS06]. Lower-Dimensional [Soh21]. Lower-Order [ACF^{+}21]. Lusztig [War11, dC02]. Lyapunov [Fou21, HK02b]. Lyons [Geb00].

M [Ano20a]. Maass [BS12, HR92a]. Machine [GHK20]. Maclaurin [CHMW15]. MacPherson [Coc05]. Madelung [Cra99a]. Magmatic [CCG21]. Mahler [Boy98, BM05, CDD^{+}16, CM21, ERSÉ19, LH23, MS05, MRW08, RSÉ03, Sil95]. Main [RW11]. Majorations [Dre93, DE93]. Maker [ST21]. Maker-Breaker [ST21]. Manifold [BFG^{+}13, Bur04, Wal03]. Manifolds [Ban06, BO17, BB22, BDSS18, CLLM07, CGKS20, Con06, CLT06, Der15, DHST08, FGG^{+}16, FMP04, GHH08, HPP08, HIK^{+}16, Lut09, MP01, Nao21, RT23, RSSS06, Saw23, SZ04, ST22, Zha22, BL00, But05, CGHN00, HW94, Jor99, KSV97, Mat98, RT00, RS05]. Manin [AKN^{+}16, Poo06]. Many [BCG10, FG23, FIK20, Gal14, RR21, VV99]. Map [ALT^{+}12, Ara07, ADDW18, BG22, CL07a, Col05, Mat01, Nil03, NSYY08, Olv01, SYY08, Sch01, Sch22b, Sha19, BÁ00, Sch92, Tom96b]. Map-Germs [Sha19]. Mapping [BS01, AT00, KSV97, Tah00]. Maps [BBvBK15, Bri04, Das23, GN09, HdJF06, HLMZ01, JJJ02, KY17, Mal22, PPSF17, RS22, BGSS22, BSTV93, FG99, Lan98, Mil92, Mil93, Mil00, Mur93, PdlL02, SL00, Tom96a]. Marching [CMST22]. Marked [BCR22, BCOS24, MT16]. Markoff [FLST23, dCIL22]. Markov [BW07, CR22, LW12, MV00]. marks [GP94, Pfe97]. Marmi [Car03]. Matching [Jon10]. materials [KK99]. Math [Ano00, Ano01a, Ano01b, Ano01c, Ano01d, Ano03a, Ano03b, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano11c]. Mathematician [Bai17]. Mathematics [GW18, PZHC09, BD09]. Mathieu [BDL96, CEM06, EOT11, SW93]. Mating [Mil04]. matings [SL00]. Matrices [AGS18, BCM21a, CILO19, EGSC18, FIK20, CR99, PS92]. Matrix [AAB^{+}21, CILO19, CKRS06, Dur11, GdlI07, Kir11, LR17, OS14, RW94, CPS^{+}01]. Matrix-Valued [GdlI07]. Matroids [GHSV23, MP15]. Maximal [Fir15, GL22, LVL22, BDL96, GR00]. Maximally [Sin15]. maximizations [Dre93, DE93]. Maximizers [BTZK23, OTZK22]. Maximum [AJK08, BU23]. McKay [Chu05]. Mean [CV03, GIKW23, LR07, May01, Ume09, Cho94, GBP97, GB97, HKS92, KMS00, RS98]. Means [JZ05]. Measure [BM05, CDD^{+}16, ERSÉ19, HV22, LH23, RSÉ03, Boy98, MS05, Sil95]. Measures [AGS18, CM21, GMK19, HOS11, MRW08, NTTY22, Bri01, JRS92, LSV98, STZ95]. MeatAxe [IL00]. Medium [GG03]. Megaforms [SSS03]. Meissel [LZ10]. Melnikov [DRR99]. memoriam [ELT^{+}97]. Menzin [LT09]. Mersenne [CFH^{+}21]. Mertens [KvdL04, LZ10]. Message [Ano92b, EL92]. Meta [DRT11]. Meta-Fibonacci [DRT11]. Method [BA02, Boo06, Gui06b, Roe07, TTS17, BHvdP95, DRR99, FG99, GPP21, SdW94, ST99]. methodology [Jor99]. Methods [DSS07, DS19, EMRSÉ14, Gal14, JZ05, RSÉ03, CPS^{+}01, SW99]. Metric [ABB22, EM02, Muk17, BFK^{+}92, Mur93]. Metrics [BGKT05, Hal14]. Metropolis [DN04]. Million [CHRR07]. minima [Bre98]. Minimal [BDdG07, BO17, BNH11, BNL17, BDS16, BMN22, BDSS18, Che19, CH08b, CLMR15, CDD^{+}16, Con18, EMRSÉ14, ESTB10, Hol04, Hye17, Jai09, Law17, LvT05, MRW08, RAG18, Tra08, VRS03, CHK93, CS93, Now97, PP93, RTW00]. Minimization [HMM05]. minimizers [He02]. Minimizing [BBC^{+}09b, HKS92]. Minimum [DJ10, HS07, Hay18, MTC14, FP92, FP93, FPDH98]. Mining [XKA14]. Minkowski [Ume09]. Minuscule [Hop22]. Mirzakhani [Bel21]. Miscellaneous [Ano01a, Ano01b, Ano01c, Ano01d, Ano11c]. Missing [HJK^{+}19, LMO13]. moats [GS97]. Möbius [CO20, DR96, Dre93, DE93, He02, KK93]. mod [AAC98]. mod- [AAC98]. Modality [GMR20, JR02]. Model [CMOS13, Hye17, KH19, ACL01]. Modeling [GG06]. Models [AD14, ABB11, BS17, BW07]. Modifications [CM16]. modulaire [Hal98, HK03]. modulaires [Lep93]. Modular [BK03, Bir21, CD09, Dem05, Dem08, Die04b, Dum13, FHL^{+}10, GJU12, GHL19, GM13, GM18, GHY13, JST16, Kat05, Kat10a, KSY08, MR21, Tak14, Wat02, BG97, Bre98, Chu05, Cre97, Hir00, Hun00, OB19]. Module [IZ14]. Modules [CH13, Hor14, LL21, LS03, LS07, Sne05, Web09, Neb00, RV99]. Moduli [ADGH20, BvBK09, CMZ21, HT19, Hye17, Tse22, Yun22]. Modulo [APS04, BHK05, Tan24, Vin01]. Moment [ACF^{+}21]. Moments [AR12, BS08a, DLN23]. Mond [Sha19]. Monic [BGMR06, EMRSÉ14]. Monochromatic [BCG10, EFMPR18, RCD21, Jun00]. Monodromy [BST20, BR17, DFH23, ST19b, Sin15]. Monogenity [GR19]. Monomial [AM22]. Monomials [AS12]. Monotone [HHdC^{+}15, PZ04]. monotonicity [FJS94]. Monster [CL04, HLM03]. MOP [JR02]. Mordell [BB18, Elk01, GPP21]. Mori [He21]. Morphic [EFMPR18]. Morse [BL14, DF00, Eng09, KKM05, LLT03]. Mosaic [Gui12]. Mosaics [DL22]. Most [HM21, MP01, McN17]. Mostow [Der05, FPUP22]. Motifs [GWW16]. Motion [Ino13, CHK93]. Motives [DPVZ22]. Motivic [CG09, Dok04]. Moussa [Car03]. Moving [Rom18, RTY21]. Mrowka [Boo21]. Mukai [Sta23]. Multidimensional [FIS^{+}14]. Multigraded [HM12, IZ14]. Multipartite [ACG^{+}21]. Multiple [GR22, BBK01]. Multiplication [CILO19, EJ22, LR17]. Multiplicative [ABL17b, Shp16]. Multiplicities [AM22]. Multiplicity [Lee23]. Multiplier [BG22]. Multipliers [DM22, JM14]. Multiply [Gir03, NR00]. multiply-connected [NR00]. Multiprecision [BM15]. Multireflection [HR14b]. Multivariate [GS05]. Multizeta [KL16]. Mumford [CLMR15]. Mutation [Law17, MO22]. Mutation-Infinite [Law17]. Mutations [BSW24].

Nagell [BOSS19]. Nash [ALP^{+}11]. Natural [Das23]. Nature [BBG^{+}13]. Near [Sch06, Mil06]. Necessarily [Roo08]. negative [Ste93]. Neighborhood [GLP16]. Neighborly [MP15]. Neighbors [CHM13]. Neighbourly [Nil03]. Néron [CT14]. Nested [Dil00]. Nets [NP23]. Neumann [HR92b]. Newforms [GJU12]. Newton [BW07, Duj01, Roe07]. Nice [Hun00]. Nilpotent [EM17, ESTB10, JN06, LMW04, LVL22, Wil23, LM97, Sch05]. No [Boe23, MTC14, Wal03, Ano03a, Ano03b, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano11c, GP94, MR95]. Nodal [ABB22, BNH11]. nodoids [Ros05]. Noetherian [DSS07]. Noisy [GMRST14]. nombres [Cos02]. Non [BR17, Chi24, DN21, Gou97, HIS16, Lam18, LM22, LM17b, MRS22, ST22, Cou94, Pfe21]. Non-Abelian [DN21]. Non-CM [Chi24]. Non-Crystallographic [Lam18]. Non-degeneracy [MRS22]. Non-Existence [LM22]. Non-geometric [HIS16]. Non-Normal [LM17b]. Non-ordinary [Gou97]. Non-Orientable [ST22]. non-realizability [Pfe21]. Non-symplectically [BR17]. Nonalternating [DERZ18]. Nonassociative [BH04]. Noncollision [Ger03]. Noncommutative [GK12]. Noncompact [Bar06, GIKW23]. nonconformal [BSTV93]. Noncongruence [FHL^{+}10]. Noncrossing [CG14]. Noncyclotomic [FK02]. Nondegenerate [BLS22]. Nonelliptic [Die04a]. Nonexistence [Nil03, DL92]. Nongeometrically [GM13]. Nonhyperbolic [Gra02]. NonIntegral [RR21]. Nonlattice [DS14]. Nonlinear [BBC09a, Neu06, FPU55]. Nonorientable [Bur07]. Nonsymmetric [Ano23, Lam21]. Nontrivial [KS01, MNS08, – Dok00]. norm [Ste93]. Normal [BC02, BS17, Bur10, LM17b, HMM98, HMM99, Jor99, Ran99]. Normality [BBC^{+}12, Lag01, Sev18, XKA14]. normalization [Rim92]. Norms [ABT96, LV15]. Norton [RW94]. Note [BBF12, Bri04, CB09, CLP15, Els14, FI18, HP18, Hen18, KMP22, Ano97]. Notes [EOT11, Wei16]. Novel [Bra21]. Nuclear [HP08]. Nucleation [GG97]. Nucleus [HP06]. Number [Aic04, Ano20a, BCIO01, BR18, BMS21, Cd18, CIO07, DdLV14, DJP^{+}05, DR08, ES22, FI18, FK09, GKT20, HST22, Hay18, HJK^{+}19, KN22, KP08, Kou19, LRZ20, Lip21, Mal10, MTC14, Mer20, Pio07, SZJ04, SP05, Shl21, Wag01, Wan01, ABG^{+}21, Ano01a, Ano01b, Ano01c, Ano01d, Cos02, EH96, FM00, GPP94, Lou98, Rog00, Ste93, WS98]. Numbers [Adz17, AS16, AMM08, AFMS17, BC02, BCM21b, BMRS19, Bri06, BDS16, BV12, CFH^{+}21, CGM15, CP10, CP12, Cha14, Das23, Dol18, Eel07, GKS22, GW18, HLM23, KB17, KM20, MT16, MU22, MV10, MGO22, Wag01, Lou98, MS05, Sil95, Thu94]. Numerical [Ave10b, AD21, BHM^{+}13, BM15, BL12, BDGS23, Ble11, BNH11, Bra95, But12, Car03, CGJM20, DD03a, DN04, DS19, DN21, GIKW23, GMK19, HP18, HP13, Hir00, JRS92, JRS03, JZ05, NR99, Now97, RW11, Tom96b, Ver00, YZ22, vB22, GH96, HMT10, Jor99, KSV97, Kob94, LR92, LR02, PZHC09]. Numerically [BMN22, ER19, Tos22]. numérique [HMT10]. numériques [LR02]. Nyman [LR02].

Observations [Bur07]. Observed [FS19]. Obstruction [Poo06]. Obtain [Gui06b]. Obtuse [Sch06, Sch09]. Occur [BU23]. Ochoa [SdW94]. octaédral [CNJ96]. Octagon [Bur10]. octahedal [CNJ96]. Octahedral [HPP08]. Octahedron [De 06b, Sch22b, De 06a]. Octic [Oma01]. Odd [ACG^{+}21, Ano20a, HM21, KB17, LRZ20, Muk22, SZ04, Tan24, Jun00]. ODEs [JZ05]. Off [kAL01, Viv94]. Off-Center [kAL01]. Once [Gar23, HT19, KSWY06]. Once-Punctured [Gar23, HT19, KSWY06]. One [CH21a, CS22a, HV19, JST20, Mat01, PW01, RT23, Rub13, Sch09, Wag01, Caz97, GH96, HW02, LP97]. One-Dimensional [Mat01, Caz97]. one-equator [LP97]. One-Relator [CH21a]. Onset [HRSS17]. Open [BCOS24, Han09, NP23]. Operad [CCG21, DZ11]. Operads [CG14]. Operator [AL03, HY22, KZ08, LN10, MS01]. Operators [BR17, DSS07, ER15, GdlI07, Sne05, AAC98, ABT96, KK99, MV00]. Opers [DN21]. OPTi [Wad06]. Optical [TTS17]. Optima [EFHY21]. Optimal [CCEK07, MP22, TS06, Nur00]. Optimality [LLT03]. Optimization [EMRSÉ14, HKPP13, NT15]. Orbifold [BK16]. Orbit [Gal14, MSV03, RS22, Lan98]. Orbits [All05, BDdG07, Col05, DS19, FLST23, Gal14, JN06, KL21, LMW04, LN01, She20, KdC05, Tom96a]. Order [ACF^{+}21, BLP09, BR17, CHRR07, De 15a, Dur11, EL14, Hor14, JM14, JZ05, KKSV16, KS05, KvdL04, LVL22, BN97, CLZ92, DL92, FJ96]. Ordered [Pec19]. Ordering [Mar20]. Orders [Aga13, GOP02, HRSS17, Kob01, GR00]. Ordinal [DKAP22]. ordinary [Gou97]. Orientable [HPP08, ST22]. Oriented [MP15]. Origami [Sch04]. Original [CS22a]. Orlik [MPY17]. Orthogonal [BST20, Dur11, DPV15, FIK20, GdlI07, HOS11, NTTY22, Sin15, Neb00]. orthonormal [CPS^{+}01]. Orthostochastic [CD21]. Oscillation [LT09]. Oscillator [ST19b]. oscillators [MP94]. Outer [BHS09, BMS21, DPPS05, Jeo15, Lut13]. Outline [Hye17]. Overconvergent [Bir21]. Overlaps [NTTY22].

Package [BRTC23, CMZ21, JZ05, MSV03]. Packard [GG06]. Packing [Ano97, AG22, CHS96, FGJ22, MP22, DS95]. Packings [BH01, DHST08, Fer22, FS11, GR22, Hen18, MVG05, TS06, Zon22, Ano97, CHS96, MMT00]. Pages [Ano01a, Ano01b, Ano01c, Ano01d]. Painlevé [Bra21, MR10]. Pair [Bur14]. Pairs [ABGS11, SS11, Sto17b, GP94]. Palindromic [Sil17]. Paper [KSV22]. Parabolic [Fou21, JR02, MOY20, RTY21]. Parameter [BCF04, BC04, GN09]. parameters [GG97]. parfaits [BM94]. Part [BDM07, CMPR21]. Partial [BFFvdW07, Gar17, KS05, Neu06, Vin01, Boy94]. Partially [GMK19, Pec19]. Partition [BHP12, DKAP22, Han09, Tan24]. Partitioned [Shl21]. Partitions [Abr22, Alm02, BO17, BNH11, BNL17, Oud11, Shl21, Wei16, FJS94]. Parts [Alm02, Tan24]. Pasta [PZHC09]. Pasting [Mil04]. Path [LL21]. Paths [KH19]. Patterns [BCG10, Cun17, Fin92]. Patters [Kho03]. Peaceable [YZ22]. Peak [BFT16]. Peeling [EHPN20]. Pell [JW00]. Penrose [DEFM23]. Pentagram [Sch01, Sch13b, Sch92]. percolation [Var98]. Perelman [Hal14]. Perfect [Aic04, BM94]. Perimeter [Oud11]. Period [EJ22, PW01]. Period-One [PW01]. Periodic [BCM21b, BL07, Con18, De 06a, De 06b, DS13, GR00, GS97, KV17, Lin04, LR05, OS14, Sch09, BC98, CHK93, CD10, HW02, KdC05, RTW00, Tom96a]. Periodicity [EM17, FS19, Lam18, Viv94, CF96]. périodiques [CD10]. Periodized [GGM11]. Periods [CGKS20, DW10, GM18, Cre97]. Peripheral [Gui18]. Permutation [CH08a, ESTB10, NSW23]. Permutations [CGP23]. Perron [Sch07]. Persistence [dFT14]. Perturbation [BFM13, Tom96b]. Perturbations [Gar21a]. PETs [Yi19]. Pezzo [CS13]. Phase [BCPV21, Guc95]. phenomenon [GP92]. Philippe [Ano20a]. Physics [BBC09a]. Pi [BBBR17, Gan17, Sev18]. Pick [EL09]. Picture [Ino13]. pieces [GM99]. Piecewise [BP07]. Pinching [GHK03]. Ping [Tsu16]. Ping-Pong [Tsu16]. Pisot [MS05, Mer20]. PL [ST22]. Planar [AF06, Guc95]. Plane [BvBK09, CR22, CP12, CM16, De 06a, De 06b, GK21, HdJF06, KY17, LRR^{+}19, MP22, Pio07, SZJ04, BSTV93, Boy99, VV99]. Planes [KV17, Ano97, CHS96, DL92]. Planimeter [LT09]. Plateaus [Ara07]. Platonic [Alp98, AAHwaabAR22]. Plücker [Pfe21]. Plumbed [Fuk11]. Pluricomplex [Wik03]. Pochhammer [Abl22]. Poincaré [BL00, HLMZ01, Sne05]. Poincaré-Betti [Sne05]. Point [BBC^{+}09b, BLP09, BRS^{+}18, BB22, De 15a, EJ10, KK12, KKM05, Mil06, Sch22b, Sha21, Ros05, Wal95]. Points [BG22, BCR22, BCOS24, Ber95, BM05, BL07, BDP16, BGKMI01, BS08b, CB09, CL22, DLS22, Dum02, Duq01, ERSÉ19, Els11, FG23, GG03, Gra02, GM13, Hop10, IC16, KS03, MP22, Poo01, RTY21, WZ21, GPP21, Gal99, Maz92, Mil00, NR00, VV99, Wal94]. Poisson [BBKL17, EGP^{+}12]. Polarization [EF18]. Polarized [Bro07]. Polars [PR21]. Poles [ST02b, Wik03]. Pólya [PE06]. Polyforms [MT16]. Polygon [KB17]. Polygonal [BP07]. Polygons [AST22, ES22]. Polyhedra [De 06a, De 06b, Fir15, Ino22, JLS20, KM18, Roe07, GH96]. Polyhedral [EM02, Nil03]. Polyhedron [Fir15]. polylogarithm [CLZ92]. Polynômes [DO99]. Polynomial [ABZ07, BGMR06, Cd18, CLMR15, CGP23, ER15, GK12, IKM02, KB19, Koh18, KMP22, MS01, Nic18, DH97, DU95, Sot00].

Polynomials [AS22, BBKL17, BCH^{+}14, BFT16, BBB^{+}18, BM00, BL21, CDV22, CLP15, CH21b, Cul19, DS13, DS15, DS14, DFJ12, Dur11, DPV15, Eel07, EMRSÉ14, ET14, Fla08, GMRST14, GXB15, GS05, GdlI07, Har11, HLLM15, HOS11, Hen23, LPU11, LH23, LM06, MSSS19, MNT23, Muk21, NTTY22, Par21, PR08, RSÉ03, Shi16, Ste21, Tha08, War11, Xu04, dC02, BB93, BC99, DF00, Sch99, SL00, DO99]. Polytope [Shl21, CR99, CRY00]. Polytopes [Ale19, BW07, BS11, DY07, Gar21a, LM17b, MP15, SZ04, IKM02]. Pong [Tsu16]. Pop [CGP23]. Pop-Stacked [CGP23]. portraits [Guc95]. Positive [BS21, KL16, LW12, MW22, Pfe21]. Positivities [HIK22]. Positivity [KZ08]. Positivity-Preserving [KZ08]. Potential [BDGS23]. Potentials [BSW24]. pour [LR02]. Power [Gaá01, GOP02, MSW22, Ola05, SS06, ACL01]. Powerful [vdP01]. Powerfully [Wil23]. Powers [BOSS19, HNR20, Ste21, Svi18, Tan24, HS05, JK95]. practical [BJP94]. Praeger [HR92b]. precision [BJL05]. Prehomogeneous [JN06]. premières [DO99]. premiers [Cos02]. prenant [DO99]. Prescribed [Kou19, SZ04, BHvdP95, BM00]. Presentations [BM04, CFH^{+}15, EHT01, ES14, HR03, IJK^{+}23, NW05, Geb00]. Presented [But11, DHJ15, ERvR15, Hul01, MW22, Roo08, HR96]. Preserving [BLS22, KZ08]. Pretzel [GK12]. Prim [GJL03]. Primality [Kid04, Kob01]. Prime [ABGS11, BDS16, DERZ18, KW08, Kim22, KS08, McN17, MU22, Shi05, Var98, Wag01, BtR96, Cos02, FM00]. prime-number [Cos02]. primeness [WPW97]. Primes [CKL^{+}23, CL07b, DT02, FS10, GKL19, GL23, KT16, Kou19, Mar22, Mül06, Gou97]. Primitive [GKL19, Lee04, Yab09, LRSST23, LLS95, CNJ96]. primitives [CNJ96]. Principal [AF06, BR18, CHLM97, DO99]. Principle [BDSS18]. Prismatoids [CS22b]. Probabilistic [HH22b]. Probability [BGMR06, GWW16]. Problem [BEZ03, BOSS19, DS19, EH04, FK09, GKR01, GHK20, HHRS21, HKPP13, HMM05, Har11, HH22a, HJL22, HH22b, MO07, MTC14, MT15, Nic18, Pau17, PZ04, Rom18, Tsu16, YZ22, dFT14, BJS^{+}21, Kra94, LSW99, Rob00, TW99, Sch13a]. Problems [Han09, NP23, Bra95, FPU55]. Product [BDM^{+}18, FIK20, HLM03, Sch99]. Products [BZ17, FS10, Frö22, LM17a, May10, OS16, Wal13, Dil00]. Profinite [Gar21b]. Program [Hye17]. Programing [HH22a]. Programming [Cd18, FK19, MV10, Ros97]. Programs [EFHY21]. Progressions [GST06, HR14a, LZ10, Mar22, PR17b, RCD21, SW12, Bre99, Jun00]. Projection [DHST08]. Projections [Alu15]. Projective [AST22, BK16, CH13, De 15b, Frö22, HT19, JK01, KY17, Lom02, MP22, OS16, DL92, MR00]. Projectivity [Sca09]. Proof [BS08a, BK22, CG09, Hol95, Pau96]. Proofs [Mel22, Guc95]. Properties [AGS18, DL22, EFHY21, GPS18, Riv10, She09, SK11, SW22b, Caz97, KK99]. Property [Bru07, He21, Hop22, KW08, NT15, PR13, Shi05, LP97, Tah00]. Propositional [MMW07]. Provers [Kon22]. Providing [RSÉ03]. Proving [Abl17a, Abl22, But11]. Prym [EF18, LM22]. Pseudo [Bri04, HS07, KKT09]. Pseudo-Anosov [Bri04]. Pseudo-Anosovs [KKT09]. Pseudo-Ansonov [HS07]. Pseudogroup [CEM06]. pseudogroups [Rim92]. Pseudoprime [AD07]. PT [CG23]. Punctured [CP10, CP12, Gar23, HT19, JLS20, KSWY06, BG99]. Pythagorean [EFMPR18].

Quadrangle [De 15a]. Quadratic [Aic04, BCIO01, BG22, BCH^{+}14, BP15, Bir14, Bre98, CDD^{+}16, CH21b, DG02a, Dem08, Els11, FK02, GN09, GXB15, GLP16, HJK^{+}19, MSSS19, Oka02, Pet04, Qua06, RS01, SP05, Sto09, Van17, BtR96, CF95, Chu05, JLW95, Mil93, RS00, SW99, Ste93, ST05]. Quadratics [PW01]. quadratiques [Lep93]. quadrature [BJL05]. Quadrilateral [Bur10, CR93]. Quadrilateral-Octagon [Bur10]. quadruply [BG99]. quadruply-punctured [BG99]. Quandle [AER^{+}08]. Quantization [BK22, KA24]. quantized [dG02b]. Quantum [CGKS20, NSW23, RS19]. Quartic [Joh11, PR21, GPP94, Ola05]. Quartics [Chi04]. Quasi [ST19b]. Quasi-Exactly [ST19b]. Quasifuchsian [Zha22]. quasiperiodic [Tom96b]. Quasiplatonic [Gir03]. Quaternary [FG23]. Quaternion [KKSV16, Oma01]. Quaternionic [CFH^{+}15, GK21]. Queens [YZ22]. Question [BZ17, Kat21]. Questions [Kat14]. Quick [GS05]. Quinn [BBC09a]. quintic [Cra01, FH99]. Quintuple [Mer22]. Quivers [Law17]. Quotient [BCOS24, DBG19, HS14, Hul01, IJK^{+}23, MSS18]. Quotients [BGBVS20, BKM24, CCG21, MSWZ93, HR96].

race [FM00]. Rademacher [Duk23]. Radii [BRS^{+}18]. Radiographie [BM94]. Radiography [BM94]. Radius [Joh08]. Rado [BMRS19]. Ramakrishnan [KT16]. Ramanujan [AG12, BCF04, BC04, DP15, Gui03, Gui12, KR15, Mer21, OS15, Rit07]. Ramanujan-like [AG12]. Ramanujan-Type [Gui03]. Ramified [Dou05]. Ramsey [MO07]. Rand [BBC09a]. Random [ACG^{+}21, AAB^{+}21, AST17, BC01, BC02, Ban06, BKN^{+}12, BL14, BBB^{+}18, BSW13, BRS^{+}18, BGKMI01, BGMR06, CKRS06, DERZ18, ES22, ERvR15, Gan14, HRSS17, KLNP20, Mar20, MOY20, MNS08, OS14, RS19, ACL01, WS98]. Random-to-Random [AST17]. Randomness [BBBR17, CDS02, Gan17]. rang [Fer96, Lep93]. Rank [Alu15, CILO19, Con21, Elk01, HIL13, HM21, Jai09, KS01, LR17, Mal22, Rog00, RS01, Rub13, Sil17, BN97, Fer96]. Rank-2 [Jai09]. Rank-Two [Sil17]. Ranks [CT14, RS00]. Rate [Bri04]. Ratio [Sha21]. Rational [BCH^{+}14, BLP09, BS08b, CB09, DK21, Dum02, Els11, FG23, Gal99, IC16, Jai09, Oha08, Pio07, RS22, Sto00, Yun22, Maz92, Mil93, Mil00, Mur93, PS96, SL00, Wal94]. Rationality [BvBK09, Mel22]. Rationals [DGJ22, Hew21, KM20]. rationnels [Ber95, Wal94, Wal95]. Ratios [Kat10b, Shp16]. Rauzy [HPR22, Mer20]. Ray [CMST22, PW01]. Ray-Marching [CMST22]. RC [BB93]. RC-graphs [BB93]. reaching [BCVZ02]. reader [Ano00]. Real [AP23, BGBVS20, BCF04, Bra21, CDD^{+}16, CL22, DG02a, De 15b, Dem08, DK21, EJ22, Esc14, FIK20, GPHH^{+}12, HHdC^{+}15, JN06, KY17, KCK22, MO22, MP22, MGO22, Oka02, RSSS06, SP05, Sot00, DF00, FP92, FP93, FG99, SW99, Ste93, ST05, Ver00]. Real-Parameter [BCF04]. Real-Time [KCK22]. realizability [Pfe21]. Realization [HLZ10]. Realizations [GKT20, Man18]. Realized [BPS12, FS12]. Reason [Bla22]. Reciprocal [Sta21]. Recognition [DPVZ22, Tar92, HR92b, Mat98]. Reconstruction [Leh22]. Recovering [BHM^{+}13]. Recovery [MSW22]. Rectangle [Sch22c]. Recurrence [BS08a]. recurrences [CF95]. Recurrent [Sch01]. Recursive [GWW16, Pin00]. Reduced [Hum02]. Reducible [PR21, HMT10]. Reduction [CL07b, HS13, Kou19, PS92, Coc05, HMM98, PR03, PR03, HMM99]. Reductions [CT14]. Redundancy [Ale21]. Ree [Kat19]. Refined [CKL^{+}23, GXB15, Mar22]. Refinements [Hen18, OS15]. Reflection [Ano20b, Bon21, CL07a]. Reflections [kAL01, GPS18, ST99]. regarding [MN05]. Regime [HRSS17]. Regina [Bur04]. Regions [AF06, CLLM07]. Regular [De 15a, Gör15, Hay18, MNS08, OS14, SW10, Sch22b, Gra96, IKM02, MR00, Smi00, Sze98]. Regularity [CLMR15, PdlL02]. Regulator [LSY03, JLW95, SW99]. Regulators [Jai09]. régulière [Gra96]. Reid [CLLM07]. Reidemeister [Ter22]. Related [DD03a, DL04, DHJ15, FS19, GdlI07, She20, Cos02, IKM02, LR92]. Relation [RS22, HdC02]. Relations [Col05, DKAP22, FHL^{+}10, Tak14, BS99]. Relative [BCOS24, DERZ18, Gaá01, Gaá15, JN06, BG99, Lou98]. Relator [CH21a]. Relators [MW22]. Remark [GJL03]. Remarkable [XZ22]. Remarks [Bax02, HSV94, Mil92, Lan98]. René [FKS19, FKS15]. Renormalization [GY07, Win22]. Rényi [FM00]. Repair [CM16]. Repeated [BK12, HK21]. Repetitive [BOSS19]. Replacement [HLM03]. Reply [Gan17]. reported [PZHC09]. Representation [Fie07, FS19, Gal14, Wag01, Tha95]. Representations [Ale21, BCM21b, BFG^{+}13, BP15, BCGY23, BH04, BOSS19, CLT06, Die02, Die04a, Dou05, Dou07, Eas20, ESTB10, FPUP22, Gor21, Gre10, Gui06a, Gui18, JRS03, AAC98, AM92, BC00, CNJ96, Cra99a, – Dok00, FS98, KSV97, LLS95, PS96, CNJ96]. Representing [Hul01]. Reproducibility [BBBR17, Gan17]. réseaux [BB92, BM94]. Reshetikhin [KM18]. Residually [Lee04]. Residue [Shp16]. Residues [ABGS11]. Resolution [BBC09a, GPP94]. Resolutions [BSW24, Da 19, DY07, DBG19]. résolvantes [RV99]. resolvents [RV99]. Resolving [ALP^{+}11]. Resonance [MP94, Olv01, Kra94]. Resonances [Bor14]. Respect [HOS11]. Restricted [BLS22, Pur18]. Resultant [DH04]. Results [BHM^{+}13, BEZ03, BDGS23, DN04, GG06, LM22, XZ22, Now97, Tom96a]. Reversible [JJJ02]. Reversing [Sil17]. Revisited [Sto17b, GPP21]. revolution [RS05]. Rewrite [CCG21]. Ribbon [Ano23, Lam21, OS22b]. Riccati [GHK03]. Ricci [RS05]. Rickard [BZ17]. Riemann [BK12, BM15, Bla22, BH18, Boo06, BFFvdW07, Bri06, Gir03, JLS20, Kaw18, LR02, LL21, Rub13, Sha21, She09, SW12, Vor20]. Right [Ino22].

Right-Angled [Ino22]. Rigid [BR17, Die04b, GKT20, LL21]. Rigidity [AD21, BFG^{+}13, Gar21b]. Rigorous [GG06, Tom96a]. Ring [CMZ21, DU95, LS02]. Rings [BCOS24, DBG19, IJK^{+}23, KKM^{+}01, LVL22, Nic18, Shp16, LS02]. Robust [AD21]. Rodrigues [Dur11]. Rogers [KR15, Mer21, OS15]. Rook [LM20]. Root [BDF^{+}22, CCEK07, Gar21a, GKL19, Lam18]. Roots [BFT16, BL21, Coq20, GHY13, MU22, Ste21, Tha08]. Ropelength [KA24]. Rote [Ano20a]. Roudneff [HOKMS24]. round [Viv94]. round-off [Viv94]. Rowmotion [Hop22]. ruler [GP94]. Rules [PR13]. Rungs [Pec19]. Running [Kim22].

SAGE [BRTC23, CMZ21]. Salem [Dol18, MS05, Shi16]. Salesman [Tsu16]. Salmon [BO11]. Samelson [Da 19]. Sampler [CR22]. Sampling [BDHZ09, CLM13, ES22, ERvR15, HS01, OSS03]. Samuel [AM22]. Sandwich [Eas20]. Satisfying [Dur11]. Sato [GJU12]. Saturation [Kie21, PR13]. Saturations [HH22b]. Scaling [BÁ00, Wan01]. Scheme [HM12, Lee23]. Schemes [BCR22, BCOS24, BPL22, BK22, BHL^{+}22, HWW05, BJL05]. Schubert [BB93, Da 19, GPHH^{+}12, HHdC^{+}15, Knu01, LW12, RSSS06, Sot00]. Schur [AS16, DM22]. Schwarz [CL07a, NSYY08, SYY08]. Schwarzian [RTY21]. Science [BBBR17, Gan17]. screen [Goe93]. screw [CHK93]. Search [Ano23, Lam21, MPS^{+}18, LM97]. Searching [BM06, CDV22, Mül06]. Secant [AB09, BDdG07, GPHH^{+}12, HHdC^{+}15, OS16]. Second [ACF^{+}21]. Secondary [JLS20]. Secrecy [OB19]. Sections [De 06a, De 06b]. Sector [Fla08, Joh08]. Sectors [CKL^{+}23]. Segment [HLZ10]. Segre [AB09, OS16]. Selberg [AAB^{+}21, BFM13]. Selection [KH19]. Self [CP10, CP12, Cha14, De 15b, DPPS05, GHSV23, HOS11, Ino19, LvF03, LN03, NTTY22, SS06, Cho94, HSV94, LSV98, STZ95]. Self-Affine [SS06, HSV94]. Self-Dual [GHSV23]. Self-Intersection [CP10, CP12, Cha14]. Self-Projective [De 15b]. Self-Similar [DPPS05, HOS11, LvF03, LN03, NTTY22, Cho94, LSV98, STZ95]. Self-Similarity [Ino19]. Selmer [Dum02]. Semi [HH22a, NW05]. Semi-Infinite [HH22a]. Semi-Presentations [NW05]. Semidefinite [Cd18, EFHY21, FK19, MV10, NT15]. Semigroup [BEN12]. Semigroups [Eas20]. Semiprimes [Mar20]. Semisimple [BJS^{+}21]. Semistable [DDG^{+}14, KK12]. Separation [BRS^{+}18, BDF^{+}22]. separatrix [DRR99]. Sequence [Bax02, CS22a, LRS02, She20, Ste17, BCVZ02, Cos02, Pin00]. Sequences [Bra21, CGJM20, DRT11, EH04, Gor21, GM08, KS08, Lam18, LS13, MVG05, Yab09, CF95, Fin92, Smi00]. Series [AG12, BGBVS20, BM15, BU23, Cha23, CS22a, CL22, DD03a, Frö22, Gui03, Gui06a, Gui06b, HS14, HHS21, Hik03, IT19, KALP22, LMW04, Mer21, Mer22, PT08, Riv10, SS06, Sne05, XZ22, dVV22, BP92, Boy94, Chu05, CVZ00, GP92, LL94]. Serre [DGJ22, Dou05, Dou07]. Set [Alm02, BGKMI01, CL07b, MTC14, SP23, ST02b, Jen00, Kra94]. Sets [ABL17b, BY13, BNH11, BRS^{+}18, HP18, Jeo15, LN03, Mil04, MVG05, Pec19, Pul23, SS06, SW22b, CF96, MPR94, Tom96b, VS00]. Sets-Ladders [Pec19]. Seven [DD03b]. Several [Abr22, Xu04]. Severi [CM13, CT14]. Sextic [ST19b, Cra99b]. sha [Ros00]. Shadow [Nao21]. Shadow-Complexity [Nao21]. Shafarevich [CM00, Dum01]. Shafarevitch [Del01]. Shahidi [KT16]. Shanks [FM00]. shape [KK97]. Shapiro [Sot00]. Sharing [CV03]. Sharp [IC16]. Sheeted [ST02b]. Shift [BMS21]. Short [CH21a, DDG^{+}14, Els14, GMRST14, GL23, SW22a, Geb00]. Shortening [EHPN20]. Shparlinski [DKRV21, Pau17]. Shuffling [AST17]. Sides [GJL03]. Siegel [CD09, EM09, GY07, HOY12, Kat10a, Tak14]. Sierpinski [DL04, OSS03, HP14]. Sieve [KL21, BtR96, EH96]. Sieving [FP99]. Sign [GN22]. Signal [Ste17]. Signatures [LvT05, RAG18]. Similar [DPPS05, HOS11, LvF03, LN03, NTTY22, Cho94, LSV98, STZ95]. Similarity [Ino19]. Simons [MMO^{+}02]. Simple [BPL22, BH04, CHRR07, Dum13, GMR20, JN06, LL21, Mar19, NW05, RAG18, Rah21, BJS^{+}21, CHLM97, Geb00, Web97]. Simplest [Duq01, Ola05]. Simplicial [BL00, CS22b, SK11]. Simultaneous [GST06, Riv04, Sto17a]. Single [HJL22]. singly [CHK93, HW02]. Singular [Ano20b, BP01, Bon21, DRR99, JST20, May01]. Singularities [DBG19, Esc14, Ger03, NSYY08, Ume09, BHvdP95, Tar92]. singularity [CS93]. Singularization [Sch07]. Sinkhorn [NSW23]. Six [CS01, Joh11, KKP^{+}19, – Dok00]. Six-Dimensional [CS01, – Dok00]. sixteenth [CLZ92]. sixteenth-order [CLZ92]. Sixty [KKP^{+}19]. Sixty-Four [KKP^{+}19]. size [JW00]. Skew [KSV22, KSV21]. Slice [BCG19, LM19]. Slicings [SK11]. Sloane [dFT14]. Slopes [BV22, Bir21]. Slow [FIS^{+}14]. Smale [Col05]. Small [BHK^{+}18, BCH^{+}14, BvBK09, BM05, Bri04, CM21, CS22b, Fab09, FG23, FMP04, Gaá15, HHMP10, Kob01, LL23, LS13, LR11, MRRTS16, RSÉ03, Shp16, Sta21, ST21, Sil95]. Smallest [EJ10, Ino22, Mer20]. Smooth [AD21, But12, CHM13, CH21b, LRR^{+}19, Sca09]. Smoothable [BCR22]. smoothness [GMSB96]. SnapPea [Bur14, Gar21b]. Snow [GG06]. Snowflake [SW22b]. Snowflakes [GG06]. soap [Bra95]. Soccer [Cra14]. Sofa [Rom18]. Software [Bur04, JZ05, ST02a]. Sol [Coi22]. solidification [Kob94]. Solids [AAHwaabAR22]. Solitar [MV00]. Solitonlike [Kus09]. solution [Bra95, MR00, MR95]. Solutions [Bra21, CR01, ER15, GP06, Gaá15, Gar17, GHK03, MR10, PPSF17, Rom18, Cho94, JW00]. Solvable [ST19b, dG05, MV00]. Solving [Cra99b, Cra01, DP15, GHK20, MTC14, dW98]. Some [AKP03, BNHLR11, Bax02, BEZ03, BDGS23, BZ16, BR17, Bon21, DL22, DPV15, Els11, FS11, Hal14, Han09, JRS03, KR15, LS02, Man18, MRRTS16, MN05, NP23, RAG18, Riv10, SW99, Wat08, Wil06, Gar02, Ros00, Tom96a, Ano20b]. sommatoire [Dre93, DE93]. Sophie [FJ93]. sous [Ber95]. sous-groupes [Ber95]. Space [ADGH20, BSW24, BCPV21, CV03, Coi22, CMZ21, DS13, GN09, GY10, HT19, HR04, He21, Hye17, KY17, KSWY06, Kus09, MPY17, Tan09, Tra08, Tse22, Ume09, ABG^{+}21, ABT96, HKS92, MR00, Now97, RS98]. Spaces [BNHLR11, BvBK09, But12, DERZ18, Frö22, GK21, JN06, JK01, Lau21, LS03, Mat01, OS16, Tan09, Yun22, Ano97, CHS96, Ter96]. Span [CDV22]. Sparse [Fie07, PS92]. Spatial [CJK^{+}19]. Specht [Web09]. Special [DPVZ22, Dok04, Kat05, KS22, Kon22, Ots15, PR08, Boy98, GM99, HR92b]. Spectra [Ter22, HW94, Ter96]. Spectral [AD21, AST17, BNL17, Joh08, KMN24, KK99, Riv16, SW22b]. Spectrum [AL03, BHS09, HS01, MRRTS16, ST19b, NR99, RT00]. Spets [Sem23]. Sphere [BRS^{+}18, HdJF06, HLMZ01, HS15, KS03, MVG05, She09, TS06, BG99, BL00]. Sphere-Hole [BRS^{+}18]. Spheres [AG22, BBC^{+}09b, BGKT05, CS22b, Fuk11, Tan09, BS99]. Spherical [Ale21, BY13, Bra16, Der15, Wan09]. Spin [Eel07, ER15, Kie21]. Spin-Euler [Eel07]. Spinor [BDM^{+}18]. Spirals [BH01, Sch13b]. Splines [GGM11]. Split [Bre08a]. Splits [BGMR06]. Splitting [DT02, DRR99]. Sporadic [DPP11, Gor21, Hol04, NW05, Rah21, CHLM97, Geb00]. Spot [DRT11]. Spot-Based [DRT11]. Sprawl [DLM13]. Spun [Ino13]. Square [Ano20a, Ben23, BNH11, DG21, Dum01, GGM11, LRZ20, MU22, SW22a]. Square-Tiled [SW22a]. squared [HKS92]. Squareful [BV12]. Squares [AS12, Dum02, vB22, Dil00, FJ96, JK95]. Squaring [LS04]. Stability [GIKW23, MS11, KdC05]. Stable [ADGH20, GOR13, GL15, KY17, Wal13]. stablement [Cou94]. Stably [BZ17]. Stacked [CGP23]. Staff [Ano04a, Ano04b]. Standard [ADDW18, BMS21, Kat05, Kat10a, Olv01, SW95]. Stanisław [PZHC09]. Stanley [IKMF16]. Stark [JRS03, Rob00]. Starscapes [HST22]. state [MPR94]. Statement [Sin03]. States [DEFM23]. Statistical [CGJM20, MR21]. Statistically [Gan14]. Statistics [AD07, Bel21, BHI^{+}11, CJK^{+}19, HS13, HJK^{+}19, SW22a, Wor20]. Status [BHK^{+}18]. Steep [ADDW18]. Stein [LS02]. Steinhaus [EH04]. Step [BSW13, Jun00]. Stepped [FIS^{+}14]. Stern [Fuk11]. Stick [ES22]. Stirling [AMM08, CS22a, Eel07]. Stochastic [AD14, HJL22, CR99]. story [Gou97]. Stover [DR17]. Strand [FS19]. Stratification [BRTC23, Lee23]. Strichartz [BKN^{+}12, WZ21]. Strict [Da 19]. Strictly [HT19]. String [AG12, OM08]. Strings [DS14, Hew21, LvF03]. Strip [BGM22]. Strogatz [BBC09a]. Strong [DP08, HK02b]. Strongly [Har02]. Structure [BV22, HM21, KH19, Lan98]. Structures [ArE11, BPS12, BdJS23, BH04, But12, CG23, CCG21, FS12, HT19, BPW95]. Studied [Ber17]. Studies [DN21, FPU55, YZ22]. Study [BBBR17, BBC^{+}09b, BLS22, Car03, Gan17, GMK19, GW18, GK21, Lau21, LN10, PPS09, SY01, TTS17, BB92, BHvdP95, Boy94, Cra99b, DRR99, Fer96, GH96, HMT10, PdlL02, Tom96b]. Sub [ER19]. Sub-dominant [ER19]. subalgebras [BJS^{+}21]. Subdivisions [BKN^{+}12]. Subgraphs [LM06, RCD21]. Subgroup [Aga13, GMR20]. Subgroups [CL04, CP03, Cum04, Der22, DFH20, EGP^{+}12, FHL^{+}10, GMY21, Hul01, Jai09, KKSV16, LR11, LT18, RD04, Shp16, BDL96, Neb96, Pfe97, Tah00, vGvdKTV97]. Subrings [Alv08]. Subset [MO07]. Subspace [BÖW18]. Substitution [Mer20]. Substitutions [ABB11]. Successive [ABG09, ABGS11]. Sudoku [MTC14]. Suffice [Ger03]. Sufficiency [BA02]. Suggested [MR21]. Suggesting [CLM13]. Suita [BZ16]. Suitable [SP05]. suite [Cos02]. Sum [Abl22, BB18, KB17, Ots15, CF96, CtR96, CtR97, TW99]. sum-free [CF96]. sum-of-divisors [CtR96, CtR97]. summation [DR96]. Summs [BFFvdW07]. Sums [Abl17a, AS12, BV12, JK95, Kat19, LZ10, LMO13, MSW22, Pat03, Pat05, Svi18, Vin01, Alm98, BBG94, Boo00, BBK01, Boy94, CB94, FS98]. Sumsets [LMO13]. Super [Sch19]. Super-Isolated [Sch19]. Superalgebras [BLS22, HM21]. Superconformal [HM21]. Supercongruences [Gui12]. Superpolynomial [DGR06]. Supersingular [Dou07, Shi16, PR03]. Supersingularland [ACNL^{+}23]. supersingulière [PR03]. Superspecial [KH22]. Supporting [Sha19]. Surface [BS01, CH08b, DK21, DR17, EOT11, EJ10, FIS^{+}14, GL22, HLZ10, iIS04, LR05, LT18, SK11, dCIL22, Bra92, BPW95, Bri00, JR99]. Surfaces [Ame08, Ano20b, AAHwaabAR22, BT18, Bar06, Bon21, Bor14, Bra16, Bro07, Bur10, ÇFM23, Cha14, Che19, CV03, CHMWS22, CLW15, Con18, CT14, Die02, Dol18, DLS22, EJ22, ET14, Esc14, FLST23, Gir03, GR22, HM02, He21, HS15, Jai09, JLS20, Kre10, LL21, MRS22, Sch19, Shi16, SW22a, Sta23, Tra08, Ume09, AR97, CHK93, CS93, GR00, GBP97, HK02a, HKS92, KMS00, MP94, Now97, PP93, Ran99, RS98, RTW00, Tah00, Tha95, Tom96a]. Surjective [Gre10]. Survey [Ter96]. Suzuki [Kat19]. Swallowtail [NSYY08]. Swinnerton [Ble11, FHL^{+}10, vB22]. Swinnerton-Dyer [Ble11, FHL^{+}10, vB22]. Sylvester [DH04]. Symbolic [YZ22, CPS^{+}01]. Symbols [Duk23, MR21]. Symmetric [AD21, BT18, BH01, BBG^{+}13, DG21, Dum01, Dum02, EGSC18, HS15, Ino13, LW12, MO07, Ter96]. Symmetries [AFMS17, Con21, HW94, Wal03]. Symmetrized [AS12]. Symmetry [Cd18, Lut09, SW22a, Rei96]. Symplectic [De 15a, DFH23, FGJ22, HS14, Kir11, MMT00, MSS18, Tom96b]. symplectically [BR17]. System [BCPV21, CCEK07, CMOS13, Joh08, Joh11, PPS09, Cos02]. système [Cos02]. Systems [CCG21, De 15b, Fou21, Lam18, MO22, Hul95, Jor99, Sot00]. Systoles [HK02a, KKSV16]. Systolic [CHMWS22]. Syzygies [BEGY20]. Szpiro [BY12, Nit93].

Table [Ano03a, Ano03b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, DGKMY15, KS05, Neb00, Pfe97]. Tables [MRRTS16]. taking [DO99]. Tammes [MT15]. Tan [Mil93]. Tangencies [FG99]. Tangle [AER^{+}08, OM08]. Tangles [OM08]. Tannaka [KK12]. Tate [CM00, Del01, DW10, Dum01, GJU12]. Taut [Par21]. Tautological [CMZ21]. Taylor [JZ05]. Techniques [MS11, SZ04, XKA14]. Teichmüller [KSWY06, Kus09, LM22, Mat01, Muk17, Muk21, Par21]. Teichons [Kus09]. Tensor [BZ17, Sch22a, Sch99]. Tentative [BP07]. Terao [MPY17]. terms [BP92]. Ternary [BP15, Chi02, Chi04, FG23, HP13, Wil06]. Tessellating [HT19]. Tessellation [Gör15]. Test [GS05, Boo00]. Testing [CJUE04, DS15, EM02]. Tests [Kid04]. Tetrahedral [FGG^{+}16]. Tetrahedron [BB22, Bur07, Zon22]. th [FI18]. Thaw [iIS04]. Their [CHRR07, DPVZ22, GdlI07, Hew21, JN06, KSV22, Muk21, Wag01, DE02, HPR22, KSV21, Leh22, PP93, Sta23]. theme [Sze98]. Theorem [CJUE04, EM09, Kon22, RSS18, FJ93, Sto00, AAB^{+}21, Bel21]. Theorems [Mer22, FJS94]. Theoretical [Lom20]. Theory [AG12, BL14, BPL22, BL22, CKRS06, EM17, HS01, LLT03, LM20, MO07, MMW07, OS14, DS95, Mar95]. There [GP94, MTC14]. Theta [Coq20, IT19, Leh22, Mer21]. Thetanullwerte [LPU11]. Thompson [Chu05, CHLM97, HHRS21]. Thomson [Sch13a]. Three [ÁCWB03, BSW13, BV12, Har02, MP01, MW22, AM92, BJL05, Cra01, HW94, KSV97, Kob94, Mat98, Now97]. Three-Dimensional [Har02, AM92, Kob94, Now97]. Three-Manifolds [MP01, HW94, KSV97, Mat98]. Three-Step [BSW13]. Three-Torus [ÁCWB03]. Threefold [Die04b, RSSS13]. Threefolds [AP23, BKZ22, Gar17]. threshold [GG97]. Thue [BHvdP95, DF00, GP06, Gaá15, SdW94, dW98]. Thurston [BGSS22, CMST22, SP23, Tos22]. Tight [AM01, BC98, BDSS18]. Tightening [ACPR11]. Tiled [SW22a]. Tiles [DPPS05]. Tiling [HPR22]. Tilings [DEFM23, HSV94]. Time [CGP23, KB19, KCK22, dVV22]. Tire [LT09]. Toeplitz [EGSC18]. Together [Mil04]. Tomography [TTS17]. too [BPL22]. Tool [iIS04]. top [Thu94]. topography [HR92a]. Topological [AJK08, CS22b, Saw23]. Topologically [KMP22]. Topology [Bur04, De 06a, De 06b, BG99, DE02, Maz92]. Tops [DDG^{+}14]. Torelli [ALT^{+}12]. Tori [BNL17, BS01, KSWY06, MMT00]. Toric [ALP^{+}11, He21, KL21, Sca09, HT00]. Tornheim [BB18]. Torsion [Aga13, BGBVS20, Chi24, GJN23, Jon10, KLNP20, KS01, MPS^{+}18, Muk22, Poo01, Sto17a]. Torus [ÁCWB03, BK21, CP10, Gar23, GOR13, GL15, GY10, HT19, KK93, BÁ00]. Total [Tra08]. totally [FP92, FP93]. Towers [BB21]. Trace [Dum13, EMRSÉ14, Fla08, Kat10b, RR21, BB92]. Traces [EGP^{+}12, LRR^{+}19]. Tracking [BL12, DLS22]. Tracks [LT09]. Trajectories [HPR22, Sch09]. Transcendence [KALP22]. Transcendental [ÇFM23]. Transform [DL04]. Transformation [Eas20, HK02b]. transformations [CD10]. Transforms [Eng09, LSV98]. Transitive [CH08a, LLS95]. Translates [ABL17b]. Translation [ÇFM23, CHMWS22]. Translative [Zon22]. transport [Viv94]. Traveling [Tsu16]. Treating [IL00]. tree [CHLM97, Pfe21]. Trees [CGM15, GWW16, AL95, Rim92]. Trefoil [Ino13]. Triangle [Ale21, DPP11, Der22, HPR22, Sch06, HK02a, Nur00, Smi00].

Triangles [Ano20a, EL09, GJL03, HP14, LRZ20, Alp98]. Triangular [Ber24, KV17, Sch06, Sch09]. Triangulated [Ame08]. Triangulation [iIS04, BL00]. Triangulations [BL14, BS01, BDSS18, CM13, HIS16, ST22, Wor20]. Tricorn [Ino19]. Trigonal [KH22]. Trilinear [BP01]. Trimming [CL98]. Triple [BDM^{+}18, Yi19]. Triples [DP08, EFMPR18]. Trisection [ST22]. Trivial [EHT01, ES14, ERvR15, HR03]. Trivializations [BM06]. trois [JR99]. Tropical [DY07, JLY16, Yun22]. Tropicalizations [CM16]. Truncated [De 06a, De 06b, FIK20, Mer21, Mer22]. tunnel [ABG^{+}21]. Tuples [ABG09, Tan24]. Turán [HMT10]. Turbulence [CMOS13]. Turing [Boo06]. Tutte [CLP15]. twelfth [Rob00]. Twenty [Joh11]. Twenty-Six [Joh11]. Twin [Che19, GKL19]. Twins [Shi05]. Twist [Ino13]. Twist-Spun [Ino13]. Twisted [DFK04, DFJ12, FK10, FK12, PT08, AAC98, Hir00]. Twists [DLN23, Qua06, RS01, RS00]. Two [ABL17b, BY13, BW07, CH08b, CLP15, DM22, FS10, Geb00, Mac08, MOY20, Mer22, RT16, Sch07, Sil17, ST02b, Wik03, XZ22, BN97, BSTV93, BÁ00, Mil00, RS98, Wil02]. Two-Dimensional [BW07, Sch07]. Two-Distance [BY13]. Two-Sheeted [ST02b]. two-torus [BÁ00]. Two-Variable [CLP15]. Type [Alv08, BNHLR11, BPL22, FLST23, GIKW23, GMR20, Gui03, Gui12, KR15, MS01, Nil03, PPS09, Sha19, BR17, CNJ96, GM18]. Type-II [GIKW23]. Types [BPL22, EL14, HS13, HdC02].

Ulam [PZHC09, Ste17]. Unavoidable [EFMPR18]. unicorns [KK97]. Unicycle [LT09]. Uniform [GLP16]. Uniformization [Der15]. Uniformly [HRSS17]. Unimodular [HM21]. Union [Pul23]. Union-Closed [Pul23]. Unipotent [Sin15, GM99]. Unique [BP15, KS08]. Unit [BL21, BGKMI01, CFH^{+}15, LPU11, Muk17]. Unitary [IJK^{+}23, KSV97]. Units [Bir14, LSY03, Tha08, Sil95, Ste93]. Unity [GHY13]. Universal [BHP12, Kat14, Kus09, RSSS13]. Universality [ABB22]. Universally [CCEK07]. Unknot [BCL^{+}22]. unknottedness [DH97]. Unknotting [BH21]. Upper [DE93, Fer22, HP18, PZ04, Dre93]. Ups [He21]. Using [CM16, DKAP22, GG03, HLLM15, Kid04, Roe07, XKA14, BPW95, CPS^{+}01, HWW05, LOS20, MPR94].

valeurs [DO99]. Valuation [AMM08, CGM15]. Valued [GdlI07, IT19, MO22]. Values [BK12, BP01, CH21b, Dok04, FK10, FK12, GMRST14, GJU12, Kat05, Kat10a, KL16, LL23, McN17, Ots15, PT08, Qua06, BDM^{+}18, BBK01, Boy98, DO99, Hir00]. Vanishing [DFK04, BM00]. Variable [CLP15]. Variables [CC07, SSS03, Xu04]. Variance [CKL^{+}23]. Variant [CO20]. Variation [CT14]. variations [BtR96, Sze98]. Varieties [AB09, ALP^{+}11, CM13, CLT06, Da 19, EF18, FGK^{+}16, GM18, PR17a, BG99, Hun00]. Variety [CD21, FM19, OS16, Sca09, GM99]. Vassiliev [Wil02]. Vector [IT19, KK12, SSS03, Guc95]. Vector-Valued [IT19]. Veech [Sch04]. Veering [HIS16, Par21, Wor20]. Verification [HP13, JRS03, vB22].

Verified [HIK^{+}16]. Veronese [AB09, BEGY20]. versal [Ste95]. Version [ADDW18, BY12, BS21, BJP94]. Versus [RTY21, GMSB96, KKT09]. Vertex [HM21, BL00]. Vertices [DD03b, Shl21]. Very [ADDW18, LM17b]. VI [MR10]. Via [BCM21b, BCR22, LGRS23, Ste21, Adz17, BCOS24, BdJS23, BS99, CD09, DHST08, DBG19, EJ22, GHK20, HMM98, HMM99, Hen18, JLY16, MTC14, NT15, PS92]. Viability [ER19]. View [BB22, KK12, MRS22]. Views [KY17]. Violating [BGM22]. Virtual [BNHLR11, BCG19]. virtually [But05]. Visualization [KCK22]. Visualizing [CM00, Muk17, RS05]. vol [Ano00, Ano03a, Ano03b, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano11c]. Volume [Gui18, Hik03, KKT09, KM18, Ano01a, Ano01b, Ano01c, Ano01d, CR99, CRY00, RT00]. Volumes [AM22]. Voronoi [Gar21a, Opg01].

W [CLLM07]. Waerden [KP08]. Walk [BSW13]. Walks [Ban06, BKN^{+}12, OS14]. Wall [ADGH20]. Waring [Chi04, Con21]. Wave [AL03, dVV22]. waveforms [HR92a]. wavelets [CPS^{+}01, JRS92]. Waves [LN10]. Way [SZ06]. Weak [BMRS19, BS12, CMOS13, RT16]. Weakly [Lee04, Nil03]. Weber [FK09]. Weierstrass [Tha95]. Weight [CKRS06, HV19, Hop10, IT19, KW08, Tak14, Van17]. Weighted [ADGH20, JK01]. Weights [Mal22]. Weil [Elk01]. Well [EVV16]. Well-Covered [EVV16]. Wendt [FJ93]. Weyl [CO02, Hum02, Mal22]. Weyl-Equivariant [Mal22]. Where [BU23]. Which [EL09]. Whitehead [HMM05, Kap07]. Whose [Tha08]. Width [GN09]. Wild [Cun17]. Wildanger [GPP21]. Wildly [Dou05]. William [LS02]. Wirsing [Har11]. Without [BPL22, CR01, KV17]. Witt [BG97]. Witten [BB18]. Word [CLM13]. Words [CH21a, ERvR15, KH19]. Worked [Mil04]. Worst [Pec19]. Worth [Sch09]. Writhe [KA24]. Wronskian [DPV15].

Yau [AP23, BKZ22, Die04b]. years [PZHC09]. Yield [CH21a]. Yoccoz [Car03].

Zagier [GM08]. Zannier [Kat21]. Zaremba [AK23]. Zariski [DFH20, LT18]. Zassenhaus [BHK^{+}18]. Zero [HOY12, Wag01, Xu04]. Zeroes [HK21]. Zeros [BFFvdW07, BFM13, Fla08, HLLM15, Mil06, SS06, DF00, Fer92, Fer92]. Zeta [BK12, BM15, BH18, BFFvdW07, BFM13, CLP15, Kat05, Kat10a, KL16, Rub13, Sha21, SW12, BBK01, Mar95]. Zeta-Like [KL16]. Ziegler [Ano20a, HR14b]. Zone [GN09]. zones [Goe93]. Zoo [CS13, RWW01]. Zwonek [ÅCL18].

References

Amzallag:2021:ERM

[AAB^{+}21] E. Amzallag, L.-P. Arguin, E. Bailey, K. Hui, and R. Rao. Evidence of random matrix corrections for the large deviations of Selberg’s Central Limit Theorem. Experimental Mathematics, 33(1):123–135, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.2011806.

Allison:1998:GRH

[AAC98] Gerald Allison, Avner Ash, and Eric Conrad. Galois representations, Hecke operators, and the mod-p cohomology of GL(3, Z) with twisted coefficients. Experimental Mathematics, 7(4):361–390, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047674153.

Athreya:2022:PSH

[AAHwaabAR22] Jayadev S. Athreya, David Aulicino, W. Patrick Hooper, and with an appendix by Anja Randecker. Platonic solids and high genus covers of lattice surfaces. Experimental Mathematics, 31(3):847–877, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2020.1712564.

Abramson:1998:EBA

[AB98] Michael Abramson and Curtis D. Bennett. Enumerating A_{3}(2) blueprints, and an application. Experimental Mathematics, 7(4):391–398, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047674154.

Abo:2009:SVS

[AB09] Hirotachi Abo and Maria Chiara Brambilla. Secant varieties of Segre–Veronese varieties P^{m} × P^{n} embedded by O(1, 2). Experimental Mathematics, 18(3):369–384, ???? 2009. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1259158472.

Arnoux:2011:GMS

[ABB11] Pierre Arnoux, Julien Bernat, and Xavier Bressaud. Geometrical models for substitutions. Experimental Mathematics, 20 (1):97–127, ???? 2011. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317924392.

Alon:2022:UNC

[ABB22] Lior Alon, Ram Band, and Gregory Berkolaiko. Universality of nodal count distribution in large metric graphs. Experimental Mathematics, 33(2):301–335, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2092565.

Ash:2009:FST

[ABG09] Avner Ash, Brandon Bate, and Robert Gross. Frequencies of successive tuples of Frobenius classes. Experimental Mathematics, 18(1):55–64, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1243430529.

Anderson:2021:SKT

[ABG^{+}21] Chris Anderson, Kenneth L. Baker, Xinghua Gao, Marc Kegel, Khanh Le, Kyle Miller, Sinem Onaran, Geoffrey Sangston, Samuel Tripp, Adam Wood, and Ana Wright. L-space knots with tunnel number > 1 by experiment. Experimental Mathematics, 32(4):600–614, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1980753.

Ash:2011:FSP

[ABGS11] Avner Ash, Laura Beltis, Robert Gross, and Warren Sinnott. Frequencies of successive pairs of prime residues. Experimental Mathematics, 20(4):400–411, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1323367154.

Ablinger:2017:DPI

[Abl17a] Jakob Ablinger. Discovering and proving infinite binomial sums identities. Experimental Mathematics, 26(1):62–71, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Abram:2017:IMT

[ABL17b] William C. Abram, Artem Bolshakov, and Jeffrey C. Lagarias. Intersections of multiplicative translates of 3-adic Cantor sets II: Two infinite families. Experimental Mathematics, 26(4):468–489, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1205532.

Ablinger:2022:DPI

[Abl22] Jakob Ablinger. Discovering and proving infinite Pochhammer sum identities. Experimental Mathematics, 31(1):309– 323, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2019.1627254.

Abrams:2022:FPS

[Abr22] Adam Abrams. Finite partitions for several complex continued fraction algorithms. Experimental Mathematics, 31 (3):755–772, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1689870.

Appel:1996:NCO

[ABT96] Matthew J. Appel, Paul S. Bourdon, and John J. Thrall. Norms of composition operators on the Hardy space. Experimental Mathematics, 5(2):111–117, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047565642.

Ash:2007:EPF

[ABZ07] Avner Ash, Jos Brakenhoff, and Theodore Zarrabi. Equality of polynomial and field discriminants. Experimental Mathematics, 16(3):367–374, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204928536.

Asada:2021:LOB

[ACF^{+}21] Megumi Asada, Ryan C. Chen, Eva Fourakis, Yujin Hong Kim, Andrew Kwon, Jared Duker Lichtman, Blake Mackall, Steven J. Miller, Eric Winsor, Karl Winsor, Jianing Yang, and Kevin Yang. Lower-order biases in the second moment of Dirichlet coefficients in families of Lfunctions. Experimental Mathematics, 32(3):431–456, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980453.

Alon:2021:AJG

[ACG^{+}21] Noga Alon, Sebastian M. Cioaba, Brandon D. Gilbert, Jack H. Koolen, and Brendan D. McKay. Addressing Johnson graphs, complete multipartite graphs, odd cycles, and random graphs. Experimental Mathematics, 30(3):372–382, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1542643.

Aiello:2001:RGM

[ACL01] William Aiello, Fan Chung, and Linyuan Lu. A random graph model for power law graphs. Experimental Mathematics, 10 (1):53–66, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/999188420.

Aahag:2018:CCB

[ÅCL18] P. Åhag, R. Czyz, and P. H. Lundow. A counterexample to a conjecture by Blocki–Zwonek. Experimental Mathematics, 27 (1):119–124, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1230913.

Arpin:2023:AS

[ACNL^{+}23] Sarah Arpin, Catalina Camacho-Navarro, Kristin Lauter, Joelle Lim, Kristina Nelson, Travis Scholl, and Jana Sotáková. Adventures in supersingularland. Experimental Mathematics, 32(2):241–268, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1926009.

Ashton:2011:KTC

[ACPR11] Ted Ashton, Jason Cantarella, Michael Piatek, and Eric J. Rawdon. Knot tightening by constrained gradient descent.

Experimental Mathematics, 20(1):57–90, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317924390.

Alvarez:2003:DBT

[ÁCWB03] Miguel Carrión Álvarez, Joseph Corneli, Genevieve Walsh, and Shabnam Beheshti. Double bubbles in the three-torus. Experimental Mathematics, 12(1):79–89, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1064858785.

Andersen:2007:PS

[AD07] Jens Kruse Andersen and Harvey Dubner. Pseudoprime statistics to 10^{19}. Experimental Mathematics, 16(2):209–214, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204905876.

Aristotelous:2014:FSG

[AD14] Andreas C. Aristotelous and Richard Durrett. Fingering in stochastic growth models. Experimental Mathematics, 23(4): 465–474, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ayub:2021:NER

[AD21] Shanza Ayub and Jacopo De Simoi. Numerical evidence of robust dynamical spectral rigidity of ellipses among smooth Z_{2}-symmetric domains. Experimental Mathematics, 32(3): 467–476, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1980456.

Arnold:2018:VSV

[ADDW18] M. Arnold, T. Dauer, M. Doucette, and S.-C. Wolf. On a very steep version of the standard map. Experimental Mathematics, 27(1):93–99, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1226993.

Ascher:2020:EHW

[ADGH20] Kenneth Ascher, Connor Dubé, Daniel Gershenson, and Elaine Hou. Enumerating Hassett’s wall and chamber decomposition of the moduli space of weighted stable curves. Experimental Mathematics, 29(1):36–53, 2020. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1428132.

Adzaga:2017:ACF

[Adz17] Nikola Adzaga. Automated conjecturing of Frobenius numbers via grammatical evolution. Experimental Mathematics, 26(2):247–252, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1175393.

Ameur:2008:TEQ

[AER^{+}08] Kheira Ameur, Mohamed Elhamdadi, Tom Rose, Masahico Saito, and Chad Smudde. Tangle embeddings and Quandle cocycle invariants. Experimental Mathematics, 17(4):487– 498, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243429961.

Antunes:2006:NBP

[AF06] Pedro Antunes and Pedro Freitas. New bounds for the principal Dirichlet eigenvalue of planar regions. Experimental Mathematics, 15(3):333–342, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1175789762.

Arnone:2017:NDB

[AFMS17] Stefano Arnone, Corrado Falcolini, Francesco Moauro, and Matteo Siccardi. On numbers in different bases: Symmetries and a conjecture. Experimental Mathematics, 26(2):197–209, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1149125.

Almkvist:1999:BBA

[AG99] Gert Almkvist and Andrew Granville. Borwein and Bradley’s Apéry-like formulae for ζ(4n + 3). Experimental Mathematics, 8(2):197–203, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047477060.

Almkvist:2012:RLS

[AG12] Gert Almkvist and Jesús Guillera. Ramanujan-like series for 1/π^{2} and string theory. Experimental Mathematics, 21(3): 223–234, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ash:2022:LPS

[AG22] Avner Ash and Robert Gross. Lattice-packing by spheres and eutactic forms. Experimental Mathematics, 31(1):302–308, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1624660.

Agashe:2013:CCO

[Aga13] Amod Agashe. Conjectures concerning the orders of the torsion subgroup, the arithmetic component groups, and the cuspidal subgroup. Experimental Mathematics, 22(4):363–366, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Alpan:2018:APJ

[AGS18] Gökalp Alpan, Alexander Goncharov, and Ahmet Nihat Simsek. Asymptotic properties of Jacobi matrices for a family of fractal measures. Experimental Mathematics, 27(1):10–21, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1209710.

Aicardi:2004:NPB

[Aic04] Francesca Aicardi. On the number of perfect binary quadratic forms. Experimental Mathematics, 13(4):451–457, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1109106437.

Alseda:2008:LBM

[AJK08] Lluís Alsedà, David Juher, and Deborah M. King. A lower bound for the maximum topological entropy of (4k+2)-cycles. Experimental Mathematics, 17(4):391–408, ???? 2008. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1243429953.

Ayadi:2023:CFE

[AK23]  Khalil Ayadi and Takao Komatsu.  Continued frac-

tion expansions towards Zaremba’s Conjecture. Experimental Mathematics, 33(4):755–767, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2023.2293285.

Amerik:2016:EDB

[AKN^{+}16] Ekaterina Amerik, Pär Kurlberg, Khoa D. Nguyen, Adam Towsley, Bianca Viray, and José Felipe Voloch. Evidence for the dynamical Brauer–Manin criterion. Experimental Mathematics, 25(1):54–65, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Almkvist:2003:SNF

[AKP03] Gert Almkvist, Christian Krattenthaler, and Joakim Petersson. Some new formulas for π. Experimental Mathematics, 12(4):441–456, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1087568020.

Applegate:1995:DT

[AL95] David Applegate and Jeffrey C. Lagarias. The distribution of 3x + 1 trees. Experimental Mathematics, 4(3):193–209, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621078.

Asch:2003:SDW

[AL03] Mark Asch and Gilles Lebeau. The spectrum of the damped wave operator for a bounded domain in R^{2}. Experimental Mathematics, 12(2):227–241, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1067634733.

Alexandersson:2019:PLC

[Ale19] Per Alexandersson. Polytopes and large counterexamples. Experimental Mathematics, 28(1):115–120, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1361367.

Alexandre:2021:RTG

[Ale21] Raphaël V. Alexandre. Redundancy of triangle groups in spherical CR representations. Experimental Mathematics, 32 (4):701–721, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1985655.

Allcock:2005:OLL

[All05] Daniel Allcock. Orbits in the Leech lattice. Experimental Mathematics, 14(4):491–509, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1136926978.

Almkvist:1998:AFG

[Alm98] Gert Almkvist. Asymptotic formulas and generalized Dedekind sums. Experimental Mathematics, 7(4):343–359, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047674152.

Almkvist:2002:PPF

[Alm02] G. Almkvist. Partitions with parts in a finite set and with parts outside a finite set. Experimental Mathematics, 11(4): 449–456, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057864654.

Alperin:1998:PTG

[Alp98] Roger C. Alperin. Platonic triangles of groups. Experimental Mathematics, 7(3):191–219, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674204.

Atanasov:2011:RTV

[ALP^{+}11] Atanas Atanasov, Christopher Lopez, Alexander Perry, Nicholas Proudfoot, and Michael Thaddeus. Resolving toric varieties with Nash blowups. Experimental Mathematics, 20(3):288–303, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317924422.

Alexeev:2012:ETM

[ALT^{+}12] Valery Alexeev, Ryan Livingston, Joseph Tenini, Maxim Arap, Xiaoyan Hu, Lauren Huckaba, Patrick McFaddin, Stacy Musgrave, Jaeho Shin, and Catherine Ulrich. Extended Torelli map to the Igusa blowup in genus 6, 7, and 8. Experimental Mathematics, 21(2):193–203, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Aluffi:2015:DPR

[Alu15] Paolo Aluffi. Degrees of projections of rank loci. Experimental Mathematics, 24(4):469–488, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Alvis:2008:SAH

[Alv08] Dean Alvis. Subrings of the asymptotic Hecke algebra of type H_{4}. Experimental Mathematics, 17(3):375–383, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227121390.

Ash:1992:EIT

[AM92] Avner Ash and Mark McConnell. Experimental indications of three-dimensional Galois representations from the cohomology of SL(3, Z). Experimental Mathematics, 1(3):209–223, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048622024.

Abbott:2001:HTH

[AM01] John Abbott and Thom Mulders. How tight is Hadamard’s bound? Experimental Mathematics, 10(3):331–336, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069786341.

Achilles:2022:GSM

[AM22]  Rüdiger Achilles and Mirella Manaresi.  Generalized

Samuel multiplicities of monomial ideals and volumes. Experimental Mathematics, 31(2):611–620, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1671919.

Amendola:2008:DET

[Ame08] Gennaro Amendola. Decomposition and enumeration of triangulated surfaces. Experimental Mathematics, 17(2):153– 166, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227118968.

Amdeberhan:2008:AVS

[AMM08] Tewodros Amdeberhan, Dante Manna, and Victor H. Moll. The 2-adic valuation of Stirling numbers. Experimental Mathematics, 17(1):69–82, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227031898.

Anonymous:1992:IA

[Ano92a] Anonymous. Instructions for authors. Experimental Mathematics, 1(2):175–177, January 1992. ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:1992:ME

[Ano92b] Anonymous. Message from the editors. Experimental Mathematics, 1(3):??, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048622021.

Anonymous:1993:IA

[Ano93] Anonymous. Instructions for authors. Experimental Mathematics, 2(2):147, January 1993. ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:1997:ENP

[Ano97] Anonymous. Editors’ note on “Packing lines, planes, etc.: packings in Grassmannian spaces”. Experimental Mathematics, 6(2):175–??, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047650003.

Anonymous:2000:REM

[Ano00] Anonymous. To the reader, Experiment. Math., vol. 9, iss. 1 (2000). Experimental Mathematics, 9(1):1, ???? 2000. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1046889586.

Anonymous:2001:MBPa

[Ano01a] Anonymous. Miscellaneous back pages, Experiment. Math., volume 10, number 3 (2001). Experimental Mathematics, 10 (3):??, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069786353.

Anonymous:2001:MBPb

[Ano01b] Anonymous. Miscellaneous back pages, Experiment. Math., volume 10, number 4 (2001). Experimental Mathematics, 10 (4):??, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069862465.

Anonymous:2001:MFPa

[Ano01c] Anonymous. Miscellaneous front pages, Experiment. Math., volume 10, number 3 (2001). Experimental Mathematics, 10 (3):??, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069786339.

Anonymous:2001:MFPb

[Ano01d] Anonymous. Miscellaneous front pages, Experiment. Math., volume 10, number 4 (2001). Experimental Mathematics, 10 (4):??, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069855246.

Anonymous:2003:TCEa

[Ano03a] Anonymous. Table of contents, Exp. Math., vol. 12, no. 3 (2003). Experimental Mathematics, 12(3):??, ???? 2003. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1087329228.

Anonymous:2003:TCEb

[Ano03b] Anonymous. Table of contents, Exp. Math., vol. 12, no. 4 (2003). Experimental Mathematics, 12(4):??, ???? 2003. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1087568014.

Anonymous:2004:ESEa

[Ano04a] Anonymous. Editorial staff, exp. math., vol. 13, no. 1 (2004). Experimental Mathematics, 13(1):??, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1086894085.

Anonymous:2004:ESEb

[Ano04b] Anonymous. Editorial staff, exp. math., vol. 13, no. 2 (2004). Experimental Mathematics, 13(2):??, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1090350927.

Anonymous:2004:TCEa

[Ano04c] Anonymous. Table of contents, Exp. Math., vol. 13, no. 1 (2004). Experimental Mathematics, 13(1):??, ???? 2004. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1086894084.

Anonymous:2004:TCEb

[Ano04d] Anonymous. Table of contents, Exp. Math., vol. 13, no. 2 (2004). Experimental Mathematics, 13(2):??, ???? 2004. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1090350926.

Anonymous:2004:TCEc

[Ano04e] Anonymous. Table of contents, Exp. Math., vol. 13, no. 3 (2004). Experimental Mathematics, 13(3):??, ???? 2004. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1103749832.

Anonymous:2004:TCEd

[Ano04f] Anonymous. Table of contents, Exp. Math., vol. 13, no. 4 (2004). Experimental Mathematics, 13(4):??, ???? 2004. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1109106430.

Anonymous:2005:TCEa

[Ano05a] Anonymous. Table of contents, Exp. Math., vol. 14, no. 1 (2005). Experimental Mathematics, 14(1):??, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1120145564.

Anonymous:2005:TCEb

[Ano05b] Anonymous. Table of contents, Exp. Math., vol. 14, no. 2 (2005). Experimental Mathematics, 14(2):??, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1128100124.

Anonymous:2005:TCEc

[Ano05c] Anonymous. Table of contents, Exp. Math., vol. 14, no. 3 (2005). Experimental Mathematics, 14(3):??, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1128371750.

Anonymous:2005:TCEd

[Ano05d] Anonymous. Table of contents, Exp. Math., vol. 14, no. 4 (2005). Experimental Mathematics, 14(4):??, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1136926968.

Anonymous:2006:TCEa

[Ano06a] Anonymous. Table of contents, Exp. Math., vol. 15, no. 1 (2006). Experimental Mathematics, 15(1):??, ???? 2006. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1150476897.

Anonymous:2006:TCEb

[Ano06b] Anonymous. Table of contents, Experiment. Math., vol. 15, no. 2 (2006). Experimental Mathematics, 15(2):??, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789735.

Anonymous:2006:TCEc

[Ano06c] Anonymous. Table of contents, Experiment. Math., vol. 15, no. 3 (2006). Experimental Mathematics, 15(3):??, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789757.

Anonymous:2006:TCEd

[Ano06d] Anonymous. Table of contents, Experiment. Math., vol. 15, no. 4 (2006). Experimental Mathematics, 15(4):??, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789774.

Anonymous:2007:TCEa

[Ano07a] Anonymous. Table of contents, Experiment. Math., vol. 16, no. 1 (2007). Experimental Mathematics, 16(1):??, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789796.

Anonymous:2007:TCEb

[Ano07b] Anonymous. Table of contents, Experiment. Math., vol. 16, no. 2 (2007). Experimental Mathematics, 16(2):??, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204905870.

Anonymous:2007:TCEc

[Ano07c] Anonymous. Table of contents, Experiment. Math., vol. 16, no. 3 (2007). Experimental Mathematics, 16(3):??, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204928527.

Anonymous:2007:TCEd

[Ano07d] Anonymous. Table of contents, Experiment. Math., vol. 16, no. 4 (2007). Experimental Mathematics, 16(4):??, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204836511.

Anonymous:2008:TCEa

[Ano08a] Anonymous. Table of contents, Experiment. Math., vol. 17, no. 1 (2008). Experimental Mathematics, 17(1):??, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227031892.

Anonymous:2008:TCEb

[Ano08b] Anonymous. Table of contents, Experiment. Math., vol. 17, no. 2 (2008). Experimental Mathematics, 17(2):??, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227118965.

Anonymous:2008:TCEc

[Ano08c] Anonymous. Table of contents, Experiment. Math., vol. 17, no. 3 (2008). Experimental Mathematics, 17(3):??, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227121380.

Anonymous:2008:TCEd

[Ano08d] Anonymous. Table of contents, Experiment. Math., vol. 17, no. 4 (2008). Experimental Mathematics, 17(4):??, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1243429951.

Anonymous:2009:TCEa

[Ano09a] Anonymous. Table of contents, Experiment. Math., vol. 18, no. 1 (2009). Experimental Mathematics, 18(1):??, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1243430525.

Anonymous:2009:TCEb

[Ano09b] Anonymous. Table of contents, Experiment. Math., vol. 18, no. 2 (2009). Experimental Mathematics, 18(2):??, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158424.

Anonymous:2009:TCEc

[Ano09c] Anonymous. Table of contents, Experiment. Math., vol. 18, no. 3 (2009). Experimental Mathematics, 18(3):??, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158464.

Anonymous:2009:TCEd

[Ano09d] Anonymous. Table of contents, Experiment. Math., vol. 18, no. 4 (2009). Experimental Mathematics, 18(4):??, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158503.

Anonymous:2010:TCEa

[Ano10a] Anonymous. Table of contents, Experiment. Math., vol. 19, no. 1 (201009). Experimental Mathematics, 19(1):??, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1268404800.

Anonymous:2010:TCEb

[Ano10b] Anonymous. Table of contents, Experiment. Math., vol. 19, no. 2 (2010). Experimental Mathematics, 19(2):??, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1276784784.

Anonymous:2011:C

[Ano11a] Anonymous. Corrigendum. Experimental Mathematics, 20 (3):371–??, ???? 2011. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317924427.

Anonymous:2011:EBE

[Ano11b] Anonymous. Editorial Board EOV. Experimental Mathematics, 20(4):ebi–??, 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:2011:MBE

[Ano11c] Anonymous. Miscellaneous backmatter, Experiment. Math., vol. 20, no. 4 (2011). Experimental Mathematics, 20(4):??, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1323367160.

Anonymous:2012:C

[Ano12a] Anonymous. Corrigendum. Experimental Mathematics, 21 (4):418, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:2012:EBE

[Ano12b] Anonymous. Editorial Board EOV. Experimental Mathematics, 21(4):ebi, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:2013:EEB

[Ano13] Anonymous. EOV editorial board. Experimental Mathematics, 22(4):ebi, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:2014:EBE

[Ano14] Anonymous. Editorial board EOV. Experimental Mathematics, 23(4):ebi, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:2015:EBE

[Ano15] Anonymous. Editorial board EOV. Experimental Mathematics, 24(4):506, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Anonymous:2017:EBE

[Ano17] Anonymous. Editorial board EOV. Experimental Mathematics, 26(4):??, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1356170.

Anonymous:2018:EEB

[Ano18] Anonymous. EOV editorial board. Experimental Mathematics, 27(4):507, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1500849.

Anonymous:2020:CAD

[Ano20a] Anonymous. Correction: “Area Difference Bounds for Dissections of a Square into an Odd Number of Triangles”, by Jean-Philippe Labbé, Günter Rote and Günter M. Ziegler. Experimental Mathematics, 29(1):123, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1519233.

Anonymous:2020:CSS

[Ano20b] Anonymous. Correction: Some Singular Curves and Surfaces Arising from Invariants of Complex Reflection Groups. Experimental Mathematics, 29(3):360, 2020. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1697490.

Anonymous:2023:CSN

[Ano23] Anonymous. Correction: [The search for nonsymmetric ribbon knots]. Experimental Mathematics, 32(3):566, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2223097. See [Lam21].

Arguz:2023:CGR

[AP23] Hülya Argüz and Thomas Prince. On the cohomology groups of real Lagrangians in Calabi–Yau threefolds. Experimental Mathematics, 32(1):169–190, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1926006.

Ash:2004:EAC

[APS04] Avner Ash, David Pollack, and Dayna Soares. SL_{3}(F_{2})- Extensions of Q and arithmetic cohomology modulo 2. Experimental Mathematics, 13(3):297–307, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1103749838.

Aitchison:1997:GSK

[AR97] Iain R. Aitchison and J. Hyam Rubinstein. Geodesic surfaces in knot complements. Experimental Mathematics, 6(2):137– 150, ???? 1997. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047650000.

Alderson:2012:CEC

[AR12] Matthew W. Alderson and Michael O. Rubinstein. Conjectures and experiments concerning the moments of L(1/2, χ_{d}). Experimental Mathematics, 21(3):307–328, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Arai:2007:HPH

[Ara07] Zin Arai. On hyperbolic plateaus of the Hénon map. Experimental Mathematics, 16(2):181–188, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1204905874.

Abrahamsen:2011:AEL

[ArE11] Mikkel Abrahamsen and Søren ren Eilers. On the asymptotic enumeration of LEGO structures. Experimental Mathematics, 20(2):145–152, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317924406.

Alexandersson:2012:DSG

[AS12] Per Alexandersson and Boris Shapiro. Discriminants, symmetrized graph monomials, and sums of squares. Experimental Mathematics, 21(4):353–361, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Ahmed:2016:GSN

[AS16] Tanbir Ahmed and Daniel J. Schaal. On generalized Schur numbers. Experimental Mathematics, 25(2):213–218, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Assaf:2022:KP

[AS22] Sami Assaf and Dominic Searles. Kohnert polynomials. Experimental Mathematics, 31(1):93–119, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1588180.

Ayyer:2017:SGR

[AST17] Arvind Ayyer, Anne Schilling, and Nicolas M. Thiéry. Spectral gap for random-to-random shuffling on linear extensions. Experimental Mathematics, 26(1):22–30, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Arnold:2022:PEP

[AST22] Maxim Arnold, Richard Evan Schwartz, and Serge Tabachnikov. On projective evolutes of polygons. Experimental Mathematics, 33(3):347–356, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2102095.

Ahara:2000:CCH

[AT00] Kazushi Ahara and Mitsuhiko Takasawa. Conjugacy classes of the hyperelliptic mapping class group of genus 2 and 3. Experimental Mathematics, 9(3):383–396, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1045604673.

Avelin:2010:CGF

[Ave10a] Helen Avelin. Computations of Green’s function and its Fourier coefficients on Fuchsian groups. Experimental Mathematics, 19(3):317–334, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317758094.

Avelin:2010:NCG

[Ave10b] Helen Avelin. Numerical computations of Green’s function and its Fourier coefficients on PSL(2, Z). Experimental Mathematics, 19(3):335–343, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317758095.

Briggs:2000:SMT

[BÁ00] Keith M. Briggs and Gonzalo Álvarez. Scaling in a map of the two-torus. Experimental Mathematics, 9(2):301–307, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045952353.

Bivia-Ausina:2002:MED

[BA02] Carles Bivià-Ausina. A method to estimate the degree of C0-Sufficiency of analytic functions. Experimental Mathematics, 11(1):81–85, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057860316.

Badziahin:2016:CIL

[Bad16] Dzmitry Badziahin. Computation of the infimum in the Littlewood conjecture. Experimental Mathematics, 25(1):100– 105, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bailey:2017:JBE

[Bai17] David H. Bailey. Jonathan Borwein: Experimental mathematician. Experimental Mathematics, 26(2):125–129, 2017.

CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1284625.

Banagl:2006:CRW

[Ban06] Markus Banagl. Combinatorial random walks on 3-manifolds. Experimental Mathematics, 15(3):367–381, ???? 2006. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1175789765.

Baragar:2006:FBE

[Bar06] Arthur Baragar. Fractals and the base eigenvalue of the Laplacian on certain noncompact surfaces. Experimental Mathematics, 15(1):33–42, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1150476901.

Baxa:2002:SRD

[Bax02] Christoph Baxa. Some remarks on the distribution of a sequence connected with ζ(\frac{1}{2}). Experimental Mathematics, 11(4):465–468, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057864656.

Bachoc:1992:EAR

[BB92] Christine Bachoc and Christian Batut. Étude algorithmique de réseaux construits avec la forme trace. (French) [Algorithmic study of lattices constructed with the trace from]. Experimental Mathematics, 1(3):183–190, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1048622022.

Bergeron:1993:RGS

[BB93] Nantel Bergeron and Sara Billey. RC-graphs and Schubert polynomials. Experimental Mathematics, 2(4):257–269, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048516036.

Borwein:1997:EDA

[BB97] Jonathan Borwein and David Bradley. Empirically determined Apéry-like formulae for ζ(4n+3). Experimental Mathematics, 6(3):181–194, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047920419.

Bailey:2018:CEE

[BB18] D. H. Bailey and J. M. Borwein. Computation and experimental evaluation of Mordell–Tornheim–Witten sum derivatives. Experimental Mathematics, 27(3):370–376, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1295687.

Boston:2021:HCT

[BB21] Nigel Boston and Michael R. Bush. Heuristics for 2-class towers of cyclic cubic fields. Experimental Mathematics, 33 (1):111–122, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.2011492.

Budney:2022:EMP

[BB22] Ryan Budney and Benjamin A. Burton. Embeddings of 3manifolds in S^{4} from the point of view of the 11-tetrahedron census. Experimental Mathematics, 31(3):988–1013, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1740836.

Bailey:2006:EDA

[BBB06] David H. Bailey, Jonathan M. Borwein, and David M. Bradley. Experimental determination of Apéry-like identities for ζ(2n + 2). Experimental Mathematics, 15(3):281–289, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789759.

Borst:2018:IRP

[BBB^{+}18] Christian Borst, Evan Boyd, Claire Brekken, Samantha Solberg, Melanie Matchett Wood, and Philip Matchett Wood. Irreducibility of random polynomials. Experimental Mathematics, 27(4):498–506, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1325790.

Bailey:2007:HFI

[BBBC07] D. H. Bailey, D. Borwein, J. M. Borwein, and R. E. Crandall. Hypergeometric forms for Ising-class integrals. Experimental Mathematics, 16(3):257–276, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1204928528.

Bailey:2017:RCS

[BBBR17] David H. Bailey, Jonathan M. Borwein, Richard P. Brent, and Mohsen Reisi. Reproducibility in computational science: A case study: Randomness of the digits of pi. Experimental Mathematics, 26(3):298–305, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1163755. See reply [Gan17].

Bailey:2009:RQR

[BBC09a] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Resolution of the Quinn–Rand–Strogatz constant of nonlinear physics. Experimental Mathematics, 18(1):107–116, ???? 2009. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1243430534.

Ballinger:2009:ESE

[BBC^{+}09b] Brandon Ballinger, Grigoriy Blekherman, Henry Cohn, Noah Giansiracusa, Elizabeth Kelly, and Achill Schürmann. Experimental study of energy-minimizing point configurations on spheres. Experimental Mathematics, 18(3):257–283, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158465.

Bailey:2012:EAN

[BBC^{+}12] David H. Bailey, Jonathan M. Borwein, Cristian S. Calude, Michael J. Dinneen, Monica Dumitrescu, and Alex Yee. An empirical approach to the normality of π. Experimental Mathematics, 21(4):375–384, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Barker:2012:NBG

[BBF12] Nathan Barker, Nigel Boston, and Ben Fairbairn. A note on Beauville p-groups. Experimental Mathematics, 21(3):298– 306, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bailey:1994:EEE

[BBG94] David H. Bailey, Jonathan M. Borwein, and Roland Girgensohn. Experimental evaluation of Euler sums. Experimental Mathematics, 3(1):17–30, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062621000.

Brumbaugh:2013:GNS

[BBG^{+}13] J. L. Brumbaugh, Madeleine Bulkow, Luis Alberto Garcia German, Stephan Ramon Garcia, Matt Michal, and Andrew P. Turner. The graphic nature of the symmetric group. Experimental Mathematics, 22(4):421–442, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Borwein:2001:CBS

[BBK01] Jonathan Michael Borwein, David J. Broadhurst, and Joel Kamnitzer. Central binomial sums, multiple Clausen values, and zeta values. Experimental Mathematics, 10(1):25– 34, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://arxiv.org/abs/

hep-th/0004153;   http://projecteuclid.org/euclid.
em/999188418.

Bailey:2017:CDA

[BBKL17] David H. Bailey, Jonathan M. Borwein, Jason S. Kimberley, and Watson Ladd. Computer discovery and analysis of large Poisson polynomials. Experimental Mathematics, 26 (3):349–363, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1180565.

Bartholdi:2015:ACH

[BBvBK15] Laurent Bartholdi, Xavier Buff, Hans-Christian Graf von Bothmer, and Jakob Kröker. Algorithmic construction of Hurwitz maps. Experimental Mathematics, 24(1):76–92, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Borner:2017:FEF

[BBW17] Michel Börner, Irene I. Bouw, and Stefan Wewers. The functional equation for L-functions of hyperelliptic curves. Experimental Mathematics, 26(4):396–411, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1189860.

Buckingham:1998:TBP

[BC98] David J. Buckingham and Paul B. Callahan. Tight bounds on periodic cell configurations in Life. Experimental Mathematics, 7(3):221–241, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047674205.

Borwein:1999:CPC

[BC99] Peter Borwein and Kwok-Kwong Stephen Choi. On cyclotomic polynomials with ±1 coefficients. Experimental Mathematics, 8(4):399–407, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047262361.

Borwein:2000:R

[BC00] Jonathan Borwein and Kwok-Kwong Stephen Choi. On the representations of xy + yz + zx. Experimental Mathematics, 9(1):153–158, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1046889597.

Bailey:2001:RCF

[BC01] David H. Bailey and Richard E. Crandall. On the random character of fundamental constant expansions. Experimental Mathematics, 10(2):175–190, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/999188630.

Bailey:2002:RGN

[BC02] David H. Bailey and Richard E. Crandall. Random generators and normal numbers. Experimental Mathematics, 11(4):527– 546, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057864662.

Borwein:2004:RAFb

[BC04] J. Borwein and R. Crandall. On the Ramanujan AGM fraction, II: The complex-parameter case. Experimental Mathematics, 13(3):287–295, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1103749837.

Borwein:2010:HDB

[BCC10] Jonathan M. Borwein, O-Yeat Chan, and R. E. Crandall. Higher-dimensional box integrals. Experimental Mathematics, 19(4):431–446, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317758103.

Borwein:2004:RAFa

[BCF04] J. Borwein, R. Crandall, and G. Fee. On the Ramanujan AGM fraction, I: The real-parameter case. Experimental Mathematics, 13(3):275–285, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1103749836.

Butler:2010:FPA

[BCG10] Steve Butler, Kevin P. Costello, and Ron Graham. Finding patterns avoiding many monochromatic constellations. Experimental Mathematics, 19(4):399–411, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317758100.

Boden:2019:VKC

[BCG19] Hans U. Boden, Micah Chrisman, and Robin Gaudreau. Virtual knot cobordism and bounding the slice genus. Experimental Mathematics, 28(4):475–491, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1422160.

Bibby:2023:HRC

[BCGY23] Christin Bibby, Melody Chan, Nir Gadish, and Claudia He Yun. Homology representations of compactified configurations on graphs applied to M_{2,n}. Experimental Mathematics, 33(4):611–623, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2023.2209749.

Baumslag:2004:EIG

[BCH04] Gilbert Baumslag, Sean Cleary, and George Havas. Experimenting with infinite groups, I. Experimental Mathematics, 13(4):495–502, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1109106441.

Benedetto:2014:SDH

[BCH^{+}14] Robert L. Benedetto, Ruqian Chen, Trevor Hyde, Yordanka Kovacheva, and Colin White. Small dynamical heights for quadratic polynomials and rational functions. Experimental Mathematics, 23(4):433–447, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Baeza:2001:HCQ

[BCIO01] Ricardo Baeza, Renaud Coulangeon, Maria Ines Icaza, and Manuel O’Ryan. Hermite’s constant for quadratic number fields. Experimental Mathematics, 10(4):543–552, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069855254.

Burton:2022:HDU

[BCL^{+}22] Benjamin A. Burton, Hsien-Chih Chang, Maarten Löffler, Clément Maria, Arnaud de Mesmay, Saul Schleimer, Eric Sedgwick, and Jonathan Spreer. Hard diagrams of the unknot. Experimental Mathematics, 33(3):482–500, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2161676.

Borwein:2017:CLA

[BCLM17] Jonathan M. Borwein, Neil J. Calkin, Scott B. Lindstrom, and Andrew Mattingly. Continued logarithms and associated continued fractions. Experimental Mathematics, 26(4): 412–429, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1195307.

Barbero:2021:AAI

[BCM21a] Stefano Barbero, Umberto Cerruti, and Nadir Murru. Approximations of algebraic irrationalities with matrices. Experimental Mathematics, 30(2):200–208, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1526722.

Barbero:2021:PRA

[BCM21b] Stefano Barbero, Umberto Cerruti, and Nadir Murru. Periodic representations and approximations of p-adic numbers via continued fractions. Experimental Mathematics, 33(1): 100–110, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.2011491.

Bertone:2024:OCL

[BCOS24] Cristina Bertone, Francesca Cioffi, Matthias Orth, and Werner M. Seiler. Open covers and Lex points of Hilbert schemes over quotient rings via relative marked bases. Experimental Mathematics, 34(1):38–52, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2024.2309517.

Bunimovich:2021:FIA

[BCPV21] Leonid A. Bunimovich, Giulio Casati, Tomaz Prosen, and Gregor Vidmar. Few islands approximation of Hamiltonian system with divided phase space. Experimental Mathematics, 30(4):459–468, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1559777.

Bertone:2022:SGP

[BCR22] Cristina Bertone, Francesca Cioffi, and Margherita Roggero. Smoothable Gorenstein points via marked schemes and double-generic initial ideals. Experimental Mathematics, 31 (1):120–137, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1592034.

Benito:2002:ASE

[BCVZ02] Manuel Benito, Wolfgang Creyaufmüller, Juan L. Varona, and Paul Zimmermann. Aliquot sequence 3630 ends after reaching 100 digits. Experimental Mathematics, 11(2):201– 206, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621215.

Borwein:2009:CCI

[BD09] Jonathan M. Borwein and Keith J. Devlin. The computer as crucible: an introduction to experimental mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-343-0. xi + 158 pp. LCCN QA8.7 .B67 2009. URL http://

www.loc.gov/catdir/toc/fy0904/2008022180.html. With

illustrations by Karl H. Hofmann.

Baur:2007:SDM

[BDdG07] Karin Baur, Jan Draisma, and Willem A. de Graaf. Secant dimensions of minimal orbits: Computations and conjectures. Experimental Mathematics, 16(2):239–251, ???? 2007. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1204905879.

Bugeaud:2022:ARS

[BDF^{+}22] Yann Bugeaud, Andrej Dujella, Wenjie Fang, Tomislav Pejković, and Bruno Salvy. Absolute root separation. Experimental Mathematics, 31(3):806–813, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1699480.

BenAv:2023:FPC

[BDGS23] Radel Ben Av, Giora Dula, Assaf Goldberger, and Yossi Strassler. Frame potential in CP^{n} some numerical and analytical results. Experimental Mathematics, 32(1):26–32, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1753597.

Boocher:2009:SLF

[BDHZ09] Adam Boocher, Jay Daigle, Jim Hoste, and Wenjing Zheng. Sampling Lissajous and Fourier knots. Experimental Mathematics, 18(4):481–497, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1259158509.

Bloch:2023:CHL

[BdJS23] Spencer Bloch, Robin de Jong, and Emre Can Sertöz. Computing heights via limits of Hodge structures. Experimental Mathematics, 33(4):547–558, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2023.2188318.

Buekenhout:1996:AGM

[BDL96] Francis Buekenhout, Michel Dehon, and Dimitri Leemans. All geometries of the Mathieu group M_{11} based on maximal subgroups. Experimental Mathematics, 5(2):101–110, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565641.

Bonifant:2007:ECA

[BDM07] Araceli Bonifant, Marius Dabija, and John Milnor. Elliptic curves as attractors in P^{2}, Part 1: Dynamics. Experimental Mathematics, 16(4):385–420, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1204836512.

Bergstrom:2018:ECS

[BDM^{+}18] Jonas Bergström, Neil Dummigan, Thomas Mégarbané, Tomoyoshi Ibukiyama, and Hidenori Katsurada. Eisenstein congruences for SO(4, 3), SO(4, 4), spinor, and triple product L-values. Experimental Mathematics, 27(2):230–250, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1251861.

Brambilla:2016:ECB

[BDP16] Maria Chiara Brambilla, Olivia Dumitrescu, and Elisa Postinghel. On the effective cone of blown-up at n + 3 points. Experimental Mathematics, 25(4):452–465, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bright:2016:MEP

[BDS16] Curtis Bright, Raymond Devillers, and Jeffrey Shallit. Minimal elements for the prime numbers. Experimental Mathematics, 25(3):321–331, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Burton:2018:CPT

[BDSS18] Benjamin A. Burton, Basudeb Datta, Nitin Singh, and Jonathan Spreer. A construction principle for tight and minimal triangulations of manifolds. Experimental Mathematics, 27(1):22–36, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1212747.

Bruce:2020:CCA

[BEGY20] Juliette Bruce, Daniel Erman, Steve Goldstein, and Jay Yang. Conjectures and computations about Veronese syzygies. Experimental Mathematics, 29(4):398–413, 2020. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1474506.

Bell:2021:ESM

[Bel21] Mark C. Bell. Experimental statistics for Mirzakhani’s Theorem. Experimental Mathematics, 32(4):683–691, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1982426.

Bohm:2012:DSA

[BEN12] Janko Böhm, David Eisenbud, and Max J. Nitsche. Decomposition of semigroup algebras. Experimental Mathematics, 21 (4):385–394, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Benoist:2023:CSA

[Ben23] Yves Benoist. Convolution and square in Abelian groups I. Experimental Mathematics, 33(3):518–528, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2174212.

Bertrand:1995:PRS

[Ber95] Daniel Bertrand. Points rationnels sur les sous-groupes compacts des groupes algébriques. Experimental Mathematics, 4(2):145–151, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047931623.

Berg:2017:FFS

[Ber17] Christian Berg. On a family of functions studied by Bhatia and Jain. Experimental Mathematics, 26(2):241–246, 2017.

CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1159624.

Bergeron:2024:TDH

[Ber24] F. Bergeron. Triangular diagonal harmonics conjectures. Experimental Mathematics, 34(1):126–139, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2024.2336074.

Beck:2003:SER

[BEZ03] Matthias Beck, David Einstein, and Shelemyahu Zacks. Some experimental results on the Frobenius problem. Experimental Mathematics, 12(3):263–270, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1087329230.

Bressaud:2016:CB

[BF16] Xavier Bressaud and Marie-Claire Fournier. Casse-briques. Experimental Mathematics, 25(4):357–378, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Borwein:2007:ZPS

[BFFvdW07] Peter Borwein, Greg Fee, Ron Ferguson, and Alexa van der Waal. Zeros of partial summs of the Riemann zeta function. Experimental Mathematics, 16(1):21–40, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789799.

Bergeron:2013:LRR

[BFG^{+}13] Nicolas Bergeron, Elisha Falbel, Antonin Guilloux, Pierre-Vincent Koseleff, and Fabrice Rouillier. Local rigidity for PGL(3, C)-representations of 3-manifold groups. Experimental Mathematics, 22(4):410–420, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Blank:1992:KMC

[BFK^{+}92] Brian E. Blank, Da Shan Fan, David Klein, Steven G. Krantz, Daowei Ma, and Myung-Yull Pang. The Kobayashi metric of a complex ellipsoid in C^{2}. Experimental Mathematics, 1(1): 47–55, ???? 1992. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048709115.

Bruggeman:2013:PZS

[BFM13] Roelof Bruggeman, Markus Fraczek, and Dieter Mayer. Perturbation of zeros of the Selberg zeta function for Γ_{0}(4). Experimental Mathematics, 22(3):217–242, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Billey:2016:CRP

[BFT16] Sara Billey, Matthew Fahrbach, and Alan Talmage. Coefficients and roots of peak polynomials. Experimental Mathematics, 25(2):165–175, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bruner:1995:BLC

[BG95] Robert Bruner and John Greenlees. The Bredon–Löffler conjecture. Experimental Mathematics, 4(4):289–297, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047674389.

Bayer:1997:HWI

[BG97] Pilar Bayer and Josep González. On the Hasse–Witt invariants of modular curves. Experimental Mathematics, 6(1): 57–76, ???? 1997. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047565284.

Benedetto:1999:TRC

[BG99] Robert L. Benedetto and William M. Goldman. The topology of the relative character varieties of a quadruplypunctured sphere. Experimental Mathematics, 8(1):85–103, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047477115.

Belova:2022:CPM

[BG22] Anna Belova and Igors Gorbovickis. Critical points of the multiplier map for the quadratic family. Experimental Mathematics, 31(1):337–345, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1627682.

Bartolo:2020:FGR

[BGBVS20] Enrique Artal Bartolo, Benoît Guerville-Ballé, and Juan Viu-Sos. Fundamental groups of real arrangements and torsion in the lower central series quotients. Experimental Mathematics, 29(1):28–35, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1428131.

Breimer:2001:HRS

[BGKMI01] Eric Breimer, Mark Goldberg, Brian Kolstad, and Malik Magdon-Ismail. On the height of a random set of points in a d-dimensional unit cube. Experimental Mathematics, 10(4): 583–598, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069855258.

Boyer:2005:EME

[BGKT05] Charles P. Boyer, Krzysztof Galicki, János Kollár, and Evan Thomas. Einstein metrics on exotic spheres in dimensions 7, 11, and 15. Experimental Mathematics, 14(1):59–64, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1120145570.

Belmans:2022:EVG

[BGM22] Pieter Belmans, Sergey Galkin, and Swarnava Mukhopadhyay. Examples violating Golyshev’s canonical strip hypotheses. Experimental Mathematics, 31(1):233–237, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1602571.

Buhler:2006:PRM

[BGMR06] Joe Buhler, Daniel Goldstein, David Moews, and Joel Rosenberg. The probability that a random monic p-adic polynomial splits. Experimental Mathematics, 15(1):21–32, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1150476900.

Bachman:2022:CFC

[BGSS22] David Bachman, Matthias Goerner, Saul Schleimer, and Henry Segerman. Cohomology fractals, Cannon–Thurston maps, and the geodesic flow. Experimental Mathematics, 31 (4):1047–1085, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1994059.

Bobenko:2001:CSC

[BH01] Alexander I. Bobenko and Tim Hoffmann. Conformally symmetric circle packings: a generalization of Doyle’s spirals. Experimental Mathematics, 10(1):141–150, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/999188429.

Bremner:2004:INA

[BH04] Murray Bremner and Irvin Hentzel. Invariant nonassociative algebra structures on irreducible representations of simple Lie algebras. Experimental Mathematics, 13(2):231–256, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1090350937.

Baig:2012:EDG

[BH12] Salman Baig and Chris Hall. Experimental data for Goldfeld’s Conjecture over function fields. Experimental Mathematics, 21(4):362–374, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Bober:2018:NCR

[BH18] Jonathan W. Bober and Ghaith A. Hiary. New computations of the Riemann zeta function on the critical line. Experimental Mathematics, 27(2):125–137, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1233083.

Brittenham:2021:CBJ

[BH21] Mark Brittenham and Susan Hermiller. A counterexample to the Bernhard–Jablan unknotting conjecture. Experimental Mathematics, 30(4):547–556, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1580633.

Buff:2001:FC

[BHH01] Xavier Buff, Christian Henriksen, and John H. Hubbard. Farey curves. Experimental Mathematics, 10(4):481–486, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069855247.

Bruns:2011:CCH

[BHI^{+}11] Winfried Bruns, Raymond Hemmecke, Bogdan Ichim, Matthias Köppe, and Christof Söger. Challenging computations of Hilbert bases of cones associated with algebraic statistics. Experimental Mathematics, 20(1):25–33, ???? 2011. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1317924388.

Brzezinski:2005:CBM

[BHK05] J. Brzeziński, W. Holsztyński, and P. Kurlberg. On the congruence ax + by ≡ 1 modulo xy. Experimental Mathematics, 14(4):391–401, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1136926970.

Bachle:2018:SZC

[BHK^{+}18] Andreas Bächle, Allen Herman, Alexander Konovalov, Leo Margolis, and Gurmail Singh. The status of the Zassenhaus conjecture for small groups. Experimental Mathematics, 27 (4):431–436, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1306814.

Buzzard:2022:SL

[BHL^{+}22] Kevin Buzzard, Chris Hughes, Kenny Lau, Amelia Livingston, Ramon Fernández Mir, and Scott Morrison. Schemes in lean. Experimental Mathematics, 31(2):355–363, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1983489.

Bates:2013:RER

[BHM^{+}13] Daniel J. Bates, Jonathan D. Hauenstein, Timothy M. Mc-Coy, Chris Peterson, and Andrew J. Sommese. Recovering exact results from inexact numerical data in algebraic geometry. Experimental Mathematics, 22(1):38–50, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Bogart:2012:UGB

[BHP12] Tristram Bogart, Ray Hemmecke, and Sonja Petrović. Universal Gröbner bases of colored partition identities. Experimental Mathematics, 21(4):395–401, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Berry:2009:OAS

[BHS09] Tyrus Berry, Steven M. Heilman, and Robert S. Strichartz. Outer approximation of the spectrum of a fractal Laplacian. Experimental Mathematics, 18(4):449–480, ???? 2009. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1259158508.

Bombieri:1995:DST

[BHvdP95] Enrico Bombieri, David C. Hunt, and Alfred J. van der Poorten. Determinants in the study of Thue’s method and curves with prescribed singularities. Experimental Mathematics, 4(2):87–96, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047931619.

Bircan:2014:CUQ

[Bir14] Nihal Bircan. A conjecture on units of quadratic fields. Experimental Mathematics, 23(2):95–98, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Birkbeck:2021:SOH

[Bir21] Christopher Birkbeck. Slopes of overconvergent Hilbert modular forms. Experimental Mathematics, 30(3):295–314, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1538909.

Birklbauer:2022:FCH

[Bir22] Philipp Birklbauer. The Fuglede conjecture holds in Z_{5}^{3}. Experimental Mathematics, 31(2):456–460, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1636427.

Bailey:2005:CTH

[BJL05] David H. Bailey, Karthik Jeyabalan, and Xiaoye S. Li. A comparison of three high-precision quadrature schemes. Experimental Mathematics, 14(3):317–329, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1128371757.

Buchmann:1994:PVG

[BJP94] Johannes Buchmann, Max Jüntgen, and Michael Pohst. A practical version of the generalized Lagrange algorithm. Experimental Mathematics, 3(3):199–207, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1048515871.

Bochenski:2021:SSS

[BJS^{+}21] Maciej Bocheński, Piotr Jastrzebski, Anna Szczepkowska, Aleksy Tralle, and Artur Woike. Semisimple subalgebras in simple Lie algebras and a computational approach to the compact Clifford–Klein forms problem. Experimental Mathematics, 30(1):86–94, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1492475.

Baier:2003:HCC

[BK03] Harald Baier and Günter Köhler. How to compute the coefficients of the elliptic modular function j(z). Experimental Mathematics, 12(1):115–126, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1064858788.

Banks:2012:RVR

[BK12] William D. Banks and Sarah Kang. On repeated values of the Riemann zeta function on the critical line. Experimental Mathematics, 21(2):132–140, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Brown:2016:FDP

[BK16] Gavin Brown and Alexander Kasprzyk. Four-dimensional projective orbifold hypersurfaces. Experimental Mathematics, 25(2):176–193, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Berchenko-Kogan:2021:EAT

[BK21] Yakov Berchenko-Kogan. The entropy of the Angenent torus is approximately 1.85122. Experimental Mathematics, 30(4): 587–594, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1583616.

Buring:2022:EMH

[BK22] R. Buring and A. V. Kiselev. The expansion ⋆ mod ō(h̄^{4}) and computer-assisted proof schemes in the Kontsevich deformation quantization. Experimental Mathematics, 31(3):701–754, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1680463.

Blair:2024:CQK

[BKM24] Ryan Blair, Alexandra Kjuchukova, and Nathaniel Morrison. Coxeter quotients of knot groups through 16 crossings. Experimental Mathematics, 34(1):27–37, 2024. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2024.2309507.

Begue:2012:RWB

[BKN^{+}12] Matthew Begue, Daniel J. Kelleher, Aaron Nelson, Hugo Panzo, Ryan Pellico, and Alexander Teplyaev. Random walks on barycentric subdivisions and the Strichartz hexacarpet. Experimental Mathematics, 21(4):402–417, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

Brown:2022:GFC

[BKZ22] Gavin Brown, Alexander M. Kasprzyk, and Lei Zhu. Gorenstein formats, canonical and Calabi–Yau threefolds. Experimental Mathematics, 31(1):146–164, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1592036.

Bjorner:2000:SMB

[BL00] Anders Björner and Frank H. Lutz. Simplicial manifolds, bistellar flips and a 16-vertex triangulation of the Poincaré homology 3-sphere. Experimental Mathematics, 9(2):275–289, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045952351.

Boyle:2007:JPP

[BL07] Mike Boyle and Bryant Lee. Jointly periodic points in cellular automata: Computer explorations and conjectures. Experimental Mathematics, 16(3):293–302, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1204928530.

Beltran:2012:CNH

[BL12] Carlos Beltrán and Anton Leykin. Certified numerical homotopy tracking. Experimental Mathematics, 21(1):69–83, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Benedetti:2014:RDM

[BL14] Bruno Benedetti and Frank H. Lutz. Random discrete Morse theory and a new library of triangulations. Experimental Mathematics, 23(1):66–94, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Braun:2021:PRU

[BL21] Benjamin Braun and Fu Liu. h^{⋆}-polynomials with roots on the unit circle. Experimental Mathematics, 30(3):332–348, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1538912.

Browning:2022:FGT

[BL22] Thomas Browning and Patrick Lutz. Formalizing Galois theory. Experimental Mathematics, 31(2):413–424, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1986176.

Blanc:2022:NRD

[Bla22] Philippe Blanc. A new reason for doubting the Riemann Hypothesis. Experimental Mathematics, 31(1):88–92, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1586600.

Bley:2011:NEE

[Ble11] Werner Bley. Numerical evidence for the equivariant Birch and Swinnerton-Dyer conjecture. Experimental Mathematics, 20(4):426–456, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1323367156.

Bernard:2009:JGC

[BLP09] Nicolas Bernard, Franck Leprévost, and Michael Pohst. Jacobians of genus-2 curves with a rational point of order 11. Experimental Mathematics, 18(1):65–70, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1243430530.

Bouarroudj:2022:CAS

[BLS22] Sofiane Bouarroudj, Dimitry Leites, and Jin Shang. Computeraided study of double extensions of restricted Lie superalgebras preserving the nondegenerate closed 2-forms in characteristic 2. Experimental Mathematics, 31(2):676–688, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1683102.

Batut:1994:RRP

[BM94] Christian Batut and Jacques Martinet. Radiographie des réseaux parfaits. (French) [Radiography of perfect lattices]. Experimental Mathematics, 3(1):39–49, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1062621002.

Belaga:1998:ECC

[BM98] Edward G. Belaga and Maurice Mignotte. Embedding the 3x + 1 conjecture in a 3x + d context. Experimental Mathematics, 7(2):145–151, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048515662.

Borwein:2000:PHP

[BM00] Peter Borwein and Michael J. Mossinghoff. Polynomials with height 1 and prescribed vanishing at 1. Experimental Mathematics, 9(3):425–433, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045604677.

Bessis:2004:EPE

[BM04] David Bessis and Jean Michel. Explicit presentations for exceptional braid groups. Experimental Mathematics, 13(3): 257–266, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1103749833.

Boyd:2005:SLP

[BM05] David W. Boyd and Michael J. Mossinghoff. Small limit points of Mahler’s measure. Experimental Mathematics, 14(4):403–414, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1136926971.

Bowman:2006:FSA

[BM06] R. Sean Bowman and Stephen B. McCaul. Fast searching for Andrews–Curtis trivializations. Experimental Mathematics, 15(2):193–197, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1175789740.

Beliakov:2015:ARZ

[BM15] Gleb Beliakov and Yuri Matiyasevich. Approximation of Riemann’s zeta function by finite Dirichlet series: A multiprecision numerical approach. Experimental Mathematics, 24(2): 150–161, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Brubaker:2022:NDM

[BMN22] Nicholas Brubaker, Thomas Murphy, and K. Oskar Negron. Numerically destabilizing minimal discs. Experimental Mathematics, 31(3):794–805, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1692261.

Boza:2019:CWR

[BMRS19] L. Boza, J. M. Marín, M. P. Revuelta, and M. I. Sanz. On the n-color weak Rado numbers for the equation x_{1} + x_{2} + ⋯ + x_{k} + c = x_{k+1}. Experimental Mathematics, 28(2):194–208, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1382403.

Bialy:2021:OBD

[BMS21] Misha Bialy, Andrey E. Mironov, and Lior Shalom. Outer billiards with the dynamics of a standard shift on a finite number of invariant curves. Experimental Mathematics, 30 (4):469–474, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1563514.

Bachoc:1997:CTG

[BN97] Christine Bachoc and Gabriele Nebe. Classification of two genera of 32-dimensional lattices of rank 8 over the Hurwitz order. Experimental Mathematics, 6(2):151–162, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047650001.

Bonnaillie-Noel:2011:NAN

[BNH11] V. Bonnaillie-Noël and B. Helffer. Numerical analysis of nodal sets for eigenvalues of Aharonov–Bohm Hamiltonians on the square with application to minimal partitions. Experimental Mathematics, 20(3):304–322, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1317924423.

Bar-Natan:2011:SDS

[BNHLR11] Dror Bar-Natan, Iva Halacheva, Louis Leung, and Fionntan Roukema. Some dimensions of spaces of finite type invariants of virtual knots. Experimental Mathematics, 20(3):282–287, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317924421.

Bonnaillie-Noel:2017:SMP

[BNL17] Virginie Bonnaillie-Noël and Corentin Léna. Spectral minimal partitions for a family of tori. Experimental Mathematics, 26(4):381–395, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1183154.

Bates:2011:TSC

[BO11] Daniel J. Bates and Luke Oeding. Toward a Salmon conjecture. Experimental Mathematics, 20(3):358–370, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317924426.

Bogosel:2017:PML

[BO17] Beniamin Bogosel and Édouard Oudet. Partitions of minimal length on manifolds. Experimental Mathematics, 26 (4):496–508, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1223570.

Boege:2023:NEC

[Boe23] Tobias Boege. No eleventh conditional Ingleton inequality. Experimental Mathematics, 33(4):808–816, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2294827.

Bonnafe:2021:SSC

[Bon21] Cédric Bonnafé. Some singular curves and surfaces arising from invariants of complex reflection groups. Experimental Mathematics, 30(3):429–440, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1555778.

Booker:2000:TIF

[Boo00] Andrew R. Booker. A test for identifying Fourier coefficients of automorphic forms and application to Kloosterman sums. Experimental Mathematics, 9(4):571–581, October 2000. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1045759522.

Booker:2006:ACT

[Boo06] Andrew R. Booker. Artin’s conjecture, Turing’s method, and the Riemann hypothesis. Experimental Mathematics, 15(4): 385–407, ???? 2006. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1175789775.

Boozer:2021:CBK

[Boo21] David Boozer. Computer bounds for Kronheimer–Mrowka foam evaluation. Experimental Mathematics, 32(4):615–630, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1982078.

Borthwick:2014:DRH

[Bor14] David Borthwick. Distribution of resonances for hyperbolic surfaces. Experimental Mathematics, 23(1):25–45, 2014. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bridy:2019:GNL

[BOSS19] Andrew Bridy, Robert J. Lemke Oliver, Arlo Shallit, and Jeffrey Shallit. The generalized Nagell–Ljunggren problem: Powers with repetitive representations. Experimental Mathematics, 28(4):428–439, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1419391.

Braun:2018:NLB

[BÖW18] Michael Braun, Patric R. J. Östergård, and Alfred Wassermann. New lower bounds for binary constant-dimension subspace codes. Experimental Mathematics, 27(2):179–183, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1239145.

Boyd:1994:ASP

[Boy94] David W. Boyd. A p-adic study of the partial sums of the harmonic series. Experimental Mathematics, 3(4):287–302, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048515811.

Boyd:1998:MMS

[Boy98] David W. Boyd. Mahler’s measure and special values of L-functions. Experimental Mathematics, 7(1):37–82, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047674271.

Boyd:1999:BFC

[Boy99] John P. Boyd. The Blasius function in the complex plane. Experimental Mathematics, 8(4):381–394, ???? 1999. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1047262359.

Bergeron:1992:CGF

[BP92] François Bergeron and Simon Plouffe. Computing the generating function of a series given its first few terms. Experimental Mathematics, 1(4):307–312, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048610118.

Bullett:1994:GIC

[BP94] Shaun Bullett and Christopher Penrose. A gallery of iterated correspondences. Experimental Mathematics, 3(2):85– 105, ???? 1994. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062620904.

Bernhardsson:2001:SVT

[BP01] Bo Bernhardsson and Jaak Peetre. Singular values of trilinear forms. Experimental Mathematics, 10(4):509–518, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069855250.

Bressaud:2007:TCB

[BP07] X. Bressaud and G. Poggiaspalla. A tentative classification of bijective polygonal piecewise isometries. Experimental Mathematics, 16(1):77–100, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1175789803.

Berkovich:2015:EUR

[BP15] Alexander Berkovich and Frank Patane. Essentially unique representations by certain ternary quadratic forms. Experimental Mathematics, 24(1):8–22, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Berg:2021:FHB

[BP21] Christian Berg and Henrik L. Pedersen. A family of Horn– Bernstein functions. Experimental Mathematics, 32(3):505– 513, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2021.1980460.

Bordg:2022:STT

[BPL22] Anthony Bordg, Lawrence Paulson, and Wenda Li. Simple type theory is not too simple: Grothendieck’s schemes without dependent types. Experimental Mathematics, 31 (2):364–382, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2062073.

Banks:2012:GSR

[BPS12] William D. Banks, Francesco Pappalardi, and Igor E. Shparlinski. On group structures realized by elliptic curves over arbitrary finite fields. Experimental Mathematics, 21(1):11– 25, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Brakke:1995:CEF

[BPW95] Kenneth Brakke, Robert Phelan, and Denis Weaire. Computation of equilibrium foam structures using the surface evolver. Experimental Mathematics, 4(3):181–192, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621077.

Batut:1995:CCL

[BQS95] Christian Batut, Heinz-Georg Quebbemann, and Rudolf Scharlau. Computations of cyclotomic lattices. Experimental Mathematics, 4(3):177–179, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062621076.

Bogner:2017:SFO

[BR17] Michael Bogner and Stefan Reiter. Some fourth order CYtype operators with non-symplectically rigid monodromy. Experimental Mathematics, 26(1):98–113, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Baker:2018:PCL

[BR18] M. D. Baker and A. W. Reid. Principal congruence links: Class number greater than 1. Experimental Mathematics, 27 (4):395–413, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1289867.

Brakke:1992:SE

[Bra92] Kenneth A. Brakke. The surface evolver. Experimental Mathematics, 1(2):141–165, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048709050.

Brakke:1995:NSS

[Bra95] Kenneth A. Brakke. Numerical solution of soap film dual problems. Experimental Mathematics, 4(4):269–287, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047674388.

Brander:2016:SS

[Bra16] David Brander. Spherical surfaces. Experimental Mathematics, 25(3):257–272, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bracken:2021:SPI

[Bra21] Anthony J. Bracken. Solutions of Painlevé II on real intervals: Novel approximating sequences. Experimental Mathematics, 30(2):147–156, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1507852.

Brent:1998:QMM

[Bre98] Barry Brent. Quadratic minima and modular forms. Experimental Mathematics, 7(3):257–274, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047674207.

Bremner:1999:APE

[Bre99] Andrew Bremner. On arithmetic progressions on elliptic curves. Experimental Mathematics, 8(4):409–413, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047262362.

Bremmer:2008:WCS

[Bre08a] Andrew Bremmer. When can (((X^{2} − P)^{2} − Q)^{2} − R)^{2} − S^{2} split into linear factors? Experimental Mathematics, 17(4): 385–390, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243429952.

Bremner:2008:E

[Bre08b] Andrew Bremner. On the equation Y^{2} = X^{5} + k. Experimental Mathematics, 17(3):371–374, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1227121389.

Brinkmann:2000:IBH

[Bri00] Peter Brinkmann. An implementation of the Bestvina– Handel algorithm for surface homeomorphisms. Experimental Mathematics, 9(2):235–240, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1045952347.

Brisebarre:2001:IM

[Bri01] Nicolas Brisebarre. Irrationality measures of log 2 and π/\sqrt{3}. Experimental Mathematics, 10(1):35–52, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/999188419.

Brinkmann:2004:NPA

[Bri04] Peter Brinkmann. A note on pseudo-Anosov maps with small growth rate. Experimental Mathematics, 13(1):49–54, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1086894089.

Briggs:2006:ANR

[Bri06] Keith Briggs. Abundant numbers and the Riemann hypothesis. Experimental Mathematics, 15(2):251–256, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789744.

Brown:2007:DPK

[Bro07] Gavin Brown. A database of polarized K3 surfaces. Experimental Mathematics, 16(1):7–20, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1175789798.

Brauchart:2018:RPS

[BRS^{+}18] J. S. Brauchart, A. B. Reznikov, E. B. Saff, I. H. Sloan, Y. G. Wang, and R. S. Womersley. Random point sets on the sphere-hole radii, covering, and separation. Experimental Mathematics, 27(1):62–81, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1226209.

Behn:2023:SPG

[BRTC23] Antonio Behn, Anita M. Rojas, and Miguel Tello-Carrera. A SAGE package for n-gonal equisymmetric stratification of M_{g}. Experimental Mathematics, 32(1):54–69, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1763872.

Bruns:2007:ICP

[Bru07] Winfried Bruns. On the integral Carathéodory property. Experimental Mathematics, 16(3):359–365, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1204928535.

Bobenko:1999:ASD

[BS99] Alexander I. Bobenko and Wolfgang K. Schief. Affine spheres: discretization via duality relations. Experimental Mathematics, 8(3):261–280, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047262407.

Brinkmann:2001:CTM

[BS01] Peter Brinkmann and Saul Schleimer. Computing triangulations of mapping tori of surface homeomorphisms. Experimental Mathematics, 10(4):571–582, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1069855257.

Borwein:2008:PRB

[BS08a] Jonathan M. Borwein and Bruno Salvy. A proof of a recurrence for Bessel moments. Experimental Mathematics, 17(2): 223–230, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227118973.

Bruin:2008:DER

[BS08b] Nils Bruin and Michael Stoll. Deciding existence of rational points on curves: An experiment. Experimental Mathematics, 17(2):181–189, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227118970.

Bremner:2011:EGD

[BS11] David Bremner and Lars Schewe. Edge-graph diameter bounds for convex polytopes with few facets. Experimental Mathematics, 20(3):229–237, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1317924417.

Bruinier:2012:CHW

[BS12] Jan H. Bruinier and Fredrik Strömberg. Computation of harmonic weak Maass forms. Experimental Mathematics, 21(2): 117–131, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bernstein:2017:NBH

[BS17] Daniel Irving Bernstein and Seth Sullivant. Normal binary hierarchical models. Experimental Mathematics, 26(2):153–164, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1142911.

Bopp:2021:VGC

[BS21] Christian Bopp and Frank-Olaf Schreyer. A version of Green’s conjecture in positive characteristic. Experimental Mathematics, 30(4):475–480, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1576082.

Bajpai:2020:CAE

[BST20] Jitendra Bajpai, Sandip Singh, and Scott Thomson. Commensurability and arithmetic equivalence for orthogonal hypergeometric monodromy groups. Experimental Mathematics, 29(2):222–233, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2018.1453424.

Bielefeld:1993:DCN

[BSTV93] Ben Bielefeld, Scott Sutherland, Folkert Tangerman, and J. J. P. Veerman. Dynamics of certain nonconformal degreetwo maps of the plane. Experimental Mathematics, 2(4):281– 300, ???? 1993. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048516038.

Borwein:2013:TSF

[BSW13] Jonathan M. Borwein, Armin Straub, and James Wan. Three-step and four-step random walk integrals. Experimental Mathematics, 22(1):1–14, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Barker:2024:CRM

[BSW24] Mary Barker, Benjamin Standaert, and Ben Wormleighton. Crepant resolutions, mutations, and the space of potentials. Experimental Mathematics, 34(1):13–26, 2024. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2024.2308246.

Bagley:2018:CAB

[BT18] Zachary Bagley and Ray Treinen. On the classification and asymptotic behavior of the symmetric capillary surfaces. Experimental Mathematics, 27(2):215–229, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1245641.

Boender:1996:FIL

[BtR96] Henk Boender and Herman J. J. te Riele. Factoring integers with large-prime variations of the quadratic sieve. Experimental Mathematics, 5(4):257–273, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047565445.

Barker:2023:BLM

[BTZK23] James Barker, Christoph Thiele, and Pavel Zorin-Kranich. Band-limited maximizers for a Fourier extension inequality on the circle, II. Experimental Mathematics, 32(2):280–293, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926011.

Berkovich:2023:WDMb

[BU23] Alexander Berkovich and Ali Kemal Uncu. Where do the maximum absolute q-series coefficients of (1 − q)(1 − q^{2})(1 − q^{3}) ⋯ (1 − q^{n−1})(1 − q^{n}) occur? Experimental Mathematics, 32(1):82–87, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1776177.

Burton:2004:IRM

[Bur04] Benjamin A. Burton. Introducing Regina, the 3-manifold topology software. Experimental Mathematics, 13(3):267– 272, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1103749834.

Burton:2007:OTN

[Bur07] Benjamin A. Burton. Observations from the 8-tetrahedron nonorientable census. Experimental Mathematics, 16(2):129– 144, ???? 2007. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1204905871.

Burton:2010:QOC

[Bur10] Benjamin A. Burton. Quadrilateral-octagon coordinates for almost normal surfaces. Experimental Mathematics, 19(3): 285–315, ???? 2010. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317758093.

Burton:2014:DPS

[Bur14] Benjamin A. Burton. A duplicate pair in the SnapPea census. Experimental Mathematics, 23(2):170–173, 2014. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Button:2005:FVF

[But05] J. O. Button. Fibred and virtually fibred hyperbolic 3manifolds in the censuses. Experimental Mathematics, 14(2): 231–255, ???? 2005. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1128100134.

Button:2011:PFP

[But11] J. O. Button. Proving finitely presented groups are large by computer. Experimental Mathematics, 20(2):153–168, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317924407.

Butler:2012:SSE

[But12] Leo T. Butler. Smooth structures on Eschenburg spaces: Numerical computations. Experimental Mathematics, 21(1):57– 64, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Browning:2012:STS

[BV12] T. D. Browning and K. Van Valckenborgh. Sums of three squareful numbers. Experimental Mathematics, 21(2):204– 211, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bandini:2022:SSD

[BV22] Andrea Bandini and Maria Valentino. On the structure and slopes of Drinfeld cusp forms. Experimental Mathematics, 31 (2):637–651, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1671921.

Bohning:2009:RMS

[BvBK09] Christian Böhning, Hans-Christian Graf von Bothmer, and Jakob Kröker. Rationality of moduli spaces of plane curves of small degree. Experimental Mathematics, 18(4):499–508, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158510.

Boston:2007:NPT

[BW07] Nigel Boston and Jay Wierer. Newton polytopes of twodimensional hidden Markov models. Experimental Mathematics, 16(2):227–238, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204905878.

Bennett:2012:LVS

[BY12] Michael A. Bennett and Soroosh Yazdani. A local version of Szpiro’s conjecture. Experimental Mathematics, 21(2):103– 116, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Barg:2013:NBS

[BY13] Alexander Barg and Wei-Hsuan Yu. New bounds for spherical two-distance sets. Experimental Mathematics, 22(2):187–194, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Blocki:2016:SCS

[BZ16] Zbigniew Blocki and Wlodzimierz Zwonek. On the Suita conjecture for some convex ellipsoids in C^{2}. Experimental Mathematics, 25(1):8–16, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Bouc:2017:QRT

[BZ17] Serge Bouc and Alexander Zimmermann. On a question of Rickard on tensor products of stably equivalent algebras. Experimental Mathematics, 26(1):31–44, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Carletti:2003:CBF

[Car03] Timoteo Carletti. The 1/2-complex Bruno function and the Yoccoz function: a numerical study of the Marmi–Moussa– Yoccoz conjecture. Experimental Mathematics, 12(4):491– 506, ???? 2003. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1087568025.

Cazals:1997:CPO

[Caz97] Frédéric Cazals. Combinatorial properties of one-dimensional arrangements. Experimental Mathematics, 6(1):87–94, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565286.

Crandall:1994:EES

[CB94] Richard E. Crandall and Joe P. Buhler. On the evaluation of Euler sums. Experimental Mathematics, 3(4):275–285, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048515810.

Castano-Bernard:2009:NRP

[CB09] Carlos Castaño-Bernard. A note on the rational points of X_{0}^{+}(N). Experimental Mathematics, 18(2):129–135, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158425.

Cheon:2007:CBB

[CC07] Younhwan Cheon and Thomas W. Cusick. Counting balanced Boolean functions in n variables with bounded degree. Experimental Mathematics, 16(1):101–106, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789804.

Cohn:2007:RSU

[CCEK07] Henry Cohn, John H. Conway, Noam D. Elkies, and Abhinav Kumar. The D_{4} root system is not universally optimal. Experimental Mathematics, 16(3):313–320, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1204928532.

Chenavier:2021:QMO

[CCG21] Cyrille Chenavier, Christophe Cordero, and Samuele Giraudo. Quotients of the magmatic operad: Lattice structures and convergent rewrite systems. Experimental Mathematics, 30(4):513–530, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1577768.

Cunningham:2009:CGH

[CD09] Clifton Cunningham and Lassina Dembélé. Computing genus-2 Hilbert–Siegel modular forms over Q(\sqrt{5}) via the Jacquet–Langlands correspondence. Experimental Mathematics, 18(3):337–345, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1259158470.

Cerveau:2010:FTP

[CD10] Dominique Cerveau and Julie Déserti. Feuilletages et transformations périodiques. (French) [Branching and periodic transformations]. Experimental Mathematics, 19(4):447–464, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758104.

CaluzaMachado:2018:ISP

[Cd18] Fabrício Caluza Machado and Fernando Mário de Oliveira Filho. Improving the semidefinite programming bound for the kissing number by exploiting polynomial symmetry. Experimental Mathematics, 27(3):362–369, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1286273.

Chen:2021:OV

[CD21] Justin Chen and Papri Dey. The 4 × 4 orthostochastic variety. Experimental Mathematics, 32(4):692–700, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1982427.

Cochrane:2016:MMM

[CDD^{+}16] Todd Cochrane, R. M. S. Dissanayake, Nicholas Donohoue, M. I. M. Ishak, Vincent Pigno, Chris Pinner, and Craig Spencer. Minimal Mahler measure in real quadratic fields. Experimental Mathematics, 25(2):107–115, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Calude:2002:CGR

[CDS02] Cristian S. Calude, Michael J. Dinneen, and Chi-Kou Shu. Computing a glimpse of randomness. Experimental Mathematics, 11(3):361–370, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057777428.

Capparelli:2022:SHP

[CDV22] Stefano Capparelli, Alberto Del Fra, and Andrea Vietri. Searching for hyperbolic polynomials with span less than 4. Experimental Mathematics, 31(3):830–842, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1706670.

Conway:2006:MGP

[CEM06] John H. Conway, Noam D. Elkies, and Jeremy L. Martin. The Mathieu group M_{12} and its pseudogroup extension M_{13}. Experimental Mathematics, 15(2):223–236, ???? 2006. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1175789742.

Cassaigne:1995:CAS

[CF95] Julien Cassaigne and Steven R. Finch. A class of 1-additive sequences and quadratic recurrences. Experimental Mathematics, 4(1):49–60, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1062621142.

Calkin:1996:CPS

[CF96] Neil J. Calkin and Steven R. Finch. Conditions on periodicity for sum-free sets. Experimental Mathematics, 5(2):131–137, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565644.

Chinburg:2015:PQU

[CFH^{+}15] Ted Chinburg, Holley Friedlander, Sean Howe, Michiel Kosters, Bhairav Singh, Matthew Stover, Ying Zhang, and Paul Ziegler. Presentations for quaternionic S-unit groups. Experimental Mathematics, 24(2):175–182, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cai:2021:LDM

[CFH^{+}21] Zhaodong Cai, Matthew Faust, A. J. Hildebrand, Junxian Li, and Yuan Zhang. Leading digits of Mersenne numbers. Experimental Mathematics, 30(3):405–421, 2021. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1551162.

Celik:2023:CTD

[ÇFM23] Türkü Özlüm Çelik, Samantha Fairchild, and Yelena Mandelshtam. Crossing the transcendental divide: From translation surfaces to algebraic curves. Experimental Mathematics, 33(4):569–587, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2023.2203413.

Cunningham:2009:MPC

[CG09] Clifton Cunningham and Julia Gordon. Motivic proof of a character formula for SL(2). Experimental Mathematics, 18 (1):11–44, ???? 2009. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243430527.

Chapoton:2014:EOB

[CG14] Frédéric Chapoton and Samuele Giraudo. Enveloping operads and bicolored noncrossing configurations. Experimental Mathematics, 23(3):332–349, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Chandler:2023:SHP

[CG23] Alex Chandler and Eugene Gorsky. Structures in HOMFLY-PT homology. Experimental Mathematics, 33(4):817–842, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2300412.

Coulson:2000:CAI

[CGHN00] David Coulson, Oliver A. Goodman, Craig D. Hodgson, and Walter D. Neumann. Computing arithmetic invariants of 3-manifolds. Experimental Mathematics, 9(1):127–152, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1046889596.

Chum:2020:NSA

[CGJM20] K. Chum, R. K. Guy, M. J. Jacobson, Jr., and A. S. Mosunov. Numerical and statistical analysis of aliquot sequences. Experimental Mathematics, 29(4):414–425, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1477077.

Coates:2020:QPC

[CGKS20] Tom Coates, Sergey Galkin, Alexander Kasprzyk, and Andrew Strangeway. Quantum periods for certain fourdimensional Fano manifolds. Experimental Mathematics, 29 (2):183–221, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1448018.

Castro:2015:AVE

[CGM15] Francis N. Castro, Oscar E. González, and Luis A. Medina. The p-adic valuation of Eulerian numbers: Trees and Bernoulli numbers. Experimental Mathematics, 24(2):183– 195, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Claesson:2023:CPS

[CGP23] Anders Claesson, Bjarki Ágúst Gumundsson, and Jay Pantone. Counting pop-stacked permutations in polynomial time. Experimental Mathematics, 32(1):97–104, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926001.

Cannon:2008:TPG

[CH08a] John J. Cannon and Derek F. Holt. The transitive permutation groups of degree 32. Experimental Mathematics, 17(3): 307–314, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227121384.

Cho:2008:MDG

[CH08b] Jin-Hwan Cho and Ji-Young Ham. The minimal dilatation of a genus-two surface. Experimental Mathematics, 17(3): 257–267, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227121381.

Cannon:2013:CPI

[CH13] John J. Cannon and Derek F. Holt. Computing projective indecomposable modules and higher cohomology groups. Experimental Mathematics, 22(1):51–59, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Cashen:2021:SHI

[CH21a] Christopher H. Cashen and Charlotte Hoffmann. Short, highly imprimitive words yield hyperbolic one-relator groups. Experimental Mathematics, 32(4):631–640, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1982079.

Conrey:2021:SVQ

[CH21b] J. B. Conrey and M. A. Holmstrom. Smooth values of quadratic polynomials. Experimental Mathematics, 30(4): 447–452, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1559775.

Chas:2014:SIN

[Cha14] Moira Chas. Self-intersection numbers of length-equivalent curves on surfaces. Experimental Mathematics, 23(3):271– 276, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Chae:2023:CKB

[Cha23] John Chae. A cable knot and BPS-series II. Experimental Mathematics, 34(1):1–12, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2023.2300432.

Chen:2019:MTS

[Che19] Hao Chen. Minimal twin surfaces. Experimental Mathematics, 28(4):404–419, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2017.1413455.

Chipalkatti:2002:DTC

[Chi02] Jaydeep V. Chipalkatti. Decomposable ternary cubics. Experimental Mathematics, 11(1):69–80, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1057860315.

Chipalkatti:2004:WLT

[Chi04] Jaydeep Chipalkatti. The Waring loci of ternary quartics. Experimental Mathematics, 13(1):93–102, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1086894091.

Chiloyan:2024:AGI

[Chi24] Garen Chiloyan. 2-adic Galois images of non-CM isogenytorsion graphs defined over Q. Experimental Mathematics, 34(1):87–125, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2024.2336060.

Callahan:1993:FSP

[CHK93] Michael Callahan, David Hoffman, and Hermann Karcher. A family of singly periodic minimal surfaces invariant under a screw motion. Experimental Mathematics, 2(3):157–182, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062620829.

Cooperman:1997:BTP

[CHLM97] Gene Cooperman, Gerhard Hiss, Klaus Lux, and Jürgen Müller. The Brauer tree of the principal 19-block of the sporadic simple Thompson group. Experimental Mathematics, 6(4):293–300, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047047190.

Conrey:2013:SN

[CHM13] J. B. Conrey, M. A. Holmstrom, and T. L. McLaughlin. Smooth neighbors. Experimental Mathematics, 22(2):195– 202, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Cellarosi:2015:CFD

[CHMW15] Francesco Cellarosi, Doug Hensley, Steven J. Miller, and Jake L. Wellens. Continued fraction digit averages and Maclaurin’s inequalities. Experimental Mathematics, 24(1): 23–44, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Columbus:2022:SGT

[CHMWS22] Tobias Columbus, Frank Herrlich, Bjoern Muetzel, and Gabriela Weitze-Schmithüsen. Systolic geometry of translation surfaces. Experimental Mathematics, 33(3):357–378, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2108946.

Chopp:1994:CSS

[Cho94] David L. Chopp. Computation of self-similar solutions for mean curvature flow. Experimental Mathematics, 3(1):1–15, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062620999.

Campbell:2007:ESG

[CHRR07] Colin M. Campbell, George Havas, Colin Ramsay, and Edmund F. Robertson. On the efficiency of the simple groups of order less than a million and their covers. Experimental Mathematics, 16(3):347–358, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1204928534.

Conway:1996:PLP

[CHS96] John H. Conway, Ronald H. Hardin, and Neil J. A. Sloane. Packing lines, planes, etc.: packings in Grassmannian spaces. Experimental Mathematics, 5(2):139–159, ???? 1996. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1047565645.

Chua:2005:EML

[Chu05] Kok Seng Chua. Extremal modular lattices, McKay Thompson series, quadratic iterations, and new series for π. Experimental Mathematics, 14(3):343–357, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1128371759.

Chiantini:2019:GRD

[CILO19] Luca Chiantini, Christian Ikenmeyer, J. M. Landsberg, and Giorgio Ottaviani. The geometry of rank decompositions of matrix multiplication I: 2 × 2 matrices. Experimental Mathematics, 28(3):322–327, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1403981.

Coulangeon:2007:LCE

[CIO07] R. Coulangeon, M. I. Icaza, and M. O’Ryan. Lenstra’s constant and extreme forms in number fields. Experimental Mathematics, 16(4):455–462, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1204836515.

Chen:2019:SSA

[CJK^{+}19] Weiru Chen, Mo Jiao, Calvin Kessler, Amita Malik, and Xin Zhang. Spatial statistics of Apollonian gaskets. Experimental Mathematics, 28(3):263–270, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1385037.

Castro-Jimenez:2004:TLC

[CJUE04] F. J. Castro-Jiménez and J. M. Ucha-Enríquez. Testing the logarithmic comparison theorem for free divisors. Experimental Mathematics, 13(4):441–449, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1109106436.

Cushing:2022:GCC

[CKL^{+}22] David Cushing, Riikka Kangaslampi, Valtteri Lipiäinen, Shiping Liu, and George W. Stagg. The graph curvature calculator and the curvatures of cubic graphs. Experimental Mathematics, 31(2):583–595, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1660740.

Chen:2023:RCV

[CKL^{+}23] Ryan C. Chen, Yujin H. Kim, Jared D. Lichtman, Steven J. Miller, Alina Shubina, Shannon Sweitzer, Ezra Waxman, Eric Winsor, and Jianing Yang. A refined conjecture for the variance of Gaussian primes across sectors. Experimental Mathematics, 32(1):33–53, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1753598.

Conrey:2006:RMT

[CKRS06] J. B. Conrey, J. P. Keating, M. O. Rubinstein, and N. C. Snaith. Random matrix theory and the Fourier coefficients of half-integral-weight forms. Experimental Mathematics, 15 (1):67–82, ???? 2006. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1150476905.

Cooper:1998:T

[CL98] Daryl Cooper and Darren D. Long. Trimming. Experimental Mathematics, 7(4):325–332, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674150.

Chua:2004:CSA

[CL04] Kok Seng Chua and Mong Lung Lang. Congruence subgroups associated to the monster. Experimental Mathematics, 13(3): 343–360, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1103749842.

Calini:2007:SRG

[CL07a] Annalisa Calini and Joel Langer. Schwarz reflection geometry II: Local and global behavior of the exponential map. Experimental Mathematics, 16(3):321–346, ???? 2007. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1204928533.

Cremona:2007:FAE

[CL07b] J. E. Cremona and M. P. Lingham. Finding all elliptic curves with good reduction outside a given set of primes. Experimental Mathematics, 16(3):303–312, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1204928531.

Cotterill:2022:RIP

[CL22] Ethan Cotterill and Cristhian Garay López. Real inflection points of real linear series on an elliptic curve. Experimental Mathematics, 31(2):506–517, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1655815.

Champanerkar:2007:ERA

[CLLM07] Abhijit Champanerkar, Jacob Lewis, Max Lipyanskiy, and Scott Meltzer. Exceptional regions and associated exceptional hyperbolic 3-manifolds, with an appendix by Alan W. Reid. Experimental Mathematics, 16(1):106–118, ???? 2007. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1175789805.

Chas:2013:ESD

[CLM13] Moira Chas, Keren Li, and Bernard Maskit. Experiments suggesting that the distribution of the hyperbolic length of closed geodesics sampling by word length is Gaussian. Experimental Mathematics, 22(4):367–371, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cioffi:2015:MCM

[CLMR15] F. Cioffi, P. Lella, M. G. Marinari, and M. Roggero. Minimal Castelnuovo–Mumford regularity for a given Hilbert polynomial. Experimental Mathematics, 24(4):424–437, 2015. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Clancy:2015:NJT

[CLP15] Julien Clancy, Timothy Leake, and Sam Payne. A note on Jacobians, Tutte polynomials, and two-variable zeta functions of graphs. Experimental Mathematics, 24(1):1–7, 2015. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cooper:2006:CVR

[CLT06] Daryl Cooper, Darren Long, and Morwen Thistlethwaite. Computing varieties of representations of hyperbolic 3manifolds into SL(4, R). Experimental Mathematics, 15(3): 291–305, ???? 2006. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1175789760.

Connor:2015:CEH

[CLW15] Peter Connor, Kevin Li, and Matthias Weber. Complete embedded harmonic surfaces in R^{3}. Experimental Mathematics, 24(2):196–224, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cohen:1992:SOP

[CLZ92] Henri Cohen, Leonard Lewin, and Don Zagier. A sixteenthorder polylogarithm ladder. Experimental Mathematics, 1(1): 25–34, ???? 1992. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048709113.

Cremona:2000:VES

[CM00] John E. Cremona and Barry Mazur. Visualizing elements in the Shafarevich–Tate group. Experimental Mathematics, 9 (1):13–28, ???? 2000. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1046889588.

Chapoton:2013:TSV

[CM13] F. Chapoton and L. Manivel. Triangulations and Severi varieties. Experimental Mathematics, 22(1):60–73, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Cueto:2016:HRT

[CM16] Maria Angelica Cueto and Hannah Markwig. How to repair tropicalizations of plane curves using modifications. Experimental Mathematics, 25(2):130–164, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Coyston:2021:SMM

[CM21] Joshua Coyston and James McKee. Small Mahler measures from digraphs. Experimental Mathematics, 32(3):527–539, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980462.

Colliander:2013:BMD

[CMOS13] James E. Colliander, Jeremy L. Marzuola, Tadahiro Oh, and Gideon Simpson. Behavior of a model dynamical system with applications to weak turbulence. Experimental Mathematics, 22(3):250–264, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cochrane:2021:GGK

[CMPR21] Todd Cochrane, Michael J. Mossinghoff, Chris Pinner, and C. J. Richardson. A generalization of the Goresky–Klapper Conjecture, Part II. Experimental Mathematics, 30(2):209– 220, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2018.1526723.

Coulon:2022:RMT

[CMST22] Rémi Coulon, Elisabetta A. Matsumoto, Henry Segerman, and Steve J. Trettel. Ray-marching Thurston geometries. Experimental Mathematics, 31(4):1197–1277, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2030262.

Costantini:2021:DSP

[CMZ21] Matteo Costantini, Martin Möller, and Jonathan Zachhuber. diffstrata — a Sage package for calculations in the tautological ring of the moduli space of Abelian differentials. Experimental Mathematics, 32(3):545–565, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980750.

Chongchitmate:2013:ALK

[CN13] Wutichai Chongchitmate and Lenhard Ng. An atlas of Legendrian knots. Experimental Mathematics, 22(1):26–37, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Cassou-Noguas:1996:FPR

[CNJ96] Philippe Cassou-Noguàs and Arnaud Jehanne. Formes primitives et représentations galoisiennes de type octaédral. (French) [Primitive forms and Galois representations of octahedal type]. Experimental Mathematics, 5(4):275–290, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565446.

Chmutova:2002:CCD

[CO02] Tanya Chmutova and Viktor Ostrik. Calculating canonical distinguished involutions in the affine Weyl groups. Experimental Mathematics, 11(1):99–117, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1057860319.

Cardinal:2020:VMI

[CO20] Jean-Paul Cardinal and Marius Overholt. A variant of Möbius inversion. Experimental Mathematics, 29(3):247–252, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1459960.

Cochet:2005:ERG

[Coc05] Charles Cochet. Effective reduction of Goresky–Kottwitz– MacPherson graphs. Experimental Mathematics, 14(2):133– 144, ???? 2005. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1128100126.

Coiculescu:2022:ISH

[Coi22] Matei P. Coiculescu. An interpolation from Sol to hyperbolic space. Experimental Mathematics, 31(4):1150–1176, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980454.

Collins:2005:FRH

[Col05] Pieter Collins. Forcing relations for homoclinic orbits of the Smale horseshoe map. Experimental Mathematics, 14(1):75– 86, ???? 2005. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1120145572.

Constantino:2006:CM

[Con06] Francesco Constantino. Complexity of 4-manifolds. Experimental Mathematics, 15(2):237–249, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789743.

Connor:2018:HGD

[Con18] Peter Connor. Higher genus doubly periodic minimal surfaces. Experimental Mathematics, 27(1):47–61, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1220339.

Conner:2021:RWD

[Con21] Austin Conner. A rank 18 Waring decomposition of sM_{⟨3⟩} with 432 symmetries. Experimental Mathematics, 30(3): 383–385, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1547231.

Cooper:2008:CCC

[Coo08] Shaun Cooper. The cubic Chan–Chua conjecture. Experimental Mathematics, 17(4):439–442, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1243429956.

Coquereaux:2020:TFL

[Coq20] Robert Coquereaux. Theta functions for lattices of SU(3) hyper-roots. Experimental Mathematics, 29(2):137–162, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1446062.

Coste:2002:SFL

[Cos02] Alain Costé. Sur un système fibré lié à la suite des nombres premiers. (French) [On a fiber system related to the prime-number sequence]. Experimental Mathematics, 11(3): 383–405, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057777430.

Cougnard:1994:AES

[Cou94] Jean Cougnard. Un anneau d’entiers stablement libre et non libre. Experimental Mathematics, 3(2):129–136, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062620906.

Cummins:2003:CSG

[CP03] C. J. Cummins and S. Pauli. Congruence subgroups of PSL(2, Z) of genus less than or equal to 24. Experimental Mathematics, 12(2):243–255, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1067634734.

Chas:2010:SIN

[CP10] Moira Chas and Anthony Phillips. Self-intersection numbers of curves on the punctured torus. Experimental Mathematics, 19(2):129–148, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1276784785.

Chas:2012:SIN

[CP12] Moira Chas and Anthony Phillips. Self-intersection numbers of curves in the doubly punctured plane. Experimental Mathematics, 21(1):26–37, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Chyzak:2001:COW

[CPS^{+}01] Frédéric Chyzak, Peter Paule, Otmar Scherzer, Armin Schoisswohl, and Burkhard Zimmermann. The construction of orthonormal wavelets using symbolic methods and a matrix analytical approach for wavelets on the interval. Experimental Mathematics, 10(1):67–86, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/999188421.

Charter:1993:DQF

[CR93] Kevin Charter and Thomas Rogers. The dynamics of quadrilateral folding. Experimental Mathematics, 2(3):209–222, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062620831.

Chan:1999:VPD

[CR99] Clara S. Chan and David P. Robbins. On the volume of the polytope of doubly stochastic matrices. Experimental Mathematics, 8(3):291–300, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047262409.

Coutinho:2001:HFA

[CR01] S. C. Coutinho and Bruno F. M. Ribeiro. On holomorphic foliations without algebraic solutions. Experimental Mathematics, 10(4):529–536, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069855252.

Chapman:2022:MCS

[CR22] Harrison Chapman and Andrew Rechnitzer. A Markov chain sampler for plane curves. Experimental Mathematics, 31(2): 552–582, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1660739.

Crandall:1999:NRM

[Cra99a] Richard E. Crandall. New representations for the Madelung constant. Experimental Mathematics, 8(4):367–379, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047262358.

Crass:1999:SSI

[Cra99b] Scott Crass. Solving the sextic by iteration: a study in complex geometry and dynamics. Experimental Mathematics, 8(3):209–240, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047262404.

Crass:2001:SQI

[Cra01] Scott Crass. Solving the quintic by iteration in three dimensions. Experimental Mathematics, 10(1):1–24, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
999188417.

Crass:2014:DSB

[Cra14] Scott Crass. Dynamics of a soccer ball. Experimental Mathematics, 23(3):261–270, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cremona:1997:CPC

[Cre97] John E. Cremona. Computing periods of cusp forms and modular elliptic curves. Experimental Mathematics, 6(2):97– 107, ???? 1997. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047649997.

Chan:2000:VCP

[CRY00] Clara S. Chan, David P. Robbins, and David S. Yuen. On the volume of a certain polytope. Experimental Mathematics, 9 (1):91–99, ???? 2000. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1046889594.

Chopp:1993:FUC

[CS93] David L. Chopp and James A. Sethian. Flow under curvature: singularity formation, minimal surfaces, and geodesics. Experimental Mathematics, 2(4):235–255, ???? 1993. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048516035.

Cid:2001:CFS

[CS01] Carlos Cid and Tilman Schulz. Computation of five- and six-dimensional Bieberbach groups. Experimental Mathematics, 10(1):109–115, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/999188425.

Cheltsov:2013:PZ

[CS13] Ivan Cheltsov and Constantin Shramov. Del Pezzo zoo. Experimental Mathematics, 22(3):313–326, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cloninger:2021:DGL

[CS21] Alexander Cloninger and Stefan Steinerberger. On the dual geometry of Laplacian eigenfunctions. Experimental Mathematics, 30(2):283–293, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2018.1538911.

Corless:2022:SOA

[CS22a] Robert M. Corless and Leili Rafiee Sevyeri. Stirling’s original asymptotic series from a formula like one of Binet’s and its evaluation by sequence acceleration. Experimental Mathematics, 31(1):184–191, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1593898.

Criado:2022:TPS

[CS22b] Francisco Criado and Francisco Santos. Topological prismatoids and small simplicial spheres of large diameter. Experimental Mathematics, 31(2):461–473, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1641766.

Costa:2014:VNS

[CT14] Edgar Costa and Yuri Tschinkel. Variation of Néron–Severi ranks of reductions of K3 surfaces. Experimental Mathematics, 23(4):475–481, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Cohen:1996:ISD

[CtR96] Graeme L. Cohen and Herman J. J. te Riele. Iterating the sum-of-divisors function. Experimental Mathematics, 5(2): 91–100, ???? 1996. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047565640. See errata [CtR97].

Cohen:1997:EIS

[CtR97] Graeme L. Cohen and Herman J. J. te Riele. Errata to: “Iterating the sum-of-divisors function”. Experimental Mathematics, 6(2):177, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047650004. See [CtR96].

Cullinan:2019:JCD

[Cul19] John Cullinan. On the Jacobians of curves defined by the generalized Laguerre polynomials. Experimental Mathematics, 28 (2):223–232, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1382405.

Cummins:2004:CSG

[Cum04] C. J. Cummins. Congruence subgroups of groups commensurable with PSL(2, Z) of genus 0 and 1. Experimental Mathematics, 13(3):361–382, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1103749843.

Cuntz:2017:WFP

[Cun17] Michael Cuntz. On wild Frieze patterns. Experimental Mathematics, 26(3):342–348, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2016.1172526.

Chopp:2003:FHS

[CV03] David Chopp and John A. Velling. Foliations of hyperbolic space by constant mean curvature surfaces sharing ideal boundary. Experimental Mathematics, 12(3):339–350, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1087329236.

Cohen:2000:CAA

[CVZ00] Henri Cohen, Fernando Rodriguez Villegas, and Don Zagier. Convergence acceleration of alternating series. Experimental Mathematics, 9(1):3–12, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1046889587.

DaSilva:2019:SBS

[Da 19] Sergio Da Silva. Strict Bott–Samelson resolutions of Schubert varieties. Experimental Mathematics, 28(3):313–321, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1398695.

Das:2023:FIM

[Das23] Angsuman Das. A family of iterated maps on natural numbers. Experimental Mathematics, 33(4):776–780, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2294184.

Donten-Bury:2019:CRD

[DBG19] Maria Donten-Bury and Maksymilian Grab. Crepant resolutions of 3-Dimensional quotient singularities via Cox rings. Experimental Mathematics, 28(2):161–180, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1368049.

duCloux:2002:CKL

[dC02] Fokko du Cloux. Computing Kazhdan–Lusztig polynomials for arbitrary Coxeter groups. Experimental Mathematics, 11(3):371–381, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057777429.

deCourcy-Ireland:2022:EMS

[dCIL22] Matthew de Courcy-Ireland and Seungjae Lee. Experiments with the Markoff surface. Experimental Mathematics, 31 (3):814–829, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1702123.

Delaunay:2003:NIR

[DD03a] C. Delaunay and S. Duquesne. Numerical investigations related to the derivatives of the L-series of certain elliptic curves. Experimental Mathematics, 12(3):311–318, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1087329234.

Deza:2003:HCS

[DD03b] Michel Deza and Mathieu Dutour. The hypermetric cone on seven vertices. Experimental Mathematics, 12(4):433–440, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1087568019.

Davis:2014:STS

[DDG^{+}14] Ryan Davis, Charles Doran, Adam Gewiss, Andrey Novoseltsev, Dmitri Skjorshammer, Alexa Syryczuk, and Ursula Whitcher. Short tops and semistable degenerations. Experimental Mathematics, 23(4):351–362, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

DeCorte:2014:FAF

[DdLV14] Evan DeCorte, David de Laat, and Frank Vallentin. Fourier analysis on finite groups and the Lovász θ-number of Cayley graphs. Experimental Mathematics, 23(2):146–152, 2014. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

DeLeo:2006:ATP

[De 06a] Roberto De Leo. Appendix to Topology of Plane Sections of Periodic Polyhedra with an Application to the Truncated Octahedron. Experimental Mathematics, 15(1):??, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/

1150476909. See [De 06b].

DeLeo:2006:TPS

[De 06b] Roberto De Leo. Topology of plane sections of periodic polyhedra with an application to the truncated octahedron. Experimental Mathematics, 15(1):109–124, ???? 2006. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1150476908.

See appendix [De 06a].

DeBruyn:2015:EGQ

[De 15a] Bart De Bruyn. Every generalized quadrangle of order 5 having a regular point is symplectic. Experimental Mathematics, 24(1):45–52, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

DeLeo:2015:CHD

[De 15b] Roberto De Leo. A conjecture on the Hausdorff dimension of attractors of real self-projective iterated function systems. Experimental Mathematics, 24(3):270–288, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Dress:1993:FSFb

[DE93] François Dress and Mohamed El Marraki. Fonction sommatoire de la fonction de Möbius. 2. Majorations asymptotiques élémentaires. (French) [upper bounds on the Möbius function. 2. Elementary asymptotic maximizations]. Experimental Mathematics, 2(2):99–112, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048516215.

Dekimpe:2002:CAG

[DE02] Karel Dekimpe and Bettina Eick. Computational aspects of group extensions and their applications in topology. Experimental Mathematics, 11(2):183–200, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1062621214.

Damanik:2023:DID

[DEFM23] David Damanik, Mark Embree, Jake Fillman, and May Mei. Discontinuities of the integrated density of states for Laplacians associated with Penrose and Ammann–Beenker tilings. Experimental Mathematics, 33(4):588–610, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2206589.

Deleglise:1998:BDA

[Del98] Marc Deléglise. Bounds for the density of abundant integers. Experimental Mathematics, 7(2):137–143, ???? 1998. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048515661.

Delaunay:2001:HTS

[Del01] Christophe Delaunay. Heuristics on Tate–Shafarevitch groups of elliptic curves defined over Q. Experimental Mathematics, 10(2):191–196, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/999188631.

Dembele:2005:ECH

[Dem05] Lassina Dembélé. Explicit computations of Hilbert modular forms on Q(\sqrt{5}). Experimental Mathematics, 14(4):457–466, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1136926976.

Dembele:2008:AME

[Dem08] Lassina Dembélé. An algorithm for modular elliptic curves over real quadratic fields. Experimental Mathematics, 17(4): 427–438, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243429955.

Deraux:2005:DDM

[Der05] Martin Deraux. Dirichlet domains for the mostow lattices. Experimental Mathematics, 14(4):467–490, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1136926977.

Deraux:2015:SCU

[Der15] Martin Deraux. On spherical CR uniformization of 3manifolds. Experimental Mathematics, 24(3):355–370, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Deraux:2022:SFI

[Der22] Martin Deraux. On subgroups finite index in complex hyperbolic lattice triangle groups. Experimental Mathematics, 33 (3):456–481, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2158969.

Diao:2018:RFA

[DERZ18] Yuanan Diao, Claus Ernst, Eric J. Rawdon, and Uta Ziegler. Relative frequencies of alternating and nonalternating prime knots and composite knots in random knot spaces. Experimental Mathematics, 27(4):454–471, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1320239.

Doche:2000:IGR

[DF00] Christophe Doche and Michel Mendès France. Integral geometry and real zeros of Thue–Morse polynomials. Experimental Mathematics, 9(3):339–350, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1045604669.

Detinko:2020:AEZ

[DFH20] A. S. Detinko, D. L. Flannery, and A. Hulpke. Algorithms for experimenting with Zariski dense subgroups. Experimental Mathematics, 29(3):296–305, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1466217.

Detinko:2023:ESH

[DFH23] A. S. Detinko, D. L. Flannery, and A. Hulpke. Experimenting with symplectic hypergeometric monodromy groups. Experimental Mathematics, 32(1):88–96, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1780516.

Dunfield:2012:TAP

[DFJ12] Nathan M. Dunfield, Stefan Friedl, and Nicholas Jackson. Twisted Alexander polynomials of hyperbolic knots. Experimental Mathematics, 21(4):329–352, 2012. ISSN 1058-6458 (print), 1944-950X (electronic).

David:2004:VTF

[DFK04] Chantal David, Jack Fearnley, and Hershy Kisilevsky. On the vanishing of twisted L-functions of elliptic curves. Experimental Mathematics, 13(2):185–198, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1090350933.

deFaria:2014:SPP

[dFT14] Edson de Faria and Charles Tresser. On Sloane’s persistence problem. Experimental Mathematics, 23(4):363–382, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Durrett:1993:ABE

[DG93] Richard Durrett and David Griffeath. Asymptotic behavior of excitable cellular automata. Experimental Mathematics, 2(3): 183–208, ???? 1993. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062620830.

Darmon:2002:ECC

[DG02a] Henri Darmon and Peter Green. Elliptic curves and class fields of real quadratic fields: Algorithms and evidence. Experimental Mathematics, 11(1):37–55, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1057860313.

deGraaf:2002:CCB

[dG02b] Willem A. de Graaf. Constructing canonical bases of quantized enveloping algebras. Experimental Mathematics, 11(2): 161–170, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621212.

deGraaf:2005:CSL

[dG05] W. A. de Graaf. Classification of solvable Lie algebras. Experimental Mathematics, 14(1):15–25, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1120145567.

Delbourgo:2021:CIS

[DG21] Daniel Delbourgo and Hamish Gilmore. Computing Linvariants for the symmetric square of an elliptic curve. Experimental Mathematics, 30(1):32–55, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1490936.

Daniels:2022:SCE

[DGJ22] Harris B. Daniels and Enrique González-Jiménez. Serre’s constant of elliptic curves over the rationals. Experimental Mathematics, 31(2):518–536, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1655816.

Donnelly:2015:TEC

[DGKMY15] Steve Donnelly, Paul E. Gunnells, Ariah Klages-Mundt, and Dan Yasaki. A table of elliptic curves over the cubic field of discriminant-23. Experimental Mathematics, 24(4):375–390, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Dunfield:2006:SKH

[DGR06] Nathan M. Dunfield, Sergei Gukov, and Jacob Rasmussen. The superpolynomial for knot homologies. Experimental Mathematics, 15(2):129–159, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1175789736.

Dasbach:1997:DJP

[DH97] Oliver T. Dasbach and Stefan Hougardy. Does the Jones polynomial detect unknottedness? Experimental Mathematics, 6 (1):51–56, ???? 1997. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047565283.

DAndrea:2004:HSR

[DH04] Carlos D’Andrea and Kevin G. Hare. On the height of the Sylvester resultant. Experimental Mathematics, 13(3):331– 341, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1103749841.

Downarowicz:2022:EAC

[DH22] Tomasz Downarowicz and Dawid Huczek. Empirical approach to the ×2, ×3 conjecture. Experimental Mathematics, 31(1):252–268, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1608603.

Dykema:2015:FPG

[DHJ15] Ken Dykema, Timo Heister, and Kate Juschenko. Finitely presented groups related to Kaplansky’s direct finiteness conjecture. Experimental Mathematics, 24(3):326–338, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Dhillon:2008:CPG

[DHST08] I. S. Dhillon, R. W. Heath, T. Strohmer, and J. A. Tropp. Constructing packings in Grassmannian manifolds via alternating projection. Experimental Mathematics, 17(1):9–35, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227031894.

Dieulefait:2002:EDI

[Die02] Luis V. Dieulefait. Explicit determination of the images of the Galois representations attached to Abelian surfaces with End(A) = Z. Experimental Mathematics, 11(4):503–512, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1057864660.

Dieulefait:2004:ENG

[Die04a] Luis Dieulefait. Existence of nonelliptic modℓ Galois representations for every ℓ > 5. Experimental Mathematics, 13(3): 327–329, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1103749840.

Dieulefait:2004:CLM

[Die04b] Luis V. Dieulefait. Computing the level of a modular rigid Calabi–Yau threefold. Experimental Mathematics, 13(2):165– 170, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1090350931.

Dilcher:2000:NSE

[Dil00] Karl Dilcher. Nested squares and evaluations of integer products. Experimental Mathematics, 9(3):369–372, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604671.

Driver:2010:MDI

[DJ10] Eric D. Driver and John W. Jones. Minimum discriminants of imprimitive decic fields. Experimental Mathematics, 19(4): 475–479, ???? 2010. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317758106.

DiazYDiaz:2005:NAC

[DJP^{+}05] Francisco Diaz Y.Diaz, Jean-François Jaulent, Sebastian Pauli, Michael Pohst, and Florence Soriano-Gafiuk. A new algorithm for the computation of logarithmic ℓ-class groups of number fields. Experimental Mathematics, 14(1):65–74, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1120145571.

Diller:2021:ERR

[DK21] Jeffrey Diller and Kyounghee Kim. Entropy of real rational surface automorphisms. Experimental Mathematics, 30 (2):172–190, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1516581.

Dzamonja:2022:FOP

[DKAP22] Mirna Dzamonja, Angeliki Koutsoukou-Argyraki, and Lawrence C. Paulson. Formalizing ordinal partition relations using Isabelle/HOL. Experimental Mathematics, 31(2):383–400, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980464.

Dupuy:2021:CCA

[DKRV21] Taylor Dupuy, Kiran Kedlaya, David Roe, and Christelle Vincent. Counterexamples to a conjecture of Ahmadi and Shparlinski. Experimental Mathematics, 32(3):540–544, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980463.

Drake:1992:NCP

[DL92] David A. Drake and Jean A. Larson. The nonexistence of certain projective planes of order 15. Experimental Mathematics, 1(1):65–69, ???? 1992. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048709117.

Dong:2004:IRC

[DL04] Xin-Han Dong and Ka-Sing Lau. An integral related to the Cauchy transform on the Sierpinski gasket. Experimental Mathematics, 13(4):415–419, ???? 2004. CODEN ????

ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1109106432.

Domokos:2022:SAP

[DL22] Gábor Domokos and Zsolt Lángi. On some average properties of convex mosaics. Experimental Mathematics, 31(3): 783–793, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1691090.

Duchin:2013:SCC

[DLM13] Moon Duchin, Samuel Lelièvre, and Christopher Mooney. The sprawl conjecture for convex bodies. Experimental Mathematics, 22(2):113–122, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

David:2023:CMA

[DLN23] Chantal David, Matilde Lalín, and Jungbae Nam. Conjectures for moments associated with cubic twists of elliptic curves. Experimental Mathematics, 32(1):105–132, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926002.

Domokos:2022:TCP

[DLS22] Gábor Domokos, Zsolt Lángi, and András A. Sipos. Tracking critical points on evolving curves and surfaces. Experimental Mathematics, 31(1):1–20, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1556136.

DeLoera:2006:CCG

[DM06] Jesús A. De Loera and Tyrrell B. McAllister. On the computation of Clebsch–Gordan coefficients and the dilation effect. Experimental Mathematics, 15(1):7–19, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1150476899.

Dietrich:2022:TGF

[DM22] Heiko Dietrich and Primoz Moravec. On two group functors extending Schur multipliers. Experimental Mathematics, 31 (4):1332–1344, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1796857.

Diaconis:2004:NRM

[DN04] Persi Diaconis and J. W. Neuberger. Numerical results for the Metropolis algorithm. Experimental Mathematics, 13(2): 207–214, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1090350935.

Dumas:2021:ONA

[DN21] David Dumas and Andrew Neitzke. Opers and non-Abelian Hodge: Numerical studies. Experimental Mathematics, 33 (1):27–68, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1988006.

Dress:1999:PPV

[DO99] François Dress and Michel Olivier. Polynômes prenant des valeurs premières. (French) [Polynomials taking principal values]. Experimental Mathematics, 8(4):319–338, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047262355.

Dokchitser:2004:CSV

[Dok04] Tim Dokchitser. Computing special values of Motivic Lfunctions. Experimental Mathematics, 13(2):137–150, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1090350929.

Dolgachev:2018:SNE

[Dol18] Igor Dolgachev. Salem numbers and Enriques surfaces. Experimental Mathematics, 27(3):287–301, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1261743.

Doud:2005:WRG

[Dou05] Darrin Doud. Wildly ramified Galois representations and a generalization of a conjecture of Serre. Experimental Mathematics, 14(1):119–127, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1120145575.

Doud:2007:SGR

[Dou07] Darrin Doud. Supersingular Galois representations and a generalization of conjecture of Serre. Experimental Mathematics, 16(1):119–128, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1175789806.

Dujella:2008:SDT

[DP08] Andrej Dujella and Vinko Petričcević. Strong Diophantine triples. Experimental Mathematics, 17(1):83–89, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227031899.

Dudek:2015:SCI

[DP15] Adrian W. Dudek and David J. Platt. On solving a curious inequality of Ramanujan. Experimental Mathematics, 24(3): 289–294, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Deraux:2011:CCH

[DPP11] Martin Deraux, John R. Parker, and Julien Paupert. Census of the complex hyperbolic sporadic triangle groups. Experimental Mathematics, 20(4):467–486, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1323367158.

Drenning:2005:OBS

[DPPS05] Shawn Drenning, Judith Palagallo, Thomas Price, and Robert S. Strichartz. Outer boundaries of self-similar tiles. Experimental Mathematics, 14(2):199–209, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1128100132.

Duran:2015:SCW

[DPV15] Antonio J. Durán, Mario Pérez, and Juan L. Varona. Some conjectures on Wronskian and Casorati determinants of orthogonal polynomials. Experimental Mathematics, 24(1):123– 132, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Dembele:2022:SHM

[DPVZ22] Lassina Dembélé, Alexei Panchishkin, John Voight, and Wadim Zudilin. Special hypergeometric motives and their Lfunctions: Asai recognition. Experimental Mathematics, 31 (4):1278–1290, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1737990.

Deleglise:1996:CSM

[DR96] Marc Deléglise and Joöl Rivat. Computing the summation of the Möbius function. Experimental Mathematics, 5(4):291– 295, ???? 1996. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047565447.

Dudek:2008:FN

[DR08] Andrzej Dudek and Vojtĕech Rödl. On the Folkman number f(2, 3, 4). Experimental Mathematics, 17(1):63–67, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227031897.

Dzambic:2017:CSS

[DR17] Amir Dzambić and Xavier Roulleau. On the cohomology of the Stover surface. Experimental Mathematics, 26(4): 490–495, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1205533.

Dress:1993:FSFa

[Dre93] François Dress. Fonction sommatoire de la fonction de Möbius. 1. Majorations expérimentales. (French) [Upper bounds on the Möbius function. 1. Experimental maximizations]. Experimental Mathematics, 2(2):89–98, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048516214.

Delshams:1999:SSS

[DRR99] Amadeu Delshams and Rafael Ramírez-Ros. Singular separatrix splitting and the Melnikov method: an experimental study. Experimental Mathematics, 8(1):29–48, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047477110.

Dalton:2011:SBG

[DRT11] Barnaby Dalton, Mustazee Rahman, and Stephen Tanny. Spot-based generations for meta-Fibonacci sequences. Experimental Mathematics, 20(2):129–137, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317924404.

Dubejko:1995:CPE

[DS95] Tomasz Dubejko and Kenneth Stephenson. Circle packing: experiments in discrete analytic function theory. Experimental Mathematics, 4(4):307–348, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674391.

DeMarco:2013:GCP

[DS13] Laura De Marco and Aaron Schiff. The geometry of the critically periodic curves in the space of cubic polynomials. Experimental Mathematics, 22(1):99–111, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Dubon:2014:CDN

[DS14] E. Dubon and J. M. Sepulcre. On the complex dimensions of nonlattice fractal strings in connection with Dirichlet polynomials. Experimental Mathematics, 23(1):13–24, 2014. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Dubickas:2015:CTD

[DS15] Arturas Dubickas and Min Sha. Counting and testing dominant polynomials. Experimental Mathematics, 24(3):312–325, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Dods:2019:NMC

[DS19] Victor Dods and Corey Shanbrom. Numerical methods and closed orbits in the Kepler–Heisenberg problem. Experimental Mathematics, 28(4):420–427, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1416709.

Dummigan:2021:AFF

[DS21] Neil Dummigan and Sebastian Schönnenbeck. Automorphic forms on Feit’s Hermitian lattices. Experimental Mathematics, 30(4):557–574, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1581857.

Damiano:2007:CMC

[DSS07] Alberto Damiano, Irene Sabadini, and Daniele C. Struppa. Computational methods for the construction of a class of Noetherian operators. Experimental Mathematics, 16(1):41– 55, ???? 2007. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1175789800.

DeSchepper:2009:BFD

[DSV09] Hennie De Schepper, Frank Sommen, and Liesbet Van de Voorde. A basic framework for discrete Clifford analysis. Experimental Mathematics, 18(4):385–395, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1259158504.

Duke:2002:SPD

[DT02] W. Duke and Á. Tóth. The splitting of primes in division fields of elliptic curves. Experimental Mathematics, 11(4): 555–565, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057864664.

DiBiase:1995:ACK

[DU95] Fausto Di Biase and Rüdiger Urbanke. An algorithm to calculate the kernel of certain polynomial ring homomorphisms. Experimental Mathematics, 4(3):227–234, ???? 1995. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1062621080.

Dujella:2001:NFC

[Duj01] Andrej Dujella. Newton’s formula and the continued fraction expansion of \sqrt{d}. Experimental Mathematics, 10(1):125–131, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
999188427.

Duke:2023:HRS

[Duk23] W. Duke. Higher Rademacher symbols. Experimental Mathematics, 33(4):624–643, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2023.2219071.

Dummigan:2001:SSF

[Dum01]  Neil Dummigan.  Symmetric square L-functions and

Shafarevich–Tate groups. Experimental Mathematics, 10(3): 383–400, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069786346.

Dummigan:2002:SSE

[Dum02] Neil Dummigan. Symmetric squares of elliptic curves: Rational points and Selmer groups. Experimental Mathematics, 11(4):457–464, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057864655.

Dummigan:2013:STF

[Dum13] Neil Dummigan. A simple trace formula for algebraic modular forms. Experimental Mathematics, 22(2):123–131, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Duquesne:2001:IPE

[Duq01] Sylvain Duquesne. Integral points on elliptic curves defined by simplest cubic fields. Experimental Mathematics, 10(1): 91–102, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/999188423.

Duran:2011:RFO

[Dur11] Antonio J. Durán. Rodrigues’s formulas for orthogonal matrix polynomials satisfying higher-order differential equations. Experimental Mathematics, 20(1):15–24, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317924387.

deVerdiere:2022:LTS

[dVV22] Yves Colin de Verdière and David Vicente. Large-time series expansion of the wave front length in the Euclidean disk. Experimental Mathematics, 33(2):254–260, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2063208.

deWeger:1998:SED

[dW98] Benjamin M. M. de Weger. Solving elliptic Diophantine equations avoiding Thue equations and elliptic logarithms. Experimental Mathematics, 7(3):243–256, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674206.

Delbourgo:2010:GCP

[DW10] Daniel Delbourgo and Thomas Ward. The growth of CM periods over false Tate extensions. Experimental Mathematics, 19(2):195–210, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1276784790.

Develin:2007:TPC

[DY07] Mike Develin and Josephine Yu. Tropical polytopes and cellular resolutions. Experimental Mathematics, 16(3):277–292, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204928529.

Dzhumadildaev:2011:AOK

[DZ11] Askar Dzhumadil’daev and Pasha Zusmanovich. The alternative operad is not Koszul. Experimental Mathematics, 20(2): 138–144, ???? 2011. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317924405.

East:2020:TRS

[Eas20] James East. Transformation representations of sandwich semigroups. Experimental Mathematics, 29(3):291–295, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1459963.

Eelbode:2007:SNS

[Eel07] D. Eelbode. Stirling numbers and spin-Euler polynomials. Experimental Mathematics, 16(1):55–66, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789801.

Eick:2009:CCG

[EF09] Bettina Eick and Dörte Feichtenschlager. Constructing CF groups by coclass. Experimental Mathematics, 18(2):205–211, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158431.

Enolski:2018:ADJ

[EF18] V. Z. Enolski and Yu. N. Fedorov. Algebraic description of Jacobians isogeneous to certain Prym varieties with polarization (1, 2). Experimental Mathematics, 27(2):147–178, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1236357.

Evert:2021:EPO

[EFHY21] Eric Evert, Yi Fu, J. William Helton, and John Yin. Empirical properties of optima in free semidefinite programs. Experimental Mathematics, 32(3):477–501, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980457.

Eliahou:2018:MPT

[EFMPR18] Shalom Eliahou, Jean Fromentin, Virginie Marion-Poty, and Denis Robilliard. Are monochromatic Pythagorean triples unavoidable under morphic colorings? Experimental Mathematics, 27(4):419–425, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1306465.

Etingof:2012:CAP

[EGP^{+}12] Pavel Etingof, Sherry Gong, Aldo Pacchiano, Qingchun Ren, and Travis Schedler. Computational approaches to Poisson traces associated to finite subgroups of Sp_{2n}((C)). Experimental Mathematics, 21(2):141–170, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ekstrom:2018:EBS

[EGSC18] Sven-Erik Ekström, Carlo Garoni, and Stefano Serra-Capizzano. Are the eigenvalues of banded symmetric Toeplitz matrices known in almost closed form? Experimental Mathematics, 27(4):478–487, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1320241.

Elkenbracht-Huizing:1996:INF

[EH96] Marije Elkenbracht-Huizing. An implementation of the number field sieve. Experimental Mathematics, 5(3):231–253, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047915103.

Eliahou:2004:PSC

[EH04] Shalom Eliahou and Delphine Hachez. On a problem of Steinhaus concerning binary sequences. Experimental Mathematics, 13(2):215–230, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1090350936.

Eppstein:2020:GPA

[EHPN20] David Eppstein, Sariel Har-Peled, and Gabriel Nivasch. Grid peeling and the affine curve-shortening flow. Experimental Mathematics, 29(3):306–316, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1466379.

Edjvet:2001:CPT

[EHT01] Martin Edjvet, Paul Hammond, and Nathan Thomas. Cyclic presentations of the trivial group. Experimental Mathematics, 10(2):303–306, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/999188639.

Epstein:1996:GFA

[EIFZ96] David B. A. Epstein, Anthony R. Iano-Fletcher, and Uri Zwick. Growth functions and automatic groups. Experimental Mathematics, 5(4):297–315, ???? 1996. CODEN ????

ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047565448.

Elsenhans:2010:SPD

[EJ10] Andreas-Stephan Elsenhans and Jörg Jahnel. On the smallest point on a diagonal cubic surface. Experimental Mathematics, 19(2):181–193, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1276784789.

Elsenhans:2022:RCM

[EJ22] Andreas-Stephan Elsenhans and Jörg Jahnel. Real and complex multiplication on K3 surfaces via period integration. Experimental Mathematics, 33(2):193–224, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2061649.

Epstein:1992:ME

[EL92] David Epstein and Silvio Levy. Message from the Editors. Experimental Mathematics, 1(3):181, January 1992. ISSN 1058-6458 (print), 1944-950X (electronic).

Eisermann:2009:WTP

[EL09] Michael Eisermann and Christoph Lamm. For which triangles is Pick’s formula almost correct? Experimental Mathematics, 18(2):187–191, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1259158428.

Ellis:2014:HTL

[EL14] Graham Ellis and Luyen Van Le. Homotopy 2-types of low order. Experimental Mathematics, 23(4):383–389, 2014. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Epstein:1992:AJ

[ELdlL92] David Epstein, Silvio Levy, and Rafael de la Llave. About this journal. Experimental Mathematics, 1(1):??, ???? 1992. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048709111.

Elkies:2001:MWL

[Elk01] Noam D. Elkies. Mordell–Weil lattices in charactistic 2, III: a Mordell–Weil lattice of rank 128. Experimental Mathematics, 10(3):467–474, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069786351.

Elsenhans:2011:RPS

[Els11] Andreas-Stephan Elsenhans. Rational points on some Fano quadratic bundles. Experimental Mathematics, 20(4):373– 379, ???? 2011. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1323367152.

Elsenhans:2014:NSC

[Els14] Andreas-Stephan Elsenhans. A note on short cosets. Experimental Mathematics, 23(4):411–413, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Epstein:1997:MFJ

[ELT^{+}97] David Epstein, Elliott H. Lieb, Jean Taylor, Robert Almgren, Robert Kusner, Frank Morgan, John Sullivan, and Albert Marden. In memoriam: Frederick J. Almgren Jr., 1933–1997. Experimental Mathematics, 6(1):1–12, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047565280.

Elder:2002:CTD

[EM02] Murray Elder and Jon McCammond. Curvature testing in 3-dimensional metric polyhedral complexes. Experimental Mathematics, 11(1):143–160, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057860322.

Everest:2009:GST

[EM09] Graham Everest and Valéry Mahé. A generalization of Siegel’s theorem and Hall’s conjecture. Experimental Mathematics, 18(1):1–10, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1243430526.

Eick:2017:CTF

[EM17] Bettina Eick and Tobias Moede. Coclass theory for finite nilpotent associative algebras: Algorithms and a periodicity conjecture. Experimental Mathematics, 26(3):267–274, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1162229.

ElOtmani:2014:FDM

[EMRSÉ14] S. El Otmani, A. Maul, G. Rhin, and J.-M. Sac-Épée. Finding degree-16 monic irreducible integer polynomials of minimal trace by optimization methods. Experimental Mathematics, 23(1):1–5, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Engstrom:2009:DMF

[Eng09] Alexander Engström. Discrete Morse functions from Fourier transforms. Experimental Mathematics, 18(1):45–54, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1243430528.

Eguchi:2011:NSM

[EOT11] Tohru Eguchi, Hirosi Ooguri, and Yuji Tachikawa. Notes on the K3 surface and the Mathieu group M_{2}4. Experimental Mathematics, 20(1):91–96, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1317924391.

Eelbode:2015:PSA

[ER15] D. Eelbode and T. Raeymaekers. Polynomial solutions for arbitrary higher spin Dirac operators. Experimental Mathematics, 24(3):339–354, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Elder:2019:SDC

[ER19] Murray Elder and Cameron Rogers. Sub-dominant cogrowth behavior and the viability of deciding amenability numerically. Experimental Mathematics, 28(1):67–80, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1353452.

ElOtmani:2019:FNL

[ERSÉ19] S. El Otmani, G. Rhin, and J.-M. Sac-Épée. Finding new limit points of Mahler’s measure by genetic algorithms. Experimental Mathematics, 28(2):129–131, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1357511.

Elder:2015:RST

[ERvR15] Murray Elder, Andrew Rechnitzer, and Esaias J. Janse van Rensburg. Random sampling of trivial words in finitely presented groups. Experimental Mathematics, 24(4):391–409, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Edjvet:2014:ICP

[ES14] Martin Edjvet and Jerry Swan. On irreducible cyclic presentations of the trivial group. Experimental Mathematics, 23 (2):181–189, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Eddy:2022:NSN

[ES22] Thomas D. Eddy and Clayton Shonkwiler. New stick number bounds from random sampling of confined polygons. Experimental Mathematics, 31(4):1373–1395, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926000.

Escudero:2014:ARL

[Esc14] Juan García Escudero. Arrangements of real lines and surfaces with A and D singularities. Experimental Mathematics, 23(4):482–491, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Elias:2010:MPR

[ESTB10] Ben Elias, Lior Silberman, and Ramin Takloo-Bighash. Minimal permutation representations of nilpotent groups. Experimental Mathematics, 19(1):121–128, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1268404806.

Enge:2014:CCP

[ET14] Andreas Enge and Emmanuel Thomé. Computing class polynomials for Abelian surfaces. Experimental Mathematics, 23 (2):129–145, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Earl:2016:ICW

[EVV16] Jonathan Earl, Kevin N. Vander Meulen, and Adam Van Tuyl. Independence complexes of well-covered circulant graphs. Experimental Mathematics, 25(4):441–451, 2016. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Everest:1998:DIG

[EW98] Graham Everest and Thomas Ward. A dynamical interpretation of the global canonical height on an elliptic curve. Experimental Mathematics, 7(4):305–316, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047674148.

Faber:2009:GBC

[Fab09] X. W. C. Faber. The geometric Bogomolov conjecture for curves of small genus. Experimental Mathematics, 18(3):347– 367, ???? 2009. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1259158471.

Fermigier:1992:ZFC

[Fer92] Stéfane Fermigier. Zéros des fonctions L de courbes elliptiques. (French) [Zeros of L functions on elliptic curves]. Experimental Mathematics, 1(2):167–173, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1048709051.

Fermigier:1996:EED

[Fer96] Stéfane Fermigier. Étude expérimentale du rang de familles de courbes elliptiques sur Q. (French) [Experimental study of the rank of families of elliptic curves on Q]. Experimental Mathematics, 5(2):119–130, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047565643.

Fernique:2022:DBD

[Fer22] Thomas Fernique. Density of binary disc packings: Lower and upper bounds. Experimental Mathematics, 33(2):235–246, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2061651.

Fornaess:1999:TRC

[FG99] John Erik Fornæss and Estela A. Gavosto. Tangencies for real and complex Hénon maps: an analytic method. Experimental Mathematics, 8(3):253–260, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047262406.

Faber:2023:TQC

[FG23] Xander Faber and Jon Grantham. Ternary and quaternary curves of small fixed genus and gonality with many rational points. Experimental Mathematics, 32(2):337–349, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926015.

Fominykh:2016:CTH

[FGG^{+}16] Evgeny Fominykh, Stavros Garoufalidis, Matthias Goerner, Vladimir Tarkaev, and Andrei Vesnin. A census of tetrahedral hyperbolic manifolds. Experimental Mathematics, 25(4):466– 481, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Fischer:2022:ASP

[FGJ22] Greta Fischer, Jean Gutt, and Michael Jünger. Algorithmic symplectic packing. Experimental Mathematics, 33(2): 175–192, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2041135.

Falbel:2016:CVF

[FGK^{+}16] E. Falbel, A. Guilloux, P.-V. Koseleff, F. Rouillier, and M. Thistlethwaite. Character varieties for SL(3, C): The figure eight knot. Experimental Mathematics, 25(2):219–235, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Finis:2010:CL

[FGT10] Tobias Finis, Fritz Grunewald, and Paulo Tirao. The cohomology of lattices in SL(2, C). Experimental Mathematics, 19 (1):29–63, ???? 2010. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1268404802.

Freitag:1999:DIQ

[FH99] Eberhard Freitag and Bruce Hunt. The dual of the invariant quintic. Experimental Mathematics, 8(2):151–155, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047477058.

Fang:2010:MFN

[FHL^{+}10] Liqun Fang, J. William Hoffman, Benjamin Linowitz, Andrew Rupinski, and Helena Verrill. Modular forms on noncongruence subgroups and Atkin–Swinnerton-Dyer relations. Experimental Mathematics, 19(1):1–27, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1268404801.

Fujima:2018:NCN

[FI18] Shoichi Fujima and Humio Ichimura. Note on the class number of the p th cyclotomic field, II. Experimental Mathematics, 27(1):111–118, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1230528.

Fieker:2007:SRC

[Fie07] Claus Fieker. Sparse representation for cyclotomic fields. Experimental Mathematics, 16(4):493–500, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1204836517.

Forrester:2020:HME

[FIK20] P. J. Forrester, J. R. Ipsen, and S. Kumar. How many eigenvalues of a product of truncated orthogonal matrices are real? Experimental Mathematics, 29(3):276–290, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1459962.

Finch:1992:PAS

[Fin92] Steven R. Finch. Patterns in 1-additive sequences. Experimental Mathematics, 1(1):57–63, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048709116.

Firsching:2015:CMC

[Fir15] Moritz Firsching. Computing maximal copies of polyhedra contained in a polyhedron. Experimental Mathematics, 24 (1):98–105, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Furukado:2014:NMS

[FIS^{+}14] Maki Furukado, Shunji Ito, Asaki Saito, Jun ichi Tamura, and Shin ichi Yasutomi. A new multidimensional slow continued fraction algorithm and stepped surface. Experimental Mathematics, 23(4):390–410, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ford:1993:WDS

[FJ93] David Ford and Vijay Jha. On Wendt’s determinant and Sophie Germain’s theorem. Experimental Mathematics, 2(2): 113–120, ???? 1993. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048516216.

Ford:1996:DLS

[FJ96] David Ford and Kenneth W. Johnson. Determinants of Latin squares of order 8. Experimental Mathematics, 5(4):317–325, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565449.

Friedman:1994:MMT

[FJS94] Jane Friedman, James T. Joichi, and Dennis Stanton. More monotonicity theorems for partitions. Experimental Mathematics, 3(1):31–37, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1062621001.

Fukuda:2002:NEI

[FK02] Takashi Fukuda and Keiichi Komatsu. Noncyclotomic Z_{p}- extensions of imaginary quadratic fields. Experimental Mathematics, 11(4):469–475, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057864657.

Fukuda:2009:WCN

[FK09] Takashi Fukuda and Keiichi Komatsu. Weber’s class number problem in the cyclotomic Z_{2}-extension of Q. Experimental Mathematics, 18(2):213–222, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1259158432.

Fearnley:2010:CVD

[FK10] Jack Fearnley and Hershy Kisilevsky. Critical values of derivatives of twisted elliptic L-functions. Experimental Mathematics, 19(2):149–160, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1276784786.

Fearnley:2012:CVH

[FK12] Jack Fearnley and Hershy Kisilevsky. Critical values of higher derivatives of twisted elliptic L-functions. Experimental Mathematics, 21(3):213–222, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Fujiwara:2019:CKC

[FK19] Koji Fujiwara and Yuichi Kabaya. Computing Kazhdan constants by semidefinite programming. Experimental Mathematics, 28(3):301–312, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1396509.

Forsgaard:2015:CRD

[FKS15] Jens Forsgård, Vladimir P. Kostov, and Boris Z. Shapiro. Could René Descartes have known this? Experimental Mathematics, 24(4):438–448, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). See corrigendum [FKS19].

Forsgaard:2019:CCR

[FKS19] Jens Forsgård, Vladimir Petrov Kostov, and Boris Shapiro. Corrigendum: “Could René Descartes have known this?”. Experimental Mathematics, 28(2):255–256, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1417775. See [FKS15].

Flammang:2008:ATP

[Fla08] V. Flammang. On the absolute trace of polynomials having all zeros in a sector. Experimental Mathematics, 17(4):443– 450, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243429957.

Fuchs:2023:OKS

[FLST23] Elena Fuchs, Matthew Litman, Joseph H. Silverman, and Austin Tran. Orbits on K3 surfaces of Markoff type. Experimental Mathematics, 33(4):663–700, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2023.2239265.

Feuerverger:2000:BSR

[FM00] Andrey Feuerverger and Greg Martin. Biases in the Shanks–Rényi prime number race. Experimental Mathematics, 9(4):535–570, October 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045759521.

Freitag:2019:GV

[FM19] Eberhard Freitag and Riccardo Salvati Manni. The Göpel variety. Experimental Mathematics, 28(3):284–291, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1389322.

Frigerio:2004:SHM

[FMP04] Roberto Frigerio, Bruno Martelli, and Carlo Petronio. Small hyperbolic 3-manifolds with geodesic boundary. Experimental Mathematics, 13(2):171–184, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1090350932.

Fougeron:2021:PDL

[Fou21]  Charles Fougeron.  Parabolic degrees and Lyapunov

exponents for hypergeometric local systems. Experimental Mathematics, 30(4):531–546, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1580632.

Ford:1992:TRE

[FP92] David Ford and Michael Pohst. The totally real A_{5} extension of degree 6 with minimum discriminant. Experimental Mathematics, 1(3):231–235, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048622026.

Ford:1993:TRE

[FP93] David Ford and Michael Pohst. The totally real A_{6} extension of degree 6 with minimum discriminant. Experimental Mathematics, 2(3):231–232, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062620833.

Flassenberg:1999:SFF

[FP99] Ralf Flassenberg and Sachar Paulus. Sieving in function fields. Experimental Mathematics, 8(4):339–349, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047262356.

Ford:1998:EDM

[FPDH98] David Ford, Michael Pohst, Mario Daberkow, and Nasser Haddad. The S_{5} extensions of degree 6 with minimum discriminant. Experimental Mathematics, 7(2):121–124, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048515659.

Fermi:1955:SNP

[FPU55] Enrico Fermi, John Pasta, and Stanisław M. Ulam. Studies of nonlinear problems. I. Technical Report LA-1940, Los Alamos Scientific Laboratory, Los Alamos, NM,

USA, May 1955. 22 pp. URL http://www.osti.gov/
accomplishments/documents/fullText/ACC0041.pdf. Also

in Enrico Fermi: Collected Papers, volume 2, edited by Edoardo Amaldi, Herbert L. Anderson, Enrico Persico, Emilio Segré, and Albedo Wattenberg. Chicago: University of Chicago Press, 1965.

Falbel:2022:RDM

[FPUP22] E. Falbel, I. Pasquinelli, and A. Ucan-Puc. Representations of Deligne-mostow lattices into PGL(3, C). Experimental Mathematics, 33(2):336–345, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2093802.

Froberg:2022:HSG

[Frö22] Ralf Fröberg. Hilbert series of generic ideals in products of projective spaces. Experimental Mathematics, 31(4):1370– 1372, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2021.1925999.

Flajolet:1998:ESC

[FS98] Philippe Flajolet and Bruno Salvy. Euler sums and contour integral representations. Experimental Mathematics, 7 (1):15–35, ???? 1998. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047674270.

Ford:2010:CBP

[FS10] Kevin Ford and Jason Sneed. Chebyshev’s bias for products of two primes. Experimental Mathematics, 19(4):385– 398, ???? 2010. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317758099.

Fuchs:2011:SEI

[FS11] Elena Fuchs and Katherine Sanden. Some experiments with integral Apollonian circle packings. Experimental Mathematics, 20(4):380–399, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1323367153.

Farashahi:2012:GSR

[FS12] Reza Rezaeian Farashahi and Igor E. Shparlinski. On group structures realized by elliptic curves over a fixed finite field. Experimental Mathematics, 21(1):1–10, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Fullarton:2019:OPR

[FS19] Neil J. Fullarton and Richard Shadrach. Observed periodicity related to the four-strand Burau representation. Experimental Mathematics, 28(1):57–66, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1335625.

Fukumoto:2011:IFS

[Fuk11] Yoshihiro Fukumoto. w-invariants and the Fintushel–Stern invariants for plumbed homology 3-spheres. Experimental Mathematics, 20(1):1–14, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317924386.

Floyd:2002:AH

[FWW02] William Floyd, Brian Weber, and Jeffrey Weeks. The Achilles’ Heel of O(3, 1)? Experimental Mathematics, 11(1): 91–97, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057860318.

Gaal:2001:PIB

[Gaá01] István Gaál. Power integral bases in cubic relative expansions. Experimental Mathematics, 10(1):133–139, ???? 2001. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/999188428.

Gaal:2015:CSS

[Gaá15] István Gaál. Calculating “small” solutions of relative Thue equations. Experimental Mathematics, 24(2):142–149, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Galbraith:1999:RP

[Gal99] Steven D. Galbraith. Rational points on X_{0}^{+}(p). Experimental Mathematics, 8(4):311–318, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047262354.

Galetto:2014:CMO

[Gal14] Federico Galetto. Computational methods for orbit closures in a representation with finitely many orbits. Experimental Mathematics, 23(3):310–321, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ganz:2014:DES

[Gan14] Reinhard E. Ganz. The decimal expansion of π is not statistically random. Experimental Mathematics, 23(2):99–104, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ganz:2017:RRC

[Gan17] Reinhard E. Ganz. Reply to “Reproducibility in Computational Science: A Case Study: Randomness of the Digits of Pi” [Bailey et al. 17]. Experimental Mathematics, 26(3):306– 307, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://www.tandfonline.com/doi/
full/10.1080/10586458.2016.1173413. See [BBBR17].

Garunkstis:2002:SIC

[Gar02] R. Garunkštis. On some inequalities concerning π(x). Experimental Mathematics, 11(2):297–301, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1062621222.

GarciaEscudero:2017:TSP

[Gar17] Juan García Escudero. Threefolds from solutions of a partial differential equation. Experimental Mathematics, 26(2): 189–196, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1149124.

Garber:2021:CVP

[Gar21a]  Alexey Garber.  On combinatorics of Voronoi poly-

topes for perturbations of the dual root lattices. Experimental Mathematics, 33(1):86–99, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1994488.

Gardam:2021:PRS

[Gar21b] Giles Gardam. Profinite rigidity in the SnapPea census. Experimental Mathematics, 30(4):489–498, 2021. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1577766.

Garden:2023:EOP

[Gar23] Grace S. Garden. Earthquakes on the once-punctured torus. Experimental Mathematics, 33(4):781–807, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2294491.

Grosse-Brauckmann:1997:GCM

[GB97] Karsten Große-Brauckmann. Gyroids of constant mean curvature. Experimental Mathematics, 6(1):33–50, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565282.

Grosse-Brauckmann:1997:CCM

[GBP97] Karsten Große-Brauckmann and Konrad Polthier. Compact constant mean curvature surfaces with low genus. Experimental Mathematics, 6(1):13–32, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047565281.

Grunenbaum:2007:MVO

[GdlI07] F. Alberto Grünenbaum and Manuel D. de la Iglesia. Matrixvalued orthogonal polynomials related to SU(N +1), their algebras of differential operators, and the corresponding curves. Experimental Mathematics, 16(2):189–208, ???? 2007. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1204905875.

Gebhardt:2000:TSP

[Geb00] Volker Gebhardt. Two short presentations for Lyons’ sporadic simple group. Experimental Mathematics, 9(3):333–338, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604668.

Gerver:2003:NSD

[Ger03] Joseph L. Gerver. Noncollision singularities: Do four bodies suffice? Experimental Mathematics, 12(2):187–198, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1067634730.

Gravner:1997:NPD

[GG97] Janko Gravner and David Griffeath. Nucleation parameters for discrete threshold growth on Z^{2}. Experimental Mathematics, 6(3):207–220, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047920421.

Gaudry:2003:CPM

[GG03] Pierrick Gaudry and Nicolas Gürel. Counting points in medium characteristic using Kedlaya’s algorithm. Experimental Mathematics, 12(4):395–402, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1087568016.

Gravner:2006:MSC

[GG06] Janko Gravner and David Griffeath. Modeling snow crystal growth I: Rigorous results for Packard’s digital snowflakes. Experimental Mathematics, 15(4):421–444, ???? 2006. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1175789778.

Gine:2011:PSC

[GGM11] E. Giné, C. S. Güntürk, and W. R. Madych. On the periodized square of L^{2} cardinal splines. Experimental Mathematics, 20(2):177–188, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317924409.

Grunewald:1996:NSE

[GH96] Fritz Grunewald and Wolfgang Huntebrinker. A numerical study of eigenvalues of the hyperbolic Laplacian for polyhedra with one cusp. Experimental Mathematics, 5(1):57–80, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047591148.

Goodman:2008:CCH

[GHH08] Oliver Goodman, Damian Heard, and Craig Hodgson. Commensurators of cusped hyperbolic manifolds. Experimental Mathematics, 17(3):283–306, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227121383.

Gerber:2003:REP

[GHK03] Marlies Gerber, Boris Hasselblatt, and Daniel Keesing. The Riccati equation: Pinching of forcing and solutions. Experimental Mathematics, 12(2):129–134, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1067634727.

Gryak:2020:SCD

[GHK20] Jonathan Gryak, Robert M. Haralick, and Delaram Kahrobaei. Solving the conjugacy decision problem via machine learning. Experimental Mathematics, 29(1):66–78, 2020. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1434704.

Gorgen:2019:CMC

[GHL19] Lukas Görgen, Gerhard Hiss, and Klaus Lux. The computation of the 3-modular characters of the Fischer group Fi_{23}. Experimental Mathematics, 28(2):185–193, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1382402.

Geiger:2023:SDM

[GHSV23] Alheydis Geiger, Sachi Hashimoto, Bernd Sturmfels, and Raluca Vlad. Self-dual matroids from canonical curves. Experimental Mathematics, 33(4):701–722, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2023.2239282.

Gunnells:2013:MFE

[GHY13] Paul E. Gunnells, Farshid Hajir, and Dan Yasaki. Modular forms and elliptic curves over the field of fifth roots of unity. Experimental Mathematics, 22(2):203–216, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Garfinkle:2023:NSA

[GIKW23] David Garfinkle, James Isenberg, Dan Knopf, and Haotian Wu. A numerical stability analysis of mean curvature flow of noncompact hypersurfaces with type-II curvature blowup: II. Experimental Mathematics, 33(4):559–568, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2201958.

Girondo:2003:MQR

[Gir03] Ernesto Girondo. Multiply quasiplatonic Riemann surfaces. Experimental Mathematics, 12(4):463–476, ???? 2003. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1087568022.

Gaal:2003:RPD

[GJL03] István Gaál, István Járási, and Florian Luca. A remark on prim divisors of lengths of sides of heron triangles. Experimental Mathematics, 12(3):303–310, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1087329233.

Gonzalez-Jimenez:2023:ADT

[GJN23] Enrique González-Jiménez and Filip Najman. An algorithm for determining torsion growth of elliptic curves. Experimental Mathematics, 32(1):70–81, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2020.1771638.

Gonzalez:2012:STD

[GJU12] Josep González and Jorge Jiménez-Urroz. The Sato–Tate distribution and the values of Fourier coefficients of modular newforms. Experimental Mathematics, 21(1):84–102, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Garoufalidis:2012:NPP

[GK12] Stavros Garoufalidis and Christoph Koutschan. The noncommutative A-polynomial of (−2, 3, n) pretzel knots. Experimental Mathematics, 21(3):241–251, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Guilloux:2021:DSD

[GK21] Antonin Guilloux and Inkang Kim. Deformation spaces of discrete groups of SU(2, 1) in quaternionic hyperbolic plane: a case study. Experimental Mathematics, 30(4):453–458, 2021.

CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1559776.

Garcia:2019:PRB

[GKL19] Stephan Ramon Garcia, Elvis Kahoro, and Florian Luca. Primitive root bias for twin primes. Experimental Mathematics, 28(2):151–160, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2017.1360809.

Gneiting:2001:EAK

[GKR01] Tilmann Gneiting, Kjell Konis, and Donald Richards. Experimental approaches to Kuttner’s problem. Experimental Mathematics, 10(1):117–124, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/999188426.

Garonzi:2022:ICN

[GKS22] Martino Garonzi, Luise-Charlotte Kappe, and Eric Swartz. On integers that are covering numbers of groups. Experimental Mathematics, 31(2):425–443, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1636425.

Grasegger:2020:LBN

[GKT20] Georg Grasegger, Christoph Koutschan, and Elias Tsigaridas. Lower bounds on the number of realizations of rigid graphs. Experimental Mathematics, 29(2):125–136, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1437851.

Gorsky:2015:SHT

[GL15] Eugene Gorsky and Lukas Lewark. On stable ∫↕_{3}-homology of torus knots. Experimental Mathematics, 24(2):162–174, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Greicius:2022:EAS

[GL22] Quinn Greicius and Aaron Landesman. An explicit Abelian surface with maximal Galois action. Experimental Mathematics, 31(2):689–693, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1691087.

Granville:2023:PSI

[GL23] Andrew Granville and Allysa Lumley. Primes in short intervals: Heuristics and calculations. Experimental Mathematics, 32(2):378–404, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1927256.

Golmakani:2016:UEO

[GLP16] Ali Golmakani, Stefano Luzzatto, and Pawel Pilarczyk. Uniform expansivity outside a critical neighborhood in the quadratic family. Experimental Mathematics, 25(2):116–124, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Geck:1999:SPU

[GM99] Meinolf Geck and Gunter Malle. On special pieces in the unipotent variety. Experimental Mathematics, 8(3):281–290, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047262408.

Geller:2005:IL

[GM05] William Geller and Michał Misiurewicz. Irrational life. Experimental Mathematics, 14(3):271–275, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1128371752.

Grunberg:2008:SEG

[GM08] Daniel B. Grünberg and Pieter Moree. Sequences of enumerative geometry: Congruences and asymptotics, with an appendix by Don Zagier. Experimental Mathematics, 17(4): 409–426, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243429954.

Guitart:2013:CAD

[GM13] Xavier Guitart and Marc Masdeu. Computation of ATR Darmon points on nongeometrically modular elliptic curves. Experimental Mathematics, 22(1):85–98, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Guitart:2018:PMT

[GM18] Xavier Guitart and Marc Masdeu. Periods of modular GL_{2}-type Abelian varieties and p-adic integration. Experimental Mathematics, 27(3):344–361, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1284624.

Gogolev:2019:NSG

[GMK19] Andrey Gogolev, Itai Maimon, and Aleksey N. Kolmogorov. A numerical study of Gibbs u-measures for partially hyperbolic diffeomorphisms on T^{3}. Experimental Mathematics, 28 (3):271–283, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1389321.

Goodwin:2020:MBS

[GMR20] Simon M. Goodwin, Peter Mosch, and Gerhard Röhrle. The modality of a Borel subgroup in a simple algebraic group of type E_{8}. Experimental Mathematics, 29(3):326–327, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1468289.

Garcia-Morchon:2014:IAP

[GMRST14] Oscar Garcia-Morchon, Ronald Rietman, Igor E. Shparlinski, and Ludo Tolhuizen. Interpolation and approximation of polynomials in finite fields over a short interval from noisy values. Experimental Mathematics, 23(3):241–260, 2014. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Gutierrez:1996:CCA

[GMSB96] Carlos Gutierrez, Francesco Mercuri, and Federico Sánchez-Bringas. On a conjecture of Carathéodory: analyticity versus smoothness. Experimental Mathematics, 5(1):33–37, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047591145.

Gunnells:2021:CCS

[GMY21] Paul E. Gunnells, Mark McConnell, and Dan Yasaki. On the cohomology of congruence subgroups of GL_{3} over the Eisenstein integers. Experimental Mathematics, 30(4):499–512, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1577767.

Gelfreich:2009:WHZ

[GN09] V. Gelfreich and V. Naudot. Width of the homoclinic zone in the parameter space for quadratic maps. Experimental Mathematics, 18(4):409–427, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1259158506.

Gekhtman:2022:ASC

[GN22] Michael Gekhtman and Tomoki Nakanishi. Asymptotic sign coherence conjecture. Experimental Mathematics, 31(2):497– 505, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2019.1650401.

Goetgheluck:1993:FZS

[Goe93] Pierre Goetgheluck. Fresnel zones on the screen. Experimental Mathematics, 2(4):301–309, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048516039.

Gaal:2002:PIB

[GOP02] István Gaál, Péter Olajos, and Michael Pohst. Power integral bases in orders of composite fields. Experimental Mathematics, 11(1):87–90, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057860317.

Gorsky:2013:SKH

[GOR13] Eugene Gorsky, Alexei Oblomkov, and Jacob Rasmussen. On stable Khovanov homology of torus knots. Experimental Mathematics, 22(3):265–281, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Gorner:2015:RTL

[Gör15] Matthias Görner. Regular tessellation link complements. Experimental Mathematics, 24(2):225–246, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Gorodetsky:2021:NRA

[Gor21] Ofir Gorodetsky. New representations for all sporadic Apéry-like sequences, with applications to congruences. Experimental Mathematics, 32(4):641–656, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1982080.

Gouvea:1997:NOP

[Gou97] Fernando Q. Gouvêa. Non-ordinary primes: a story. Experimental Mathematics, 6(3):195–205, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047920420.

Gray:1992:CGN

[GP92] Alfred Gray and Mark A. Pinsky. Computer graphics and a new Gibbs phenomenon for Fourier–Bessel series. Experimental Mathematics, 1(4):313–316, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048610119.

Gilbert:1994:TNN

[GP94] Peter Gilbert and Eric Postpischil. There are no new homometric Golomb ruler pairs with 12 marks or less. Experimental Mathematics, 3(2):147–152, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062620908.

Gaal:2006:DEG

[GP06] István Gaál and Michael Pohst. Diophantine equations over global functions fields II: R-integral solutions of Thue equations. Experimental Mathematics, 15(1):1–6, ???? 2006. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1150476898.

Garcia-Puente:2012:SCR

[GPHH^{+}12] Luis D. García-Puente, Nickolas Hein, Christopher Hillar, Abraham Martín del Campo, James Ruffo, Frank Sottile, and Zach Teitler. The secant conjecture in the real Schubert calculus. Experimental Mathematics, 21(3):252–265, 2012. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Gaal:1994:RIF

[GPP94] István Gaál, Attila Pethő, and Michael Pohst. On the resolution of index form equations in dihedral quartic number fields. Experimental Mathematics, 3(3):245–254, ???? 1994. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048515874.

Gaal:2021:CIP

[GPP21] István Gaál, Maximilian C. Pohst, and Michael E. Pohst. On computing integral points of a Mordell curve — the method of Wildanger revisited. Experimental Mathematics, 30(1): 127–134, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1502700.

Gomez-Perez:2018:API

[GPS18] Domingo Gómez-Pérez and Igor E. Shparlinski. Arithmetic properties of integers in chains and reflections of g-ary expansions. Experimental Mathematics, 27(2):184–192, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1239146.

Geiges:2000:PAS

[GR00] Hansjörg Geiges and Diego Rattaggi. Periodic automorphisms of surfaces: invariant circles and maximal orders. Experimental Mathematics, 9(1):75–84, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1046889592.

Gaal:2019:IBM

[GR19] István Gaál and László Remete. Integral bases and monogenity of composite fields. Experimental Mathematics, 28 (2):209–222, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1382404.

Girondo:2022:MED

[GR22] Ernesto Girondo and Cristian Reyes. Multiple extremal disc-packings in compact hyperbolic surfaces. Experimental Mathematics, 33(2):276–300, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2075491.

Granboulan:1996:CER

[Gra96] Louis Granboulan. Construction d’une extension régulière de Q(T) de groupe de Galois M_{24}. (French) [Construction of a regular extension of Q(T) for the Galois group M_{24}]. Experimental Mathematics, 5(1):3–14, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047591143.

Graca:2002:CNF

[Gra02] Mário M. Graça. On the computation of nonhyperbolic fixed points. Experimental Mathematics, 11(4):477–485, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1057864658.

Greicius:2010:ECS

[Gre10] Aaron Greicius. Elliptic curves with surjective adelic Galois representations. Experimental Mathematics, 19(4):495–507, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758108.

Gethner:1997:PGM

[GS97] Ellen Gethner and H. M. Stark. Periodic Gaussian moats. Experimental Mathematics, 6(4):289–292, ???? 1997. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1047047189.

GrafvonBothmer:2005:QDI

[GS05] H.-C. Graf von Bothmer and F.-O. Schreyer. A quick and dirty irreducibility test for multivariate polynomials over F_{q}. Experimental Mathematics, 14(4):415–422, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1136926972.

Garoufalidis:2016:ECL

[GSS16] Stavros Garoufalidis, Eric Sabo, and Shane Scott. Exact computation of the n-loop invariants of knots. Experimental Mathematics, 25(2):125–129, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Garcia-Selfa:2006:SAP

[GST06] Irene García-Selfa and José M. Tornero. On simultaneous arithmetic progressions on elliptic curves. Experimental Mathematics, 15(4):471–478, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1175789781.

Guckenheimer:1995:PPP

[Guc95] John Guckenheimer. Phase portraits of planar vector fields: computer proofs. Experimental Mathematics, 4(2):153–165, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047931624.

Guillera:2003:ANK

[Gui03] Jesús Guillera. About a new kind of Ramanujan-type series. Experimental Mathematics, 12(4):507–510, ???? 2003. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1087568026.

Guillera:2006:CCS

[Gui06a] Jesús Guillera. A class of conjectured series representations for 1/π. Experimental Mathematics, 15(4):409–414, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789776.

Guillera:2006:NMO

[Gui06b] Jesús Guillera. A new method to obtain series for 1/π and 1/π^{2}. Experimental Mathematics, 15(1):83–89, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1150476906.

Guillera:2012:MSR

[Gui12] Jesús Guillera. Mosaic supercongruences of Ramanujan type. Experimental Mathematics, 21(1):65–68, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Guilloux:2018:VRB

[Gui18] Antonin Guilloux. Volume of representations and birationality of peripheral holonomy. Experimental Mathematics, 27 (4):472–477, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1320240.

Gunnells:2000:CHE

[Gun00] Paul E. Gunnells. Computing Hecke eigenvalues below the cohomological dimension. Experimental Mathematics, 9(3): 351–367, ???? 2000. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1045604670.

Gronau:2018:BEA

[GW18] Quentin F. Gronau and Eric-Jan Wagenmakers. Bayesian evidence accumulation in experimental mathematics: A case study of four irrational numbers. Experimental Mathematics, 27(3):277–286, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1256006.

Gopaladesikan:2016:APF

[GWW16] Mohan Gopaladesikan, Stephan Wagner, and Mark Daniel Ward. On the asymptotic probability of forbidden motifs on the fringe of recursive trees. Experimental Mathematics, 25 (3):237–245, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Goksel:2015:RCF

[GXB15] Vefa Goksel, Shixiang Xia, and Nigel Boston. A refined conjecture for factorizations of iterates of quadratic polynomials over finite fields. Experimental Mathematics, 24(3):304–311, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Gaidashev:2007:CRS

[GY07] Denis Gaidashev and Michael Yampolsky. Cylinder renormalization of Siegel disks. Experimental Mathematics, 16(2): 215–226, ???? 2007. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1204905877.

Goshima:2010:DSH

[GY10] Masato Goshima and Masakazu Yamagishi. On the dimension of the space of harmonic functions on a discrete torus.

Experimental Mathematics, 19(4):421–429, ???? 2010. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1317758102.

Halberstadt:1998:CM

[Hal98] Emmanuel Halberstadt. Sur la courbe modulaire Xndép(11). Experimental Mathematics, 7(2):163–174, ???? 1998. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048515664.

Hall:2014:PES

[Hal14] Stuart J. Hall. Perelman’s entropy for some families of canonical metrics. Experimental Mathematics, 23(3):277–284, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Han:2009:SCO

[Han09] Guo-Niu Han. Some conjectures and open problems on partition hook lengths. Experimental Mathematics, 18(1):97–106, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1243430533.

Hardcastle:2002:TDG

[Har02] D. M. Hardcastle. The three-dimensional Gauss algorithm is strongly convergent almost everywhere. Experimental Mathematics, 11(1):131–141, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057860321.

Hare:2011:GGW

[Har11] Kevin G. Hare. Generalized Gorshkov–Wirsing polynomials and the integer Chebyshev problem. Experimental Mathematics, 20(2):189–200, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317924410.

Hare:2015:BED

[Har15] Kevin G. Hare. Base-d expansions with digits 0 to q − 1. Experimental Mathematics, 24(3):295–303, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Haythorpe:2018:MNH

[Hay18] M. Haythorpe. On the minimum number of Hamiltonian cycles in regular graphs. Experimental Mathematics, 27(4): 426–430, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1306813.

Hall:2002:FRH

[HdC02] Toby Hall and André de Carvalho. The forcing relation for horseshoe braid types. Experimental Mathematics, 11(2):271– 288, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621220.

Hacon:2006:FMS

[HdJF06] D. Hacon, C. Mendes de Jesus, and M. C. Romero Fuster. Fold maps from the sphere to the plane. Experimental Mathematics, 15(4):491–497, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1175789783.

He:2002:MMC

[He02] Zheng-Xu He. On the minimizers of the Möbius cross energy of links. Experimental Mathematics, 11(2):244–248, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621218.

He:2021:BUT

[He21] Zhuang He. Blow-ups of toric surfaces and the Mori dream space property. Experimental Mathematics, 30(3):386–404, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1547232.

Henk:2018:NLP

[Hen18] Martin Henk. A note on lattice packings via lattice refinements. Experimental Mathematics, 27(1):1–9, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1208595.

Hennecart:2023:ABK

[Hen23] Lucien Hennecart. Asymptotic behavior of Kac polynomials. Experimental Mathematics, 32(2):294–312, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926012.

Hew:2021:CDR

[Hew21] Patrick Chisan Hew. Collatz on the dyadic rationals in [0.5, 1) with fractals: How bit strings change their length under 3x + 1. Experimental Mathematics, 30(4):481–488, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1577765.

Hare:2022:AIS

[HH22a] Kevin G. Hare and Philip W. Hodges. Applications of integer semi-infinite programing to the integer Chebyshev problem. Experimental Mathematics, 31(2):694–700, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1691089.

Hauenstein:2022:PSA

[HH22b] Jonathan D. Hauenstein and Martin Helmer. Probabilistic saturations and Alt’s problem. Experimental Mathematics, 31(3):975–987, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1740835.

Hein:2015:MSC

[HHdC^{+}15] Nickolas Hein, Christopher J. Hillar, Abraham Martín del Campo, Frank Sottile, and Zach Teitler. The monotone secant conjecture in the real Schubert calculus. Experimental Mathematics, 24(3):261–269, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Heard:2010:HGS

[HHMP10] Damian Heard, Craig Hodgson, Bruno Martelli, and Carlo Petronio. Hyperbolic graphs of small complexity. Experimental Mathematics, 19(2):211–236, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1276784791.

Haagerup:2021:CET

[HHRS21] Søren Haagerup, Uffe Haagerup, and Maria Ramirez-Solano. Computational explorations of the Thompson group T for the amenability problem of F. Experimental Mathematics, 30(1):105–126, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1502699.

Herbig:2021:LCH

[HHS21] Hans-Christian Herbig, Daniel Herden, and Christopher Seaton. The Laurent coefficients of the Hilbert series of a Gorenstein algebra. Experimental Mathematics, 30(1):56–75, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1492473.

Hikami:2003:VCA

[Hik03] Kazuhiro Hikami. Volume conjecture and asymptotic expansion of q-series. Experimental Mathematics, 12(3):319–338, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1087329235.

Hoffman:2016:VCH

[HIK^{+}16] Neil Hoffman, Kazuhiro Ichihara, Masahide Kashiwagi, Hidetoshi Masai, Shin’ichi Oishi, and Akitoshi Takayasu. Verified computations for hyperbolic 3-manifolds. Experimental Mathematics, 25(1):66–78, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hamer:2022:PKL

[HIK22] Jesse Hamer, Tetsuya Ito, and Keiko Kawamuro. Positivities of knots and links and the defect of bennequin inequality. Experimental Mathematics, 31(1):199–225, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1596848.

Hauenstein:2013:ELB

[HIL13] Jonathan D. Hauenstein, Christian Ikenmeyer, and J. M. Landsberg. Equations for lower bounds on border rank. Experimental Mathematics, 22(4):372–383, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hiraoka:2000:NCT

[Hir00] Yoshio Hiraoka. Numerical calculation of twisted adjoint L-values attached to modular forms. Experimental Mathematics, 9(1):67–73, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1046889591.

Hodgson:2016:NGV

[HIS16] Craig D. Hodgson, Ahmad Issa, and Henry Segerman. Nongeometric veering triangulations. Experimental Mathematics, 25(1):17–45, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Holmin:2019:MCG

[HJK^{+}19] S. Holmin, N. Jones, P. Kurlberg, C. McLeman, and K. Petersen. Missing class groups and class number statistics for imaginary quadratic fields. Experimental Mathematics, 28 (2):233–254, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1383952.

Harlev:2022:DSS

[HJL22] Amit Harlev, Charles R. Johnson, and Derek Lim. The doubly stochastic single eigenvalue problem: a computational approach. Experimental Mathematics, 31(3):936–945, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1727799.

Hamenstadt:2002:SFT

[HK02a] Ursula Hamenstädt and Roman Koch. Systoles of a family of triangle surfaces. Experimental Mathematics, 11(2):249–270, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621219.

Hardcastle:2002:DGT

[HK02b] D. M. Hardcastle and K. Khanin. The d-dimensional Gauss transformation: Strong convergence and Lyapunov exponents. Experimental Mathematics, 11(1):119–129, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1057860320.

Halberstadt:2003:CM

[HK03] Emmanuel Halberstadt and Alain Kraus. Sur la courbe modulaire X_{E}(7). Experimental Mathematics, 12(1):27–40, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1064858782.

Hoskins:2021:DZU

[HK21] Jeremy Hoskins and Zakhar Kabluchko. Dynamics of zeroes under repeated differentiation. Experimental Mathematics, 32(4):573–599, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1980752.

Hajdu:2013:OPL

[HKPP13] L. Hajdu, T. Kovács, A. Pethő, and M. Pohst. An optimization problem for lattices. Experimental Mathematics, 22(4): 443–455, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hsu:1992:MSM

[HKS92] Lucas Hsu, Rob Kusner, and John Sullivan. Minimizing the squared mean curvature integral for surfaces in space forms. Experimental Mathematics, 1(3):191–207, ???? 1992. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048622023.

Hauenstein:2015:EZH

[HLLM15] J. D. Hauenstein, A. Lerario, E. Lundberg, and D. Mehta. Experiments on the zeros of harmonic polynomials using certified counting. Experimental Mathematics, 24(2):133–141, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Holmes:2003:PRM

[HLM03] Petra E. Holmes, Stephen A. Linton, and Scott H. Murray. Product replacement in the monster. Experimental Mathematics, 12(1):123–126, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1064858789.

Hamalainen:2023:INF

[HLM23] Rimma Hämäläinen, Jason Lo, and Edward Morales. Intersection numbers on fibrations and Catalan numbers. Experimental Mathematics, 33(4):768–775, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2023.2293292.

Hayat-Legrand:2001:CCD

[HLMZ01] Claude Hayat-Legrand, Sergei Matveev, and Heiner Zeischang. Computer calculation of the degree of maps into the Poincaré homology sphere. Experimental Mathematics, 10(4): 497–508, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069855249.

Hougardy:2010:SRI

[HLZ10] Stefan Hougardy, Frank H. Lutz, and Mariano Zelke. Surface realization with the intersection segment functional. Experimental Mathematics, 19(1):79–92, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1268404804.

Hart:2002:ASF

[HM02] Enric Hart and Daniel Maisner. Abelian surfaces over finite fields as Jacobians. Experimental Mathematics, 11(3):321– 337, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057777425.

Hering:2012:GMH

[HM12] Milena Hering and Diane Maclagan. The T -graph of a multigraded Hilbert scheme. Experimental Mathematics, 21(3): 280–297, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hohn:2021:MVS

[HM21] Gerald Höhn and Geoffrey Mason. Most vertex superalgebras associated to an odd unimodular lattice of rank 24 have an N = 4 superconformal structure. Experimental Mathematics, 32(3):502–504, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1980458.

Havas:1998:EGH

[HMM98] George Havas, Bohdan S. Majewski, and Keith R. Matthews. Extended GCD and Hermite normal form algorithms via lattice basis reduction. Experimental Mathematics, 7(2):125– 136, ???? 1998. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/

euclid.em/1048515660. See addenda and errata [HMM99].

Havas:1999:AEE

[HMM99] George Havas, Bohdan S. Majewski, and Keith R. Matthews. Addenda and errata: “Extended GCD and Hermite normal form algorithms via lattice basis reduction. Experimental Mathematics, 8(2):205, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047477061. See

[HMM98].

Haralick:2005:HWM

[HMM05] R. M. Haralick, A. D. Miasnikov, and A. G. Myasnikov. Heuristics for the Whitehead minimization problem. Experimental Mathematics, 14(1):7–14, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1120145566.

Hanrot:2010:CTK

[HMT10] Guillaume Hanrot, Bruno Martin, and Gérald Tenenbaum. Constantes de Turán–Kubilius friables: une étude numérique. (French) [Reducible Turán–Kubilius constants: a numerical study]. Experimental Mathematics, 19(3):345–361, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758096.

Heim:2020:IPD

[HNR20] Bernhard Heim, Markus Neuhauser, and Florian Rupp. Imaginary powers of the Dedekind eta function. Experimental Mathematics, 29(3):317–325, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1468288.

Hernandez-Ortiz:2024:RCD

[HOKMS24] Rangel Hernández-Ortiz, Kolja Knauer, Luis Pedro Montejano, and Manfred Scheucher. Roudneff’s Conjecture in dimension 4. Experimental Mathematics, 34(1):67–71, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2024.2334379.

Holt:1995:APF

[Hol95] Derek F. Holt. An alternative proof that the Fibonacci group F (2, 9) is infinite. Experimental Mathematics, 4(2):97–100, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047931620.

Holmes:2004:MFS

[Hol04] P. E. Holmes. On minimal factorisations of sporadic groups. Experimental Mathematics, 13(4):435–440, ???? 2004. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1109106435.

Hopkins:2010:HWH

[Hop10] Kimberly Hopkins. Higher-weight Heegner points. Experimental Mathematics, 19(3):257–266, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317758090.

Hopkins:2022:MDN

[Hop22] Sam Hopkins. Minuscule doppelgängers, the coincidental down-degree expectations property, and rowmotion. Experimental Mathematics, 31(3):946–974, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2020.1731881.

Horn:2014:CFH

[Hor14] Peter D. Horn. On computing the first higher-order Alexander modules of knots. Experimental Mathematics, 23(2):153– 169, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Heilman:2011:OPR

[HOS11] Steven M. Heilman, Philip Owrutsky, and Robert S. Strichartz. Orthogonal polynomials with respect to selfsimilar measures. Experimental Mathematics, 20(3):238–259, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317924418.

Hayata:2012:ZCS

[HOY12] Takahiro Hayata, Takayuki Oda, and Tomoki Yatougo. Zero cells of the Siegel–Gottschling fundamental domain of degree 2. Experimental Mathematics, 21(3):266–279, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hentzel:2006:NFA

[HP06] I. R. Hentzel and L. A. Peresi. The nucleus of the free alternative algebra. Experimental Mathematics, 15(4):445–454, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789779.

Hentzel:2008:NED

[HP08] I. R. Hentzel and L. A. Peresi. Nuclear elements of degree 6 in the free alternative algebra. Experimental Mathematics, 17(2):245–255, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227118975.

Helfgott:2013:NVT

[HP13] H. A. Helfgott and David J. Platt. Numerical verification of the ternary Goldbach conjecture up to 8.875 · 10^{30}. Experimental Mathematics, 22(4):406–409, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hasselblatt:2014:IFS

[HP14] Boris Hasselblatt and Donald Plante Jr. On the interior of “fat” Sierpiński triangles. Experimental Mathematics, 23(3): 285–309, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Habsieger:2018:NNU

[HP18] Laurent Habsieger and Alain Plagne. A numerical note on upper bounds for B_{2}[g] sets. Experimental Mathematics, 27 (2):208–214, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1245640.

Heard:2008:OOM

[HPP08] Damian Heard, Ekaterina Pervova, and Carlo Petronio. The 191 orientable octahedral manifolds. Experimental Mathematics, 17(4):473–486, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1243429960.

Hubert:2022:TTB

[HPR22] Pascal Hubert and Olga Paris-Romaskevich. Triangle tiling billiards and the exceptional family of their escaping trajectories: Circumcenters and Rauzy gasket. Experimental Mathematics, 31(1):58–87, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1583615.

Hejhal:1992:TMW

[HR92a] Dennis A. Hejhal and Barry N. Rackner. On the topography of Maass waveforms for PSL(2, Z). Experimental Mathematics, 1(4):275–305, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048610117.

Holt:1992:INP

[HR92b] Derek F. Holt and Sarah Rees. An implementation of the Neumann–Praeger algorithm for the recognition of special linear groups. Experimental Mathematics, 1(3):237–242, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048622027.

Holt:1996:FQF

[HR96] Derek F. Holt and Sarah Rees. Free quotients of finitely presented groups. Experimental Mathematics, 5(1):49–56, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047591147.

Havas:2003:ICP

[HR03] George Havas and Edmund F. Robertson. Irreducible cyclic presentations of the trivial group. Experimental Mathematics, 12(4):487–490, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1087568024.

Hausmann:2004:SCE

[HR04] Jean-Claude Hausmann and Eugenio Rodriguez. The space of clouds in Euclidean space. Experimental Mathematics, 13 (1):31–48, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1086894088.

Hare:2014:KCA

[HR14a] Kathryn E. Hare and L. Thomas Ramsey. Kronecker constants of arithmetic progressions. Experimental Mathematics, 23(4):414–422, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hoge:2014:ZMA

[HR14b] Torsten Hoge and Gerhard Röhrle. Ziegler’s multireflection arrangements are free. Experimental Mathematics, 23(4):448– 451, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Henson:2017:OAR

[HRSS17] Joe Henson, David Rideout, Rafael D. Sorkin, and Sumati Surya. Onset of the asymptotic regime for (uniformly random) finite orders. Experimental Mathematics, 26(3):253– 266, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2016.1158134.

Huang:2001:STF

[HS01] Nina N. Huang and Robert S. Strichartz. Sampling theory for functions with fractal spectrum. Experimental Mathematics, 10(4):619–640, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069855261.

Hermiller:2005:CFP

[HS05] Susan Hermiller and Irena Swanson. Computations with Frobenius powers. Experimental Mathematics, 14(2):161–173, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1128100129.

Ham:2007:MDP

[HS07] Ji-Young Ham and Won Taek Song. The minimum dilatation of pseudo-Ansonov 5-braids. Experimental Mathematics, 16(2):167–180, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204905873.

Harvey:2013:SDR

[HS13] David Harvey and Igor E. Shparlinski. Statistics of different reduction types of Fermat curves. Experimental Mathematics, 22(3):243–249, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Herbig:2014:HSL

[HS14] Hans-Christian Herbig and Christopher Seaton. The Hilbert series of a linear symplectic circle quotient. Experimental Mathematics, 23(1):46–65, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Heller:2015:DSC

[HS15] Sebastian Heller and Nicholas Schmitt. Deformations of symmetric CMC surfaces in the 3-sphere. Experimental Mathematics, 24(1):65–75, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Hare:2018:LBD

[HS18] Kevin G. Hare and Nikita Sidorov. A lower bound for the dimension of Bernoulli convolutions. Experimental Mathematics, 27(4):414–418, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1301841.

Harriss:2022:ANS

[HST22] Edmund Harriss, Katherine E. Stange, and Steve Trettel. Algebraic number starscapes. Experimental Mathematics, 31 (4):1098–1149, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2102094.

Hacon:1994:RSA

[HSV94] Derek Hacon, Nicolau C. Saldanha, and J. J. P. Veerman. Remarks on self-affine tilings. Experimental Mathematics, 3(4): 317–327, ???? 1994. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048515813.

Huber:2000:CGF

[HT00] Birkett Huber and Rekha R. Thomas. Computing Gröbner fans of toric ideals. Experimental Mathematics, 9(3):321–331, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604667.

Haraway:2019:TMS

[HT19] Robert C. Haraway III and Stephan Tillmann. Tessellating the moduli space of strictly convex projective structures on the once-punctured torus. Experimental Mathematics, 28 (3):369–384, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1409671.

Hulpke:1995:BSG

[Hul95] Alexander Hulpke. Block systems of a Galois group. Experimental Mathematics, 4(1):1–9, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062621138.

Hulpke:2001:RSF

[Hul01] Alexander Hulpke. Representing subgroups of finitely presented groups by quotient subgroups. Experimental Mathematics, 10(3):369–382, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069786345.

Humphreys:2002:AWF

[Hum02] J. E. Humphreys. Analogues of Weyl’s formula for reduced enveloping algebras. Experimental Mathematics, 11(4):567– 573, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057864665.

Hunt:2000:NMV

[Hun00] Bruce Hunt. Nice modular varieties. Experimental Mathematics, 9(4):613–622, October 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045759526.

Harris:2019:DHA

[HV19] Michael Harris and Akshay Venkatesh. Derived Hecke algebra for weight one forms. Experimental Mathematics, 28 (3):342–361, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1409144.

Hiary:2022:CIM

[HV22] Ghaith Hiary and Joseph Vandehey. Calculations of the invariant measure for Hurwitz continued fractions. Experimental Mathematics, 31(1):324–336, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1627255.

Hodgson:1994:SIL

[HW94] Craig D. Hodgson and Jeffrey R. Weeks. Symmetries, isometries and length spectra of closed hyperbolic three-manifolds. Experimental Mathematics, 3(4):261–274, ???? 1994. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048515809.

Hoffman:2002:DSP

[HW02] David Hoffman and Fusheng Wei. Deforming the singly periodic genus-one helicoid. Experimental Mathematics, 11(2): 207–218, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621216.

Harkins:2005:ESU

[HWW05] Ryan Harkins, Eric Weber, and Andrew Westmeyer. Encryption schemes using finite frames and Hadamard arrays. Experimental Mathematics, 14(4):423–433, ???? 2005. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1136926973.

Hoffmann:2022:DEI

[HY22] Tim Hoffmann and Zi Ye. A discrete extrinsic and intrinsic Dirac operator. Experimental Mathematics, 31(3):920–935, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1727798.

Hyeon:2017:OLM

[Hye17] Donghoon Hyeon. An outline of the log minimal model program for the moduli space of curves. Experimental Mathematics, 26(1):114–124, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ibrahimoglu:2016:SBL

[IC16] B. Ali Ibrahimoglu and Annie Cuyt. Sharp bounds for Lebesgue constants of barycentric rational interpolation at equidistant points. Experimental Mathematics, 25(3):347– 354, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Itoh:2004:TTA

[iIS04] Jin ichi Itoh and Robert Sinclair. Thaw: a tool for approximating cut loci on a triangulation of a surface. Experimental Mathematics, 13(3):309–325, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1103749839.

Ingalls:2023:QGA

[IJK^{+}23] Colin Ingalls, Bruce W. Jordan, Allan Keeton, Adam Logan, and Yevgeny Zaytman. Quotient graphs and amalgam presentations for unitary groups over cyclotomic rings. Experimental Mathematics, 32(1):191–217, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926007.

Iwasaki:2002:PIH

[IKM02] Katsunori Iwasaki, Atsufumi Kenma, and Keiji Matsumoto. Polynomial invariants and harmonic functions related to exceptional regular polytopes. Experimental Mathematics, 11(2):313–319, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1062621224.

Ichim:2016:LLS

[IKMF16] Bogdan Ichim, Lukas Katthän, and Julio José Moyano-Fernández. LCM lattices and Stanley depth: A first computational approach. Experimental Mathematics, 25(1):46– 53, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ivanyos:2000:TEC

[IL00] Gábor Ivanyos and Klaus Lux. Treating the exceptional cases of the MeatAxe. Experimental Mathematics, 9(3):373–381, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604672.

Inoue:2013:SMP

[Ino13] Ayumu Inoue. A symmetric motion picture of the twist-spun trefoil. Experimental Mathematics, 22(1):15–25, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Inou:2019:SST

[Ino19] Hiroyuki Inou. Self-similarity for the tricorn. Experimental Mathematics, 28(4):440–455, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1420502.

Inoue:2022:ELS

[Ino22] Taiyo Inoue. Exploring the list of smallest right-angled hyperbolic polyhedra. Experimental Mathematics, 31(1):165–183, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1593897.

Ibukiyama:2019:CTS

[IT19] Tomoyoshi Ibukiyama and Sho Takemori. Construction of theta series of any vector-valued weight and applications to lifts and congruences. Experimental Mathematics, 28 (1):95–114, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1353454.

Ichim:2014:ACM

[IZ14] Bogdan Ichim and Andrei Zarojanu. An algorithm for computing the multigraded Hilbert depth of a module. Experimental Mathematics, 23(3):322–331, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Jain:2009:MRR

[Jai09] Sonal Jain. Minimal regulators for rank-2 subgroups of rational and K3 elliptic surfaces. Experimental Mathematics, 18(4):429–447, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1259158507.

Jenkinson:2000:FLB

[Jen00] Oliver Jenkinson. Frequency locking on the boundary of the barycentre set. Experimental Mathematics, 9(2):309–317, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045952354.

Jeong:2015:OBC

[Jeo15] In-Jee Jeong. Outer billiards with contraction: Attracting Cantor sets. Experimental Mathematics, 24(1):53–64, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Jordan:2002:RCH

[JJJ02] C. R. Jordan, D. A. Jordan, and J. H. Jordan. Reversible complex Hénon maps. Experimental Mathematics, 11(3):339– 347, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057777426.

Jagy:1995:SSC

[JK95] William C. Jagy and Irving Kaplansky. Sums of squares, cubes, and higher powers. Experimental Mathematics, 4(3): 169–173, ???? 1995. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621075.

Johnson:2001:FHW

[JK01] Jennifer M. Johnson and János Kollár. Fano hypersurfaces in weighted projective 4-spaces. Experimental Mathematics, 10(1):151–158, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/999188430.

Joswig:2020:SFS

[JLS20] Michael Joswig, Robert Löwe, and Boris Springborn. Secondary fans and secondary polyhedra of punctured Riemann surfaces. Experimental Mathematics, 29(4):426–442, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1477078.

Jacobson:1995:IBR

[JLW95] Michael J. Jacobson, Jr., Richard F. Lukes, and Hugh C. Williams. An investigation of bounds for the regulator of quadratic fields. Experimental Mathematics, 4(3):211–225, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621079.

Jensen:2016:CTC

[JLY16] Anders Jensen, Anton Leykin, and Josephine Yu. Computing tropical curves via homotopy continuation. Experimental Mathematics, 25(1):83–93, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Jezernik:2014:BMG

[JM14] Urban Jezernik and Primoz Moravec. Bogomolov multipliers of groups of order 128. Experimental Mathematics, 23(2):174– 180, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Jackson:2006:PSA

[JN06] Steven Glenn Jackson and Alfred G. Noël. Prehomogeneous spaces associated with nilpotent orbits in simple real Lie algebras E_{6(6)} and E_{6(−26)} and their relative invariants. Experimental Mathematics, 15(4):455–469, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789780.

Johnson:2008:SRE

[Joh08] Tomas Johnson. L^{p} spectral radius estimates for the Lamé system on an infinite sector. Experimental Mathematics, 17(3):333–339, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227121386.

Johnson:2011:QST

[Joh11] Tomas Johnson. A quartic system with twenty-six limit cycles. Experimental Mathematics, 20(3):323–328, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317924424.

Jonsson:2010:MTH

[Jon10] Jakob Jonsson. More torsion in the homology of the matching complex. Experimental Mathematics, 19(3):363–383, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758097.

Jorba:1999:MNC

[Jor99] Ángel Jorba. A methodology for the numerical computation of normal forms, centre manifolds and first integrals of Hamiltonian systems. Experimental Mathematics, 8(2):155– 195, ???? 1999. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047477059.

Joets:1999:CSE

[JR99] Alain Joets and Roland Ribotta. Caustique de la surface ellipsoïdale à trois dimensions. Experimental Mathematics, 8 (1):49–55, ???? 1999. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047477111.

Jurgens:2002:MAM

[JR02] Ulf Jürgens and Gerhard Röhrle. MOP — algorithmic modality analysis for parabolic group actions. Experimental Mathematics, 11(1):57–67, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057860314.

Janardhan:1992:NEF

[JRS92] Prem Janardhan, David Rosenblum, and Robert S. Strichartz. Numerical experiments in Fourier asymptotics of Cantor measures and wavelets. Experimental Mathematics, 1(4):249–273, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048610116.

Jehanne:2003:NVS

[JRS03] Arnaud Jehanne, Xavier-François Roblot, and Jonathan Sands. Numerical verification of the Stark–Chinburg conjuncture for some icosahedral representations. Experimental Mathematics, 12(4):419–432, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1087568018.

Jorgenson:2016:KLF

[JST16] Jay Jorgenson, Lejla Smajlović, and Holger Then. Kronecker’s limit formula, holomorphic modular functions, and qexpansions on certain arithmetic groups. Experimental Mathematics, 25(3):295–320, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Jorgenson:2020:ESI

[JST20] Jay Jorgenson, Lejla Smajlović, and Holger Then. On the evaluation of singular invariants for canonical generators of certain genus one arithmetic groups. Experimental Mathematics, 29(1):1–27, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2017.1422161.

Jungic:2000:MAP

[Jun00] Veselin Jungic. On monochromatic arithmetic progressions having odd step. Experimental Mathematics, 9(3):467–471, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604680.

Jacobson:2000:SFS

[JW00] Michael J. Jacobson and Hugh C. Williams. The size of the fundamental solutions of consecutive Pell equations. Experimental Mathematics, 9(4):631–640, October 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1045759528.

Jorba:2005:SPN

[JZ05] Àngel Jorba and Maorong Zou. A software package for the numerical integration of ODEs by means of high-order Taylor methods. Experimental Mathematics, 14(1):99–117, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1120145574.

Klotz:2024:RWQ

[KA24] Alexander R. Klotz and Caleb J. Anderson. Ropelength and writhe quantization of 12-crossing knots. Experimental Mathematics, 34(1):59–66, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2024.2334378.

Au:2001:CRC

[kAL01] Thomas Kwok keung Au and Xiao-Song Lin. Off-center reflections: Caustics and chaos. Experimental Mathematics, 10(2): 287–302, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/999188638.

Koutsoukou-Argyraki:2022:ITC

[KALP22] Angeliki Koutsoukou-Argyraki, Wenda Li, and Lawrence C. Paulson. Irrationality and transcendence criteria for infinite series in Isabelle/HOL. Experimental Mathematics, 31 (2):401–412, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1980465.

Kapovich:2007:CCW

[Kap07] Ilya Kapovich. Clusters, currents, and Whitehead’s algorithm. Experimental Mathematics, 16(1):67–76, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789802.

Katsurada:2005:SVS

[Kat05] Hidenori Katsurada. Special values of the standard zeta functions for elliptic modular forms. Experimental Mathematics, 14(1):27–45, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1120145568.

Katsurada:2010:ESZ

[Kat10a] Hidenori Katsurada. Exact standard zeta values of Siegel modular forms. Experimental Mathematics, 19(1):65–77, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1268404803.

Katz:2010:LTR

[Kat10b] Nicholas M. Katz. 2, 3, 5, Legendre: ± trace ratios in families of elliptic curves. Experimental Mathematics, 19(3):267–277, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758091.

Katz:2014:EQU

[Kat14] Nicholas M. Katz. Equidistribution questions for universal extensions. Experimental Mathematics, 23(4):452–464, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Katz:2019:ESR

[Kat19] Nicholas M. Katz. Exponential sums, Ree groups and Suzuki groups: Conjectures. Experimental Mathematics, 28(1):49– 56, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2017.1334246.

Katz:2021:QZ

[Kat21] Nicholas M. Katz. On a question of Zannier. Experimental Mathematics, 30(3):422–428, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1551818.

Kawahira:2018:RHH

[Kaw18] Tomoki Kawahira. The Riemann hypothesis and holomorphic index in complex dynamics. Experimental Mathematics, 27(1):37–46, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1217443.

Kim:2017:PNA

[KB17] Daeyeoul Kim and Abdelmejid Bayad. Polygon numbers associated with the sum of odd divisors function. Experimental Mathematics, 26(3):287–297, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1162231.

Kapovich:2019:DFI

[KB19] Ilya Kapovich and Mark C. Bell. Detecting fully irreducible automorphisms: A polynomial time algorithm. Experimental Mathematics, 28(1):24–38, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1326326.

Kopczynski:2022:RTV

[KCK22] Eryk Kopczyński and Dorota Celińska-Kopczyńska. Realtime visualization in anisotropic geometries. Experimental Mathematics, 31(4):1177–1196, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2050324.

Kamphorst:2005:FBC

[KdC05] Sylvie Oliffson Kamphorst and Sônia Pinto de Carvalho. The first Birkhoff coefficient and the stability of 2-periodic orbits on billiards. Experimental Mathematics, 14(3):299–306, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1128371755.

Kjos-Hanssen:2019:FPF

[KH19] Bjørn Kjos-Hanssen. Few paths, fewer words: Model selection with automatic structure functions. Experimental Mathematics, 28(1):121–127, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1368048.

Kudo:2022:STC

[KH22] Momonari Kudo and Shushi Harashita. Superspecial trigonal curves of genus 5. Experimental Mathematics, 31(3):908–919, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1723745.

Khovanov:2003:PKC

[Kho03] Mikhail Khovanov. Patters in knot cohomology, I. Experimental Mathematics, 12(3):365–374, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1087329238.

Kida:2004:PTU

[Kid04] Masanari Kida. Primality tests using algebraic groups. Experimental Mathematics, 13(4):421–427, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1109106433.

Kiers:2021:SCS

[Kie21] Joshua Kiers. On the saturation conjecture for Spin(2n). Experimental Mathematics, 30(2):258–267, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1537866.

Kim:2022:PRF

[Kim22] Jaeyoon Kim. Prime running functions. Experimental Mathematics, 31(4):1291–1313, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2020.1786863.

Kirschmer:2011:FSM

[Kir11] Markus Kirschmer. Finite symplectic matrix groups. Experimental Mathematics, 20(2):217–228, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317924412.

Kim:1993:TKE

[KK93] Denise Kim and Rob Kusner. Torus knots extremizing the Möbius energy. Experimental Mathematics, 2(1):1–9, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062620733.

Krauskopf:1997:HCL

[KK97] Bernd Krauskopf and Hartje Kriete. Hausdorff convergence and the limit shape of the unicorns. Experimental Mathematics, 6(2):117–135, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047649999.

Kuchment:1999:SPH

[KK99] Peter Kuchment and Leonid A. Kunyansky. Spectral properties of high contrast band-gap materials and operators on graphs. Experimental Mathematics, 8(1):1–28, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047477109.

Kaid:2012:SVB

[KK12] Almar Kaid and Ralf Kasprowitz. Semistable vector bundles and tannaka duality from a computational point of view. Experimental Mathematics, 21(2):171–188, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Kemper:2001:DIR

[KKM^{+}01] Gregor Kemper, Elmar Körding, Gunter Malle, B. Heinrich Matzat, Denis Vogel, and Gabor Wiese. A database of invariant rings. Experimental Mathematics, 10(4):537–542, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069855253.

King:2005:GDM

[KKM05] Henry King, Kevin Knudson, and Neža Mramor. Generating discrete Morse functions from point data. Experimental Mathematics, 14(4):435–444, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1136926974.

Kaihnsa:2019:SFC

[KKP^{+}19] Nidhi Kaihnsa, Mario Kummer, Daniel Plaumann, Mahsa Sayyary Namin, and Bernd Sturmfels. Sixty-four curves of degree six. Experimental Mathematics, 28(2):132–150, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1360808.

Katz:2016:BQO

[KKSV16] Karin Katz, Mikhail Katz, Michael M. Schein, and Uzi Vishne. Bolza quaternion order and asymptotics of systoles along congruence subgroups. Experimental Mathematics, 25 (4):399–415, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Kin:2009:EVV

[KKT09] E. Kin, S. Koijima, and M. Takasawa. Entropy versus volume for pseudo-Anosovs. Experimental Mathematics, 18(4):397– 407, ???? 2009. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1259158505.

Kuan:2016:CDZ

[KL16] Yen-Liang Kuan and Yi-Hsuan Lin. Criterion for deciding zeta-like multizeta values in positive characteristic. Experimental Mathematics, 25(3):246–256, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Kontorovich:2021:TOA

[KL21] Alex Kontorovich and Jeffrey Lagarias. On toric orbits in the affine sieve. Experimental Mathematics, 30(4):575–586, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1616337.

Kahle:2020:CLH

[KLNP20] Matthew Kahle, Frank H. Lutz, Andrew Newman, and Kyle Parsons. Cohen–Lenstra heuristics for torsion in homology of random complexes. Experimental Mathematics, 29 (3):347–359, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1473821.

Kluners:2002:AFF

[Klü02] Jürgen Klüners. Algorithms for function fields. Experimental Mathematics, 11(2):171–181, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062621213.

Kolpakov:2018:CDK

[KM18] Alexander Kolpakov and Jun Murakami. Combinatorial decompositions, Kirillov–Reshetikhin invariants, and the volume conjecture for hyperbolic polyhedra. Experimental Mathematics, 27(2):193–207, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1242441.

Kogoshi:2020:CNC

[KM20] Yuta Kogoshi and Takayuki Morisawa. On the class numbers in the cyclotomic Z_{29} and Z_{31}-extensions of the field of rationals. Experimental Mathematics, 29(4):470–474, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1482481.

Kaluba:2024:SGC

[KMN24] Marek Kaluba, Piotr Mizerka, and Piotr W. Nowak. Spectral gap for the cohomological Laplacian of SL_{3}(Z). Experimental Mathematics, 34(1):53–58, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2024.2333722.

Kostic:2022:NCP

[KMP22] Aleksandra Kostić, Nela Milosević, and Zoran Z. Petrović. Note on the cyclotomic polynomial topologically. Experimental Mathematics, 31(2):669–675, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1680464.

Kilian:2000:NCM

[KMS00] Martin Kilian, Ian McIntosh, and Nicholas Schmitt. New constant mean curvature surfaces. Experimental Mathematics, 9(4):595–611, October 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045759525.

Kirschmer:2022:BHL

[KN22] Markus Kirschmer and Gabriele Nebe. Binary Hermitian lattices over number fields. Experimental Mathematics, 31 (1):280–301, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1618756.

Knutson:2001:DCS

[Knu01] Allen Knutson. Descent-cycling in Schubert calculus. Experimental Mathematics, 10(3):345–354, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1069786343.

Kako:1995:CCG

[KO95] Fujio Kako and Mitsuyuki Ochiai. Computational construction of W -graphs of Hecke algebras H(q, n) for n up to 15. Experimental Mathematics, 4(1):61–67, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1062621143.

Kobayashi:1994:NAT

[Kob94] Ryo Kobayashi. A numerical approach to three-dimensional dendritic solidification. Experimental Mathematics, 3(1):59– 81, ???? 1994. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621004.

Koblitz:2001:APG

[Kob01] Neal Koblitz. Almost primality of group orders of elliptic curves defined over small finite fields. Experimental Mathematics, 10(4):553–558, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069855255.

Kohli:2018:LGI

[Koh18] Ben-Michael Kohli. The Links–Gould invariant as a classical generalization of the Alexander polynomial? Experimental Mathematics, 27(3):251–264, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1255860.

Kontorovich:2022:FSI

[Kon22] Alex Kontorovich. Foreword to: Special issue on interactive theorem provers. Experimental Mathematics, 31(2):347–348, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2088982.

Koutsianas:2019:CAE

[Kou19] Angelos Koutsianas. Computing all elliptic curves over an arbitrary number field with prescribed primes of bad reduction. Experimental Mathematics, 28(1):1–15, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1325791.

Kouril:2008:VWN

[KP08] Michal Kouril and Jerome L. Paul. The van der Waerden number W (2, 6) is 1132. Experimental Mathematics, 17(1): 53–61, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227031896.

Kanade:2015:ISN

[KR15] Shashank Kanade and Matthew C. Russell. IdentityFinder and some new identities of Rogers–Ramanujan type. Experimental Mathematics, 24(4):419–423, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Krauskopf:1994:BSR

[Kra94] Bernd Krauskopf. The bifurcation set for the 1: 4 resonance problem. Experimental Mathematics, 3(2):107–128, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062620905.

Kraus:1998:EFE

[Kra98] Alain Kraus. Sur l’équation a^{3} + b^{3} = c^{p}. (French) [On the equation a^{3}+b^{3} = c^{p} ]. Experimental Mathematics, 7(1):1–13, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047674269.

Kresch:2010:CCC

[Kre10] Andrew Kresch. CW complexes for complex algebraic surfaces. Experimental Mathematics, 19(4):413–419, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758101.

Kulesz:2001:ECH

[KS01] Leopoldo Kulesz and Colin Stahlke. Elliptic curves of high rank with nontrivial torsion group over Q. Experimental Mathematics, 10(3):475–480, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069786352.

Katanforoush:2003:DPS

[KS03] Ali Katanforoush and Mehrdad Shahshahani. Distributing points on the sphere, I. Experimental Mathematics, 12(2): 199–209, ???? 2003. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1067634731.

Kitano:2005:POK

[KS05] Teruaki Kitano and Masaaki Suzuki. A partial order in the knot table. Experimental Mathematics, 14(4):385–390, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1136926969.

Kurokawa:2008:EPS

[KS08] Nobushige Kurokawa and Takakazu Satoh. Euclid prime sequences over unique. Experimental Mathematics, 17(2):145– 152, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227118967.

Kolesnikov:2022:SIG

[KS22] P. S. Kolesnikov and B. K. Sartayev. On the special identities of Gelfand–Dorfman algebras. Experimental Mathematics, 33 (1):165–174, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2041134.

Karowski:1997:ITM

[KSV97] Michael Karowski, Robert Schrader, and Elmar Vogt. Invariants of three-manifolds, unitary representations of the mapping class group, and numerical calculations. Experimental Mathematics, 6(4):317–352, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047047192.

Konovalov:2021:SBT

[KSV21] Alexander Konovalov, Agata Smoktunowicz, and Leandro Vendramin. On skew braces and their ideals. Experimental Mathematics, 30(1):95–104, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1492476. See erratum [KSV22].

Konovalov:2022:EPS

[KSV22] A. Konovalov, A. Smoktunowicz, and L. Vendramin. Erratum to the paper “On Skew Braces and Their Ideals”. Experimental Mathematics, 31(1):346, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980466. See [KSV21].

Kigami:2001:CLD

[KSW01] Jun Kigami, Robert S. Strichartz, and Katharine C. Walker. Constructing a Laplacian on the diamond fractal. Experimental Mathematics, 10(3):437–448, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1069786349.

Komori:2006:DBE

[KSWY06] Yohei Komori, Toshiyuki Sugawa, Masaaki Wada, and Yasushi Yamashita. Drawing Bers embeddings of the Teichmüller space of once-punctured tori. Experimental Mathematics, 15(1):51–60, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1150476903.

Khan:2008:CCG

[KSY08] Mizan Khan, Igor E. Shparlinski, and Christian L. Yankov. On the convex closure of the graph of modular inversions. Experimental Mathematics, 17(1):91–104, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1227031900.

Katsurada:2016:CPK

[KT16] Hidenori Katsurada and Sho Takemori. Congruence primes of the Kim–Ramakrishnan–Shahidi lift. Experimental Mathematics, 25(3):332–346, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Kushnarev:2009:TSG

[Kus09] Sergey Kushnarev. Teichons: Solitonlike geodesics on universal Teichmüller space. Experimental Mathematics, 18(3): 325–336, ???? 2009. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1259158469.

Kangaslampi:2017:HTB

[KV17] Riikka Kangaslampi and Alina Vdovina. Hyperbolic triangular buildings without periodic planes of genus 2. Experimental Mathematics, 26(1):54–61, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Kotnik:2004:OMF

[KvdL04] Tadej Kotnik and Jan van de Lune. On the order of the Mertens function. Experimental Mathematics, 13(4):473–481, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1109106439.

Kilford:2008:FGP

[KW08] L. J. P. Kilford and Gabor Wiese. On the failure of the Gorenstein property for Hecke algebras of prime weight. Experimental Mathematics, 17(1):37–52, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1227031895.

Kundu:2023:HAC

[KW23] Debanjana Kundu and Lawrence C. Washington. Heuristics for anti-cyclotomic Z_{p}-extensions. Experimental Mathematics, 33(4):644–662, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2023.2221866.

Kobayashi:2017:VRP

[KY17] Mahito Kobayashi and Minoru Yamamoto. Views of real projective 3-space by stable maps into the plane. Experimental Mathematics, 26(2):138–152, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1142910.

Kim:2020:CEG

[KY20] Hyun Kyu Kim and Masahito Yamazaki. Comments on exchange graphs in cluster algebras. Experimental Mathematics, 29(1):79–100, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1437849.

Kauers:2008:EPP

[KZ08] Manuel Kauers and Doron Zeilberger. Experiments with a positivity-preserving operator. Experimental Mathematics, 17(3):341–345, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227121387.

Kowalski:2012:CIF

[KZ12] Emmanuel Kowalski and David Zywina. The Chebotarev invariant of a finite group. Experimental Mathematics, 21(1): 38–56, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Lagarias:2001:NAC

[Lag01] Jeffrey C. Lagarias. On the normality of arithmetical constants. Experimental Mathematics, 10(3):355–368, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069786344.

Lampe:2018:APS

[Lam18] Philipp Lampe. On the approximate periodicity of sequences attached to non-crystallographic root systems. Experimental Mathematics, 27(3):265–271, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1255861.

Lamm:2021:SNR

[Lam21] Christoph Lamm. The search for nonsymmetric ribbon knots. Experimental Mathematics, 30(3):349–363, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1540313. See correction [Ano23].

Lanford:1998:IRO

[Lan98] Oscar E. Lanford III. Informal remarks on the orbit structure of discrete approximations to chaotic maps. Experimental Mathematics, 7(4):317–324, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674149.

Lauret:2021:CSL

[Lau21] Emilio A. Lauret. A computational study on lens spaces isospectral on forms. Experimental Mathematics, 30(2):268– 282, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2018.1538908.

Lawson:2017:MMI

[Law17] John W. Lawson. Minimal mutation-infinite quivers. Experimental Mathematics, 26(3):308–323, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1166353.

Leemans:2004:RWP

[Lee04] Dimitri Leemans. The residually weakly primitive geometries of J_{3}. Experimental Mathematics, 13(4):429–433, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1109106434.

Lee:2023:CMC

[Lee23] Cheolgyu Lee. A construction of multiplicity class of hypersurfaces from Hesselink stratification of a Hilbert scheme. Experimental Mathematics, 32(2):423–430, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1978117.

Lehavi:2022:ERG

[Leh22] David Lehavi. Effective reconstruction of generic genus 5 curves from their theta hyperplanes. Experimental Mathematics, 33(1):150–164, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2022.2041133.

Leprevost:1993:CMR

[Lep93] Franck Leprévost. Courbes modulaires et 11-rang de corps quadratiques. Experimental Mathematics, 2(2):137–146, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048516218.

Lombardo:2023:DJG

[LGRS23] Davide Lombardo, Elisa Lorenzo García, Christophe Ritzenthaler, and Jeroen Sijsling. Decomposing Jacobians via Galois covers. Experimental Mathematics, 32(1):218–240, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926008.

Liu:2023:MMF

[LH23] Hang Liu and Qin Hourong. Mahler measure of families of polynomials defining genus 2 and 3 curves. Experimental Mathematics, 32(2):321–336, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1926014.

Linner:2004:PGG

[Lin04] Anders Linnér. Periodic geodesics generator. Experimental Mathematics, 13(2):199–206, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1090350934.

Lipnowski:2021:BHN

[Lip21] Michael Lipnowski. On Bhargava’s heuristics for GL_{2}(F_{p})- number fields and the number of elliptic curves of bounded conductor. Experimental Mathematics, 30(2):221–234, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1537864.

Lewis:1994:ICD

[LL94] Robert H. Lewis and Sal Liriano. Isomorphism classes and derived series of certain almost-free groups. Experimental Mathematics, 3(3):255–258, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048515875.

Lee:2021:CBR

[LL21] Kyu-Hwan Lee and Kyungyong Lee. A correspondence between rigid modules over path algebras and simple curves on Riemann surfaces. Experimental Mathematics, 30(3): 315–331, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1538910.

Lamzouri:2023:SV

[LL23] Youness Lamzouri and Alessandro Languasco. Small values of |L′/L(1, χ)|. Experimental Mathematics, 32(2):362–377, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1927255.

Linton:1995:PDT

[LLS95] Stephen A. Linton, Klaus Lux, and Leonard H. Soicher. The primitive distance-transitive representations of the Fischer groups. Experimental Mathematics, 4(3):235–253, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621081.

Lewiner:2003:TOD

[LLT03] Thomas Lewiner, Hélio Lopes, and Geovan Tavares. Toward optimality in discrete Morse theory. Experimental Mathematics, 12(3):271–286, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1087329231.

Lewis:1997:CSN

[LM97] Robert H. Lewis and Guy D. Moore. Computer search for nilpotent complexes. Experimental Mathematics, 6(3):239– 246, ???? 1997. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047920423.

Lundow:2006:BCF

[LM06] P. H. Lundow and K. Markström. Broken-cycle-free subgraphs and the log-concavity conjecture for chromatic polynomials. Experimental Mathematics, 15(3):343–353, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789763.

Lagarias:2017:PFF

[LM17a] Jeffrey C. Lagarias and Harsh Mehta. Products of Farey fractions. Experimental Mathematics, 26(1):1–21, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Lason:2017:NNV

[LM17b] Michał Lasoń and Mateusz Michałek. Non-normal very ample polytopes — constructions and examples. Experimental Mathematics, 26(2):130–137, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2015.1128370.

Lewark:2019:CSG

[LM19] Lukas Lewark and Duncan McCoy. On calculating the slice genera of 11- and 12-crossing knots. Experimental Mathematics, 28(1):81–94, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1353453.

Lewis:2020:RTF

[LM20] Joel Brewster Lewis and Alejandro H. Morales. Rook theory of the finite general linear group. Experimental Mathematics, 29(3):328–346, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1470045.

Lanneau:2022:NEF

[LM22] Erwan Lanneau and Martin Möller. Non-existence and finiteness results for Teichmüller curves in Prym loci. Experimental Mathematics, 31(2):621–636, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1671920.

Lazarev:2013:DMS

[LMO13] Oleg Lazarev, Steven J. Miller, and Kevin O’Bryant. Distribution of missing sums in sumsets. Experimental Mathematics, 22(2):132–156, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Landsberg:2004:SNO

[LMW04] J. M. Landsberg, Laurent Manivel, and Bruce W. Westbury. Series of nilpotent orbits. Experimental Mathematics, 13(1): 13–30, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1086894087.

Lubeck:2001:ELO

[LN01] Frank Lübeck and Max Neunhöffer. Enumerating large orbits and direct condensation. Experimental Mathematics, 10(2): 197–206, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/999188632.

Lau:2003:DBS

[LN03] Ka-Sing Lau and Sze-Man Ngai. Dimensions of the boundaries of self-similar sets. Experimental Mathematics, 12(1): 13–26, ???? 2003. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1064858781.

Lebeau:2010:ESH

[LN10] Gilles Lebeau and Maëlle Nodet. Experimental study of the HUM control operator for linear waves. Experimental Mathematics, 19(1):93–120, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1268404805.

Lomont:2002:YMP

[Lom02] Chris Lomont. Yet more projective curves over F_{2}. Experimental Mathematics, 11(4):547–554, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1057864663.

Lombard:2020:HHT

[Lom20] John Lombard. Honey from the hives: A theoretical and computational exploration of combinatorial hives. Experimental Mathematics, 29(4):361–382, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1473822.

Lobb:2020:KHC

[LOS20] Andrew Lobb, Patrick Orson, and Dirk Schütz. Khovanov homotopy calculations using flow category calculus. Experimental Mathematics, 29(4):475–500, 2020. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1482805.

Louboutin:1998:CRC

[Lou98] Stéphane Louboutin. Computation of relative class numbers of imaginary abelian number fields. Experimental Mathematics, 7(4):293–303, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047674147.

Lakner:1997:OEP

[LP97] Mitja Lakner and Peter Petek. The one-equator property. Experimental Mathematics, 6(2):109–115, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047649998.

Lackenby:2014:GCB

[LP14] Marc Lackenby and Jessica S. Purcell. Geodesics and compression bodies. Experimental Mathematics, 23(2):218–240, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Leprevost:2011:CCP

[LPU11] Franck Leprévost, Michael Pohst, and Osmanbey Uzunkol. On the computation of class polynomials with “Thetanullwerte” and its applications to the unit group computation. Experimental Mathematics, 20(3):271–281, ???? 2011. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1317924420.

Lafferty:1992:FFA

[LR92] John D. Lafferty and Daniel Rockmore. Fast Fourier analysis for SL_{2} over a finite field and related numerical experiments. Experimental Mathematics, 1(2):115–139, ???? 1992. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048709049.

Landreau:2002:CBN

[LR02] Bernard Landreau and Florent Richard. Le critère de Beurling et Nyman pour l’hypothèse de Riemann: aspects numériques. (French) [The Beurling and Nyman criteria for the Riemann Hypothesis: numerical aspects]. Experimental Mathematics, 11(3):349–360, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1057777427.

Linner:2005:DPG

[LR05] Anders Linnér and Robert Renka. Discrete periodic geodesics in a surface. Experimental Mathematics, 14(2):145–152, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1128100127.

Lehavi:2007:EFA

[LR07] D. Lehavi and C. Ritzenthaler. An explicit formula for the arithmetic-geometric mean in genus 3. Experimental Mathematics, 16(4):421–440, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204836513.

Long:2011:SS

[LR11] D. D. Long and A. W. Reid. Small subgroups of SL(3, Z). Experimental Mathematics, 20(4):412–425, ???? 2011. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1323367155.

Landsberg:2017:GBR

[LR17] J. M. Landsberg and Nicholas Ryder. On the geometry of border rank algorithms for n × 2 by 2 × 2 matrix multiplication. Experimental Mathematics, 26(3):275–286, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1162230.

Lercier:2019:DTF

[LRR^{+}19] Reynald Lercier, Christophe Ritzenthaler, Florent Rovetta, Jeroen Sijsling, and Benjamin Smith. Distributions of traces of Frobenius for smooth plane curves over finite fields. Experimental Mathematics, 28(1):39–48, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1328321.

Lagarias:2002:ES

[LRS02] J. C. Lagarias, E. M. Rains, and N. J. A. Sloane. The EKG sequence. Experimental Mathematics, 11(3):437–446, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1057777433.

Lee:2023:GCI

[LRSST23] Destine Lee, Iris Rosenblum-Sellers, Jakwanul Safin, and Anda Tenie. Graph coverings and (im)primitive homology: Examples of low degree. Experimental Mathematics, 32(2): 313–320, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1926013.

Labbe:2020:ADB

[LRZ20] Jean-Philippe Labbé, Günter Rote, and Günter M. Ziegler. Area difference bounds for dissections of a square into an odd number of triangles. Experimental Mathematics, 29(3): 253–275, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1459961.

Lario:2002:SCH

[LS02] Joan-C. Lario and René Schoof. Some computations with Hecke rings and deformation ring, with an appendix by Amod Agashe and William Stein. Experimental Mathematics, 11(2): 303–311, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621223.

Lux:2003:CHS

[LS03] Klaus M. Lux and Magdolna Szőke. Computing homomorphism spaces between modules over finite dimensional algebras. Experimental Mathematics, 12(1):91–98, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1064858786.

Lagarias:2004:AS

[LS04] J. C. Lagarias and N. J. A. Sloane. Approximate squaring. Experimental Mathematics, 13(1):113–128, ???? 2004. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1086894093.

Lux:2007:CDM

[LS07] Klaus M. Lux and Magdolna Szőke. Computing decompositions of modules over finite-dimensional algebras. Experimental Mathematics, 16(1):1–6, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1175789797.

Leong:2013:CSS

[LS13] Alexander Leong and Jeffrey Shallit. Counting sequences with small discrepancies. Experimental Mathematics, 22(1): 74–84, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Lund:1998:CTS

[LSV98] John-Peter Lund, Robert S. Strichartz, and Jade P. Vinson. Cauchy transforms of self-similar measures. Experimental Mathematics, 7(3):177–190, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674203.

Letherman:1999:PHD

[LSW99] Simon Letherman, Dierk Schleicher, and Reg Wood. The 3n + 1-problem and holomorphic dynamics. Experimental Mathematics, 8(3):241–251, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047262405.

Lee:2003:CFU

[LSY03] Y. Lee, R. Scheidler, and C. Yarrish. Computation of the fundamental units and the regulator of a cyclic cubic function field. Experimental Mathematics, 12(2):211–225, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1067634732.

Levi:2009:BTT

[LT09] Mark Levi and Serge Tabachnikov. On bicycle tire tracks geometry, hatchet planimeter, Menzin’s conjecture, and oscillation of unicycle tracks. Experimental Mathematics, 18(2): 173–186, ???? 2009. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1259158427.

Long:2018:ZDS

[LT18] D. D. Long and M. B. Thistlethwaite. Zariski dense surface subgroups in SL(4, Z). Experimental Mathematics, 27 (1):82–92, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1226992.

Lutowski:2009:SFM

[Lut09] Rafał Lutowski. On symmetry of flat manifolds. Experimental Mathematics, 18(2):201–204, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1259158430.

Lutowski:2013:FOA

[Lut13] Rafal Lutowski. Finite outer automorphism groups of crystallographic groups. Experimental Mathematics, 22(4):456–464, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Loring:2015:ENC

[LV15] Terry A. Loring and Fredy Vides. Estimating norms of commutators. Experimental Mathematics, 24(1):106–122, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Lapidus:2003:CDS

[LvF03] Michel L. Lapidus and Machiel van Frankenhuysen. Complex dimensions of self-similar fractal strings and Diophantine approximation. Experimental Mathematics, 12(1):41–69, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1064858783.

Lee:2022:GNL

[LVL22] Seungjai Lee and Michael Vaughan-Lee. The groups and nilpotent Lie rings of order p^{8} with maximal class. Experimental Mathematics, 33(2):247–253, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2062074.

Lempken:2005:MLS

[LvT05] Wolfgang Lempken and Tran van Trung. On minimal logarithmic signatures of finite groups. Experimental Mathematics, 14(3):257–269, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1128371751.

Lam:2012:MCS

[LW12] Thomas Lam and Lauren Williams. A Markov chain on the symmetric group that is Schubert positive? Experimental Mathematics, 21(2):189–192, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Languasco:2010:CMM

[LZ10] Alessandro Languasco and Alessandro Zaccagnini. Computing the Mertens and Meissel–Mertens constants for sums over arithmetic progressions. Experimental Mathematics, 19(3): 279–284, ???? 2010. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317758092.

Macasieb:2008:DAF

[Mac08] Melissa L. Macasieb. Derived arithmetic Fuchsian groups of genus two. Experimental Mathematics, 17(3):347–369, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227121388.

Malle:2004:DGG

[Mal04] Gunter Malle. On the distribution of Galois groups, II. Experimental Mathematics, 13(2):129–136, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1090350928.

Malle:2010:DCG

[Mal10] Gunter Malle. On the distribution of class groups of number fields. Experimental Mathematics, 19(4):465–474, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758105.

Malkoun:2022:WWE

[Mal22] Joseph Malkoun. Weights, Weyl-equivariant maps and a rank conjecture. Experimental Mathematics, 31(3):843–846, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1712272.

Manneville:2018:FRS

[Man18] Thibault Manneville. Fan realizations for some 2-associahedra. Experimental Mathematics, 27(4):377–394, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1289866.

Martin:1995:CTL

[Mar95] Roland Martin. A counterexample in the theory of local zeta functions. Experimental Mathematics, 4(4):299–305, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047674390.

Martins:2019:SCF

[Mar19] Rogério Martins. The (not so simple!) chain fountain. Experimental Mathematics, 28(4):398–403, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1413454.

Marschner:2020:ROS

[Mar20] Ian C. Marschner. Random ordering of semiprimes. Experimental Mathematics, 29(4):383–397, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1474505.

Martin:2022:RGC

[Mar22] Kimball Martin. Refined Goldbach Conjectures with primes in progressions. Experimental Mathematics, 31(1):226–232, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1596849.

Matveev:1998:CRT

[Mat98] Sergei V. Matveev. Computer recognition of three-manifolds. Experimental Mathematics, 7(2):153–161, ???? 1998. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048515663.

Matthews:2001:AMB

[Mat01] Charles A. Matthews. Approximation of a map between one-dimensional Teichmüller spaces. Experimental Mathematics, 10(2):247–266, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/999188635.

Mayer:2001:SEA

[May01] Uwe F. Mayer. A singular example for the averaged mean curvature flow. Experimental Mathematics, 10(1):103–107, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
999188424.

Mayer:2010:CHB

[May10] Sebastian Mayer. Calculation of Hilbert Borcherds products. Experimental Mathematics, 19(2):243–256, ???? 2010. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1276784793.

Mazur:1992:TRP

[Maz92] Barry Mazur. The topology of rational points. Experimental Mathematics, 1(1):35–45, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048709114.

McNew:2017:MFV

[McN17] Nathan McNew. The most frequent values of the largest prime divisor function. Experimental Mathematics, 26(2): 210–224, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1155188.

Mellit:2022:RPC

[Mel22] Anton Mellit. Rationality proofs by curve counting. Experimental Mathematics, 31(3):773–782, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1691088.

Mercat:2020:RFS

[Mer20] Paul Mercat. Rauzy fractal of the smallest substitution associated with the smallest Pisot number. Experimental Mathematics, 29(2):163–182, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1446063.

Merca:2021:TTS

[Mer21]  Mircea Merca.  Truncated theta series and Rogers–

Ramanujan functions. Experimental Mathematics, 30(3): 364–371, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1542642.

Merca:2022:TTQ

[Mer22] Mircea Merca. On two truncated quintuple series theorems. Experimental Mathematics, 31(2):606–610, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1660742.

Morier-Genoud:2022:DRN

[MGO22] Sophie Morier-Genoud and Valentin Ovsienko. On qdeformed real numbers. Experimental Mathematics, 31(2): 652–660, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1671922.

Milnor:1992:RIC

[Mil92] John Milnor. Remarks on iterated cubic maps. Experimental Mathematics, 1(1):5–24, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048709112.

Milnor:1993:GDQ

[Mil93] John Milnor. Geometry and dynamics of quadratic rational maps, with an appendix by the author and Lei Tan. Experimental Mathematics, 2(1):37–83, ???? 1993. CODEN ????

ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062620736.

Milnor:2000:RMT

[Mil00] John Milnor. On rational maps with two critical points. Experimental Mathematics, 9(4):481–522, October 2000. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1045759519.

Miller:2001:GKF

[Mil01] Sally M. Miller. Geodesic knots in the figure-eight knot complement. Experimental Mathematics, 10(3):419–436, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1069786348.

Milnor:2004:PTJ

[Mil04] John Milnor. Pasting together Julia sets: a worked out example of mating. Experimental Mathematics, 13(1):55–92, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1086894090.

Miller:2006:IZN

[Mil06] Steven J. Miller. Investigations of zeros near the central point of elliptic curve L-functions, with an appendix by Eduardo Dueñez. Experimental Mathematics, 15(3):257–279, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789758.

Maurer:2001:FEE

[MM01] Markus Maurer and Volker Müller. Finding the eigenvalue in Elkies’ algorithm. Experimental Mathematics, 10(2):275–286, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
999188637.

Murakami:2002:KCC

[MMO^{+}02] Hitoshi Murakami, Jun Murakami, Miyuki Okamoto, Toshie Takata, and Yoshiyuki Yokota. Kashaev’s conjecture and the Chern–Simons invariants of knots and links. Experimental Mathematics, 11(3):427–435, ???? 2002. CODEN ????

ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1057777432.

Maley:2000:SPC

[MMT00] F. Miller Maley, Jean Mastrangeli, and Lisa Traynor. Symplectic packings in cotangent bundles of tori. Experimental Mathematics, 9(3):435–455, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1045604678.

McCluskey:2007:PTC

[MMW07] A. E. McCluskey, D. W. McIntyre, and W. S. Watson. The propositional theory of closure. Experimental Mathematics, 16(4):501–512, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204836518.

Muller:2005:SCR

[MN05] Jurgen Müller and Max Neunhöffer. Some computations regarding Foulkes’ conjecture. Experimental Mathematics, 14(3):277–283, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1128371753.

Miller:2014:ADI

[MN14] Jonathan Miller and Ramin Naimi. An algorithm for detecting intrinsically knotted graphs. Experimental Mathematics, 23(1):6–12, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Miller:2008:DLN

[MNS08] Steven J. Miller, Tim Novikoff, and Anthony Sabelli. The distribution of the largest nontrivial eigenvalues in families of random regular graphs. Experimental Mathematics, 17(2): 231–244, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227118974.

Minac:2023:AGF

[MNT23] Ján Minác, Tung T. Nguyen, and Nguy˜ ên Duy Tân. On the arithmetic of generalized Fekete polynomials. Experimental Mathematics, 33(4):723–754, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2023.2293283.

Martin:2007:SSP

[MO07] Greg Martin and Kevin O’Bryant. The symmetric subset problem in continuous Ramsey theory. Experimental Mathematics, 16(2):145–166, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204905872.

Machacek:2022:DDS

[MO22] John Machacek and Nicholas Ovenhouse. Discrete dynamical systems from real valued mutation. Experimental Mathematics, 33(2):261–275, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2065555.

Martin:2020:RKG

[MOY20] Gaven Martin, Graeme O’Brien, and Yasushi Yamashita. Random Kleinian groups, II two parabolic generators. Experimental Mathematics, 29(4):443–451, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2018.1477079.

McGehee:1994:RSF

[MP94] Richard P. McGehee and Bruce B. Peckham. Resonance surfaces for forced oscillators. Experimental Mathematics, 3(3): 221–244, ???? 1994. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048515873.

Martelli:2001:TMH

[MP01] Bruno Martelli and Carlo Petronio. Three-manifolds having complexity at most 9. Experimental Mathematics, 10(2):207– 236, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/999188633.

Miyata:2015:ENP

[MP15] Hiroyuki Miyata and Arnau Padrol. Enumerating neighborly polytopes and oriented matroids. Experimental Mathematics, 24(4):489–505, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Mixon:2022:OPE

[MP22] Dustin G. Mixon and Hans Parshall. The optimal packing of eight points in the real projective plane. Experimental Mathematics, 31(2):474–485, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1641767.

McShane:1994:DLS

[MPR94] Greg McShane, John R. Parker, and Ian Redfern. Drawing limit sets of Kleinian groups using finite state automata. Experimental Mathematics, 3(2):153–170, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1062620909.

Mukherjee:2018:STK

[MPS^{+}18] Sujoy Mukherjee, Józef H. Przytycki, Marithania Silvero, Xiao Wang, and Seung Yeop Yang. Search for torsion in Khovanov homology. Experimental Mathematics, 27(4):488–497, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1320242.

Moseley:2017:OTA

[MPY17] Daniel Moseley, Nicholas Proudfoot, and Ben Young. The Orlik–Terao algebra and the cohomology of configuration space. Experimental Mathematics, 26(3):373–380, 2017. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1187097.

Mignotte:1995:CEN

[MR95] Maurice Mignotte and Yves Roy. Catalan’s equation has no new solution with either exponent less than 10651. Experimental Mathematics, 4(4):259–268, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047674387.

Matsumoto:2000:RPS

[MR00] Saburo Matsumoto and Richard Rannard. The regular projective solution space of the figure-eight knot complement.

Experimental Mathematics, 9(2):221–234, ???? 2000. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1045952346.

Movasati:2010:PVE

[MR10] Hossein Movasati and Stefan Reiter. Painlevé VI equations with algebraic solutions and family of curves. Experimental Mathematics, 19(2):161–173, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1276784787.

Mazur:2021:ACS

[MR21] Barry Mazur and Karl Rubin. Arithmetic conjectures suggested by the statistical behavior of modular symbols. Experimental Mathematics, 32(4):657–672, 2021. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1982424.

Martin:2016:ESA

[MRRTS16] P. Martín, R. Ramírez-Ros, and A. Tamarit-Sariol. Exponentially small asymptotic formulas for the length spectrum in some billiard tables. Experimental Mathematics, 25(4):416– 440, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Moschetti:2022:CVN

[MRS22] Riccardo Moschetti, Franco Rota, and Luca Schaffler. A computational view on the non-degeneracy invariant for Enriques surfaces. Experimental Mathematics, 33(3):400–421, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2113576.

Mossinghoff:2008:MMM

[MRW08] Michael J. Mossinghoff, Georges Rhin, and Qiang Wu. Minimal Mahler measures. Experimental Mathematics, 17(4):451– 458, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1243429958.

Masson:2001:PEH

[MS01] Gisli Másson and Boris Shapiro. On polynomial eigenfunctions of a hypergeometric-type operator. Experimental Mathematics, 10(4):609–618, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069855260.

McKee:2005:SNP

[MS05] James McKee and Chris Smyth. Salem numbers, Pisot numbers, Mahler measure, and graphs. Experimental Mathematics, 14(2):211–229, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1128100133.

Morrison:2011:GTL

[MS11] Ian Morrison and David Swinarski. Gröbner techniques for low-degree Hilbert stability. Experimental Mathematics, 20 (1):34–56, ???? 2011. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1317924389.

Morita:2015:IEC

[MSS15] Shigeyuki Morita, Takuya Sakasai, and Masaaki Suzuki. Integral Euler characteristic of out F_{11}. Experimental Mathematics, 24(1):93–97, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Morita:2018:AQS

[MSS18] Shigeyuki Morita, Takuya Sakasai, and Masaaki Suzuki. An Abelian quotient of the symplectic derivation Lie algebra of the free Lie algebra. Experimental Mathematics, 27(3): 302–315, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1266975.

Mans:2019:FGQ

[MSSS19] Bernard Mans, Min Sha, Igor E. Shparlinski, and Daniel Sutantyo. On functional graphs of quadratic polynomials. Experimental Mathematics, 28(3):292–300, 2019. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1391725.

Magaard:2003:GPB

[MSV03] Kay Magaard, Sergey Shpectorov, and Helmut Völklein. A GAP package for braid orbit computation and applications.

Experimental Mathematics, 12(4):385–394, ???? 2003. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1087568015.

Melanova:2022:RPS

[MSW22] Hana Melánová, Bernd Sturmfels, and Rosa Winter. Recovery from power sums. Experimental Mathematics, 33(2): 225–234, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2061650.

Mestre:1993:QHG

[MSWZ93] Jean-François Mestre, René Schoof, Lawrence Washington, and Don Zagier. Quotients homophones des groupes libres. Experimental Mathematics, 2(3):153–155, ???? 1993. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1062620828.

Musin:2015:TP

[MT15] Oleg R. Musin and Alexey S. Tarasov. The Tammes problem for N = 14. Experimental Mathematics, 24(4):460–468, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Mann:2016:HNE

[MT16] Casey Mann and B. Charles Thomas. Heesch numbers of edge-marked polyforms. Experimental Mathematics, 25(3): 281–294, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

McGuire:2014:TNC

[MTC14] Gary McGuire, Bastian Tugemann, and Gilles Civario. There is no 16-clue Sudoku: Solving the Sudoku minimum number of clues problem via hitting set enumeration. Experimental Mathematics, 23(2):190–217, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Miska:2022:CCF

[MU22] Piotr Miska and Maciej Ulas. On consecutive 1’s in continued fractions expansions of square roots of prime numbers. Experimental Mathematics, 31(1):238–251, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1605319.

Mukamel:2017:VUB

[Muk17] Ronen E. Mukamel. Visualizing the unit ball for the Teichmüller metric. Experimental Mathematics, 26(1):93–97, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Mukamel:2021:PDT

[Muk21] Ronen E. Mukamel. Polynomials defining Teichmüller curves and their factorizations modp. Experimental Mathematics, 30(1):19–31, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1488156.

Mukherjee:2022:OTE

[Muk22] Sujoy Mukherjee. On odd torsion in even Khovanov homology. Experimental Mathematics, 31(3):1039–1045, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1743215.

Muller:2006:SLE

[Mül06] Tom Müller. Searching for large elite primes. Experimental Mathematics, 15(2):183–186, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1175789738.

Murray:1993:IMR

[Mur93] Michael K. Murray. The information metric on rational maps. Experimental Mathematics, 2(4):271–279, ???? 1993. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1048516037.

Martin:2000:MOS

[MV00] Florian Martin and Alain Valette. Markov operators on the solvable Baumslag–Solitar groups. Experimental Mathematics, 9(2):291–300, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045952352.

Mittelmann:2010:HAS

[MV10] Hans D. Mittelmann and Frank Vallentin. High-accuracy semidefinite programming bounds for kissing numbers. Experimental Mathematics, 19(2):174–178, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1276784788.

Muraz:2005:LBD

[MVG05] G. Muraz and J.-L. Verger-Gaugry. On lower bounds of the density of Delone sets and holes in sequences of sphere packings. Experimental Mathematics, 14(1):47–57, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1120145569.

Mohamed:2022:ICP

[MW22] Esamaldeen Mohamed and Gerald Williams. An investigation into the cyclically presented groups with length three positive relators. Experimental Mathematics, 31(2):537–551, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1655817.

Naoe:2021:CLS

[Nao21] Hironobu Naoe. Corks with large shadow-complexity and exotic four-manifolds. Experimental Mathematics, 30(2): 157–171, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1514332.

Nebe:1996:FS

[Neb96] Gabriele Nebe. Finite subgroups of GL_{24}(Q). Experimental Mathematics, 5(3):163–195, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047915100.

Nebe:2000:IOG

[Neb00] Gabriele Nebe. Invariants of orthogonal G-modules from the character table. Experimental Mathematics, 9(4):623– 629, October 2000. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1045759527.

Neuberger:2006:NEP

[Neu06] John M. Neuberger. Nonlinear elliptic partial difference equations on graphs. Experimental Mathematics, 15(1):91–107, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1150476907.

Newkirk:2016:BB

[New16] Edward Newkirk. Billiards with bombs. Experimental Mathematics, 25(2):194–212, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Nichols:2018:APP

[Nic18] Daniel Nichols. Analogues of the 3x + 1 problem in polynomial rings of characteristic 2. Experimental Mathematics, 27 (1):100–110, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1227734.

Nilakantan:2003:NWN

[Nil03] Nandini Nilakantan. Nonexistence of a weakly neighbourly polyhedral map of type 6, 6. Experimental Mathematics, 12(3):257–262, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1087329229.

Nitaj:1993:AFG

[Nit93] Abderrahmane Nitaj. Algorithms for finding good examples for the abc and Szpiro conjectures. Experimental Mathematics, 2(3):223–230, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1062620832.

Nowak:1997:NFR

[Now97] Ivo Nowak. Numerical finiteness results for minimal surfaces in three-dimensional space forms. Experimental Mathematics, 6(4):301–315, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047047191.

Nabutovsky:2023:GNS

[NP23] Alexander Nabutovsky and Fabian Parsch. Geodesic nets: Some examples and open problems. Experimental Mathematics, 32(1):1–25, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1743216.

Neuberger:1999:NCE

[NR99] J. W. Neuberger and R. J. Renka. Numerical calculation of the essential spectrum of a Laplacian. Experimental Mathematics, 8(3):301–308, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047262410.

Neuberger:2000:CPG

[NR00] J. W. Neuberger and R. J. Renka. Critical points of the Ginzburg–Landau functional on multiply-connected domains. Experimental Mathematics, 9(4):523–533, October 2000. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1045759520.

Nechita:2023:SAQ

[NSW23] Ion Nechita, Simon Schmidt, and Moritz Weber. Sinkhorn algorithm for quantum permutation groups. Experimental Mathematics, 32(1):156–168, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1926005.

Noro:2008:CSS

[NSYY08] Masayuki Noro, Takeshi Sasaki, Kotaro Yamada, and Masaaki Yoshida. Confluence of swallowtail singularities of the hyperbolic Schwarz map defined by the hypergeometric differential equation. Experimental Mathematics, 17(2):191– 204, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227118971.

Netzer:2015:KPS

[NT15] Tim Netzer and Andreas Thom. Kazhdan’s property (T) via semidefinite optimization. Experimental Mathematics, 24(3): 371–374, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ngai:2022:OPD

[NTTY22] Sze-Man Ngai, Wei Tang, Anh Tran, and Shuai Yuan. Orthogonal polynomials defined by self-similar measures with overlaps. Experimental Mathematics, 31(3):1026–1038, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1743214.

Nurmela:2000:COC

[Nur00] Kari J. Nurmela. Conjecturally optimal coverings of an equilateral triangle with up to 36 equal circles. Experimental Mathematics, 9(2):241–250, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045952348.

Nickerson:2005:SPS

[NW05] S. J. Nickerson and R. A. Wilson. Semi-presentations for the sporadic simple groups. Experimental Mathematics, 14(3): 359–371, ???? 2005. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1128371760.

Oggier:2019:SGE

[OB19] Frédérique Oggier and Jean-Claude Belfiore. On the secrecy gain of extremal even l-modular lattices. Experimental Mathematics, 28(4):492–508, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1423249.

Ohashi:2008:RHG

[Oha08] Ryo Ohashi. The rational homology group of Out(F_{n}) for n ≤ 6. Experimental Mathematics, 17(2):167–179, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1227118969.

Dokovic:2000:ISD

[– Dok00] Dragomir Ž. – Doković. The irreducible six-dimensional complex representations of Aut(F_{2}) that are nontrivial on F_{2}. Experimental Mathematics, 9(3):457–465, ???? 2000. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1045604679.

Okada:2002:HER

[Oka02] Kaoru Okada. Hecke eigenvalues for real quadratic fields. Experimental Mathematics, 11(3):407–426, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1057777431.

Olajos:2005:PIB

[Ola05] Péter Olajos. Power integral bases in the family of simplest quartic fields. Experimental Mathematics, 14(2):129– 132, ???? 2005. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1128100125.

Olvera:2001:EAR

[Olv01] Arturo Olvera. Estimation of the amplitude of resonance in the general standard map. Experimental Mathematics, 10(3): 401–418, ???? 2001. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1069786347.

Ochiai:2008:BTD

[OM08] Mitsuyuki Ochiai and Noriko Morimura. Base-tangle decompositions of n-string tangles with 1 < n < 10. Experimental Mathematics, 17(1):1–8, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1227031893.

Omar:2001:AFO

[Oma01] Sami Omar. On Artin L-functions for octic quaternion fields. Experimental Mathematics, 10(2):237–246, ???? 2001. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/999188634.

Opgenorth:2001:DCV

[Opg01] Jürgen Opgenorth. Dual cones and the Voronoi algorithm. Experimental Mathematics, 10(4):599–608, ???? 2001. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1069855259.

Odlyzko:1999:JC

[ORW99] Andrew Odlyzko, Michael Rubinstein, and Marek Wolf. Jumping champions. Experimental Mathematics, 8(2):107– 118, ???? 1999. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047477055.

Oren:2014:PWL

[OS14] Idan Oren and Uzy Smilansky. Periodic walks on large regular graphs and random matrix theory. Experimental Mathematics, 23(4):492–498, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

OHara:2015:RRR

[OS15] Kathleen O’Hara and Dennis Stanton. Refinements of the Rogers–Ramanujan identities. Experimental Mathematics, 24 (4):410–418, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Oeding:2016:EFS

[OS16] Luke Oeding and Steven V. Sam. Equations for the fifth secant variety of Segrè products of projective spaces. Experimental Mathematics, 25(1):94–99, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ogievetsky:2022:ECC

[OS22a] Oleg Ogievetsky and Senya Shlosman. Extremal cylinder configurations II: Configuration O_{6}. Experimental Mathematics, 31(2):486–496, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1641768.

Owens:2022:AFR

[OS22b] Brendan Owens and Frank Swenton. An algorithm to find ribbon disks for alternating knots. Experimental Mathematics, 33(3):437–455, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2022.2158968.

Oberlin:2003:SSG

[OSS03] Richard Oberlin, Brian Street, and Robert S. Strichartz. Sampling on the Sierpinski gasket. Experimental Mathematics, 12(4):403–418, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1087568017.

Otsubo:2015:SVJ

[Ots15] Noriyuki Otsubo. On special values of Jacobi-sum Hecke Lfunctions. Experimental Mathematics, 24(2):247–259, 2015.

CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

eSilva:2022:BLM

[OTZK22] Diogo Oliveira e Silva, Christoph Thiele, and Pavel Zorin-Kranich. Band-limited maximizers for a Fourier extension inequality on the circle. Experimental Mathematics, 31(1): 192–198, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1596847.

Oudet:2011:APL

[Oud11] Édouard Oudet. Approximation of partitions of least perimeter by Γ-convergence: Around Kelvin’s conjecture. Experimental Mathematics, 20(3):260–270, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317924419.

Parlak:2021:CTV

[Par21] Anna Parlak. Computation of the taut, the veering and the Teichmüller polynomials. Experimental Mathematics, 33 (1):1–26, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1985656.

Patterson:2003:ADE

[Pat03] S. J. Patterson. The asymptotic distribution of exponential sums, I. Experimental Mathematics, 12(2):135–153, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1067634728.

Patterson:2005:ADE

[Pat05] S. J. Patterson. The asymptotic distribution of exponential sums, II. Experimental Mathematics, 14(1):87–98, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1120145573.

Paule:1996:PCK

[Pau96] Peter Paule. A proof of a conjecture of Knuth. Experimental Mathematics, 5(2):83–89, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047565639.

Pausinger:2017:AHP

[Pau17] Florian Pausinger. An accurate heuristic for a problem of Shparlinski. Experimental Mathematics, 26(1):72–76, 2017. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Petrov:2002:RCB

[PdlL02] Nikola P. Petrov and Rafael de la Llave. Regularity of conjugacies between critical circle maps: an experimental study. Experimental Mathematics, 11(2):219–241, ???? 2002. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1062621217.

Peachey:2006:DBC

[PE06] T. C. Peachey and C. M. Enticott. Determination of the best constant in an inequality of Hardy, Littlewood, and Pólya. Experimental Mathematics, 15(1):43–50, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1150476902.

Peczarski:2019:WBP

[Pec19] Marcin Peczarski. The worst balanced partially ordered setsladders with broken rungs. Experimental Mathematics, 28 (2):181–184, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1368050.

Peters:2004:DEI

[Pet04] Meinhard Peters. The Diophantine equation xy+yz +xz = n and indecomposable binary quadratic forms. Experimental Mathematics, 13(3):273–274, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1103749835.

Pfeiffer:1997:SHC

[Pfe97] Götz Pfeiffer. The subgroups of M_{24}, or how to compute the table of marks of a finite group. Experimental Mathematics, 6(3):247–270, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047920424.

Pfeifle:2021:PPC

[Pfe21] Julian Pfeifle. Positive Plücker tree certificates for nonrealizability. Experimental Mathematics, 33(1):69–85, 2021.

CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1994487.

Pinn:2000:CCC

[Pin00] Klaus Pinn. A chaotic cousin of Conway’s recursive sequence. Experimental Mathematics, 9(1):55–66, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1046889590.

Piontkowski:2007:NCR

[Pio07] Jens Piontkowski. On the number of cusps of rational cuspidal plane curves. Experimental Mathematics, 16(2):251–256, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204905880.

Poonen:2001:CTP

[Poo01] Bjorn Poonen. Computing torsion points on curves. Experimental Mathematics, 10(3):449–466, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1069786350.

Poonen:2006:HBM

[Poo06] Bjorn Poonen. Heuristics for the Brauer–Manin obstruction for curves. Experimental Mathematics, 15(4):415–420, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789777.

Pinkall:1993:CDM

[PP93] Ulrich Pinkall and Konrad Polthier. Computing discrete minimal surfaces and their conjugates. Experimental Mathematics, 2(1):15–36, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1062620735.

Petrera:2009:IHK

[PPS09] Matteo Petrera, Andreas Pfadler, and Yuri B. Suris. On integrability of Hirota–Kimura-type discretizations: Experimental study of the discrete Clebsch system. Experimental Mathematics, 18(2):223–247, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1259158433.

Petrera:2017:CES

[PPSF17] Matteo Petrera, Andreas Pfadler, Yuri B. Suris, and Yuri N. Fedorov. On the construction of elliptic solutions of integrable birational maps. Experimental Mathematics, 26(3):324–341, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1166354.

Perrin-Riou:2003:ACE

[PR03] Bernadette Perrin-Riou. Arithmétique des courbes elliptiques à réduction supersingulière en p. (French) [Arithmetic on elliptic curves with a supersingular reduction in p ]. Experimental Mathematics, 12(2):155–186, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1067634729.

Plesken:2005:CIF

[PR05] W. Plesken and D. Robertz. Constructing invariants for finite groups. Experimental Mathematics, 14(2):175–188, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1128100130.

Plesken:2008:ECS

[PR08] W. Plesken and D. Robertz. Elimination for coefficients of special characteristic polynomials. Experimental Mathematics, 17(4):499–510, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1243429962.

Pasquier:2013:SPB

[PR13] B. Pasquier and N. Ressayre. The saturation property for branching rules — examples. Experimental Mathematics, 22 (3):299–312, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Paulhus:2017:CDJ

[PR17a] Jennifer Paulhus and Anita M. Rojas. Completely decomposable Jacobian varieties in new genera. Experimental Mathematics, 26(4):430–445, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1199981.

Platt:2017:EEA

[PR17b] D. J. Platt and O. Ramaré. Explicit estimates: From Λ(n) in arithmetic progressions to Λ(n)/n. Experimental Mathematics, 26(1):77–92, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Pokora:2021:RPK

[PR21] Piotr Pokora and Joaquim Roé. The 21 reducible polars of Klein’s quartic. Experimental Mathematics, 30(1):1–18, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1488155.

Pralle:2005:H

[Pra05] Harm Pralle. The hyperplanes of DW (5, 2). Experimental Mathematics, 14(3):373–384, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1128371761.

Pomerance:2022:CC

[PRS22] Carl Pomerance and Simon Rubinstein-Salzedo. Cyclotomic coincidences. Experimental Mathematics, 31(2):596–605, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1660741.

Pomerance:1992:RHS

[PS92] Carl Pomerance and J. W. Smith. Reduction of huge, sparse matrices over finite fields via created catastrophes. Experimental Mathematics, 1(2):89–94, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048709047.

Plesken:1996:CRR

[PS96] Wilhelm Plesken and Bernd Souvignier. Constructing rational representations of finite groups. Experimental Mathematics, 5(1):39–47, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047591146.

Plesken:2000:CCG

[PS00] Wilhelm Plesken and Tilman Schulz. Counting crystallographic groups in low dimensions. Experimental Mathematics, 9(3):407–411, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045604675.

Pacetti:2008:CCV

[PT08] Ariel Pacetti and Gonzalo Tornaría. Computing central values of twisted L-series: The case of composite levels. Experimental Mathematics, 17(4):459–472, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1243429959.

Pulaj:2023:CSU

[Pul23] Jonad Pulaj. Characterizing 3-sets in union-closed families. Experimental Mathematics, 32(2):350–361, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1927254.

Purslow:2018:REA

[Pur18] Thomas Purslow. The restricted Ermolaev algebra and F_{4}. Experimental Mathematics, 27(3):272–276, 2018. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1256005.

Pinner:2001:IHR

[PW01] Christopher G. Pinner and Dan Wolczuk. On the inhomogeneous Hall’s ray of period-one quadratics. Experimental Mathematics, 10(4):487–496, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069855248.

Pfeifle:2004:MUB

[PZ04] Julian Pfeifle and Günter M. Ziegler. On the monotone upper bound problem. Experimental Mathematics, 13(1):1–12, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1086894086.

Porter:2009:FPU

[PZHC09] Mason A. Porter, Norman J. Zabusky, Bambi Hu, and David K. Campbell. Fermi, Pasta, Ulam and the birth of experimental mathematics: a numerical experiment that Enrico Fermi, John Pasta, and Stanisław Ulam reported 54 years ago continues to inspire discovery. American Scientist, 97(3):214–221, May/June 2009. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.

americanscientist.org/issues/feature/2009/3/fermi-
pasta-ulam-and-the-birth-of-experimental-mathematics.

The work referred to in the title is [FPU55].

Quattrini:2006:DAV

[Qua06] Patricia L. Quattrini. On the distribution of analytic \sqrt{|[Cyrillic − sha]|} values on quadratic twists of elliptic curves. Experimental Mathematics, 15(3):355–365, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1175789764.

Rahimipour:2018:EML

[RAG18] A. R. Rahimipour, A. R. Ashrafi, and A. Gholami. The existence of minimal logarithmic signatures for some finite simple groups. Experimental Mathematics, 27(2):138–146, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1235997.

Rahimipour:2021:EFS

[Rah21] Ali Reza Rahimipour. Exact factorizations of sporadic simple groups. Experimental Mathematics, 30(4):441–446, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1556137.

Rannard:1999:CIN

[Ran99] Richard Rannard. Computing immersed normal surfaces in the figure-eight knot complement. Experimental Mathematics, 8(1):73–84, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047477114.

Rawdon:2003:CCD

[Raw03] Eric J. Rawdon. Can computers discover ideal knots? Experimental Mathematics, 12(3):287–302, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1087329232.

Robertson:2021:DMC

[RCD21] Aaron Robertson, William Cipolli, and Maria Dascalu. On the distribution of monochromatic complete subgraphs and arithmetic progressions. Experimental Mathematics, 30(1): 135–145, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1507851.

Roney-Dougal:2004:CSG

[RD04] Colva M. Roney-Dougal. Conjugacy of subgroups of the general linear group. Experimental Mathematics, 13(2):151–164, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1090350930.

Reiter:1996:ASC

[Rei96] Clifford A. Reiter. Attractors with the symmetry of the n-cube. Experimental Mathematics, 5(4):327–336, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047565450.

Rimlinger:1992:BTN

[Rim92] Frank Rimlinger. R-trees and normalization of pseudogroups. Experimental Mathematics, 1(2):95–114, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1048709048.

Ritter:2007:HRF

[Rit07] Gordon Ritter. A Hardy–Ramanujan formula for Lie algebras. Experimental Mathematics, 16(3):375–384, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1204928537.

Rivoal:2004:SGK

[Riv04] T. Rivoal. Simultaneous generation of Koecher and Almkvist-Granville’s Apéry-like formulae. Experimental Mathematics, 13(4):503–508, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1109106442.

Rivoal:2010:EPS

[Riv10] Tanguy Rivoal. Extremality properties of some Diophantine series. Experimental Mathematics, 19(4):481–494, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1317758107.

Rivin:2016:SE

[Riv16] Igor Rivin. Spectral experiments+. Experimental Mathematics, 25(4):379–388, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Roblot:2000:SCH

[Rob00] Xavier-François Roblot. Stark’s conjectures and Hilbert’s twelfth problem. Experimental Mathematics, 9(2):251–260, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045952349.

Roeder:2007:CHP

[Roe07] Roland K. W. Roeder. Constructing hyperbolic polyhedra using Newton’s method. Experimental Mathematics, 16(4): 463–492, ???? 2007. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1204836516.

Rogers:2000:RCC

[Rog00] Nicholas F. Rogers. Rank computations for the congruent number elliptic curves. Experimental Mathematics, 9(4):591– 594, October 2000. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1045759524.

Romik:2018:DEE

[Rom18] Dan Romik. Differential equations and exact solutions in the moving sofa problem. Experimental Mathematics, 27 (3):316–330, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1270858.

Roos:2008:HLA

[Roo08] Jan-Erik Roos. The homotopy Lie algebra of a complex hyperplane arrangement is not necessarily finitely presented.

Experimental Mathematics, 17(2):129–143, ???? 2008. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1227118966.

Rostand:1997:CLC

[Ros97] Jérémie Rostand. Computing logarithmic capacity with linear programming. Experimental Mathematics, 6(3):221–238, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047920422.

Rose:2000:SEC

[Ros00] Harvey E. Rose. On some elliptic curves with large sha. Experimental Mathematics, 9(1):85–89, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1046889593.

Rossman:2005:FBP

[Ros05] Wayne Rossman. The first bifurcation point for Delaunay nodoids. Experimental Mathematics, 14(3):331–342, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1128371758.

Reid:2021:IMK

[RR21] A. W. Reid and N. Rouse. Infinitely many knots with NonIntegral trace. Experimental Mathematics, 32(3):514–526, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980461.

Rubinstein:1994:CB

[RS94] Michael Rubinstein and Peter Sarnak. Chebyshev’s bias. Experimental Mathematics, 3(3):173–197, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1048515870.

Rossman:1998:CMC

[RS98] Wayne Rossman and Katsunori Sato. Constant mean curvature surfaces with two ends in hyperbolic space. Experimental Mathematics, 7(2):101–119, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1048515658.

Rubin:2000:REC

[RS00] Karl Rubin and Alice Silverberg. Ranks of elliptic curves in families of quadratic twists. Experimental Mathematics, 9(4):583–590, October 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045759523.

Rubin:2001:RFQ

[RS01] Karl Rubin and Alice Silverberg. Rank frequencies for quadratic twists of elliptic curves. Experimental Mathematics, 10(4):559–570, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1069855256.

Rubinstein:2005:VRF

[RS05] J. Hyam Rubinstein and Robert Sinclair. Visualizing Ricci flow of manifolds of revolution. Experimental Mathematics, 14(3):285–298, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1128371754.

Rivin:2019:QCR

[RS19] Igor Rivin and Naser T. Sardari. Quantum chaos on random Cayley graphs of SL_{2}[Z/pZ]. Experimental Mathematics, 28 (3):328–341, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1403982.

Ramadas:2022:EIC

[RS22] Rohini Ramadas and Rob Silversmith. Equations at infinity for critical-orbit-relation families of rational maps. Experimental Mathematics, 33(3):379–399, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2022.2113575.

Rhin:2003:NMP

[RSÉ03] G. Rhin and J.-M. Sac-Épée. New methods providing high degree polynomials with small Mahler measure. Experimental Mathematics, 12(4):457–462, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1087568021.

Richey:2018:EBA

[RSS18] Joseph Richey, Noah Shutty, and Matthew Stover. Explicit bounds from the Alon–Boppana theorem. Experimental Mathematics, 27(4):444–453, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1311813.

Ruffo:2006:ECR

[RSSS06] Jim Ruffo, Yuval Sivan, Evgenia Soprunova, and Frank Sottile. Experimentation and conjectures in the real Schubert calculus for flag manifolds. Experimental Mathematics, 15(2): 199–221, ???? 2006. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1175789741.

Ren:2013:UKT

[RSSS13] Qingchun Ren, Steven V. Sam, Gus Schrader, and Bernd Sturmfels. The universal Kummer threefold. Experimental Mathematics, 22(3):327–362, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ratcliffe:2000:VSH

[RT00] John G. Ratcliffe and Steven T. Tschantz. The volume spectrum of hyperbolic 4-manifolds. Experimental Mathematics, 9(1):101–125, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1046889595.

Roure:2016:TIC

[RT16] Eduard Roure and Artur Travesa. Two independent checkings of the weak Goldbach conjecture up to 10^{27}. Experimental Mathematics, 25(1):79–82, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Ratcliffe:2023:HCM

[RT23] John G. Ratcliffe and Steven T. Tschantz. Hyperbolic 24-cell 4-manifolds with one cusp. Experimental Mathematics, 32 (2):269–279, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1926010.

Rossman:2000:EDP

[RTW00] Wayne Rossman, Edward C. Thayer, and Meinhard Wohlgemuth. Embedded, doubly periodic minimal surfaces. Experimental Mathematics, 9(2):197–219, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1045952345.

Rugh:2021:SVF

[RTY21] Hans Henrik Rugh, Lei Tan, and Fei Yang. Schwarzian versus a family of moving parabolic points. Experimental Mathematics, 33(1):136–149, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.2011807.

Rubinstein:2013:ECH

[Rub13] Michael O. Rubinstein. Elliptic curves of high rank and the Riemann zeta function on the one line. Experimental Mathematics, 22(4):465–480, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Rennert:1999:CRA

[RV99] Nicolas Rennert and Annick Valibouze. Calcul de résolvantes avec les modules de Cauchy. (French) [Computing resolvents with Cauchy modules]. Experimental Mathematics, 8(4):351– 366, ???? 1999. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047262357.

Ryba:1994:MGH

[RW94] Alexander J. E. Ryba and Robert A. Wilson. Matrix generators for the Harada–Norton group. Experimental Mathematics, 3(2):137–145, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1062620907.

Roblot:2011:NET

[RW11] Xavier-François Roblot and Alfred Weiss. Numerical evidence toward a 2-adic equivariant “main conjecture”. Experimental Mathematics, 20(2):169–176, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1317924408.

Renze:2001:GZ

[RWW01] John Renze, Stan Wagon, and Brian Wick. The Gaussian zoo. Experimental Mathematics, 10(2):161–174, ???? 2001. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/999188629.

Sawon:2023:TBH

[Saw23] Justin Sawon. Topological bounds on Hyperkähler manifolds. Experimental Mathematics, 33(3):501–517, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2172630.

Scaramuzza:2009:APE

[Sca09] Anna Scaramuzza. Algorithms for projectivity and extremal classes of a smooth toric variety. Experimental Mathematics, 18(1):71–84, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1243430531.

Schwartz:1992:PM

[Sch92] Richard Schwartz. The pentagram map. Experimental Mathematics, 1(1):71–81, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1048709118.

Schwingel:1999:TPP

[Sch99] Ruth Schwingel. The tensor product of polynomials. Experimental Mathematics, 8(4):395–397, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1047262360.

Schwartz:2001:PMR

[Sch01] Richard Evan Schwartz. The pentagram map is recurrent. Experimental Mathematics, 10(4):519–528, ???? 2001. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1069855251.

Schmithusen:2004:AFV

[Sch04] Gabriela Schmithüsen. An algorithm for finding the veech group of an origami. Experimental Mathematics, 13(4):459– 472, ???? 2004. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1109106438.

Schneider:2005:CBA

[Sch05] Csaba Schneider. A computer-based approach to the classification of nilpotent Lie algebras. Experimental Mathematics, 14(2):153–160, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1128100128.

Schwartz:2006:OTB

[Sch06] Richard Evan Schwartz. Obtuse triangular billiards I: Near the (2, 3, 6) triangle. Experimental Mathematics, 15(2):161– 182, ???? 2006. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1175789737.

Schratzberger:2007:STD

[Sch07] B. Schratzberger. On the singularization of the twodimensional Jacobi–Perron algorithm. Experimental Mathematics, 16(4):441–454, ???? 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1204836514.

Schwartz:2009:OTB

[Sch09] Richard Evan Schwartz. Obtuse triangular billiards II: One hundred degrees worth of periodic trajectories. Experimental Mathematics, 18(2):137–171, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1259158426.

Schwartz:2013:FEC

[Sch13a] Richard Evan Schwartz. The five-electron case of Thomson’s Problem. Experimental Mathematics, 22(2):157–186, 2013. ISSN 1058-6458 (print), 1944-950X (electronic).

Schwartz:2013:PS

[Sch13b] Richard Evan Schwartz. Pentagram spirals. Experimental Mathematics, 22(4):384–405, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Scholl:2019:SIE

[Sch19] Travis Scholl. Super-isolated elliptic curves and Abelian surfaces in cryptography. Experimental Mathematics, 28(4): 385–397, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1412371.

Scholze:2022:LTE

[Sch22a] Peter Scholze. Liquid tensor experiment. Experimental Mathematics, 31(2):349–354, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2021.1926016.

Schwartz:2022:FPM

[Sch22b] Richard Evan Schwartz. The farthest point map on the regular octahedron. Experimental Mathematics, 31(4):1086–1097, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1894269.

Schwartz:2022:FLR

[Sch22c] Richard Evan Schwartz. Four lines and a rectangle. Experimental Mathematics, 31(2):661–668, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1671923.

Stroeker:1994:EDE

[SdW94] Roel J. Stroeker and Benjamin M. M. de Weger. On elliptic Diophantine equations that defy Thue’s method: the case of the ochoa curve. Experimental Mathematics, 3(3):209–220, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048515872.

Semeraro:2023:CGS

[Sem23] Jason Semeraro. A 2-compact group as a spets. Experimental Mathematics, 32(1):140–155, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1926004.

Sengun:2011:ICB

[Şen11] Mehmet Haluk Şengün. On the integral cohomology of Bianchi groups. Experimental Mathematics, 20(4):487–505, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1323367159.

Sevcik:2018:FAP

[Sev18] Carlos Sevcik. Fractal analysis of pi normality. Experimental Mathematics, 27(3):331–343, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2017.1279092.

Sharland:2019:EFD

[Sha19] Ayse Sharland. Examples of finitely determined mapgerms of corank 3 supporting Mond’s μ ≥ τ-type conjecture. Experimental Mathematics, 28(3):257–262, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1385036.

Shanker:2021:GBG

[Sha21] O. Shanker. Good-to-bad Gram point ratio for Riemann zeta function. Experimental Mathematics, 30(1):76–85, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1492474.

Sherman:2009:CPD

[She09] Morgan Sherman. Convergence properties of Donaldson’s T -iterations on the Riemann sphere. Experimental Mathematics, 18(1):117–126, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1243430535.

Shen:2020:KSR

[She20] Bobby Shen. The Kolakoski sequence and related conjectures about orbits. Experimental Mathematics, 29(1):54–65, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1434703.

Shiu:2005:DPA

[Shi05] P. Shiu. A Diophantine property associated with prime twins. Experimental Mathematics, 14(1):1–6, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1120145565.

Shimada:2016:ASS

[Shi16] Ichiro Shimada. Automorphisms of supersingular K3 surfaces and Salem polynomials. Experimental Mathematics, 25(4): 389–398, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Shlyk:2021:NVP

[Shl21] Vladimir A. Shlyk. Number of vertices of the polytope of integer partitions and factorization of the partitioned number. Experimental Mathematics, 32(3):457–466, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980455.

Shparlinski:2016:RSI

[Shp16] Igor E. Shparlinski. Ratios of small integers in multiplicative subgroups of residue rings. Experimental Mathematics, 25 (3):273–280, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Silverman:1995:EUN

[Sil95] Joseph H. Silverman. Exceptional units and numbers of small Mahler measure. Experimental Mathematics, 4(1):69–83, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621144.

Silverio:2017:RPE

[Sil17] Andrew E. Silverio. Reversing palindromic enumeration in rank-two free groups. Experimental Mathematics, 26(3): 364–372, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2016.1180655.

Sinclair:2003:LGS

[Sin03] R. Sinclair. On the last geometric statement of Jacobi. Experimental Mathematics, 12(4):477–486, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1087568023.

Singh:2015:OHG

[Sin15] Sandip Singh. Orthogonal hypergeometric groups with a maximally unipotent monodromy. Experimental Mathematics, 24(4):449–459, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Spreer:2011:CPS

[SK11] Jonathan Spreer and Wolfgang Kühnel. Combinatorial properties of the K^{3} surface: Simplicial blowups and slicings. Experimental Mathematics, 20(2):201–216, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1317924411.

Shishikura:2000:FCR

[SL00] Mitsuhiro Shishikura and Tan Lei. A family of cubic rational maps and matings of cubic polynomials. Experimental Mathematics, 9(1):29–53, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1046889589.

Smith:2000:IRH

[Smi00] Alvy Ray Smith. Infinite regular hexagon sequences on a triangle. Experimental Mathematics, 9(3):397–406, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604674.

Snellman:2005:CPB

[Sne05] Jan Snellman. A conjecture on Poincaré-Betti series of modules of differential operators on a generic hyperplane arrangement. Experimental Mathematics, 14(4):445–456, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1136926975.

Soh:2021:EAA

[Soh21] Célestin Wafo Soh. An evolutionary approach to the automatic classification of automorphisms of lower-dimensional Lie algebras. Experimental Mathematics, 30(2):191–199, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1516582.

Sottile:2000:RSC

[Sot00] Frank Sottile. Real Schubert calculus: polynomial systems and a conjecture of Shapiro and Shapiro. Experimental Mathematics, 9(2):161–182, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045952343.

Schielzeth:2005:RQN

[SP05] Daniel Schielzeth and Michael E. Pohst. On real quadratic number fields suitable for cryptography. Experimental Mathematics, 14(2):189–197, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1128100131.

Silvestri:2023:ABT

[SP23] Stefano Silvestri and Rodrigo A. Pérez. Accessibility of the boundary of the Thurston set. Experimental Mathematics, 32(2):405–422, 2023. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2021.1974984.

Shmerkin:2006:ZPS

[SS06] Pablo Shmerkin and Boris Solomyak. Zeros of {−1, 0, 1} power series and connectedness loci for self-affine sets. Experimental Mathematics, 15(4):499–511, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789784.

Silverman:2011:APA

[SS11] Joseph H. Silverman and Katherine E. Stange. Amicable pairs and aliquot cycles for elliptic curves. Experimental Mathematics, 20(3):329–357, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1317924425.

Sabadini:2003:DCA

[SSS03] Irene Sabadini, Frank Sommen, and Daniele C. Struppa. The Dirac complex on abstract vector variables: Megaforms. Experimental Mathematics, 12(3):351–364, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1087329237.

Stewart:1995:ILC

[ST95] Iain A. Stewart and Ben Thompson. On the intersections of longest cycles in a graph. Experimental Mathematics, 4(1): 41–48, ???? 1995. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621141.

Stroeker:1999:ELM

[ST99] Roel J. Stroeker and Nikos Tzanakis. On the elliptic logarithm method for elliptic diophantine equations: reflections and an improvement. Experimental Mathematics, 8(2):135– 149, ???? 1999. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047477057.

Sinclair:2002:LSC

[ST02a] Robert Sinclair and Minoru Tanaka. Loki: Software for computing cut loci. Experimental Mathematics, 11(1):1–25, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1057860311.

Sinclair:2002:SPT

[ST02b] Robert Sinclair and Minoru Tanaka. The set poles of a twosheeted hyperboloid. Experimental Mathematics, 11(1):27– 36, ???? 2002. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1057860312.

Sumida-Takahashi:2005:IIH

[ST05] Hiroki Sumida-Takahashi. The Iwasawa invariants and the higher K-groups associated to real quadratic fields. Experimental Mathematics, 14(3):307–316, ???? 2005. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1128371756.

Schastnyy:2019:LIB

[ST19a] V. Schastnyy and D. Treschev. On local integrability in billiard dynamics. Experimental Mathematics, 28(3):362–368, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1409145.

Shapiro:2019:AMA

[ST19b] Boris Shapiro and Milos Tater. Asymptotics and monodromy of the algebraic spectrum of quasi-exactly solvable sextic oscillator. Experimental Mathematics, 28(1):16–23, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1325792.

Stojakovic:2021:HMB

[ST21] Milos Stojaković and Nikola Trkulja. Hamiltonian makerbreaker games on small graphs. Experimental Mathematics, 30(4):595–604, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1586599.

Spreer:2022:DTG

[ST22] Jonathan Spreer and Stephan Tillmann. Determining the trisection genus of orientable and non-orientable PL 4-manifolds through triangulations. Experimental Mathematics, 31(3): 897–907, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1723744.

Stankov:2021:RAI

[Sta21] Dragan Stankov. The reciprocal algebraic integers having small house. Experimental Mathematics, 32(4):673–682, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1982425.

Stagliano:2023:ECC

[Sta23] Giovanni Staglianò. Explicit computations with cubic fourfolds, Gushel–Mukai fourfolds, and their associated K3 surfaces. Experimental Mathematics, 33(4):529–546, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2023.2184882.

Stevenhagen:1993:NRQ

[Ste93] Peter Stevenhagen. The number of real quadratic fields having units of negative norm. Experimental Mathematics, 2(2): 121–136, ???? 1993. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048516217.

Stevens:1995:CVD

[Ste95] Jan Stevens. Computing versal deformations. Experimental Mathematics, 4(2):129–144, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047931622.

Steinerberger:2017:HSU

[Ste17] Stefan Steinerberger. A hidden signal in the Ulam sequence. Experimental Mathematics, 26(4):460–467, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1204638.

Steinerberger:2021:FCP

[Ste21] Stefan Steinerberger. Free convolution powers via roots of polynomials. Experimental Mathematics, 32(4):567–572, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1980751.

Stoimenow:2000:RKT

[Sto00] Alexander Stoimenow. Rational knots and a theorem of Kanenobu. Experimental Mathematics, 9(3):473–478, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045604681.

Stopple:2009:QCE

[Sto09] Jeffrey Stopple. The quadratic character experiment. Experimental Mathematics, 18(2):193–200, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1259158429.

Stoll:2017:STL

[Sto17a] Michael Stoll. Simultaneous torsion in the Legendre family. Experimental Mathematics, 26(4):446–459, 2017. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1201443.

Stopple:2017:LPR

[Sto17b] Jeffrey Stopple. Lehmer pairs revisited. Experimental Mathematics, 26(1):45–53, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Strichartz:1995:DSS

[STZ95] Robert S. Strichartz, Arthur Taylor, and Tong Zhang. Densities of self-similar measures on the line. Experimental Mathematics, 4(2):101–128, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047931621.

Svinin:2018:CIG

[Svi18] Andrei K. Svinin. Conjectures involving a generalization of the sums of powers of integers. Experimental Mathematics, 27(4):437–443, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1306815.

Suleiman:1993:CFC

[SW93] Ibrahim A. I. Suleiman and Robert A. Wilson. Construction of the fourfold cover of the Mathieu group M_{22}. Experimental Mathematics, 2(1):11–14, ???? 1993. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1062620734.

Suleiman:1995:SG

[SW95] Ibrahim A. I. Suleiman and Robert A. Wilson. Standard generators for J_{3}. Experimental Mathematics, 4(1):11–18, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621139.

Stein:1999:SME

[SW99] Andreas Stein and Hugh C. Williams. Some methods for evaluating the regulator of a real quadratic function field. Experimental Mathematics, 8(2):119–133, ???? 1999. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://projecteuclid.org/euclid.em/1047477056.

Sarveniazi:2010:ECR

[SW10] Alireza Sarveniazi and Stefan Wiedmann. Explicit construction of the (13,13)-regular hypergraph. Experimental Mathematics, 19(2):237–242, ???? 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1276784792.

Steuding:2012:RZF

[SW12] Jörn Steuding and Elias Wegert. The Riemann zeta function on arithmetic progressions. Experimental Mathematics, 21 (3):235–240, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Shrestha:2022:SST

[SW22a] Sunrose T. Shrestha and Jane Wang. Statistics of square-tiled surfaces: Symmetry and short loops. Experimental Mathematics, 31(4):1314–1331, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2020.1786864.

Strichartz:2022:SPL

[SW22b] Robert S. Strichartz and Samuel C. Wiese. Spectral properties of Laplacians on snowflake domains and filled Julia sets. Experimental Mathematics, 31(3):1014–1025, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1743213.

Sasaki:2001:GSH

[SY01] Takeshi Sasaki and Masaaki Yoshida. A geometric study of the hypergeometric function with imaginary exponents. Experimental Mathematics, 10(3):321–330, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1069786340.

Sasaki:2008:HSM

[SYY08] Takeshi Sasaki, Kotaro Yamada, and Masaaki Yoshida. The hyperbolic Schwarz map for the hypergeometric. Experimental Mathematics, 17(3):269–282, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1227121382.

Schwartz:2004:CTC

[SZ04] Alexander Schwartz and Günter M. Ziegler. Construction techniques for cubical complexes, odd cubical 4-polytopes, and prescribed dual manifolds. Experimental Mathematics, 13(4):385–413, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1109106431.

Sills:2006:DDC

[SZ06] Andrew V. Sills and Doron Zeilberger. Disturbing the Dyson conjecture (in a GOOD way). Experimental Mathematics, 15(2):187–191, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1175789739.

Szekeres:1998:AER

[Sze98] George Szekeres. Abel’s equation and regular growth: variations on a theme by Abel. Experimental Mathematics, 7(2): 85–100, ???? 1998. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1048515657.

Schaeffer:2004:ANP

[SZJ04] Gilles Schaeffer and Paul Zinn-Justin. On the asymptotic number of plane curves and alternating knots. Experimental Mathematics, 13(4):483–493, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1109106440.

Taherkhani:2000:KPM

[Tah00] Feraydoun Taherkhani. The Kazhdan property of the mapping class group of closed surfaces and the first cohomology group of its cofinite subgroups. Experimental Mathematics, 9(2):261–274, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045952350.

Takemori:2014:CRS

[Tak14] Sho Takemori. Congruence relations for Siegel modular forms of weight 47, 71, and 89. Experimental Mathematics, 23(4): 423–428, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Tange:2009:LSG

[Tan09] Motoo Tange. Lens spaces given from L-space homology 3-spheres. Experimental Mathematics, 18(3):285–301, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158466.

Tang:2024:ICM

[Tan24] Dazhao Tang. Internal congruences modulo powers of 5 for partition k-tuples with odd parts distinct. Experimental Mathematics, 34(1):72–86, 2024. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2024.2334381.

Tari:1992:RSF

[Tar92] Farid Tari. Recognition of K-singularities of functions. Experimental Mathematics, 1(3):225–229, ???? 1992. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1048622025.

Terras:1996:SSL

[Ter96] Audrey Terras. Survey of spectra of Laplacians on finite symmetric spaces. Experimental Mathematics, 5(1):15–32, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047591144.

Tertooy:2022:RSL

[Ter22] Sam Tertooy. The Reidemeister spectra of low dimensional almost-crystallographic groups. Experimental Mathematics, 31(2):444–455, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2019.1636426.

Thayer:1995:HGC

[Tha95] Edward C. Thayer. Higher-genus Chen–Gackstatter surfaces and the Weierstrass representation for surfaces of infinite genus. Experimental Mathematics, 4(1):19–39, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621140.

Thaine:2008:CFC

[Tha08] F. Thaine. On the construction of families of cyclic polynomials whose roots are units. Experimental Mathematics, 17(3): 315–331, ???? 2008. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1227121385.

Theobald:2002:CA

[The02] Thorsten Theobald. Computing amoebas. Experimental Mathematics, 11(4):513–526, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1057864661.

Thunberg:1994:FNC

[Thu94] Hans Thunberg. Feigenbaum numbers for certain flattop families. Experimental Mathematics, 3(1):51–57, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1062621003.

Tompaidis:1996:AIS

[Tom96a] Stathis Tompaidis. Approximation of invariant surfaces by periodic orbits in high-dimensional maps: some rigorous results. Experimental Mathematics, 5(3):197–209, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047915101.

Tompaidis:1996:NSI

[Tom96b] Stathis Tompaidis. Numerical study of invariant sets of a quasiperiodic perturbation of a symplectic map. Experimental Mathematics, 5(3):211–230, ???? 1996. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1047915102.

Tosic:2022:CBC

[Tos22] Ognjen Tosić. Confirming Brennan’s Conjecture numerically on a counterexample to Thurston’s K = 2 conjecture. Experimental Mathematics, 33(3):422–436, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2022.2153287.

Traizet:2008:ESE

[Tra08] Martin Traizet. Exploring the space of embedded minimal surfaces of finite total curvature. Experimental Mathematics, 17(2):205–221, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1227118972.

Torquato:2006:NCL

[TS06] S. Torquato and F. H. Stillinger. New conjectural lower bounds on the optimal density of sphere packings. Experimental Mathematics, 15(3):307–331, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1175789761.

Tseng:2022:DMS

[Tse22] Dennis Tseng. Divisors on the moduli space of curves from divisorial conditions on hypersurfaces. Experimental Mathematics, 31(4):1358–1369, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2021.1980749.

Tsukerman:2016:PPT

[Tsu16] E. Tsukerman. Ping-pong and the traveling salesman problem. Experimental Mathematics, 25(1):1–7, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Tsuchioka:2020:BFI

[Tsu20] Shunsuke Tsuchioka. BMR freeness for icosahedral family. Experimental Mathematics, 29(2):234–245, 2020. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1455072.

Tamminen:2017:DBM

[TTS17] J. Tamminen, T. Tarvainen, and S. Siltanen. The D-Bar method for diffuse optical tomography: A computational study. Experimental Mathematics, 26(2):225–240, 2017. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1157775.

Teske:1999:PCC

[TW99] Edlyn Teske and Hugh C. Williams. A problem concerning a character sum. Experimental Mathematics, 8(1):63–72, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1047477113.

Umeda:2009:CMC

[Ume09] Yuhei Umeda. Constant-mean-curvature surfaces with singularities in Minkowski 3-space. Experimental Mathematics, 18(3):311–323, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1259158468.

VanHirtum:2017:DFW

[Van17] Jasper Van Hirtum. On the distribution of Frobenius of weight 2 eigenforms with quadratic coefficient field. Experimental Mathematics, 26(2):165–188, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1142912.

Vardi:1998:PP

[Var98] Ilan Vardi. Prime percolation. Experimental Mathematics, 7(3):275–289, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047674208.

vanBommel:2022:NVB

[vB22] Raymond van Bommel. Numerical verification of the Birch and Swinnerton-Dyer conjecture for hyperelliptic curves of higher genus over Q up to squares. Experimental Mathematics, 31(1):138–145, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1592035.

vanderPoorten:2001:PD

[vdP01] Alfred J. van der Poorten. A powerful determinant. Experimental Mathematics, 10(2):307–320, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/999188640.

Verschelde:2000:NEC

[Ver00] Jan Verschelde. Numerical evidence for a conjecture in real algebraic geometry. Experimental Mathematics, 9(2):183–196, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1045952344.

vanGeemen:1997:HEC

[vGvdKTV97] Bert van Geemen, Wilberd van der Kallen, Jaap Top, and Alain Verberkmoes. Hecke eigenforms in the cohomology of congruence subgroups of SL(3, Z). Experimental Mathematics, 6(2):163–174, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047650002.

Vinson:2001:PSM

[Vin01] Jade Vinson. Partial sums of ζ(1/2) modulo 1. Experimental Mathematics, 10(3):337–344, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:

//projecteuclid.org/euclid.em/1069786342.

Vivaldi:1994:PTR

[Viv94] Franco Vivaldi. Periodicity and transport from roundoff errors. Experimental Mathematics, 3(4):303–315, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1048515812.

Voros:2020:DKL

[Vor20] André Voros. Discretized Keiper/Li approach to the Riemann hypothesis. Experimental Mathematics, 29(4):452–469, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1482480.

Vasco:2003:MLF

[VRS03] María Isabel González Vasco, Martin Rötteler, and Rainer Steinwandt. On minimal length factorizations of finite groups. Experimental Mathematics, 12(1):1–12, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1064858780.

Veerman:2000:DCI

[VS00] J. J. P. Veerman and B. D. Stošić. On the dimensions of certain incommensurably constructed sets. Experimental Mathematics, 9(3):413–423, ???? 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1045604676.

Villegas:1999:CPC

[VV99] Fernando Rodriguez Villegas and José Felipe Voloch. On certain plane curves with many integral points. Experimental Mathematics, 8(1):57–62, ???? 1999. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047477112.

Wada:2006:OAD

[Wad06] Masaaki Wada. OPTi’s algorithm for discreteness determination. Experimental Mathematics, 15(1):61–66, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1150476904.

Wagstaff:2001:PNF

[Wag01] Samuel S. Wagstaff. Prime numbers with a fixed number of one bits or zero bits in their binary representation. Experimental Mathematics, 10(2):267–274, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/999188636.

Waldschmidt:1994:DPR

[Wal94] Michel Waldschmidt. Densité des points rationnels sur un groupe algébrique. (French) [Density of rational points on an algebraic group]. Experimental Mathematics, 3(4):329–352, ???? 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/

1048515814. See errata [Wal95].

Waldschmidt:1995:EDP

[Wal95] Michel Waldschmidt. Errata to: “Densité des point rationnels sur un groupe algébrique”. Experimental Mathematics, 4(3): 255, ???? 1995. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1062621082. See [Wal94].

Waldmuller:2003:FMN

[Wal03] Reinhard Waldmüller. A flat manifold with no symmetries. Experimental Mathematics, 12(1):71–77, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1064858784.

Walker:2013:SCL

[Wal13] Alden Walker. Stable commutator length in free products of cyclic groups. Experimental Mathematics, 22(3):282–298, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Wang:2001:NDS

[Wan01] Yang Wang. On the number of Daubechies scaling functions and a conjecture of chyzak et al. Experimental Mathematics, 10(1):87–89, ???? 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/999188422.

Wang:2009:FIE

[Wan09] Jeffrey Wang. Finding and investigating exact spherical codes. Experimental Mathematics, 18(2):249–256, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158434.

Warrington:2011:ECC

[War11] Gregory S. Warrington. Equivalence classes for the μcoefficient of Kazhdan–Lusztig polynomials in S_{n}. Experimental Mathematics, 20(4):457–466, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1323367157.

Watkins:2002:CMD

[Wat02] Mark Watkins. Computing the modular degree of an elliptic curve. Experimental Mathematics, 11(4):487–502, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1057864659.

Watkins:2008:SHA

[Wat08] Mark Watkins. Some heuristics about elliptic curves. Experimental Mathematics, 17(1):105–125, ???? 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1227031901.

Weber:1997:HSF

[Web97] Hermann-Josef Weber. Hyperelliptic simple factors of J_{0}(N) with dimension at least 3. Experimental Mathematics, 6(4): 273–287, ???? 1997. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1047047188.

Weber:2009:LDC

[Web09] Christian Weber. Low-degree cohomology of integral Specht modules. Experimental Mathematics, 18(1):85–96, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1243430532.

Weintraub:2014:AD

[Wei14] Steven H. Weintraub. The adjoint of differentiation. Experimental Mathematics, 23(4):429–432, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Weinstein:2016:NFP

[Wei16] F. V. Weinstein. Notes on Fibonacci partitions. Experimental Mathematics, 25(4):482–499, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Wikstrom:2003:CPG

[Wik03] Frank Wikström. Computing the pluricomplex Green function with two poles. Experimental Mathematics, 12(3):375– 384, ???? 2003. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1087329239.

Willerton:2002:FTV

[Wil02] Simon Willerton. On the first two Vassiliev invariants. Experimental Mathematics, 11(2):289–296, ???? 2002. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://projecteuclid.org/euclid.em/1062621221.

Wilson:2006:SGC

[Wil06] P. M. H. Wilson. Some geometry and combinatorics for the Sinvariant of ternary cubics. Experimental Mathematics, 15(4):

479–490, ???? 2006. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://projecteuclid.org/
euclid.em/1175789782.

Williams:2023:PNG

[Wil23] James Williams. Powerfully nilpotent groups of class 2. Experimental Mathematics, 32(1):133–139, 2023. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2021.1926003.

Winckler:2022:LRC

[Win22] Björn Winckler. The Lorenz Renormalization Conjecture. Experimental Mathematics, 31(1):42–57, 2022. CO-DEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2019.1583145.

Worden:2020:ESV

[Wor20] William Worden. Experimental statistics of veering triangulations. Experimental Mathematics, 29(1):101–122, 2020. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2018.1437850.

Wilson:1997:BPC

[WPW97] Mark C. Wilson, Geoffrey Pritchard, and David H. Wood. Bell’s primeness criterion for W (2n+1). Experimental Mathematics, 6(1):77–85, ???? 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047565285.

Woodcock:1998:ACR

[WS98] Christopher F. Woodcock and Nigel P. Smart. p-adic chaos and random number generation. Experimental Mathematics, 7(4):333–342, ???? 1998. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1047674151.

Williams:2003:NCC

[WtR03] Hugh Williams and Herman te Riele. New computations concerning the Cohen–Lenstra heuristics. Experimental Mathematics, 12(1):99–113, ???? 2003. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://

projecteuclid.org/euclid.em/1064858787.

Wayne:2021:CPS

[WZ21] C. Eugene Wayne and Vadim Zharnitsky. Critical points of Strichartz functional. Experimental Mathematics, 30(2): 235–257, 2021. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2018.1537865.

Xylogiannopoulos:2014:EAN

[XKA14] Konstantinos F. Xylogiannopoulos, Panagiotis Karampelas, and Reda Alhajj. Experimental analysis on the normality of using advanced data-mining techniques. Experimental Mathematics, 23(2):105–128, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Xu:2004:PLD

[Xu04] Yuan Xu. On polynomials of least deviation from zero in several variables. Experimental Mathematics, 13(1):103–113, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1086894092.

Xia:2022:GHR

[XZ22] Ernest X. W. Xia and Alice X. H. Zhao. Generalizations of Hirschhorn’s results on two remarkable q-series expansions. Experimental Mathematics, 31(3):878–882, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2020.1712565.

Yabuta:2009:PDC

[Yab09] Minoru Yabuta. Primitive divisors of certain elliptic divisibility sequences. Experimental Mathematics, 18(3):303–310, ???? 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://projecteuclid.org/euclid.em/
1259158467.

Yi:2019:TLP

[Yi19] Ren Yi. The triple lattice PETs. Experimental Mathematics, 28(4):456–474, 2019. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2017.1422159.

Yun:2022:ERH

[Yun22] Claudia He Yun. The S_{n}-equivariant rational homology of the tropical moduli spaces Δ_{2,n}. Experimental Mathematics, 31 (4):1345–1357, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10586458.2020.1870179.

Yao:2022:NSS

[YZ22] Yukun Yao and Doron Zeilberger. Numerical and symbolic studies of the peaceable queens problem. Experimental Mathematics, 31(1):269–279, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2019.1616338.

Zhang:2022:CAH

[Zha22] Yongquan Zhang. Construction of acylindrical hyperbolic 3-manifolds with quasifuchsian boundary. Experimental Mathematics, 31(3):883–896, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
10586458.2020.1718566.

Zong:2022:CAD

[Zon22] Chuanming Zong. A computer approach to determine the densest translative tetrahedron packings. Experimental Mathematics, 31(1):21–41, 2022. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10586458.2019.1582378.